A Selected Bibliography of Publications by, and about, Eugene Wigner

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

23 August 2025 Version 1.125

Title word cross-reference

+ [LPN11]. 0 [OW84]. $1 [Duf46]. 1/2 [Lan87, OW84]. $15.00 [Kol60]. 2 [LR83, RI03]. 3 [Bru85, Bru87, Loc76, RY03, SG75, Win76]. 3j [Max10]. $4.00 [Bec58, Edw60]. $49.50 [Sch85]. 6 [Bru85, Bru87, RY03, SG75]. 6j [LY09, Max10]. 6 × 9 [Wig65c]. 6 × 9.25 [Bec58, Edw60]. $7.50 [Pai67a, Wei69a]. $7.75 [Wig65c]. $8 [Wig64f]. 9 [CD79, Ear81, Hol73a, Jan68, Wu72, Wu73]. 9j [Max10, Ros98, Ros99]. ^{∗} [Fro79]. ^{4} [CPCF99]. ^{8} [BW37, WB36]. _{2} [LPN11, SBR^{+}86]. _{3} [DA73]. _{n} [BGL67, Tem07]. A [Wig36a]. β [RW54, WB36, Wig39a]. C^{∗} [BG02, BG03, CIMS08]. C^{2} [Pos00]. D [Mar06a, PFG06, Nou99]. E(2) [Got71]. E = mc^{2} [KN19]. h [CMS77]. h̄ → 0 [Ara95]. ISO_{n} [Wol71]. IU_{n} [Wol72]. j [Bru85, Bru87, CD79, Ear81, Hol73a, Jan68, JF16, Loc76, RY03, SG75, Win76, Wu72, Wu73]. ≤ 80 [CM66]. m [Beh94]. C^{2} [Pos01]. }↕(1|n) [LSV08]. SU (3) [RSdG99]. SU (3) ⊃ U (2) [PD98]. SU (n) [PD98]. μ [BZ74]. N [HW35]. n ≥ 2 [SW67]. O(5) [HST74]. O(5) ⊆ O(3) [CFM78]. O(h̄8) [Sha92]. O(n) [Gou81, Gou86b]. OSp(1, 2) [ZY88]. \overline{SA}(4, R) [LNPC92]. πN → πB_{s} [KAUZ71]. R [DW64, NW71, Szm98, Szm99, Wig52b, Wig64a, Wig64h]. R_{4} [Bie61]. D [Dra86]. σ [AK00]. SO(5) [Hol73b]. SO(n) [Won71]. SO_{n} [Wol71]. SO_{n,1} [Wol71]. Sp(4) [Hol73c, Hol73b]. SU(1, 1) [CLM75]. SU(2) [CLM75, RdGS01]. SU(2) × SU(2) [Que76]. SU(3) [AD84, DLD^{+}21, RdGS01]. SU(3) ⊇ R(3) [HS83]. SU(3) ⊃ SU(2) × U(1) [HB90]. SU(4) [Hol73c, Que76]. SU(6) ⊇ SU(3) ⊗ SU(2) [SS79, Str79]. SU(N) [Kot05]. SU_{2} ⊃ ⋯ ⊃ G′′ ⊃ G′ ⊃ G [Kib76]. SU_{3} [AD73, BB65, Mos62]. su_{q}(1, 1) [Mae91]. su_{q}(2) [Mae91]. U(2Ω) ⊃ U(Ω) ⊃ ⊗SU(2) [Kot07]. U(3) [BL72, HB90]. u(n) [LB71, CCB72, Gou86a, Won76, Won79]. U(n = n_{1} + n_{2}): U(n_{1}) × u(n_{2}) [GP86]. U_{n,1} [Wol72]. U_{q}[gl(n)] [GLB92, Gou92]. U_{q→0}(sl(2)) [MS00]. W_{∞} [DV97].

-Algebras [BG02]. -coefficients [SG75]. -covariance [DV97]. -D [RI03]. -Decay [RW54]. -dimensional [Nou99]. -function [Beh94]. -functions [PFG06]. -Invariant [Tem07]. -matrix [Szm99, DW64, NW71, Szm98, Wig64a, Wig64h]. -Model [AK00]. -Modules [BG02, BG03, CIMS08]. -product [Fro79]. -Radioactivities [Wig39a]. -Ray [WB36]. -space [BZ74]. -stability [CMS77]. -Symbol [LY09]. -symbols [Ros98, Ros99].

İnönü [FB88]. İnönü-Wigner [FB88].

/Abonyi [Wig92f]. /van [Sch85]. /Vigier [Sch85].

0-521-31911-0 [Ble86, Sch85]. 0-521-34013-6 [Hen89]. 0-521-34017-9 [Hen89].

10 [Wig66f]. 11th [BW61, Ano61]. 13 [Wig85d]. 145th [Her79]. 17 [Wig88d]. 17-én [Wig88d]. 1902-1995 [Ano02]. 1930/41 [Fer68, Fer71]. 1930s [Stu79]. 1945 [Dan95]. 1952 [Ano54a, Wig53a]. 1953 [Ano54b]. 1954 [Wig55b]. 1957 [Mac59, Wig57d, Wig67b, Wig67-29, Wig80a]. 1958 [SW59]. 1960s [Mla98]. 1961 [Ano61, Wig64f, Wig64g]. 1963 [Wig64o, Wig64p, Wig64n, Wig67a, Wig78a]. 1964 [Wig64b]. 1966 [Gre69, Wig66d, Wig66f]. 1967 [Wig69s]. 1968 [Sal69, Wig68e]. 1969 [Wig69e]. 1970 [BH71, Kir73, d’E71]. 1971 [Wig71k]. 1972 [Meh73]. 1974 [MR76]. 1975 [Bug77, W^{+}75, Wig75c]. 1976 [WH77a]. 1977 [KPS78, Wig81a]. 1979 [Ano79a, WT79, Wig79j]. 1980 [Vig86]. 1982 [Wig82l]. 1983 [DGW84, GHM^{+}87]. 1984 [Zac84]. 1986 [Gre86]. 1987 [Wig88d]. 1995 [SVW98, SVW00, Tel95]. 1st [GHM^{+}87, Wig82l].

2-billion-year [Ano15]. 2-D [RI05]. 2007 [SM07]. 20th [Meh73]. 22 [Wig74a]. 25 [Ble86]. 29th [Ano79a].

3-j [Lai94]. 30.00 [Hen89]. 33 [BB93]. 39 [Szm99]. 3j [CM66].

41 [Pos01]. 46 [Maj05a]. 4th [Mac59].

50th [Gau94].

6-j [Lai94]. 60th [MF69, Wig69f, Wig96b]. 6j [CM66].

8-reference [PČH02].

9-j [Lai94]. 9j [CM66].

A. [Wig96j]. Abdus [Bel73, Hei74]. Abner [CHS97]. Abonyi [Wig92f, Wig92f]. ábrázolásainak [Wig71n]. ábrázolásairól [Wig71m]. Absolute [Beh91, Ano75, Ano77b, Ano78b, Ano79b, Ano83a, Vil85]. Absorption [CJSW55a, CJW55, CJSW55b, Dir27, WCJS55, WCJS56, Wig57a, Wig57g, Wig65l]. Abstract [NW71, WE39, Wig54e, Wig67l, Wig73b, Wig74e, Wig47b]. abstraction [CGPT21]. Abstracts [Her79]. Academic [Wig65c, Wig69s]. Academy [Gre86, MR76, Wig65c]. accensione [Wig83l]. Acceptable [KPS78]. Accident [SBR^{+}86]. Accomplishments [Wig89b, W^{+}75]. according [WW28]. Accuracy [JCT11]. Accurate [JF16, MH94, DLD^{+}21]. achievements [Woo80]. Achinstein [L.88]. action [Wig32c]. Activation [HW39]. Adalékok [Wig33d]. Adaptive [BGA07, JLC14]. Addison [Vig86]. Addison-Wesley [Vig86]. Address [Wig64b, Wig67m, Wig68b, Wig82l, Wig87b, Wig96a, Wig96d, Wig59b, Wig74c, Wig75c, Wig82e, Wig83e]. Adiabatic [vNW29a]. adiabatischen [vNW29a]. adjoint [Won71]. adjointness [OW92]. admitting [Wig73k, Wig74g]. adódó [Wig64-28]. adolescence [TW79b]. Adrien [Hen89, KW87d, Ram89, Sch88, Stu88, Wig88b]. Advance [Wig69e]. Advanced [TC10, Zic78]. adventurous [ABES12]. Advisors [Wig76c]. Advisory [MWW66]. AEC [Ano52, Wig62b]. Affairs [GR63, Wig69w]. affine [BB92, BB93]. affluence [Wig77i]. Again [Isl22, Wig71f, Wig72e]. Age [Ano15, GR63, Stu18, Wat93, Mar96, Wag98, Wig46f, Wig63f, Wig07, Wag81]. ago [TW79b, Wig62d, Wig76o]. al [Sch85]. Alain [Bac89]. alapján [NW70]. Albert [Wig96g, Ein39, Wig49a, Wig96g, Wig96l, Woo80, Wig79j]. Alberta [Wig59b]. Alfred [Wig66d]. Algebra [Tem07, Agr80, Bec72, Jan14, Kot05, Kot07, MM92, Que76, RR96]. Algebraic [JvNW34, Lai94]. Algebras [BG02, Con72, Wig73j, BW73, CCT08, CLM76, Hud78, Kol73, Maj93, Mez77, WW95]. Alisauskas [Wu73]. alkali [SM05]. alkali-doped [SM05]. allgemeinen [Wig29a]. allows [Ano19]. Almost [BY88, FKS97, Tau61b, vNW40]. Alternative [BAA^{+}75, Gir75, Hal01]. alternatives [KPS78, Wig78f]. Alvin [Gla59, Kol60, Kou59, Wig66d, Wig69t, Coh59]. Alwyn [Ble86, Tor87]. am [Wig64n]. Ambiguity [HS01, Lie90]. Ambrosino [Wig61d]. amelyek [Wig71n]. America [TW79a, WT79, Wig83l]. American [BW61, Sec62, Wig82e, Wig82l, Rom91, Wig62c]. Amin [Wig72m]. amnesia [Scu07b]. among [Ano77b, Jan68]. Amos [Wig63a]. amplitudes [Kli74]. analog [Sch96, Wig66h]. Analysis [AN92, CM80, DFS20, Gou02, Gou03, HS01, Mac04, MPPS02, MPPS03, Tau63b, WY73, CMS77, GRT91, HMW02, LNPC92, Mic85, PI02, VC01]. Analytic [Wig51f]. anatomy [Gre69]. Andreev [GS04b, GS04a]. Andrew [Bon92, WS92, Wig02c, WS03]. angles [Wig67v, Wig68c]. Angolból [Wig92e, Wig92f]. Angular [RV11, WE47a, BV65, CM66, CW71, PFG06, SG75]. anharmonic [WN90]. anni [Wig80b]. anniversaire [Wig62d]. Anniversary [Gau94, Wig62i, Wig62d]. Announcement [Ano48]. annual [Wig65u, Wig74a]. Anomalous [Wig46e]. ans [Wig62d]. answered [Wig76f]. answers [Wig83j]. anthology [MGM62]. Anthropistike [Ano77a]. Antiballistic [HW77, HW74]. Antiferromagnetic [RI05]. Antisymmetrical [CW41]. Antiunitarity [LM63]. Antiunitary [Wig60g, Wig60h, Mol99]. Anwendung [Wig31b, Wig44]. apathy [Wig77b]. Apparatus [Wig57a, WYW61, Wig61g, WOYW62]. Appendix [Wig72d, Wig02a]. Appl [Wig85d]. Applicability [Gel14]. Applicable [Wig79c]. Application [Fan60, Gut60, LPN11, Sac59, Wig31b, Wig54b, Cha01, HS83, Hol73a, Wig44, Wig59g, SCS99]. Applications [Bie61, Mos62, Wig73j, Wig76b, Kol73, Loe75a, Loe71, Loe75b, SM06]. Applied [AN92, BW61, Cut63, How62, Sec62, HMW05, SW59, TCT16]. apply [Wig79n]. Appointed [Ano52]. appraisal [Wig69m, Wig69n]. Appreciation [Wig96b, Wig69f]. Approach [For97, JW87, MH98, Tal68, CPCF99, DK04, GI03, GS04b, GP86, LLBC85, MMP94, Pal82, PD98, PMHW07, Wu72, Bie69]. Approximate [MYPL90, WWCS42]. Approximation [DW64, LY09, BT76, Nau73, TCT16]. April [BW61, Wig67t]. Äquivalenzverbot [JW28]. arbitrario [Maj32]. arbitrary [Agr80, Maj32]. architect [Wag81]. archives [Wig79j]. area [Wig79l, Wig81b]. Argument [Cla71b, Wig71j, FW73]. Arithmetic [JF16, GRT91]. Armed [MW75]. arms [CW74, Wig77b]. Army [WFS^{+}69]. Arthur [Duf46, Rom00]. Article [Wig64c, Wig82k]. Articles [GR63]. ascended [Wig82j]. ascendesse [Wig82j]. Asim [Ble86, Tor87]. Aspects [SW72, Wie53, Hei74, Bel73]. Assembly [Wig58c]. Association [Wig37a, Wig82e]. assumption [HKW68]. assured [Wig70f]. astronomy [FF91]. Astrophysics [Tau63c]. Asymptotic [Kla91, RdGS01, VW66, MM73]. átfogó [Wig88d, Wig88e]. átírta [Wig88d]. Atom [Duf46, FR13a, Wig54b, App87, DK70a, DK70b, DK71, DK73a, DK73b, DK75, DK78, Dav74, wYDK74]. Atomic [Ano04, Com56, Dah95, Duf46, Fan60, FR13e, Fin60, GR63, Gut60, Sac59, Wig31b, Wig48c, Wig48d, Wig48a, Wig48e, Wig51b, MW46, MW07, SAZ90, Wig44, Wig46f, Wig59g, Wig63f, Wig65c, Wig07, Wag81]. Atomic-Test [Fin60]. atomkor [Wag98]. atommag [EGW69, Wig69y]. Atomphysiker [Her76]. Atoms [Ano78a, SD96, Wig77b, KPM97]. Atomspektren [Wig31b, Wig44]. attack [Wig67u]. Áttekintés [Wig64y]. Attempting [Kas21]. Attila [Wig88c]. Audio [TC10]. August [Gau94, WT79, Zic78, Zic83a, Zic83b, Zic85, Zic88]. Austausch [PW32]. Austausch-Reaktionen [PW32]. Austin [WNY^{+}84]. Automata [Tau63b]. Award [LW34, Ano58, Wig62b]. Awards [Ano63, Ano78a].

B [Wig95b, Wig98, Sch85, Wig01b]. Back [Isl22]. Background [SIM93, PI02]. baked [MGM62]. balance [CW74]. Baltimore [L.88]. Ban [Fin60]. Banesh [Wig79j]. Banff [Mac59]. Bargmann [Stu03, Stu04, Cri67, HL88, HL90, LR83, MS06, RL81, Wig78k, Wu72, dDG83]. barrier [FR03]. Barriers [But88]. Barut [Ble86, Tor87, Sch85]. Based [DW64, Tal68, TC10, Wen98, HMW02]. basic [Wig79e]. Basis [HVW65, HVW82, Ker85, WW30b]. Bass [Wig56e]. be [Nar88, Wig56e, Wig71n]. BeC [TKT^{+}05]. became [Wig02a]. because [Pal88]. been [Wig83l]. before [Wig69w, Wig76d, Wig82j]. Behavior [Wig48f, vNW29a, Kla91, MM73]. Behind [Wig70e, Wig79s]. Behram [L.88, Ram89, Hen89, Sch88, Stu88]. Beilenson [Wig82k]. Bell [FW73]. Bellows [Wig61c]. Bemerkung [Wig29a]. benchmark [SNDS14]. benefits [WW84]. Berechnung [WW30b]. Berlin [Ano30, Wig79o]. Berlino [Wig79o]. beset [TW79b]. Besprechungen [WBBE30]. Beszéd [Wig68j]. Beszélgetés [Wig73q, Wig88c, Pal88]. Beta [CW41, Wig54g]. Beta-Decay [CW41, Wig54g]. Better [TW79a, GW71, WT79]. Between [FW37, MWW66, OW77, OW78b, Stu18, Wig52b, AW75, EW41, Maj05a, Maj05b, OW78a, WWCS42, Wig60h, Wig61a, Wig72f, Wig75a, Wig77a, Wig79e, ZHH60]. Beyond [Wig37c, ZWE12]. bibliográfiával [Wag98]. bibliography [Wag81, Wag98]. big [Pal88, Wig69t]. Bildung [PW25, Wig25]. billion [Ano15]. binding [Wig41a]. Biographical [Wig69a, Wig01b]. biographies [Nob72]. Biography [Wig80a, Wig96c]. biorthogonal [Kli74]. Birkhoff [Cut63, How62, Sec62]. Birth [Anoxxa, Anoxxb, Wat93]. Birthday [Ano62, Wig96b, Ano61, BvdM83, MF69, Wig63f, Wig69f]. Bitter [Wig66k]. Bizonyos [Wig71m]. Blackmail [WH77b]. Blanc [Wig61d]. Blast [Wig70a]. Blatt [Wig53a]. block [Fre93]. Bloomington [Pai67a, Wei69a]. Board [MWW66]. Bochner [Wig67p]. Bodies [Wig36c, WWY58b, Wig61a]. Body [HW01, HMW01, CvV82, CPCF99, PS97, PMHW07, Wig62f, Wig67i]. Bohr [Duf46, FR13c, Whi96]. Boltzmann [CvV82]. Bomb [Ada70, Ano04, Duf46, Lan93a, MD67, Ros70, San95, Wei69a, Wei69b, Wig69x, Wig70b, Wol67, MW46, MW07, Wig63f, Wig71f, Wig72e, Duf46]. bond [Wig29b]. Book [Ada70, Ano48, Bec58, Bel73, Bie69, Ble86, Bre72, Cho68, Coh59, CH68, Cut63, De 58, Edw60, Fan60, Fes60, Gla59, Gut60, Gut68, Hei74, Hen89, How62, Kin58, Kir73, Kol60, Kou59, L.88, Mar67, Mar69, Pai67a, Pai67b, Par65, Ram89, Rom00, Ros70, Sac59, Sch67, Sch85, Sch88, Sec62, Set58, Stu88, Tor87, Van94, Wei69a, Wei69b, Wig49b, Wig50b, Wig56c, Wig58a, Wig61d, Wig61e, Wig62c, Wig63a, Wig64f, Wig64g, Wig65c, Wig65d, Wig66a, Wig67o, Wig67n, Wig67p, Wig67q, Wig69d, Wig70b, Wig74a, Wig76c, Wig76b, Wig81a, WBBE30]. Books [Fla68, L.88, Wig62a]. Boolean [Jan14]. Bordered [Wig55a, Wig57b]. Bose [OW84, OK78]. Bose-like [OK78]. boson [LB73]. bosons [AN09]. Boston [Ano79b]. boundaries [DP02, Wig64-29]. boundary [JC10, JCT11, JLC14, LSV08, ML91, WWCS42]. Bounds [WB05, Lie90]. Bowdoin [Gre69]. Brain [Wig69e]. breaking [BH13, RNM04, Vig66, dDG83]. Breeder [FWN^{+}45]. Breit [FR13c, Flü46, FR03, Kli74, ML91, PL75, TN84]. bridge [TEEBH03]. bright [Wig86e]. Brillouin [Ahl70, Ahl70, BSW36, BSW64, BM64, CV80, Gra74, HW01, HMW01, HMW02, HMW05, LWA74, Löw64, MH98, PMHW07, PČH02, TKT^{+}05, TCT16, Wen98]. brine [SW23]. British [Gau94]. Broglie [Ble86, Sch85, Stu03, Stu04, Tor87, WBBE30, BVV84, F^{+}76, Stu04, dBDW^{+}84]. Brussels [Wig64f, Wig64g]. Bub [FW73]. Budapest [Pal88]. Budapesti [Pal88]. Builders [MD67]. Bulk [CJSW55b, Eck73, Veb60, von96, vN96]. Bulletin [GR63]. Bundling [dSFA05]. BWCC [HMW05].

C [DLD^{+}21, MWW66, Wig65k, Wig72d]. C.W [Hen89]. calculate [SAZ90]. Calculation [FW36, UW42, Wig37a, Ahl70, CM66, GP86, WW30b]. Calculations [OG07, WN90]. calculus [DMMT93, Gou86a, Gou86b, MM92]. Cambridge [Ble86, Hen89, Kir73, Sch85]. can [Wig78e]. Canada [Hoo73]. Canadian [Mac59]. Candid [HH04a]. Canonical [BGL67, KW89a, KW90a, Loc76, QM71, Bec72, BL72, KP90, LB73, SCS99]. Capital [Fin60]. Capture [BW36, Wig57i]. Carbon [dSFA05, DB08]. career [Wag81, Wag98]. Carl [Eck73]. Carlo [KLW09, SNDS14]. Carlson [Wig67t]. Case [NAH^{+}89, MACS04, Wig82b, WB05]. Case-Study [NAH^{+}89]. Casimir [LNPC92]. Cat [Tes86, Vil86]. Catalonia [GHM^{+}87]. category [Joy01, JBR02]. Causality [Wig64a, Wig64h, Wig79c, CM68, F^{+}76, Wig64x, Wig79n, Par65]. cause [PW28]. Cavities [TW53, TP01]. CCSD [HMW05]. celebrate [Woo80]. celebrated [F^{+}76]. Celebration [Hen89, Wig88d]. céljáról [Wig92d]. Cell [KN94, CV80]. cells [Fre93]. Centenary [Hen89, Fre79]. centennial [Woo80, JKM03]. centrality [Ano75]. Centre [Ano68, DGW84, Meh73, Sal69, Wig68e]. century [Gau94, Har06b, Meh73]. Ceremony [Fin60]. Certain [Wig41b, Wig58h, WY64, Wig65h, Wig71m]. Certificate [Wig58b]. chaiman [Wig79h]. Chain [Coh59, Gla59, Kou59, WW61b, WOYW62, And73, CFM78, Kib76, Kol60, WW58, Wig62d, Wig67m, Wig82l]. chaine [Wig62d]. chains [LSV08, RNM04]. Chairman [Wig79f, Wig77g, Wig83e]. changed [Har06b, Mar06b]. Changes [Wig84b, Wig79t]. Changing [SM06, Ano78b].

Chaotic [TKC^{+}08]. Characteristic [Wig55a]. Characteristics [Wig57b]. characterized [BL72]. Charge [WWW70]. charged [Mar06a]. Charging [OYWW61, WY60]. Charging/Discharging [OYWW61, WY60]. Charpie [Wig57h]. Chase [Wig67u]. che [Wig82j, Wig83l]. Chemical [WE37, PW28, Wig29b, Wig32b]. chemischen [Wig32b]. chemischer [PW28]. Chemistry [Rom00]. Chernobyl [SBR^{+}86]. chianti [Wig83l]. Chicago [Kol60, Wig46b, Wig67m]. chiral [GS04b, GS04a]. Church [EW80]. címu [Wig69y]. circuits [Jan14]. Citation [Wig66b, Wig79a, Wig83a, Wig78j]. Cities [HW77, HW74]. City [Wig68j, Gre86, Wig78a]. Civil [Ada70, BWD76b, BWD76a, CH68, KW77, Mar69, PW64a, PW64b, Ros70, TW79a, Wei69a, Wei69b, Wig64b, Wig64i, Wig64j, Wig66c, Wig67r, Wig67s, Wig69h, Wig69i, Wig69x, Wig70b, Wig70c, Wig73b, WG74a, WG74b, Wig76d, WH77b, Wig82c, Wig82d, GW71, PFW^{+}76, TWB73a, TWB73b, Wig64w, Wig65e, Wig66e, Wig66j, Wig68h, Wig68d, Wig68i, Wig69q, Wig70j, Wig71e, Wig71i, WG72, Wig73m, WT79, Wig82b, Wig82e, Wig76e]. Clark [Wig71k]. Class [Wig51f, BG03, BLS03, BB92, BB93, CLM75, MM73]. Classic [Mar03, Wig79a, Wig83a]. Classical [Ber77, LPN11, VW65, VW66, Wig66a, BW73, CMS77, MMP94, Wig54c, Wig69c, Wig69b, dDG83]. Clebsch [Rno74]. Clerical [Wei75]. Closed [Goe49]. closing [Wig75f, Wig79h]. Cloud [Wig72a, WL72]. Club [Wig02a]. Cluster [MH98, HMW05, PČH02, TKT^{+}05]. codimension [KL03]. coefficient [Loc76]. Coefficients [Bie61, BGL67, BMR65, CCB72, Mos62, RY03, AD73, AD84, Bru85, Bru87, CM66, CLM75, DA73, DLD^{+}21, Got71, Gou81, Gou86a, Gou86b, GP86, GLB92, Gou92, HS95, HS83, HB90, Hol73b, HST74, Hon76, Ker85, LB71, Mae91, ÖAÖ25, PD98, Rno74, SG75, SS79, Str79, Wig54f, Win76, Won76, Won79, Wu72, Wu73]. Coherent [But88, HB90, RR96]. cohesion [Wig54i, WS55]. Collected [Rom00, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, Wig92c, WWM98, Wig93, Wig95b, Wig96m, Wig97a, Wig97b, Wig98, Wig01b]. collection [BV65, F^{+}76, K^{+}84, NKW88, Por65]. College [Gre69, Zac84, Wig67t]. Collinear [LPN11]. Collision [DW64, KS07, GW72a, Wig62g, Wig64a, Wig64h]. Collision-Free [KS07]. Collisions [Wig51f, WvN54, Wig64y]. Collocation [Rin92, Arn94]. Colloquium [DGW84, Zac84, KR78]. Columbia [Gau94]. Come [Wig79g]. comforting [Wig66k]. Comm [Wig85d]. Commemoration [Wig67m]. Comment [CD79, Wig76e]. Commentary [Wig48e, Wig65e]. Comments [MW62, Put64, WF68, Wig81a, JWY67, NL69]. Committee [ACF^{+}42, Wig69w, Wig76d, Ano52, LK04, Wig77g, Wig82b]. Common [For97]. Commutation [Wig50c, OW92]. Como [d’E71]. Compact [BG02, FB88, GW72b, Sha73]. comparative [TKT^{+}05]. Comparison [SBR^{+}86, Eck33, HMW05]. Complementary [PD98]. complete [Bo09]. completeness [HS95]. complex [Maj05a, Maj05b, ZHH60, Wig29c]. component [CMS77, HKS86]. Compound [Wig55d, Wig67e]. comprehensive [Wag81, Wig88d, Wig88e]. Compton [Duf46]. Computation [Lai94, Wol71, DLD^{+}21, Gau94, Sha92, SM07]. Computational [SM07, Wig73j, Gau94, Kol73]. Computers [Tau63b, K^{+}84, Wig69e]. con [Maj32, WU75]. concentrate [Ano30]. conception [Meh73]. Concepts [Wig76b, Wig79e]. Conceptual [HVW65, PE77, HVW82]. Concern [Fin60]. concerning [JWY67]. conciousness [Wig79k, Wig82f]. Concluding [Wig74b, Wig78b, Wig78c, Wig96d]. conclusions [Wig64-28]. Condition [GW72b, CM68, JCT11, JLC14, MFAB85]. Conditions [BY88, Nar88, OW81a, Wig39a, LSV08, MF69, ML91, WWCS42, Wig83f, Wig95c]. Condon [Wig56d]. conductance [TP01]. Conducting [Wig59a, Wig60a, WOYW62]. Conference [Ano54b, Ano75, Ano77b, Ano78b, Ano79b, Ano83a, SM07, Vil85, Wig51a, Wig51e, Wig58a, Wig60b, Wig67l, Wig74c, Wig79h, WNY^{+}84, ZHH60, Gre86, Hoo73, Wig60i, Wig64f, Wig64g, Wig69u, Wig71l, W^{+}75, SW56a]. Conferences [Rub82]. configuration [HMW02, TKT^{+}05]. conflict [Wig79e]. conflicting [Ano19]. conformally [HL88]. Congress [Mac59, Wig76d]. Congresses [Wig64g]. Congressional [WT79]. Congruences [SCS99]. Conjecture [FP03]. Connection [Par65, Wig52b, AW75, Wig64x]. Connections [Lan93b, Lan99]. Connes [Bac89]. consciousness [Jah81, Wig72j, Wig67n]. consequence [Wig70i]. Consequences [Wig37b, Wig65o]. Conservation [VW66, Wig52c, Wig54c, Wig64t, Wig85c, DGW34, Vig64, Wig27c, Wig64s, Wig65p, Wig65q, Wig65r, Wig65s, Wig82h]. Considerations [HW39]. Consistent [OG07, Wigxx]. Constant [Wig36a, PW32]. constituents [Zic78]. Constitution [WS33, WS34]. constraints [MYPL90]. constructed [Ell92]. Construction [WO58, LR78, GS04a]. constructions [HB90]. contain [GW72b]. Containing [CW58, CW61]. Contemporary [Sal69, Hoo73, Sal69]. Contents [WY63]. continued [Gra74]. Continues [Ble86, Sch85, Tor87, BVV84, dBDW^{+}84]. Continuity [Beh91]. Continuous [Tau62]. Contraction [IW53, FB88, FS82, Maj93]. Contractions [Con72, WW95, Hud78]. contradictions [Wig69p]. contribution [Wig80d]. Contributions [Wig62b, F^{+}76, Wig33d]. Control [NWEC61, Wig64b]. Controlling [NW61, WYW61]. Convergence [BY88, Bai93, BY05, IW54, Ahl70]. Conversation [Wal73, Pal88, Wol67, Wig72m]. conversations [HH04a]. Conversion [WOY57]. Conversions [Wig59a, Wig60a, PW28]. Convocation [Wig59b]. convoluted [MO20]. convolution [Nar88]. Coolant [Wig61c]. Cooled [WC60, ACF^{+}42, LCM^{+}42]. Cooling [Wig43, WOYW59]. coordinate [WI98]. Coordinates [HW35, Poi99]. copolymers [Fre93]. Core [RW54]. corepresentations [KP90]. Corfu [SM07]. Correction [Pla25, Wig32a, Wig58b, Wig83a]. corrections [HMW02]. correlation [Mar03, PMHW07, WI98]. correlation-energy [WI98]. correlations [CHDK08, SM06]. correspondence [Poi99]. cost [Wig70a]. Coulomb [CHDK08, Ker05, Nou99]. counter [Wig70d]. counter-strike [Wig70d]. County [Wig64b]. Coupled [MH98, HMW05, LSV06, PČH02, TKT^{+}05]. Coupled-Cluster [MH98, PČH02]. Coupling [Wig39a, ÖAÖ25, PFG06, SG75, Wig54e, wYDK74]. Course [Zic83a, Zic85, Wig73o, Zic83b, Zic88, d’E71]. covariance [BB92, BB93, DV97]. Covariant [KW87c, KW87a, KW88c, KW88a, KW88b, KW89b, HS95]. coverings [FS82]. CP [Wig62i]. CP-I [Wig62i]. create [WG74c]. created [Wig73q]. creation [Ano83a]. Creativity [NAH^{+}89]. creators [Str11]. Credibility [Sau75]. Crick [Wig67x]. crisi [Wig83l]. crisis [Wig83l]. Critical [PE77]. Criticism [Wig60c, Wig65k, Wig72e]. Cronkite [CH68]. Cross [Wig46a, Wig48f, Wig61e, MM73]. Cross-Sections [Wig46a]. crossing [Wig32b]. crossings [KL03]. Crossover [AF95, FKS97, RNM04, DK04]. crustal [VC01]. Crystal [Ano06, DB08, PI02, RI03, RI05]. crystallization [Ano08, ISRK12, LM89, Mar06a]. Crystals [BSW36, BSW64, Mar03]. csökkent [Pal88]. Csoportelméleti [Wig79x]. csoportok [Wig71m, Wig71n]. Cuba [Wig64k]. Cultural [Gre69, Vil85, Wie53]. Current [WF68, JC10, LW78]. Curvature [Wig60d]. curved [APW02, APW03, Fon94]. cusps [Hod76]. Cylindrical [KW87b, RRC05]. Czechoslovak [Wig65c].

D [Bie69, RI03, Wig61e, RI05]. D. [Wig96j]. dangers [Wig67t]. Daniel [Wig61d]. dans [Dem15]. Darrow [Ano48]. Data [BGA07, HNP09, VC01]. Day [Wig79o, Wig82i]. days [Wig83c, Wig83d]. DC [W^{+}75, Ano77b, Vil85]. De-Shalit [Wig63a]. Death [Wig83b]. debate [BWD76b, BWD76a]. debated [Wig73g]. Dec [Wig80a]. Decay [CW41, Dah95, RW54, WW48, PW25, WW51, Wig25, Wig54g]. December [Wig64o, Wig64n, Wig67a, Wig64p]. Decision [Fel67]. decomposition [Lit72, RV11, Vas89]. Deconvolution [RB92]. deep [VC01]. Defeats [Hal01]. Defect [Wig33a, Wig65a]. Defects [dSFA05, TEEBH03]. defence [BWD76a]. Defended [Wig68g]. Defense [CH68, HW74, HW77, KW77, Mar69, TW79a, Wig64b, Wig67u, Wig69j, Wig69k, Wig69x, Wig70b, Wig70c, Wig76d, WH77b, BWD76b, GW71, PFW^{+}76, PW64a, PW64b, TWB73a, TWB73b, Wig64i, Wig64j, Wig64w, Wig65e, Wig66c, Wig66e, Wig66j, Wig67r, Wig67s, Wig68h, Wig68d, Wig68i, Wig69h, Wig69i, Wig69q, Wig70j, Wig71e, Wig71i, WG72, Wig73b, Wig73m, WG74a, WG74b, Wig76e, WT79, Wig82b, Wig82c, Wig82d, Wig82e, Ada70, Ros70, Wei69a, Wei69b, Wig70b]. Defining [Mla98]. Definition [BGL67]. Deformations [BW73]. Deformed [RW54]. del [Maj37, Wig79q, Wig83l]. Delano [Ein39]. Delivered [Wig96a, Wig87b]. dell [Maj37]. della [Wig80b]. demands [Wig69m, Wig69n, Wig72h]. demon [Scu07a]. Density [OG07, WYW61, BZ74, Lag84, MYPL90, Nau73, PL75, Wig72d]. departments [Ano30]. dependent [KNS04, Rob93, SP74]. Deploy [Fel67]. Derivation [CM68, JW87, Wig52d, AW75, Flü46, Jor80]. Derivations [Wig54d]. Derivative [Wig52b, Wig52e, Wig52f, Wig55c]. derived [WE39]. description [SS05]. Design [ACF^{+}42, Tau63b]. destruction [Wig70d, Wig70f]. Detection [Ore58, Wig58e, TN84]. Determination [Nou99]. Determine [Wig50c]. determinism [F^{+}76, Wig83h]. Deterministic [AN92]. Development [Meh73, Mei61, NWA72, WW61a, Wig55d, Wig67e, WA72]. Developments [Wig51b, LB73]. Device [OYWW61, WOY57, WY58, Wig01a, JLC14]. Devices [Tes86]. Dexer [Duf46]. Dezember [Wig64n]. DG [JC10]. Diagram [Joy01, SM05]. diagrams [Joy01, Wig65c]. Dialog [Jha11]. Dialogue [MWW66, Zsa76]. diatomic [WW28]. dietro [Wig79s]. different [Ano19]. Differential [BM64]. difficile [Wig83l]. difficult [Wig83l]. Diffusion [KS07, WWCS42]. díjas [Zsa76]. dilemma [Whi96]. dilettáns [Wig72m]. dilettante [Wig72m]. Dimensional [Ano06, Con72, Ste91, Wig60d, Ano08, CV80, DK70a, DK70b, DK71, DK73a, DK73b, DK75, DK78, DB08, Fro79, Hon76, Hud78, Nou99, RRC05, RV10, RV11, wYDK74, KPT86]. Dimensions [Wig55a, Wig57b, BZ74, Mar06a, SW67, Wig59f]. diodes [JCT11]. Dirac [Ble86, Hen89, Sch85, Stu88, Szm99, Tor87, Wig96a, BVV84, Bel73, For97, KW87d, L.88, LW34, LW78, Szm98, Tay87, WW30b, Wig88b, Wig87c, Wig96d, Wig96k, dBDW^{+}84, Ram89, Sch88, Stu88]. Diracschen [WW30b]. Direct [Wig73j, Lit72, Wig73i]. disagree [WB77]. Disagreeable’ [Har99]. Disarmament [Wig69w, Wig67r, Wig67z]. Disaster [Wig64b]. Discharging [OYWW61, WY60]. Disclose [Fin60]. Discovered [Wig76b]. Discovery [BHT86, Rog13, Wig79u]. Discrete [Bo09, Mac04, vNW29b, ALZ00, GRT91, LSV06]. discrete-velocity [ALZ00]. Discussion [BW48, FWN^{+}45, LCM^{+}42, NAH^{+}89, Wig56e, Wig70e, Wig88c, HWWJ^{+}38, W^{+}75, WNY^{+}84]. discussions [Wig71b]. Disintegration [BW37]. diskrete [vNW29b]. dispersing [ISRK12]. Dispersion [TW52, Wig64x, Par65]. Dispersive [Col94]. dissenting [SW56a]. Dissidents [Fre15]. Dissipation [LZ94]. Distances [SW58, SV86]. Distinction [Wig60h]. Distribution [BC62, CM80, HOSW84, KW88b, RB92, SSP^{+}91, TC10, UW42, Wig52b, Wig57c, Wig58h, Wig65f, Wig65g, Wig71a, Wig84a, ACW98, CW71, GRT91, JC90, JC10, KW89b, Lan87, MYPL90, Nar88, Nou99, OW81a, OW81b, OW84, Wig51g, Wig65h, Wig65j, Wig87a, WN90]. Distributions [Bai93, HS01, Jan81, WY63, Jan84, Lie90, LO10, MACS04, Nar88, TP01, VC01]. Djilas [Wig69l]. Dnej [Vig68]. Do [FR13a, Kas21, Wig50c]. doctorate [Wig66b]. Documenting [LK04]. Does [WL65, Wig79n]. dolgokat [Wig88c]. domestic [Frö93, Lad93]. doped [SM05, SM06]. Dr. [Cla71a, Wig82k]. Drafted [WFS^{+}69]. Drama [Her76]. Dream [Lan83]. Drehelektrons [vNW28a, vNW28b, vNW28c]. Dual [Tem07]. Dufort [MPPS02, MPPS03]. Dukas [Wig79j]. Dunworth [Wig61e]. Dutch [Wig76e]. dynamic [CPCF99, Wig54h]. Dynamics [Dav74, MDH94, Tem07, CPCF99, Ker05, KLW09, PS02, wYDK74]. Dyson [AF95, AK00, BLS03, Fol78, Fol80, FKS97, GS04b, GS04a, LLW^{+}09, RNM04, TP01, Yan10].

E. [Ada70, Fan60, Par65, Ros70, Wig86a, Wig49b, Wig49e, Wig56e, Wig56d, Wig67t, Wig74c, Wig75c, Wig76f, Wig77b, Wig83j, Wig96j]. Earth [TW79b]. East [SW59]. Eckart [Agr80, Eck73, Ell92, Gin63, MS00, Mez77]. Economical [Mei61, WW61a]. Ed [Wei69a, Hen89]. Edited [Duf46, Sch67, Sec62, Wig79j]. edition [Wig02a]. editor [Par65, BJH68, Kue61, Lap60, Mei61, WW61a, Wig56d, Wig58e, Wig69o, WA72]. Editorial [Fel67, Rom91]. editors [Tor87, Wig63a]. eds [Ble86, Hen89, Sch85]. Edsel [Ano78a]. Education [Wig82e]. Edward [BH89, Hor97, MF69, Wig69f, Wig96b]. Effect [CJSW55a, CJW55, JC10, SIM93]. Effective [LPN11, Wig65e]. Effectiveness [Gel14, HNP09, Isl22, Nic12, Omn11, Tis98, Wig60j, Wig60k, Wig60l, Hed93, Jan14, Vig68, Wig64-27, Wig66e, Wig67k, Wig70a, Wig85d, Wig90b, Wig91, Wig92b]. Effects [Col94, SW56b, Wig38a]. Efficient [RY03, PMHW07, SAZ90]. Effizienz [Hed93]. Effort [Wig68g]. egyik [Wag98]. egyszerûen [Wig71n]. Eigenschaften [vNW28a, vNW28b, vNW28c]. Eigenschwingungen [PW28, Wig20]. Eigenstates [TKC^{+}08, Stu03]. Eigenvalue [BY88, FKS97, vNW28a, vNW28b, vNW28c, JD90, LWA74]. Eigenvalues [vNW29b, vNW29a, Wig65o]. eigenvibrations [Wig64m]. Eigenwerte [vNW29b]. Eigenwerten [vNW29a]. Eighteenth [Zic83a]. eightieth [BvdM83]. Einführung [EW61, WBBE30]. Einige [Wig27a, Wig27b]. einiger [vNW28a, vNW28b, vNW28c]. Einstein [Wig71k, Wig79j, Ein39, Fre79, KN19, Vig86, Whi96, Wig29a, Wig49a, Wig79d, Wig79j, Wig79i, Wig79m, Wig80d, Wig80e, WW86, Wig96g, Wig96e, Wig96j, Wig96l, Woo80, Wig71k]. Einsteins [Wig29a]. Eisenbud [Bec58, De 58, Edw60, Kin58, Set58]. Elastic [Wig64m, Wig20]. elastischen [Wig20]. electrical [SW23]. Electrodynamics [Sch58, Sch03, dlPC79]. Electromagnetic [TW53]. Electron [OG07, RB92, SD96, WCT39, vNW28a, vNW28b, vNW28c, LM89, Maj37, Mar03, NW70, PS02, RI03, RRC05, Wig38a]. Electronic [KKF89, SD96, Ano08]. Electrons [FR13a, Wig34, ISRK12, MMP94, Wig38a]. Electrostatic [Fol78, Hal81, Hal83]. elektrische [SW23]. elektron [NW70]. Element [RW54, WSC59]. Elementary [NW49, OW77, WWW52, Wig37a, Wig52d, WNY^{+}84, FW36, GW63, Ram00, WWW64, Wig77a]. Elementary-Particles [WNY^{+}84]. Elements [QM71, SWC59, Wol71, Wol72, CLM76, Gou81, GP86, GLB92, Wig61d]. elemi [WWW64]. élete [Wig02b]. elettrone [Maj37]. Eleventh [Ano83a, Sec62]. elmélet [Wig35c, Wig88d, Wig88e]. elméletbol [Wig64-28]. elmélete [Wig29b]. elméletéhez [Wig33d]. elméletérol [Wig64y]. Elöadás [Wig88d, Wig65u]. Eloszó [Wig69y]. ELTE [Wig88d]. elvek [Wig76o]. embedded [Kot05, Kot07]. ember [Wig88c]. Emberi [Wig75g]. Emission [Dir27, DK70a, DK75, DK78]. emlékezés [Wig66k]. emlékiratai [Wig02c]. emotions [Wig75d]. Empirical [BY05]. Ends [Wig61h, Wig67g]. Energies [Wen98, WNY^{+}84, Wig41a]. Energieschwankungen [PW28]. Energy [BW35, HW39, PW67, SIM93, Seg85, SBR^{+}86, TW79b, Wig48c, Wig48d, Wig51b, Wig55c, WNY^{+}84, CMS77, DHW49, DK04, KPS78, MM73, MYPL90, PW68, PW28, WW60, Wig38a, Wig48a, Wig48e, Wig48g, Wig67d, Wig69s, Wig72c, Wig73k, Wig74g, WU75, Wig76a, Wig76n, Wig78f, WI98, Zic83a, Wig92c, Wig57h]. energy-level [DK04]. engineer [Wei02]. Engineering [SM07]. Engineers [SBR^{+}86]. England [Ano65]. English [Wig92e, Wig92f, Wig72e]. Enough [Gul75, Wig72e, Wig86e]. Enrico [d’E71, Ano58, Wig55b, Wig96f]. ensemble [Kot05, Kot07, KPM97]. entitled [Gre86, Wig71k]. Entropy [KW90b, KW90c, Wig90a]. enveloping [Que76]. Eötvös [Stu05, Wig88d]. Epistemiology [Wig69m, Wig69n]. Epistemological [Wig73c]. Epistemology [Jha11, Wig68a]. Equation [AN92, Gou02, Gou03, HW35, KNS04, MPPS02, MPPS03, NAKS04, Wig36a, Arn94, ALZ00, CIMS08, DF77, LWA74, WWCS42]. Equations [BW48, BM64, Col94, Cri67, Gar88, Stu04, Wig48b, Wig50c, Zha03, BT73, CvV82, Duf71, HL88, HL90, JCT11, JLC14, LR83, MS06, RL81, Stu03, Szm98, Szm99, Wig47b, Wig56f, Wig63b, Wig73k, Wig73l, Wig74g]. Equilibrium [Wig32a, Wig83a, DK70a, DK70b, DK71, DK73a, DK73b, DK75, DK78, wYDK74]. Equivalence [CLM75, JW28]. equivalences [ZWE12]. Eraser [Scu07b]. érdekessége [Pal88]. Ereignisse [Wig64n]. erfolgreiche [Wig70d]. Ergodic [Tau61b]. Erhaltungssätze [Wig27c, Wig85c]. Erice [Zic78, Zic83a, Zic83b, Zic85, Zic88]. Erinnerungen [Wig96g]. Erklärung [vNW28a, vNW28b, vNW28c]. Errata [MR73, WCJS56]. Erratum [BB93, Pos01, Szm99]. Error [Wei75, GRT91]. Erscheinungen [SW23]. Erster [Wig26a]. értelmezésér [NW70]. érzi [Wig88c]. escape [Mar06b]. Események [Wig65u]. Essays [Ano61, Ble86, Sch67, Sch85, Tor87, BVV84, K^{+}84, Wig67-28, Wig70k, Wig72n, Wig78l, dBDW^{+}84, Gut68, Mar67, Pai67a, Kir73]. esszéi [Wig72n]. est [Dem15]. Establishment [Cob72]. estestvennyh [Vig68]. Estimation [BGA07]. esztendeje [Wig76o]. Eszter [Wig92e, Wig92e]. Ethics [HKT77, Wig72f, Wig75a]. Euclidean [Rno74]. Eugene [Bec58, Ble86, Cho68, Coh59, CH68, De 58, Edw60, Fla68, Gla59, Goe01, Hei74, Hen89, Hor97, Kin58, Kol60, Kou59, L.88, Pai67a, Ram89, Rom00, Sac59, Sch67, Sch85, Sch88, Set58, Stu88, Tor87, Wei69a, Wei69b, Wig73q, Zsa76, Ano62, Ano83b, Ano92, Ano02, Ano04, Bac89, BvdM83, BVV84, Bie69, Bon92, BH89, Cha01, CFW78, DMS84, FR13b, Fes60, Gut68, Har99, HH04b, KW01, Mar67, Mar69, Mar95, Muk95, Nag88, Ne’04, Olt74a, Pai67b, Pal88, Rom00, SW89, Sei95, SVW98, SVW00, Stu05, Sza92, Tal68, Tel95, Tis03, Vog95, Vos93, Wag81, Wag98, Wal73, Wei97, Wei02, Wig95a, Wig70k, Wig71d, Wig72f, Wig72n, Wig72o, Wig73a, Wig78l, Wig88c, Wig92c, WWM98, WS92, Wig93, Wig95b, Wig96m, Wig97a, Wig97b, Wig98, Wig01b, Wig02c, WS03, Wol67]. Eugene [Woo95, dBDW^{+}84]. Euler [Wig67v, Wig68c]. Eulerian [HKS86].

Europe [Fer68, Fer71]. evacuation [WG74c]. Evaluation [CCB72, JF16, Str79, Ano79b, SG75]. evening [Ano68]. Events [Wig64o, Wig64p, Wig67a, Wig72b, Wig85a, Vig65, Wig64n, Wig65u]. Ever [Fra05]. Everyday [Wig73j, Wig73i]. évi [Wig65u]. Evolution [Wig82a, MFG^{+}83, SEC83, Wig65d, Wig65m]. Exact [KS82, SG75, DK73a, DK73b]. Examination [ACF^{+}42]. Excerpts [Wig74c]. Exchange [BW35, BJH68, Wig36b, PW32]. Excitation [Wen98]. Excited [TW53]. exclude [WL65]. Exhibiting [OW78b, OW78a]. Exiles [Fra05]. Existence [Wig61b, Wig79k, Wig82f]. existing [Ano79b]. expansion [Gra74, Pul06, Sha92]. Expansions [Kli74, TW53, Nau73]. expectations [Wig75c]. Expected [Bai93, WÁ68]. Expelled [MP01]. Experience [CGPT21, Ano19]. Experiences [L.88]. experiment [Ano19, Zic84]. Experimental [CHS97, PPG^{+}19, DGW34, MS83]. experimentelle [DGW34]. Explaining [Wig67n]. Explanation [vNW28a, vNW28b, vNW28c, Wig70h, Wig74f]. Explicit [BB65, HST74]. expression [WY64]. Extended [Ste25, BB92, BB93]. Extending [WNY^{+}84]. Extension [NW71, Wig79l, Wig81b]. Extensions [Lan87, Ara95]. external [Fon94]. extra [DK75, DK78]. Extraordinary [Gib19]. extremum [SCS99].

F [Wig53a, Wig67t]. Facets [BH70]. facsimile [Wig02a]. Factorization [JF16]. Factors [Wig65d, Wig65m]. Falk [Wig73g]. Fallout [Wig60c]. families [Pos00, Pos01]. Family [Wig82g]. famous [HH04a]. Fano [CGPT21]. far [Zic85]. Farkas [FW52]. Fashioned [Duf46]. Fasori [Wig73r]. Fast [JF16, SW61, Wig61e]. fears [WG74c]. Feb [Wig80a]. február [Wig02b]. February [WJH76, WH77a, Wig02b]. feel [Wig88c]. Feenberg [CFW78, FR13d]. Feld [BJH68, WF68]. féle [Wig35c, Wig64-28]. Feliu [GHM^{+}87]. Fellow [DFS20]. feltételektõl [Wig95c]. fénypontjai [Wag98]. Fermi [Ano58, Wig55b, Wig96f, d’E71, Nau73, OW84, Seg03, Wig62b]. fermion [BZ74, HKW68]. Ferromagnetic [RI03]. festschrift [CGPT21, LS76]. fiatal [Frö93, Lad93]. fiataloknak [Wig95d]. Field [Hal01, TW53, WCT39, AN09, BT73, DK70a, DK70b, DK71, DK73a, DK73b, DK75, DK78, wYDK74]. Fields [Wig64f, Wig64g, Fon94, LW78, LM89, SS05, dDG83, Rom00]. Fifth [Ano77b]. Fifty [Wig78d, Wig76o]. Figure [Vog95]. filozófus [Wig72m]. fingerprints [Mar06a]. Finite [Wig41b, ACW98, DK73a, DK75, LLBC85, Ras89, VGB87, Wig71m, DK70b]. fire [Wig82j, Wig83l]. first [Ano19, Bug77, Wig62d, Wig67m, Zic85, Stu18]. Fisher [Luo04]. Fisica [Wig77c, Wig79p]. Fisk [Ano52]. Fission [BHT86, BW39, COWY57, WW48, CS51, WW46a, WW46b, WW51, Whe09]. Fissionable [WWY58b]. fitting [TN84]. Five [Wol67, Har06b, Har06a, Hon76]. five-dimensional [Hon76]. fixed [GRT91, LSV08]. fixed-point [GRT91]. fizika [Pal88, Wig92d, Vig68]. fizikát.‘ [Wig92f]. fizike [Vig66]. fizikus [Wig02a]. FL [FR66]. fled [Mar06b]. Flow [JW87, MR98]. fluctuations [PW28, Por65, TP01]. Flügge [Wig56c]. Fluid [WOYW57]. Flux [NWEC61]. Fock [OG07]. Fodor [Wig95c, Wig95c]. fogunk [Wig75g]. Folgerungen [Wig27a, Wig27b]. following [Bug77]. fontos [Wig88c]. Force [LPN11]. Forces [BW38, Bri65a, Wig36b, WCT39, WE39, Wig70i, Zic85, Zic88]. Ford [Ano78a, Wig92e, Wig92f, Ano78a]. Foreign [Wig69w, Wig82b]. Foreword [Duf46, Wig47a, Ano48]. forgiven [Wig69l]. Form [Wig60g, FR13a, KP90, Wu73]. Formalism [JvNW34, Löw64, Maj05a, Maj05b]. Formation [PW25, Wig25]. Forms [EW41, SCS99]. formula [CFM78, Flü46, FR03, ML91, NL69, Ros98, Ros99, Win76]. Formulas [Lai94, CvV82, Wig67q]. Formulation [Stu04, BBDP08, MH94, Szm98, Szm99, Wig29a]. Formulierung [Wig29a]. Fortran [AD73, AD84, CM66]. Forum [KPS78]. foundation [WE47b]. Foundations [Fre15, GL23, Gre69, PE77, Rom00, Wig66a, Hoo73, Wie53, d’E71]. founders [Wag98]. Four [Fin60, Lan93b, Lan99, PČH02, Hud78, Lan11]. Four- [PČH02]. four-dimensional [Hud78]. fourth [Ano75]. fractions [Gra74]. Francis [Wig67x]. Francisco [Ano78b]. Frankel [MPPS02, MPPS03]. Franklin [Ein39]. Free [CCB72, KS07, TW53, Gou86a, Gou86b, SS05, Won76, Won79]. freedom [HL88, Wig83b]. French [Dem15, DMS84, Wig62d, Wig73f]. Frensley [JCT11]. Frequency [CM80, JW87, Mac04, TC10, VC01]. Friedrich [ABES12]. Friend [DFS20, Die21, Kas21, Lee22, Scu07b, Ste25, Wes21, Wig79m, Pla24, Pla25]. friends [MO20, Scu07b]. Fritz [Wig67q]. Fuel [OYWW61, SWC59, Wig58c, WSC59, WY60]. Full [Duf46, MW46, MW07, Wag98]. fullerides [SM05, SM06]. Function [Ber77, BC62, BGA07, FM03, KS07, MH94, SD96, UW42, Wig52b, Abe92, ACW98, Bo09, Beh94, CPCF99, Fon94, JC10, KPM97, MMP94, Nou99, OW81b, OW84, RRC05, SEC83, SC83, SK05, Sta10, TCT16, Wig87a, WN90, WB05, wYDK74]. Functional [OG07, WI98]. Functions [Bie69, BSW36, Dah95, HS01, KW88b, Ler90, OLBC10, Tal68, Tau61b, WB35, Wig51f, Wig67q, Wig71a, APW02, APW03, Ara95, BB92, BB93, BSW64, BW95, CUZ01, DP02, Dra86, HOSW84, Jan84, KW89b, Kla91, Kli74, KLW09, Lie90, MR98, MYPL90, MFAB85, OW81a, ÖAÖ25, PFG06, PL75, Pul06, Rno74, Rob93, RSdG99, RdGS01, SAZ90, Wig84a]. fundamental [Zic78, Zic83b]. Fundamentals [Wig84a, HOSW84]. fuoco [Wig82j, Wig83l]. Further [LB73]. Fusion [Stu04]. Future [KPS78, WPG77, Wig78e, Wig78h, WM95, Wig99].

G [Cut63, How62, L.88, Stu88]. GAC [Lap60]. Galilei [IW52]. Gallery [Pai00]. Games [Tau63c]. ganz [Hed93]. Gap [Ano15, TEEBH03]. Gatlinburg [ZHH60]. Gauge [Zic84, ZDF^{+}84, HKS86, HL88, Zic85]. Gaunt [RY03]. Gaussian [Hal81, Kot05, Kot07, PI02]. Gegen [Wig70d]. Gegen- [Wig70d]. General [EGW67, HWWJ^{+}38, Wig29a, WNY^{+}84, GP86, Wig70i, Wig79e, Wig87a]. Generalization [JvNW34, Mol99, Pos00, Pos01, Wig67v, Wig68c, Löw64]. Generalized [Col94, DP01, Duf71, HW01, Stu04, CvV82, Ker85, LWA74, MFAB85, ÖAÖ25, Stu03, WW95, Yan10]. Generating [CUZ01]. generators [Gou81]. genesis [Cha01]. Geneva [SW56a, SW59]. Genius [Pai00]. Geniuses [Lan93a]. Geometric [HVW65, GI03, HVW82]. Geometrical [FM03, Maj05a, Maj05b]. Geometry [CJSW55a, KW89c, Tau62, Wig61a, KW90d]. Georg [ABES12]. George [Wig62a, Wig72o, Stu05, Wig73r, Wig88d]. Georges [Wig61d]. German [Wig62j, Ahl70, DGW34, EW61, Hed93, Her76, JW28, MW24, PW32, PW25, PW28, SW23, WW30b, WW30a, Wig20, Wig25, Wig26a, Wig26b, Wig27a, Wig27b, Wig27c, WW28, Wig29a, WBBE30, Wig31b, Wig31a, Wig32c, Wig32b, Wig33c, Wig33b, Wig44, Wig47c, Wig48b, Wig51c, Wig52a, Wig58i, Wig62h, Wig64n, Wig64l, Wig65r, Wig70d, Wig76l, Wig79n, Wig85c, Wig96g, vNW28a, vNW28b, vNW28c, vNW29b, vNW29a]. Geschwindigkeitskonstante [PW32]. Gets [Ano58]. Géza [Wig72o]. Giant [LTW55, Wig66h]. Giants [MD67]. Gift [MP01, Wig82g]. Gilpin [Wig62c]. Gilt [Wig79n]. Gimnáziumban [Wig88c]. Gimnáziumból [Wig73r]. Ginibre [FKS97]. Giorno [Wig79o]. Gitterstruktur [MW24]. glass [BTW73, TWB73a, TWB73b]. glimpses [Wig79j]. glorious [Wig83c, Wig83d]. Go [Fra05]. Goeppert [Wig72i]. going [Wig75g]. Goldberger [Wig72d]. golden [Mar96, Seg03]. Gomolcak [Wig65c]. gondolkodtam [Wig72m]. good [Wig73r]. Gordan [Rno74]. Gordon [SS05]. Gore [Wig69w]. Government [Zic78]. graded [Mez77, WW95]. gradient [CFM78]. graphene [ISRK12]. Graphite [SWC59, TEEBH03]. graphs [GS04b, GS04a]. Gravitation [Wig49b, Wig49e]. gravity [MS83]. Great [Hen89, L.88, Ram89, Sch88, Stu88, KW87d, Mar06b]. Greater [WNY^{+}84]. Greece [SM07]. Green [JC10]. Greenberger [Lee22]. greeting [Nag88]. grid [MW24]. Griffin [Gut60, Sac59]. Groenewold [DV97]. Ground [SIM93, TKT^{+}05]. Ground-State [SIM93]. Group [BW48, Bie61, Bie69, GW63, Gre69, KR78, Loe75a, Loe71, Loe75b, Mei64, Mos62, Tal68, Tau61b, Wig39b, Wig59g, Wig79a, Zac84, Agr80, Cha01, CW71, Dir45, Ear81, GW72b, Got71, Gou81, GP86, IW52, KW87b, LR78, MR72, MR73, NKW88, PD98, Rno74, RSdG99, Vas89, Wig64e, Wig71n, Wig89a, DGW84, Fan60, Fes60, Wig31b, Wig44, Wig79x, Gut60, Sac59]. Groups [BB65, BMR65, Gin63, IW53, Wig41b, Wig66a, Wig71c, vNW40, BL72, BW73, CFM78, Ell92, Gou86a, Gou86b, HKS86, Ker85, LB73, LLBC85, MACS04, MMZ^{+}05, Sha73, Wig65i, Wig71m]. growth [Olt74b, Wig67t, WÁ68]. Grund [WW30b]. Gruppentheorie [Wig31b, Wig44]. guerra [Wig80b]. guide [AD73, AD84]. Guíxols [GHM^{+}87]. György [Wig73r, Wig88d]. Györgyi [Wig72o].

H [Duf46, LPN11, Wig63a, LPN11]. H8 [HMW05]. half [Gau94]. half-century [Gau94]. Hall [Wig78a, Wig68j]. Hamiltonian [Ell92, Wig37b]. Hamiltonians [Beh91, Beh94, RV10]. Hamming [Isl22]. Hamos [Wig73r, Wig73r]. Handbook [OLBC10, Wig71i]. Handbuch [Wig56c]. Hanford [Ban04, San95, Wei04]. Harbor [MWW66, PW64a, PW64b, Wig65k, Wig66c, Wig69h, Wig69i]. Harder [Wig69e]. Harkabi [Wig67o, Wig67y]. Harmonic [KW88c, LSV06, LSV08, KPT86, KW88a, Pal82, RV11, WW30a, dlPC79]. harmonics [Dra86]. harmonischen [WW30a]. harmony [Ano77b]. Három [Frö93, Lad93]. Hartree [OG07]. hasonmáskiadása [Wig02a]. határa [Wig64-30, Wig64-31]. határai [Wig64-29]. hatékonysága [Wig64-27]. hazai [Frö93, Lad93]. hazards [WÁ68]. He-cooled [ACF^{+}42]. Hearing [Wig76d]. heartily [WB77]. Heaven [TW79b]. Heavy [Wig52c, WOYW56a, FW36]. Heisenberg [LW34, Bri65b, Eck33, Wig76h]. Held [Meh73, BW61, DGW84, GHM^{+}87, Gre86, Hoo73, MR76, Sal69, Wig64f, Wig64g, Zic78, Zic83a, Zic83b, Zic85, Zic88]. Helen [Wig79j]. Helicity [Bec72, Jor78]. Helium [FW37, Wig33a, Wig65a, GBO04, KPM97, LCM^{+}42]. helium-like [GBO04]. Hellenike [Ano77a]. help [TW79b]. helps [Ano15]. Henry [Ano78a, LS76]. Herbert [Wig76c]. Hermitean [Wig65f]. Hermitian [FKS97]. Hetaireira [Ano77a]. heuristic [McD17]. Hevesy [Stu05]. Hevesy [?]LorándStuewer:2005:GHL. Hidden [Cla71b, Wig70g, Wig71j, Wig76i, Wig76j]. Hidden-Variable [Cla71b, Wig71j]. Higgs [dDG83]. High [Mac04, Nau73, NWEC61, Wig88c, MM73, Wig73r, Zic83a]. high-energy [Zic83a]. High-Frequency [Mac04]. Higher [WE47a, WNY^{+}84]. Highlights [Zic73, Wag81, Wag98]. Hilbert [BG02, BG03, Mol99]. Historical [GL23, Seg85, Wig95b, Wig98, Wig01b, Jan14, Rom00]. History [GHM^{+}87, San95, KHFA67, LK04, Sch96, Wig83f, Wig50b]. Hitler [Mar06b, MP01]. Hoffman [Wig79j]. hogy [Wig88c]. Hogyan [Wig02a]. Holes [TW53]. Homodyne [BGA07]. homomorphic [SW67]. honor [F^{+}76, LW78, MF69, Wig73e]. honorary [Wig66b, Wig83e]. Honoring [Fin60]. Honors [Wig62b]. Honour [Ble86, Sch85, Tor87, BVV84, dBDW^{+}84]. hope [Wig78e]. hopeful [TW79b]. Hopf [CCT08]. hopping [KLW09]. Horne [Lee22]. Hotel [BW61]. House [Duf46, Wig65c, BTW73, TWB73a, TWB73b]. Houston [Her79]. Houtermans [ABES12]. HTSC [SM05]. Hudson [Jan84]. Hughes [Wig57h]. Human [Ano48, Duf46, Dar47, Wig76l, Wig77h, Wig79j]. humanist [Rub82]. Hungarian [KW01, Wig02a, EGW69, Frö93, GW03, Har06a, Jon10, Lad93, Lan93a, Mar96, Nag88, NW70, Pal88, Stu05, Wag98, WWW64, Wig29b, Wig29c, Wig33d, Wig35c, Wig64z, Wig64-27, Wig64-28, Wig64-29, Wig64y, Wig64-30, Wig64-31, Wig65u, Wig66k, Wig67-29, Wig68j, WÁ68, Wig69y, Wig71m, Wig71n, Wig72m, Wig72n, Wig72o, Wig73q, Wig73r, Wig73s, Wig75g, Wig76o, Wig79x, Wig88c, Wig88d, Wig88e, Wig92d, Wig92e, Wig92f, Wig95c, Wig95d, Wig02a, Wig02b, Wig02c, Zsa76].

Hungarians [Lan11]. Husimi [TKC^{+}08]. Hydrocarbon [Wig59a, Wig60a]. Hydrodynamics [Tau63c]. Hydrogen [WH35, Wig54b, FW36, Wig33c]. hyperbolic [WB05]. hyperboloids [APW02]. hypergeometric [ÖAÖ25].

I. [Wig73a]. ICCMSE [SM07]. Ideals [Wig79j, Wig79d]. Ideas [GHM^{+}87, Gre86, Hor97, MGM62, Wig59d]. idempotency [MYPL90]. Identity [LB71]. idotükrözés [Wig64z]. if [GW72b]. Igal [Wig63a]. ignition [Wig83l]. ihre [Wig31b, Wig44]. II [Wig26b, Wig96m, APW03, BL72, Bru87, DK70b, GS04b, Goe49, Gou86b, Löw64, PD98, Tau61b, WS34, Wig57b, Wig82c]. III [CvV82, CM80, DK71, LB73, Tau63a, Wig97a]. Illus [Wig64f, Pai67a, Wei69a]. Illustrious [Fer68, Fer71]. Image [HS01, CGPT21]. immigrants [Fer68, Fer71]. Impact [Wig51b, Wig73d, Wig73f, Wig79u]. implementation [Zie88]. implications [LB73, Wig27a, Wig27b]. Important [TC10, Wig88c]. inaugural [Bug77]. inauguration [Bug77, Wig66f, Wig75c]. including [Nob72, SM05, Wig64e]. increased [WW62a, WW62b]. incredible [Hed93]. independence [PPG^{+}19]. Indiana [Pai67a, Wei69a]. Indispensability [Gel14]. inequivalence [BH13]. Infinite [Wig50a, Wig55a, Wig57b, DK73b, DK78]. inflow [JCT11, JLC14]. influence [Wig74d, wYDK74]. Informal [Cla71b, Wig71j]. Information [Luo04, WY63, GI03, K^{+}84, LLW^{+}09, Yan10]. informational [HS95]. informationally [Bo09]. Inhomogeneous [Wig39b, Wig79a, BW73, Duf71, Wig64e, Wig89a]. Initial [Wig83f, Wig95c]. inner [CIMS08]. Innocence [Stu18]. Inönü [FS82, Hud78, Maj93, WW95]. Inönü-Wigner [Hud78, Maj93, WW95]. Inquiry [Gib19]. Inspiration [Wig62a]. instabilities [Fol80]. Instantaneous [VW66]. Institute [Bug77, MR76, Zic78]. Instruments [L.88]. insulating [SM06]. Integer [JF16]. integrability [RNM04]. integrable [BLS03]. Integral [Lie90, Ber06]. integrals [DF77, WB05]. Integration [Win76]. integrations [CV80]. intellectual [Fer68, Fer71, Wig74d]. Interacting [VW65, VW66, KPT86, Pal82]. Interaction [CW41, EW41, LTW55, MDH94, Wig34, WE47a, HMW02, TKT^{+}05, Wig38a, Wig69c, Wig69b]. Interactions [Kas21, WNY^{+}84, ML91, SP74, Zic83b, Zic84]. Interest [L.88]. interesting [Pal88]. interface [WWCS42]. interference [PW28]. Interferenz [PW28]. Interjú [Wig73r, Wig73q]. interjút [Wig73q]. Intermediate [Wig54e]. Internal [WK72, Wig65d, Wig65m]. International [Ano54b, Ano75, Ano77b, Ano78b, Ano79b, Ano83a, BH71, DGW84, GHM^{+}87, KPS78, Meh73, Rub82, Sal69, SM07, Vil85, Wig68e, Wig69w, Zac84, Zic78, Zic83a, Zic83b, Zic85, Zic88, d’E71, F^{+}76, KR78, MR76, Ano68, Wig51a, Wig51e]. Interpretation [Hug89, LTW55, NW70, Wig54f, Wig66h, Wig83g]. Interscience [Wig64f]. Interview [DMS84, Olt74a, Wig59c, Wig71d, Wig72f, Wig75d, Wig83j, Wig96j, Wig73q, Wig73a, Wig77b, DMS84, Wig73q]. Interviews [BH89]. intracule [GBO04]. intrinseco [Maj32]. Intrinsic [WWW52, Maj32]. Introducing [Wig66d]. Introduction [Duf46, Wig73e, Wie53, Wig86a, EW61, WBBE30]. Introductory [Wig72g, Por65]. Invariance [HVW65, Wig49d, Wig57f, Wig60b, Wig64o, Wig64p, Wig72b, Wig82h, Wig85a, ZDF^{+}84, HKW68, HVW82, Vig58, Vig65, Wig56a, Wig56f, Wig56b, Wig64n, Wig64l, Wig64d, Wig65u, Wig67a, Wig84f, Wig51c]. invarianciaelvek [Wig65u]. Invariant [AW75, EW41, Fro79, Tem07, Wig52b, Wig63b, HL88, Wig67h]. invariantnost [Vig58]. invariantnosti [Vig65]. Invariants [KAUZ71]. Invarianz [Wig51c]. Invarianzprinzipien [Wig64n, Wig64l]. inventai [Wig79g]. invented [Wig79g]. inventory [KHFA67]. Involving [HW39]. ionic [Mar03]. ions [GBO04]. Iowa [Wig67t]. Iranian [Wig76f, Wig77d]. Irradiator [WSCF61]. Irreducible [Kib76, Wig71c, Ell92, GW72b, KP90, Lit72]. Isa [Wig78i]. ISBN [Ble86, Hen89, Sch85]. iskolára [Wig73s]. Isobaric [Wig66g, FR66]. isomorphism [MMZ^{+}05]. Isotope [WOY57]. Isotopic [Wig58d, Wig79g]. isotopico [Wig79g]. issue [CW74, JKM03]. István [Wig73q, Wig95c]. Italian [Maj32, Maj37, Wig77c, Wig77e, Wig79g, Wig79o, Wig79s, Wig79t, Wig79w, Wig79q, Wig79p, Wig80b, Wig82j, Wig83l]. Italy [DGW84, Meh73, Wig68e, Zic85, Zic88]. itself [Nar88]. IV [DK73a, HH04a, Tau62, Wig97b]. Ivan [Wig92f, Wig92f]. Iván/ [Wig92f]. IX [PFG06].

J [BB93, Gut60, L.88, Maj05a, Pos01, Sac59, Sch67, Sch85, Szm99, Wig61e, Wig67t, Lai94]. J.-P [Sch85]. Jacketed [SWC59, WSC59]. Jagdish [Rom00]. Jaki [Wig69e, Fla68]. James [Bie69, Wig78j]. János [Wig67-29, GW03]. January [Her79, SVW98, SVW00, Wei97, Gre86]. javlenija [Vig58]. Jean [Ble86, Tor87]. Jean-Pierre [Ble86, Tor87]. jeffektivnost [Vig68]. Jeno [Nag88, Wag98, Wig72o, KW01, Wig72n, Wig02c, Zsa76]. Jenövel [Pal88, Wig88c, Wig73q]. Jersey [Wig64b]. Jews [Mar06b]. jó [Wig73r]. Joe [Wig76b]. John [Stu88, Wig53a, Wig92f, GW57, GW03, Hor97, NAH^{+}89, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, Wig57d, Wig67b, Wig67-29, Wig80a, Wig96c, Wig96h, WG96, von96, vN96]. Joint [Wig76d]. Jones [L.88]. Jost [Kla91]. Journal [Rom91]. Journey [FR13e]. József [Wig88c]. Jr [Bie69]. Jr. [Wig78j]. jubilee [F^{+}76]. Jucys [Wu73]. July [Ano54a, Zic78, Zic83a, Zic83b, d’E71, Dan95]. Jump [SSP^{+}91]. June [Ano79a, Sal69, Wig68e, d’E71].

K. [Ano48]. Kac [Maj93]. Kahan [Wig66a]. Kampen [CM68]. Kardos [Wig73q, Wig73a]. Kargon [L.88]. Kármán [Hor97]. karrier [Wag98]. Kathlerine [Duf46]. Kausalitätsprinzip [Wig79n]. kedvezo [WÁ68]. Keldysh [AK00]. Kellner [GBO04]. Kelvin [L.88]. kémiai [Wig29b]. Kepler [Fro79]. Kernphysik [EW61]. Keseru [Wig66k]. készítette [Wig73q]. Kettenreaktion [Her76]. Key [Wig82e]. kezdõ [Wig95c].

Khubas [Wig78i]. kilátások [WÁ68]. Killians [Wig78j]. Kilmister [Hen89]. Kinds [Wig64u, Wig64v, Wig67j]. kinematic [Wig54h, Wig49e, Wig49b]. Kinetic [BT73, BT76, Fon94, Maj05a, Maj05b, MYPL90]. kiredukálják [Wig71n]. Kirkwood [Sha92]. Klein [SS05]. knew [Wig92f]. know [Wig88c]. Knowing [Wig96l, Vig86, Wig80e]. Knowledge [Ano61, Jha11, Gre69, Wig74h, Wig76l, Wig77h]. kombinierende [Wig26a, Wig26b]. Konvergenz [Ahl70]. korba [Wig75g]. Korzybski [Wig66d]. köszöntése [Nag88]. kötés [Wig29b]. következtetés [Wig64-28]. Közreadta [Wig72o]. Kramers [TCT16, Wig75c]. Kratzer [Sta10]. Kronecker [Wig71n, Wig65i, Wig71n]. Kronecker-multiplied [Wig71n]. Kronecker-szorzatát [Wig71n]. Kursunoglu [Hen89, L.88, Ram89, Sch88, Stu88]. kvantovye [Vig58]. kvantummechanikában [Wig64z, Wig79x]. kvantummechanikája [NW70]. Kyoto [Ano54b].

L [Fla68, Wig65c, Wig69e, Wig72d, Wig95c, NW70, Wig95c]. L. [FW52, Wig56e]. Laboratory [Lan93b, Wig46b, Lan99]. Lagrangian [LR83]. Lagrangians [LR78]. Lake [d’E71]. Lancelot [Wig65m, Wig65d]. Landed [Lan93b, Lan99]. landmark [Wig66f]. Langevin [Gar88]. language [Wig74h]. Large [Bai93, JD90]. Largest [BY88]. Lattice [SIM93]. lattices [CMS77, Fol78, Fol80]. Laudisa [Wes21]. Laue [Wig50b]. laureate [Zsa76]. laureates [Hoi01, Nob72]. Law [Wig52c, Wig65d, Wig65m, DGW34]. Laws [VW66, Wig64o, Wig64p, Wig64t, Wig72b, Wig85a, Wig85c, Vig64, Vig65, Wig54c, Wig54h, Wig64n, Wig64s, Wig65f, Wig65u, Wig65p, Wig65q, Wig65r, Wig65s, Wig67a, Wig79q, Wig82h, Wig83f, Wig95c]. Leakage [WSCF61]. Learned [Wig80f, Wig73o]. leaves [Wig72o]. Lecture [Wig64n, Wig68e, Wig74h, Wig64o, Wig64p, Wig67a, Wig67t, Wig73b, Wig65u, Wig88d]. lecturer [Wig66d]. lecturers [Ano30]. Lectures [L.88, Tal68, Ano68, Bug77, Nob72]. Leerstoel [Wig75c]. legacy [Tis03]. leggi [Wig79q]. Lejegyezte [Wig95c]. lejegyzésében [Wig02c]. Leó [Wig92e, Fra05, Hor97, Wig69a, Wig92e, Wig96i]. Leonard [Bec58, De 58, Edw60, Kin58, Set58]. lépni [Wig75g]. lettem [Wig02a]. Letter [Ein39, Kue61, Lap60, Mei61, WW61a, Wig58e, Wig64c, Wig67u, Wig71f, WA72, Wig72h, Wig73g, Wig75b, Wig76g, Wig79r, Wig80c, Wig96j, Dem15, Wig67-27, Wig69l, Wig76f, WW86]. Letters [BJH68, Ger75, Gir75, Gul75, MW75, Sau75, Smi75, Wei75, Wig69o]. Level [Wig49c, DK04, FS82, RNM04, Wig57c, Wig65g]. levelei [Wig72o]. levels [Wig38a, WF41, Wig51g, Wig65j, Wig69s]. Leverett [ACF^{+}42]. Lewis [Wig75b]. Li [BW37, WB36]. library [DLD^{+}21]. Lichttheorie [WW30b]. Lie [Agr80, BB65, BW73, CLM76, Con72, Gin63, GW72b, Gou81, Gou86a, Gou86b, Hud78, Kol73, Mez77, MACS04, MMZ^{+}05, WW95, Wig73j, ZY88]. Life [Pai67b, WL65, ABES12, Ano68, BDH^{+}89, DMS84, Wig70h, Wig71k, Wig74f, Wig79p, Wig02b]. Lifetimes [Wig39a]. lifting [FS82]. Light [Cox61, KW87c, Wig39a, Wig61a, Wig96j, BH13, FW36, KW87a, WW30b, WW86]. like [GBO04, OK78, SM06, WI98]. Limit [Wig55c, Zha03, AN09, Ara95, Wig64-30, Wig64-31, Zic83a]. limitation [CW74]. Limitations [SW58, Tem07, SV86, Wig82i, Wig83h, Wig89b]. limited [BWD76b, HMW02, Wig76d]. Limits [Wig50d, MMP94, RdGS01, Wig51d, Wig53b, Wig67c, Wig78h, Wig09]. Lindau [Ano79a]. Linear [CvV82, KS82, BW73, Bru85]. linewidth [WW30b, WW30a]. Linienbreite [WW30b, WW30a]. Liouville [BT76]. Liouvillian [MR98]. L’irreale [Wig79s]. little [HKS86, Vas89, Wig69h, Wig69i]. Littler [Wig57h]. locality [FW73]. localizability [Bac89]. Localization [Hal01, LO10, HS95, PS02, Ram00]. Localized [KW87c, NW49, KW87a]. Loewner [WvN54]. Logic [Ano61, Tau61a]. London [Ano65, Hoo73]. Long [WE47a, Ano19]. Longer [WW60, Wig67d]. look [dlPC79]. looks [Wig68b]. Loránd [Stu05]. Lorentz [MR73, Dir45, Ker85, KW90b, KW90c, MR72, SW67, Vas89, Wig39b, Wig64e, Wig79a, Wig89a, Wig90a]. Losev [Wig69d]. Loss [TW53]. Loss-Free [TW53]. Lost [Sau75]. Louis [Ble86, Sch85, Tor87, WBBE30, BVV84, F^{+}76, dBDW^{+}84]. Low [Con72]. Low-Dimensional [Con72]. Lower [Wig55c]. lumps [Wig57i].

M [Coh59, Gla59, Kol60, Kou59, LW78, Wig53a, Wig66d, Wig96j]. M. [Wig72d, Wig73e]. M.I.T [Kir73]. Machine [Lan83, Ano15]. Machines [Wig69g]. Madison [BH71, Wig67l]. Magic [Ne’04]. Magnetic [MW40, LM89]. Magnus [Wig67q, HW57a, HW57b]. Magyar [Wig02a]. Majorana [BW35, Bri65b]. Makers [Ano52]. Manhattan [Lan93b, Lan99, Lan11]. Manifestations [OW84]. mankind [Wig86e]. Manpower [Wig69p]. Many [HW01, HMW01, PS97, PMHW07, SD96, Ste25, CPCF99, Wig59f]. Many-and [Ste25]. Many-Body [HW01, HMW01, PS97, PMHW07, CPCF99]. Many-Electron [SD96]. map [DP01]. Maple [Lai94, PFG06]. mapped [Sta10]. March [Ano61, Bug77, FR66, MR76, Wig73q, W^{+}75, WJH76, WH77a, Wig02b]. Märchen [Wig70d]. március [Wig73q, Wig02b]. Margenau [LS76, Put64]. Maria [Wig72i]. marslakó [Har06a]. Martians [Har06b, Har06a, Lan93b, Lan99, Mar96]. Marx [Wig88d]. Maryland [Zac84]. Mass [Kir73, Vig86, Wig33a, Wig65a, Jor78, SS05, Wig56e]. Massachusetts [Ano79b]. masses [Wig41a]. Massey [Wig63a]. Massive [WSCF61, LR83]. massless [KW87b]. master [Duf71]. Masters [Duf46]. Matematica [Wig77c]. matematika [Wig64-27]. matematiki [Vig68]. Material [Wig61a]. Math [BB93, Maj05a, Pos01, Szm99, Wig85d]. Mathematic [Gel14]. Mathematical [Ban09, BW61, Mac59, MR76, Mic85, OLBC10, Sec62, Wig56c, Wig61f, Wig66a, Wig67q, WE47b, Wig50e, MR76]. Mathematics [BW61, Cut63, How62, Isl22, Nic12, Omn11, Sec62, Tis98, Wig60j, Wig60k, Wig60l, Wig67p, FF91, Gau94, Hed93, Jan14, MR76, Vig68, Wig64-27, Wig67k, Wig74h, Wig77c, Wig85d, Wig90b, Wig91, Wig92b, Gau94]. Mathematik [Hed93]. Matrices [Bai93, BY05, FP03, FKS97, LY09, Wig55a, Wig57b, Wig58h, Wig67w, Ell92, Kot05, Kot07, MYPL90, PFG06, SW67, Wig59f, Wig63d, Wig65h, Wig65i, Wig69s]. Matrix [BY88, DW64, GLB92, IW54, QM71, RW54, Wol71, Wol72, CLM76, DK04, GW72a, Gou81, GP86, JD90, Kot05, NW71, SCS99, Szm98, Szm99, Wig52e, Wig52f, Wig64a, Wig64h, WY64, Wig65f, Wig71n]. mátrixokról [Wig71n]. matter [MF69, Zic78]. Maurice [Hen89, KW87d, Ram89, Sch88, Stu88, Wig88b]. Max [Wig50b]. maximal [BL72]. Maxwell [Scu07a]. May [Wig64b, Zac84, ZHH60, Wig71n]. Mayer [Wig72i]. Mean [AN09]. Meaning [Duf46, Hor97, MW46, MW07, Wig84c]. Means [COWY57, NW61, WOYW59, WW61b, Wig61h, LLW^{+}09, ML91, Wig67g]. Meant [LB96]. Measure [MPPS02, MPPS03, Zha03, Bo09]. Measurement [Ano78c, Jha11, KPM97, SW58, Wig57a, Wig60d, Wig63e, Wig86b, Gre86, MS83, SV86, WZ83, Wig52a, Wig62h, Wig67f, WY73, Wig78g, Wig83i, Wig83k, Wig84e, Wig86f]. Measurements [CJSW55b, JWY67]. Measures [Mac04, AN09, KL03, Ker05]. Mechanical [CHS97, HW39, JvNW34, Wig50c, Wig70g, Wig71a, Wig76i, Wig86b, DK71, LM89, OW81a, OW81b, OK78, SG75, Wig52a, Wig56f, Wig62h, Wig63b, WY73, Wig76j, Wig83k, Wig84e, Wig86f]. Mechanics [Ano78c, Ber77, Cob72, Fan60, For97, Fre15, GL23, Gut60, Hug89, IÖRZ67, Sac59, Tau61a, VW65, VW66, Wes21, Wig31b, Wig47c, WL65, vNW28a, vNW28b, vNW28c, Ano19, BBDP08, Cha01, DP01, F^{+}76, JWY67, KW89a, KW90a, MR98, MACS04, MMZ^{+}05, NW70, Rob93, SP74, Wig25, Wig27c, WW28, Wig29c, Wig32c, Wig44, Wig56a, Wig56b, Wig59g, Wig64z, Wig64q, Wig68a, Wig69m, Wig69n, Wig73l, Wig79e, Wig79x, Wig83g, Wigxx, d’E71, Rom00, WBBE30]. Mechanik [Wig25]. mechanika [Wig29c]. Mechanism [BW39, Whe09]. Medal [Wig78k]. media [WWCS42]. Medium [KS07]. Meet [Ano30, Lee22]. Meeting [Ano79a, Bal19, GHM^{+}87, Wig68e, Wig82l, Wig87b, Wig96a, Her79, WU75]. meetings [LK04]. Meets [Fle00]. még [Wig88c, Wig88d, Wig88e]. megalapítója [Wag98]. meghökkento [Wig64-27]. Mehra [Rom00]. Melléklet [Wig02a]. member [Wig02b]. Memoir [Wig69a, Wig92a, Wig96e, Wig79i]. Memoirs [Wig02c, Wig02a]. Memorable [Rom91]. memorial [Ano78a, FW52, Wig66d, Wig87b, Wig96a]. memories [Wig66k, Wig96g]. men [Wig67x]. menschliches [Wig76l]. Mercer [Wig64b]. merkwürdige [vNW29b]. mert [Pal88]. Merwe [Ble86, Sch85, Tor87]. meson [WE39]. Mesoscopic [AF95]. Message [Wig95d]. Messung [Wig52a]. Messung. [Wig62h]. Metal [WWY58b]. Metallic [WS33, WS34, WH35]. Metallurgical [Wig46b]. Metals [Wig34, WB35, Wig38a, Wig54i, WS55]. Metaphysics [CHS97]. Meteorology [Tau63c]. Method [AN92, AR96, DW64, IÖRZ67, LPN11, MH94, MDH94, NAKS04, WM58, WYW61, Wig61g, WOYW62, Wol71, Arn94, Kib76, Maj05a, Maj05b, PČH02, SNDS14, SAZ90, TCT16, Wig38b, Wig77e]. Methods [Ada70, HMW01, Ler90, Ros70, SW70, SM07, Wei69a, Wei69b, Wig69x, Wig70b, Wig73j, Zac84, DGW84, GW63, JC10, Kol73, KR78, RRC05, Wig50e, Wig79x, Wig56c]. Metodo [Wig77e]. MgC [TKT^{+}05]. Michael [Sch67, Ano61, WJH76, WH77a, Wig02b]. microcosm [Wig79q]. microcosmo [Wig79q]. Microscopy [RB92]. migration [Fer68, Fer71]. Mihály [Wig02b]. Military [NWA72, WA72]. Milne [Wig49b, Wig49e]. Mind [Elv82, Wig62f, Wig67i, Wig69e, Pai67b]. mind-body [Wig62f]. Minimally [vNW40]. Minimax [BGA07]. Minnesota [Wig74c]. mint [Wig72m]. Minutes [LK04]. Miramare [Meh73]. mirror [Wig79s]. Mirroring [Wig64z]. Missiles [HW77, HW74, Wig72e]. misunderstanding [FW73]. Mixed [BW95, Arn94, LLW^{+}09]. Model [AK00, Kas21, KNS04, RW54, Wig55d, HMW05, Wig52g, Wig65c, Wig67e]. Modeling [MH94]. Modelling [Tem07]. models [BLS03]. Moderated [SWC59, WOYW56a, WOYW57]. Modern [Anoxxa, Anoxxb, Gib19, L.88, SM07, Stu04, Vog95, Wig56d, Wig81a, Bug77, Wig26a, Wig26b, Wig72j]. Modernization [Rub82]. Modification [WH35, Wig35b, Wig35c]. módosításáról [Wig35c]. módszer [Wig79x]. Modular [Ram00]. module [Jan14]. Modules [BG02, BG03, CIMS08, Mol99]. Moldauer [Wig61e]. molecular [Kib76]. molecule [WW28]. Molecules [SD96, PW25, Wig25, Wig67x]. Molekelspektren [WW28]. Molekülen [Wig25, PW25]. Momenta [WE47a, CM66, PFG06, SG75]. momento [Maj32]. Moments [MW40, Kot05]. Momentum [OW77, OW78b, AW75, BV65, CW71, Maj32, OW78a, RV11, Wig77a]. mondo [Wig79s]. Monographs [Sch85]. monopole [Dra86]. Monovalent [WB35]. Monte [KLW09, SNDS14]. Moody [Maj93]. Moore [Sch67, ACF^{+}42]. Morse [SK05, Sta10]. MOSFET [JC10]. Moshinsky [Wig73e]. Motion [Wig50c, BT73, RL81, Wig63b]. Mott [CHDK08]. move [Wig75g]. Moyal [DV97, SS05]. MPI [Zic78]. MR [HMW05]. MR-BWCC [HMW05]. MRST [Zic78]. Much [Ger75, WNY^{+}84]. muhöz [Wig69y]. multidimensional [LO10, SC83]. multilevel [Dav74]. Multiple [Fra05, Wig54a]. Multiplicity [CCB72, Gou86a, Gou86b, PD98, Won76, Won79]. Multiplicity-Free [CCB72, Gou86a, Gou86b, Won76, Won79]. multiplied [Wig71n]. Multireference [HMW01, HMW05, MH98, Wen98, HMW02, PMHW07, TKT^{+}05]. multispinor [LR78]. multispinors [RL81]. Multiword [JF16]. munkássága [GW03]. Must [Wig56e]. Mutamenti [Wig79t]. Muth [Wig76e]. My [Wig85b, Wig79p]. Mystery [FR13e, Scu07a]. mystics [Scu07a]. myth [Mar96, Wig70d, Wig70f].

N [Hen89, L.88, Ram89, Sch88, Stu88]. N.J [Vig86]. nach [WW28]. nagyra [Pal88]. named [Ano30]. nano [JC10]. nano-scale [JC10]. nanostructures [RRC05]. Nanotubes [dSFA05, DB08]. narrative [Com56]. Narushenie [Vig66]. Nashih [Vig68]. National [GR63, Lan93b, Her79, Wig66f, Lan99]. NATO [Zic78]. NATO-MPI-MRST [Zic78]. Natural [Tis98, Wig60j, Wig60k, Wig60l, Vig68, WW30b, WW30a, Wig64-27, Wig64l, Wig64d, Wig67k, Wig82h, Wig85d, Wig86d, Wig90b, Wig91, Wig92b]. Naturalism [Ban09]. Nature [Wig64o, Wig64p, Wig72b, Wig76b, Wig85a, Wig86b, Elv82, Meh73, Vig65, Wig64n, WY64, Wig65u, Wig67a, Wig70i, Wig72k, Wig83f, Wig86f, Wig95c, Zic88]. Naturgesetze [Wig64n]. natürliche [WW30a]. natürlichen [WW30b]. Naturphilosophie [Wig64l]. Naturwissenschaften [Hed93]. naukah [Vig68]. Nazi [MP01]. Near [Wig48f, Kla91]. Necessary [BY88, MFAB85, Wig71e]. Need [Smi75, TW79b]. Needs [TW79a, WU75, Wig76a, WT79]. negative [OW81b]. neglected [CW74]. néhány [NW70, Wig64-28]. nella [Wig79t, Wig79p]. nelle [Wig79q]. nem [Wig88c]. Nepostizhimaja [Vig68]. neuer [Wig29a]. neueren [Wig26a, Wig26b]. Neumann [Beh91, Wig57d, Wig67b, Wig80a, Wig96h, Beh94, Cla71b, Frö93, GW57, GW03, Hor97, Kla91, Lad93, Lan93b, Lan99, NAH^{+}89, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, Wig67-29, Wig71j, Wig96c, WG96, vN96]. Neumannról [Frö93, Lad93]. Neutrality [MW75]. Neutron [Coh59, Gla59, Kou59, Rog13, SW61, Wig57a, WYW61, Kol60, WW58, Wig33d, Wig57c, Wig65g, Wig79u]. Neutronen [Wig33b]. Neutronic [COW^{+}59, OWWY58, OWWY59, WOYW56a, Wig57e, Wig58c, WO58, WWY58a, Wig58f, Wig58g, WC60, Wig60e, Wig60f, Wig62e, Wig63c, Wig01a]. Neutrons [BW36, CJSW55a, CJW55, CJSW55b, FR13a, WCJS55, WCJS56, Wig65b, Wig33b]. News [Ano52, Wig51e]. News-Makers [Ano52]. Newton [Fle00, Hal01, Jor80]. nicht [Hed93, Wig26a, Wig26b]. Niels [Duf46]. Nincs [Wig88d, Wig88e]. nine [Mar06b]. nineteenth [Zic83b]. NIST [OLBC10]. NITM [Ell92]. NMR [Tem07]. No [BAA^{+}75, Gir75, Wig82d, Ano19, GW72b, Jor78, Wig88d, Wig88e]. Nobel [Ano79a, Wig64n, Wig65u, Hoi01, LW34, Wig64o, Wig64p, Wig65u, Wig67a, Wig82g, Wol67, Zsa76]. Nobel-díjas [Zsa76]. Nobel-eloadás [Wig65u]. Nobel-Vortrag [Wig64n]. Nobelmen [Ano63]. Nobelwoman [Ano63]. Nodal [SD96]. Noise [Gar88]. Noisy [BGA07]. non [OW81b, Wig79w, Wig86f]. non-negative [OW81b]. non-relativistic [Wig86f]. Noncanonical [Pal82]. Noncircular [Fre93]. noncombining [Wig26a, Wig26b]. noncommutative [BBDP08]. None [Duf46, MW46, MW07]. nonlinear [Bru87]. nonmagnetic [RI03]. nonnegative [Jan84, SC83]. Nonrelativistic [Wig86b, CM68]. Nontrivial [Bru85, Bru87]. nonzero [Jor78]. Normal [Wig60g]. Normalized [Ell92]. not-so-incredible [Hed93]. Note [Bar64, BW35, Maj05a, Wig54g, Jan84, Pos00, Pos01, Wig29a]. Notes [Ban04, Wei04, Wig51e, Wig70e]. notions [Bac89]. nött [Pal88]. Nov [Wig82l]. növekedése [WÁ68]. November [Ano75, Ano77b, Ano78b, Ano79b, Ano83a, KPS78, SVW98, SVW00, Wei97, WNY^{+}84, Wig88d, Wig02a]. nucléaire [Wig61d]. Nuclear [BHT86, BAA^{+}75, BW39, BW38, Bri65a, COWY57, COW^{+}61, Cut63, EW41, EW58, EW61, FBR^{+}41, How62, KPS78, MR76, Mla98, MW79, OYWW61, Ore58, Seg85, Stu18, TW52, Wat93, Wig37b, WCT39, Wig41a, Wig49c, Wig51a, Wig51e, Wig58a, Wig58e, Wig59a, Wig60a, WY60, Wig62c, Wig62b, WOYW62, Wig67o, Wig68h, WU75, WH77b, Wig78f, Wig79r, Wig79v, Wig79b, Wig82l, Wig92c, Wig96m, And73, BH71, Eck33, EGW67, FR66, Her76, Hod76, SW59, Wag98, WW60, Wei02, WF41, Wig48g, Wig51g, Wig53a, Wig54e, Wig57h, Wig59d, Wig61f, Wig65j, Wig66g, Wig67u, Wig67d, Wig67y, Wig69y, Wig70i, Wig74a, Wig76d, Wig76g, WH76, Wig82j, Wig83l, Wig85b, BW61, Her76, KPS78, Set58, Stu79, Wig63a, Bec58, De 58, Edw60, Kin58, Wig67o, Wig67y]. nucleare [Wig82j, Wig83l]. Nuclei [FW37, Goe49, MW40, Wig37b, Wig37c, Wig39a, Wig52g, Wig58d, ZHH60]. Nucleon [LTW55]. Nucleus [FR13e, LTW55, Wig55d, EGW69, Wig67e, Wig65c]. null [BL72]. number [Wig58d]. numbers [Ne’04]. Numerical [AN92, RRC05, Tau63b, CMS77, GS04a, SAZ90].

O [Ble86, Sch85, Tor87]. Oak [Wig59c]. Oberhettinger [Wig67q]. Obituary [CFW78, Sei95, Wig72i, WJH76, Wig76h]. Objections [Cla71a]. objective [Ano19]. objectives [Wig69q, Wig58i]. observability [GW72a]. observables [HS95, Pos00, Pos01]. Observation [Ste25]. observer [PPG^{+}19]. observer-independence [PPG^{+}19]. observers [Ano19]. Obstacles [BAB^{+}82]. occasion [F^{+}76]. octahedral [Ell92]. October [Wig64f, Wig64g, Wig74c, Wig75c, Wig82e]. Odd [MW40]. Oesper [Wig50b]. offer [WU75]. Okrent [Wig61e]. Old [Zic88, Wig68f, Zic88]. once [GW72b]. One [Ano06, CvV82, CHS97, DFS20, Gib19, MW46, MW07, Ste91, CMS77, DK70a, DK70b, DK71, DK73a, DK73b, DK75, DK78, DB08, Fro79, RV10, Wag98, Wig78e, wYDK74, Wig26a, Duf46]. One-body [CvV82]. One-Dimensional [Ano06, Ste91, DK70a, DK70b, DK71, DK73a, DK73b, DK75, DK78, DB08, Fro79, RV10, wYDK74]. ones [Wig66k]. only [Wig73k, Wig74g]. Onsager [Wig54d]. onslaught [K^{+}84]. Ontario [Hoo73]. onto [Sta10]. opening [Wig79f]. Operation [MW79, Wig32c, Wig79b, Wig64q]. Operations [Bar64]. Operator [AR96, NAKS04, Szm98, Szm99, Arn94, Bo09, Jor78, Jor80, Sha73, ZY88]. Operator-Split [NAKS04]. Operatoren [Wig52a]. Operators [BB65, BG02, KS82, Mac04, OW78b, Tau63a, WK72, Wig60g, Wig60h, Bec72, BL72, LNPC92, LB73, MS00, MM73, Mez77, OW78a, OW92, Wig52a, Wig78d, Won71, Tau61b]. Opinion [Fel67, SW56a]. opinions [Frö93, Lad93]. Oppenheimer [Wig76c]. Optical [JW87, Wig65c]. Optics [FM03, Maj05a, Maj05b, MS83]. options [Wig76n]. Oral [San95]. order [PMHW07, Ras89, VGB87]. Organization [Wig64b, Wig69w]. organized [MR76, Sal69]. original [BV65, Por65]. origins [TW79b]. ortho [Wig33c]. orthogonal [FB88]. Orthogonalization [SW70]. Orthowasserstoff [Wig33c]. Ortvay [Wig72o]. oscillations [Wig20]. oscillator [LSV08, OK78, RV11, SK05, Sta10, WW30a, dlPC79]. Oscillators [KW88c, CCT08, KW88a, LSV06, WN90]. Oscillatory [Dah95]. Osipov [Wig69d]. Összetett [Wig29c]. Oswald [Veb60]. Oszillators [WW30a]. öt [Har06a]. Other [L.88, Tau62]. ötven [Wig76o]. Our [DHW49, Wig76a, Wig82d, WNY^{+}84, Wig71b, Wig86d]. ours [GW71]. Outcomes [Kas21]. outer [PD98]. outlook [WÁ68]. Overall [NWEC61]. Overlapping [KW88b, KW89b]. Overview [Wig64y]. Oxford [Wig51a, Wig51e]. Oxide [CJW55]. Oxygen [FW37, Wig37c, PČH02].

P [Ada70, Ano48, Bec58, Bel73, Bie69, BH89, Cha01, Cho68, Coh59, CH68, Cut63, De 58, DMS84, Edw60, Fan60, FR13b, Fes60, Fla68, Gla59, Gut60, Har99, HH04b, Hen89, How62, Kin58, Kir73, Kol60, Kou59, L.88, LW78, Mar69, Mar95, Pai67a, Pai67b, Ram89, Ros70, Sac59, Sch67, Sch85, Sch88, Set58, Stu88, Stu05, Tal68, Vos93, Wag81, Wei69a, Wei69b, Wig86a, Wig61e, Wig67q, Wig67t, Wig67-28, Wig70k, Wig72f, Wig78l, Wig83j, WS92, Wig96j, WS03, Wol67]. P/b [Sch85]. pace [Wig80b]. pages [Bec58, Edw60, Wig65c]. Pál [Wag98]. Pamphlet [Wig73n]. Panel [SBR^{+}86]. Panorama [Tru51]. paper [Wig56e]. Papers [Rom00, Sch58, Sch03, vN96, BV65, Eck73, Gre86, Gre69, Her79, Kol73, Mei64, NKW88, Por65, Rom91, Veb60, Wig92c, Wig93, Wig95b, Wig96m, Wig97a, Wig97b, Wig98, Wig01b, von96]. para [Wig33c, Wig33c]. para-ortho [Wig33c]. Para-Orthowasserstoff [Wig33c]. parabolic [Ker85]. paradox [Vil86]. Parallel [AC09]. paramagnetic [Wig33c]. paramagnetische [Wig33c]. parameters [Loc76, ÖAÖ25]. parametrization [HKS86]. Párbeszéd [Zsa76]. Parity [WWW52, WWW64]. Park [Zac84]. Part [Rom00, Bai93, Wig26a, Wig26b, Wig92c, Wig93, Wig95b, Wig96m, Wig97a, Wig97b, Wig98, Wig01b]. particelle [Maj32]. Participants [MWW66]. Participation [Wig72o]. Particle [AN92, KNS04, LB96, BT73, GW63, Zic73]. Particles [EW41, HW35, KW89c, KS07, Rom00, VW65, VW66, WWW52, Wig52c, WNY^{+}84, F^{+}76, KPT86, KW87b, KW90d, Lan87, LR83, Maj32, Mar06a, OW84, Pal82, WWW64, Wig66i, Wig69c, Wig69b]. Particular [IW54]. partitioning [GP86]. partly [MGM62]. partly-baked [MGM62]. path [Ber06]. Paths [Wig61a]. pattern [Gou86a, Gou86b]. Paul [Ano62, BvdM83, Ble86, Hen89, Hor97, L.88, Muk95, Ram89, SW89, Sch85, Sch88, SVW98, SVW00, Stu88, Tel95, Tor87, Vog95, Wag98, Wei97, Wig95a, Wig71d, Wig74c, Wig88b, Wig96a, BVV84, Goe01, KW87d, Tay87, Wig87c, Wig96k, dBDW^{+}84]. Pauli [JW28, Wig73p]. Paulische [JW28]. Peace [Ano78a, BAB^{+}82, Wig67s, Wig80b, Wig67o, Wig67y]. Peaks [TC10]. Peierls [Wig81a]. Pennsylvania [Ano83a]. People [Wig88c, Wig95d]. perfect [Wig79w]. perfette [Wig79w]. performed [Ano19]. Periodic [Tau61b, vNW40, MMP94, Wig72d, LSV08]. Periphery [Cha01]. permutation [LLBC85]. Personal [Ano61, Com56]. Perspective [Nic12, Seg85, Wig81a, Wig73c]. perspectives [KN19]. perturbáció [Wig35c]. perturbáció-elmélet [Wig35c]. Perturbation [BM64, CIMS08, FP03, HW01, Löw64, Wen98, Wig54b, Ahl70, LWA74, PMHW07, Wig35b, Wig35c]. perturbations [Gra74]. perturbative [HMW02]. perturbed [BT73]. Peter [L.88]. Petition [Dan95]. Phase [Ber77, BC62, KW87c, KW88c, KW88b, Wig55c, Ara95, KW87a, KW88a, KW89b, Lag84, RI05, SM05]. phase-space [KW87a, KW88a, KW89b, Lag84]. phases [RI03]. Phenomena [Wig57f, BH71, SW23, Vig58, Wig67h]. Phenomenological [Wig60h]. Philadelphia [Ano83a, Wig02a, Wig02a]. philosopher [Wig72m]. Philosophical [Rom00, WMW95, Wig97c, Wig71h, Wig95b, Wig98, Wig01b]. philosophy [Hoo73, Wie53, Wig64l, Wig64d, Wig82h]. Photo [Ano83b]. photon [DK75, DK78, Stu03, Wig56e]. Phys [BB93, Maj05a, Pos01, Szm99]. physic [Wig73h, Wig74e, Wig76k]. Physical [Coh59, Dar47, Gla59, GL23, Kou59, Rom00, Wig49d, Wig62j, Jah81, Kol60, LS76, MR76, Mic85, WW58, Wig50e, Wig51c, Wig71h, Wig74h, Wig77f, Wig77g, Wig82h, Ano48]. Physicist [Hen89, Ram89, Sch88, Stu88, ABES12, Her76, KW87d, Meh73, Wig68b, Wig02a, L.88]. Physicists [Ano19, HH04a, Har06b, Jon10, Wig82g]. Physics [Ano54b, Anoxxa, Anoxxb, EW61, Meh73, Mla98, Sch85, Stu18, Vog95, Wig46b, Wig65t, Wig66a, Wig67w, Wig68e, Wig70h, Wig74f, Wig77h, WNY^{+}84, Zac84, Zic78, Zic83a, Zic83b, Zic85, Zic88, d’E71, Ano54a, Ano68, BDH^{+}89, DGW84, FBR^{+}41, FF91, FR66, GHM^{+}87, GW63, HOSW84, Kib76, KR78, KHFA67, MR76, Mei64, Pal88, Ros82, Sal69, SW59, Str11, Stu79, Vig66, Wig54c, Wig66g, Wig68f, Wig69r, Wig71g, Wig72j, Wig74d, Wig79p, Wig82i, Wig83c, Wig83d, Wig83f, Wig84a, Wig89b, Wig92d, WM95, Wig99, Zic73, Ano68, L.88, LW34, Rom00, Vig68, Wig51a, Wig51e, Wig76l, Wig77c, Wig81a, BH70, Wig96m, Fla68, Pai67b, Wig50b, Wig67q, Wig53a, Rom91, Wig56d, Wig81a]. physics.‘ [Wig92f]. Physik [Wig56c, Wig76l]. physikalischen [Wig51c]. Physique [Wig61d]. Picture [MDH94, CLM76]. Pierre [Ble86, Tor87]. Pile [FWN^{+}45, NWEC61]. Piles [LK04]. piu [Wig83l]. place [Wig72j]. plan [Wig72l]. plans [WG74c]. Plant [ACF^{+}42, LCM^{+}42, Wig69y, Wig79r]. plasma [CMS77]. Pleasantly [Har99]. Pluhar [Wig65c]. Plutonium [WW61b]. Poi [Wig80b]. Poincaré [MR73, Bec72, MR72, NKW88]. Point [VW65, VW66, GRT91]. points [CV80, Kla91]. Poisson [AF95, Arn94, AR96, CHH05, DK04, LZ94, RNM04, Rin92, Ste91, TP01, Zha03]. Polányi [WJH76, WH77a, Wig02b, Jha11, Ano61, Wig02b]. Polanyian [Jha11]. Polarization [BH71]. Poles [Wig52b]. Policing [Fin60]. Policy [Wig62c]. Polish [MR76]. Political [Rom00, Wig95b, Wig98, Wig01b]. polyhedra [SAZ90]. Polymath [Hen89]. polynomials [LB73]. Polytechnic [Bug77]. Pondering [Wig61h, Tis03, Wig67g]. Population [CW74]. pörg [NW70]. Portrait [Pai00]. Position [OW78b, OW78a, WK72, Jor78, Jor80]. Positive [PW67, SIM93, Bo09, BW95, PI02, PW68, SCS99, Wig63d, WY64, WÁ68, Wig73k, Wig74g]. Positivity [Jan81, MFAB85]. Positron [WCT39, Maj37]. positrone [Maj37]. Possibilities [WNY^{+}84]. Possibility [WH35]. possible [BH13, Wig66e]. posteriori [HMW02]. potential [BT73, KPT86, Ker05, MMP94, Pal82, Wig32b]. Potentials [Beh91, Beh94, Kla91]. Potentialschwellen [Wig32b]. Power [BAA^{+}75, Wig61g, LCM^{+}42]. Pp [Hen89, Wig79j, Duf46, Kol60, Pai67a, Wei69a, Wig64f]. Precalculated [RY03]. predict [Wig83l]. Preface [Wig71i, Wig69y]. prefer [Wig83b]. Preis [Ble86]. Premature [Wig72h]. premier [Wig62d]. Prepared [Wig82b]. preparedness [Wig76d]. presence [DK75, DK78]. Present [Wig86b, TW79b, Wig82i, Wig86f]. present-day [Wig82i]. presentation [Nob72]. presented [Ano61, Gre86, Gre69, Wig67t]. president [Bug77, Dan95, Ein39]. Press [Bec58, Ble86, Edw60, Hen89, Kir73, Kol60, Pai67a, Sch85, Wei69a, Wig65c, Wig69s]. prevedere [Wig83l]. Price [Bec58, Edw60, Wig65c]. Prima [Wig82j]. Prime [JF16]. primitive [Wig74h]. Princeton [Bec58, Edw60, Vig86, Wig79j, Ano30, Ano58]. Principle [Wig79c, SCS99, Wig29a, Wig79n]. Principles [HVW65, Wig64o, Wig64p, Wig72b, Wig85a, EGW67, HVW82, Vig65, Wig64n, Wig64l, Wig64d, Wig67a, Wig68f, Wig76o, Wig80d, Wig82h, Wig84b, Wig84f]. principy [Vig65]. prirody [Vig65]. Prize [Ano79a, Wol67]. Prize-winner [Wol67]. Prizes [LW34]. Pro [WU75]. Probabilities [Wig70g, Wig76i, Wig76j]. Probability [Wig61b]. Problem [AR96, Cox61, HW01, Jha11, Rin92, Ste91, Wig54a, Wig63e, Fro79, Nou99, PD98, PMHW07, Wig48g, Wig67f, Wig67i, Wig71h, Wig78g, Wig83i, Wig83k, Wig84e]. Probleme [Wig58i]. problems [Rub82, TW79b, Wig58i, Wig61f, Wig69r, Wig71g, Wig73h, Wig74e, Wig76k, Wig78h, Wig86d]. procedures [PFG06]. Proceedings [SM07, Wig64f, Wig64g, Zic83a, Zic85, Zic88, d’E71, Ano75, Ano77b, Ano78b, Ano79b, Ano83a, BH71, BW61, DGW84, FR66, GHM^{+}87, Hoo73, KPS78, MR76, Rub82, Sal69, Stu79, Vil85, Zic78, Zic83b, ZHH60, Ano54b, Ano78a, Mac59, WNY^{+}84, Cut63, How62, Sec62, Wig58a]. Process [BY05, WY73]. Processes [vNW29a]. Producing [Wig61g, WW61b]. Product [DF77, CIMS08, Ear81, Fro79]. Production [Wig76d]. Products [WW48, Wig73j, CS51, Lit72, WW46a, WW46b, WW51, Wig65i, Wig73i]. Professional [WFS^{+}69]. Professor [MW62, MW64, Zsa76]. Professors [LW34]. professzorral [Zsa76]. Program [TW79a, GW71, TN84, WT79]. programs [AD73, AD84, CM66]. Progress [Wig57h]. prohibition [JW28]. Project [Wig92a, Lan93b, Lan99, Lan11, PW64a, PW64b, Wig65k, Wig66c]. Projection [Sha73, Joy01]. Proliferation [FR13c]. promise [Rub82, Wig67t, Wig77j]. Proof [HKW68, Jor78, Ras89, Ros99]. propagation [AN09]. Properties [BSW36, KS82, MF69, Wig52d, BB92, BB93, BSW64, NW70, OW81b, Stu03, WF41, Wig59f, Wig87a]. property [CMS77]. proportion [Woo80, Vig86]. Proposal [SM05]. Protection [Wei69b]. Protonen [Wig33b]. Protons [FR13a, Wig65b, Wig33b]. proves [Ano19]. provided [Wig71n]. Prozessen [vNW29a]. Prüfung [DGW34]. Public [Duf46, MW46, MW07]. Publisher [Maj05a]. Publishers [Wig71k]. Publishing [Wig65c]. Pure [SW59, Wig85d]. purpose [Wig84d, Wig88a, Wig92d]. Putnam [MW62, MW64]. Puzzle [Ban09, McD17]. Puzzling [Gel14]. Pythagorean [McD17, Scu07a]. Pythagoreans [Ne’04].

QBism [DFS20]. quadratically [ISRK12]. Qualitative [Wig54i, WS55]. Quando [Wig79w]. quanta [Jor78]. quantal [Mar03]. Quantenmechanik [Wig31b, Wig44, vNW28a, vNW28b, vNW28c, Wig27c, WW28, Wig32c]. quantenmechanischen [Wig62h]. quantenmechanischer [Wig52a]. Quantentheorie [Wig26a, Wig26b]. Quantization [BC62, HL88, CCT08, DV97, Fro79, Pal82, Poi99, RV10]. quantized [BT73]. Quantum [Ano78c, BVV84, Bel73, BV65, BGA07, Cob72, CHS97, Dir27, Fan60, Fer32, F^{+}76, For97, Fre15, Gar88, GL23, Gut60, Hal01, Hei74, HW39, Hug89, IÖRZ67, JvNW34, KKF89, KS07, LM89, LPN11, Luo04, MH94, Mar06a, MS83, OW81a, PE77, Rom00, Sac59, SW58, SV86, Sch85, Sch58, Sch03, Scu07b, SM05, Tau61a, Wes21, WZ83, Wig31b, Wig32a, Wig50c, Wig51f, WvN54, Wig57f, Wig64f, Wig64g, WL65, Wig66a, Wig70g, Wig71a, Wig73j, Wig76i, Wig83a, Wig86b, dBDW^{+}84, vNW28a, vNW28b, vNW28c, Ano19, BBDP08, Ber06, Cha01, CPCF99, DK70a, DK70b, DK71, DK73a, DK73b, DK75, DK78, DP01, FB88, Fon94, GI03, Gre86, Hoo73, JWY67, JC10, JLC14, KPT86, KW89a, KW90a, KLW09, KHFA67, LSV06, LSV08, MR98, MACS04, MMZ^{+}05, NW70, NKW88]. quantum [OW81b, OW84, PS97, PPG^{+}19, SW72, Sch96, SG75, Scu07a, Str11, TP01, Vig58, Whi96, Wig26a, Wig26b, Wig27c, WW28, Wig29c, Wig32c, Wig44, Wig52a, Wig54c, Wig56a, Wig56f, Wig56b, Wig58d, Wig59g, Wig62h, Wig63b, Wig64z, Wig64q, Wig67h, Wig68a, Wig69m, Wig69n, WY73, Wig73c, Wig73i, Wig73l, Wig76j, Wig79e, Wig79x, Wig83g, Wig83k, Wig84e, Wig86f, Wigxx, wYDK74, ZWE12, d’E71, dlPC79, Ble86, Tor87]. Quantum-Mechanical [HW39, LM89, OW81a, DK71, OW81b, SG75, Wig52a, Wig56f, Wig62h, Wig84e]. quantum-mechanika [Wig29c]. quantum-statistical [DK70a, DK71, DK73a, DK73b, DK75, wYDK74]. quasi [Ano08]. quasi-three-dimensional [Ano08]. quasispin [Hon76]. query [Wig76f]. Quest [Ble86, Sch85, Tor87, BVV84, Com56, dBDW^{+}84]. Question [Wig75b, Wig62f]. Questions [Wig83j, And73, Wig71h]. Qui [Dem15].

R [L.88, Wig67q, Wig78j]. R. [Wig65i]. Racah [AD73, AD84, Bru85, Bru87, DMMT93, DA73, DLD^{+}21, Joy01, JBR02, Kot05, Kot07, LB71, LLBC85, Mae91, MM92, RR96, Tem07, Wig54f, Won79, ZY88]. Race [Duf46]. radar [Lie90]. radial [CLM76]. Radiation [Dir27, Fer32, SW56b, WW46a, WW46b, WY58, DK70a, DK70b, DK71, DK73a, DK73b, DK75, DK78, WW30a, wYDK74]. radiation-field [DK73b]. radiations [Wig66i]. Radioactivities [Wig39a]. Radioactivity [Wig43]. Radiochemical [CS51]. Ralph [Wig50b]. Random [Bai93, FKS97, Web73, Wig67w, DK04, JD90, Kot05, Kot07, Wig65f, Wig69s]. random-sign [JD90]. randomly [BT73]. Range [WE47a, WW60, Wig67d, Wig70i]. Rate [Bai93, WW48, WW51, Wig37a, WE37, FW36, PW32, Wig25]. Rates [Wig39c]. Ray [WB36]. Rayleigh [Wig35b, Wig35c, Wig54b]. Rayleigh-Schrödinger-féle [Wig35c]. Re [MH94, Ano79b]. re-evaluation [Ano79b]. Re-formulation [MH94]. Reaction [BHT86, LPN11, Wig39c, Wig46a, WOYW62, Wig63c, And73, Her76, Wig25, Wig62d, Wig67m, Wig82l, Wig62d]. Reactions [HW39, NW61, TW52, Wig37a, WE37, Wig46c, Wig46e, Wig46d, WE47a, Wig49c, WW61b, ZHH60, BH71, FW36, Hod76, PW32, Wig32b, Wig54e, Wig67l, ZHH60]. Reactor [COW^{+}59, Cut63, How62, Mei61, MW79, OWYW56, OWWY58, OWWY59, OYWW61, SW61, WW61a, WOYW56a, WOYW56b, WOYW57, Wig57e, Wig58c, WO58, WWY58a, Wig58f, Wig59e, WC60, WY60, Wig60e, Wig61e, Wig61c, Wig62e, Wig79b, Wig01a, Ban04, BW61, Wei04, Wig61f, Wig66f]. Reactors [COW^{+}61, SBR^{+}86, SWC59, Wig57h, Wig58g, Wig59a, Wig60a, WW58, Wig58i, Wig59d, Wig60f, WU75, Wig85b, Coh59, Gla59, Kou59, Kol60]. Reading [Vig86]. Readings [Wie53, Rub82]. Ready [Fra05]. Reaktionen [PW32, Wig32b]. Reaktionsgeschwindigkeit [Wig25]. Reaktoren [Wig58i]. Real [Con72, Hud78, Wig59f]. realities [Ano19]. Reality [Wig64u, Wig64v, Ano19, Hor97, LS76, Wig67j]. Realization [Tes86]. really [GW71]. Reasonable [Tis98]. Reasoning [Ste25]. Rebuilding [Fre15]. Rebuttal [Wig67z]. Recall [Wig61h, Wig67g]. receipt [Wig66b]. Reciprocal [Wig54d]. recognition [Wig80d]. Recollections [Bon92, Vos93, Wig75c, WS92, WS03, Wig73s]. Reconstruction [Mac04]. Record [WT79]. recordings [Wig88d]. Recoupling [PD98, Joy01]. Recursion [BMR65]. Recursive [Wol71, SG75]. reduce [Wig65i]. Reduced [Gou92, GP86, Pal88, PD98]. reducibilis [Wig71n]. reducible [Wig71n]. Reduction [MR72, MR73, NAKS04, Wig73j, Wig73i]. Reeh [Fle00, Hal01]. reference [PČH02]. Reflected [Cox61]. Reflection [Wig84d, Wig88a, VC01]. Reflections [Gut68, Mar67, Pai67a, Rom00, Sch67, Wig62d, Wig63f, Wig72n, WMW95, Kir73, Wig64e, Wig67-28, Wig70k, Wig78l, Wig97c, Wig01b, Fla68, Cho68, Wig69t]. reflexiók [Wig72n]. Réflexions [Wig62d]. Refugee [Seg85, Jon10]. Regge [Hol73a]. Regime [AF95, MP01]. Regional [Zic78]. Regular [Gra74, MACS04]. Rehovoth [Wig58a]. rejection [PPG^{+}19]. Rejoinder [Wig71j]. related [Rno74]. Relation [OW77, OW78b, Ste91, LLW^{+}09, OW78a, Wig72c, Wig77h, Wig77a, Yan10]. Relational [Wes21]. Relations [BMR65, Par65, Wig50c, Wig54d, Wig69w, Wig82b, OW92, Wig64x]. relationship [Maj05a, Maj05b, Wig72f, Wig75a]. Relationships [Jan68]. Relativistic [BW48, Hal01, KW89c, ML91, OW78b, VW65, VW66, Wig47b, Wig56a, Wig56f, Wig56b, Wig57f, Wig67h, Wig69c, Wig69b, Wig73k, Wig73l, Wig74g, KW90d, OW78a, Wig86f, Maj32, Vig58, Wig47c, Wig48b]. relativistica [Maj32]. Relativistische [Wig47c, Wig48b]. Relativitätsprinzips [Wig29a]. relativity [Wig49b, Wig49e, Wig79e, Wigxx, NKW88, Wig29a]. Relaxation [DK70b, DK71, DK70a, DK73a, DK73b, DK75, DK78, wYDK74, Wig69d]. Relevance [Fla68]. Reljativistskaja [Vig58]. Remarks [CHH05, Wig39c, Wig50a, Wig62f, Wig66h, Wig66g, Wig66f, Wig67u, Wig67i, Wig82k, Wig96d, JD90, Wig74b, Wig78b, Wig78c, Wig78i, Wig87a]. Remembered [Woo95]. Remembering [Tis03, Wig87c, Wig96k, Wig88b, Wig71k]. Remembers [Har99, Wig79m]. Remind [Duf46]. Reminiscences [Hen89, KW87d, L.88, Sch88, Stu88, Ram89]. rendszerek [Wig29c]. replies [Wig74i]. Reply [Cla71a, MW64, Put64, Wes21, Wig65k, Wig78i, Wig72e]. Report [ACF^{+}42, Duf46, PW64a, PW64b, KHFA67, MW46, MW07, Wig69h, Wig69i, Lap60]. Reporter [Wig92f]. Reports [Wig68g]. Representation [KW87c, KW88b, TKC^{+}08, Wol71, Wol72, GW72b, Hol73c, KW87a, KW88a, KW89b, MACS04, OK78, Seg03, Wig71n, Wig73i]. Representations [BB65, IW52, IW53, RSdG99, Wig39b, Wig41b, Wig71c, Wig73j, Wig79a, Dir45, GW72b, Lit72, LLBC85, Wig64e, Wig65i, Wig71m, Wig89a]. reprints [BV65, Por65, Ros82]. Reproducing [Wig61b]. Reprresentation [KW88c]. Republican [Wig72l]. repulsive [CHH05]. Requirement [Wig82d]. Requirements [BW38, WI98]. Research [MR76, NWA72, Ano54a, HKT77, Hoo73, WA72]. Residues [Wig52b]. Resolution [RB92, PD98]. RESON [TN84]. Resonance [CJSW55a, CJW55, CJSW55b, LTW55, Wig46c, Wig46e, Wig46d, WE47a, WCJS55, WCJS56, Kla91, WE47b, Wig51g, Wig57c, Wig57i, Wig57g, Wig65g, Wig65l, Wig65j, Wig66h, Wig67l]. resonances [TN84]. resonant [JCT11]. resources [DHW49]. respect [MR72, MR73]. Respecting [DFS20]. Response [Kue61, Wig78j, Wig86c, CvV82, Rub82, Wig66b]. responsibilities [WW62a, WW62b]. Restraint [Smi75]. restricted [GW72b]. Restriction [Wig71c, Wig73j, Wig73i]. result [Ras89]. Results [Wig57g, Wig65l, GS04a]. részek [WWW64]. retaliation [Wig69j, Wig69k]. retentissement [Wig73f]. retrospect [Stu79]. reversal [HKW68, Wig32c, Wig64q]. Review [Ada70, Bec58, Bel73, Bie69, Ble86, Bre72, Cho68, Coh59, CH68, CJSW55b, Cut63, De 58, Duf46, Edw60, Fan60, Fes60, Gla59, Gut60, Gut68, Hei74, Hen89, How62, Kin58, Kir73, Kol60, Kou59, L.88, Mar67, Mar69, Pai67a, Pai67b, Par65, PE77, Ram89, Rom00, Ros70, Sac59, Sch67, Sch85, Sch88, Sec62, SBR^{+}86, Set58, Stu88, Tor87, Van94, Wei69a, Wei69b, Wig49b, Wig49e, Wig50b, Wig53a, Wig56c, Wig57h, Wig57i, Wig58a, Wig60i, Wig61d, Wig61e, Wig62c, Wig62g, Wig63a, Wig64f, Wig64g, Wig65c, Wig65d, Wig65m, Wig66a, Wig67o, Wig67n, Wig67p, Wig67q, Wig67x, Wig67y, Wig69d, Wig69t, Wig69s, Wig70b, Wig71k, Wig76c, Wig76b, Wig79j, Wig81a, Wig83k, Wig84e, Por65, Wig67l, Wig69e, Wig74a]. Reviews [WBBE30, Wig56d]. Revisited [Omn11, Wig71a, CvV82, DMMT93, dlPC79]. revolution [Vil85]. rhombic [MW24]. rhombischen [MW24]. Richard [Wig72e]. Richardson [Wig36a]. Ridger [Wig59c]. Rings [Tau63a, PS02]. Riporter [Wig92f]. Rise [Wig67p]. risks [Wig79v]. road [Wig67z]. Roadblocks [Wig68d, Wig73m]. Robert [L.88, Wig62c]. robust [PMHW07]. Role [KKF89, Wig67p, Jah81, Wig50e, Wig64l, Wig64d, Wig80d, Wig82h, Wig84b, Wig84d, Wig88a, Wig92d, ZWE12, Mar03]. roles [Wig66j, Wig74h]. Rolle [Wig64l]. Ronald [Wig71k]. Roosevelt [Dem15, Ein39]. Root [MH98]. Roots [Wig46f, Wig58h, Wig07, Wig65f, Wig65h]. Rosen [Wig76b]. rotating [SK05, Sta10]. Rotation [AC09, LY09, DMMT93, MM92, PFG06]. Rotational [HW35]. rotations [HKS86]. rotten [Wig67-27]. Round [MFG^{+}83]. Roundoff [GRT91]. Roundtable [W^{+}75]. Royal [Wig02b]. rpm [Wig92e, Wig92f]. Rudolf [Wig72o]. Rudolfhoz [Wig72o]. Rudolp [Wig81a]. Rule [WWW70, HKW68, Seg03]. Rules [TW52, CM68, Joy01]. Russian [Vig58, Vig64, Vig65, Vig66, Vig68, WG74c].

S [Rom00, Sch85, Wig56c, Wig63a, Wig69e, Wig69d, Wig82b, MWW66, Tem07, Wig65k]. S. [Wig61e, Wig65i]. S/C [MWW66, Wig65k]. sa [DMS84]. Safe [Gul75]. safest [WU75]. Safety [Wig01a, Wig79r]. sajátparitása [WWW64]. Salam [Bel73, Hei74]. Salomon [Wig67p]. SALT [Wig70e, WG72]. Sampling [Mac04, KLW09]. San [Ano78b]. Sandia [Ano15]. Sant [GHM^{+}87]. Sarebbe [Wig83l]. satisfaction [Wig77i]. satisfied [LB71]. satisfy [Wig76a]. Saturation [BW38, Wig36b, WE39]. Saturn [Ano15]. says [Wig75d]. scalars [Que76]. scale [JC10]. scaling [WI98]. Scattering [Wig46e, Wig46a, Wig54a, Wig55c, Wig65b, CPCF99, MM73, Wig33b, Wig52e, Wig52f]. scenarios [Pla24, Pla25]. Scheme [HP69, MPPS02, MPPS03, RY03]. Schemes [Hal01]. Schlieder [Fle00, Hal01]. School [Zic78, Zic83a, Zic83b, Zic85, Zic88, d’E71, Wig73r, Wig73s, Wig88c]. Schrödinger [Hen89, LW34, Szm99, Wig64-28, CLM76, DF77, HW35, MPPS02, MPPS03, Ste91, Szm98, Tes86, Vil86, Wig27a, Wig27b, Wig35b, Wig35c, Wig54b, Wig56e, Wig64-28, Wig65o, Zha03]. Schrödinger-féle [Wig64-28]. Schrödingerschen [Wig27a, Wig27b]. Schwefels [MW24]. Schweinler [SCS99]. Science [Ano48, Gib19, K^{+}84, Pai00, SM07, Web73, Wig50d, Wig62b, Wig64r, Wig65n, Wig66i, Wig73n, Wig75d, Wig77i, Wig82a, Ano65, Ano75, Dar47, GM80, HH04a, Hed93, Mar96, Tru51, Wie53, Wig51d, Wig53b, Wig64-29, Wig64-30, Wig64-31, Wig67t, Wig67c, WÁ68, Wig72f, Wig73d, Wig73f, Wig74a, Wig75a, Wig75e, Wig75f, Wig76m, WPG77, Wig77j, Wig78e, Wig78h, Wig79l, Wig79t, Wig81b, Wig82g, Wig84d, WW84, Wig88a, Wig09, Bug77, Har06b, Zic78, Wig62a, Wig67p, Wig76b, Wig69t]. Sciences [Ano75, Ano78b, Ano79b, Ano83a, MR76, Tis98, Vil85, Wig51b, Wig60j, Wig60k, Wig60l, Wig65c, Ano77b, Vig68, Wie53, Wig64-27, Wig67k, Wig74h, Wig77f, Wig77g, Wig80c, Wig85d, Wig86d, Wig90b, Wig91, Wig92b, Gre86, Rub82, Ano79b]. Scientific [Cob72, GHM^{+}87, Gut68, KPS78, Kir73, Mar67, MWW66, NAH^{+}89, Pai67a, Rom00, Sch67, WG96, CGPT21, DMS84, GW57, HKT77, Wig67-28, Wig70k, Wig72n, Wig78l, Wig92c, Wig93, Wig96m, Wig97a, Wig97b, GW03]. scientifique [DMS84]. Scientist [Mar95, Stu05, MGM62, WW62a, WW62b, Wig68e, Wig89c]. Scientists [BAA^{+}75, Duf46, Fin60, GR63, MP01, Seg85, SBR^{+}86, Wig62c, Str11, GR63]. scienza [Wig79t]. scope [Wig77j]. Scriven [Sch67]. SE [SBR^{+}86]. Seal [Wig61c]. search [Ano77b, Ano78b, Ano79b]. Second [BC62, Wig48g, Wig76d, ZHH60, CCT08, Kot05, PMHW07, Stu18]. second-order [PMHW07]. section [MM73]. Sections [Wig46a, Wig48f, Wig61e]. Secure [SBR^{+}86]. Seeding [Wig72a, WL72]. segregated [Fre93]. seismic [VC01]. Seitz [CV80, Fre93, KN94, SAZ90]. Select [SBR^{+}86]. Selected [Sch58, Sch03, Wig79j, Ros82, Rub82, Vil85, WN90]. selection [Mei64]. Selections [Wig81a]. Self [OG07, OW92, Wig61b]. Self-adjointness [OW92]. Self-Consistent [OG07]. Self-Reproducing [Wig61b]. Semi [Ber77, MMP94]. Semi-Classical [Ber77]. Semiclassical [FM03, LY09, Pul06, Rob93, Zha03, Ara95, BZ74, CPCF99, Poi99, GS04b]. semiconductor [PS02]. Semiconductors [KKF89]. semidefinite [WY64]. Semidiscrete [Gou02, Gou03]. Semisimple [BB65, Gou81, Gou86a, Gou86b]. Senate [WT79, Wig69w, Wig82b]. Sense [For97]. Separating [CW58, WM58]. Separation [CW61, HW35]. September [Ano54b, DGW84, GHM^{+}87, Meh73, SM07, Vil85, Wig73q, Wig66f]. sequel [Wig49b, Wig49e]. Sequential [But88]. Serbian [Wig65s]. Series [But88]. Session [ZDF^{+}84, Wig76d]. set [Wig31a]. Sets [Tau61a, Tem07]. Setting [Mac04]. seventh [Ano79b]. seventieth [Ano61]. Seventy [Wig49a]. Shalit [Wig63a]. Shannon [Jan14]. Shape [JC90, SAZ90]. Shaped [Gib19, WWY58b]. shaping [Har06a]. Shell [Wig63a, Wig52g]. Shells [FR13a, Goe49]. shelter [Wig70a]. Shenkman [Wig67o]. Shield [WO58, Wig59e]. Shielding [WY58, WOYW59]. Shift [Wig55c]. Shimony [CHS97]. Shock [Wig69d]. shocking [Wig77d]. short [Wig70i]. short-range [Wig70i]. Shot [SSP^{+}91]. should [Wig75d]. si [Wig82j]. Sicilian [Zic78]. Sicily [Zic78, Zic83b, Zic85, Zic88, Zic83a]. side [Wig79j]. sign [JD90]. Signal [CM80, LB96]. Signals [Cox61]. Significance [WvN54, W^{+}75]. Similarity [KP90]. simmetrica [Maj37]. simmetrie [Wig79w, Wig79q]. simmetrii [Vig66]. Simmetrija [Vig64]. Simple [FB88, Gin63, Jor78, Jor80, Ras89]. Simplified [Wig52d]. simply [Wig71n]. simulating [JCT11, MR98]. Simulation [KKF89, RRC05]. Single [MH98, Ste25]. Single-Root [MH98]. singlet [PČH02]. Singular [IW54]. Sir [Wig62a, Wig81a]. Sitter [Cox61, HL90, PW67, PW68, Wig50a]. six [ÖAÖ25]. Sixth [Ano78b, KR78]. Sixtieth [Ano62]. Sixty [FR13e]. Skew [Luo04, Yan10]. Slow [BW36]. Smoothing [Col94]. Sobytija [Vig65]. Societies [Wig62j]. Society [BW61, Gib19, Wig82l, Wig02b, Bug77, Wig68e, Wig72f, Wig73n, Wig75a, Wig89c, Sec62]. Socio [Rom00, Wig95b, Wig98, Wig01b]. Socio-Political [Rom00, Wig95b, Wig98, Wig01b]. Sodium [WS33, WS34]. software [Zie88]. sohranenija [Vig64]. Solid [Rom00, Wig36c, Hal81, Hal83, Mei64]. Solids [CW58, CW61, SW56b, Mar03, SM05]. Solution [KNS04, Rin92, DK71, DK73a, DK73b]. Solutions [NAKS04, Bru85, Bru87, LSV06, LSV08]. Solvay [WNY^{+}84, Wig64g]. solve [Ano15]. Some [Ara95, Bie61, HS83, HW39, JWY67, JD90, KS82, L.88, NL69, OW81b, Wig39c, Wig50a, Wig65o, Wig70i, Wig86d, Woo80, vNW28a, vNW28b, vNW28c, Mos62, NW70, RV10, Wig64-28, Wig69r, Wig71g, Wig73h, Wig74e, Wig76k, BB65, Bru87, Wig27a, Wig27b, Vig86]. Soni [Wig67q]. sono [Wig79w]. soul [Wig68b]. Sources [KHFA67]. Soviet [Wig71i, Ano65, GW71, KW77, SBR^{+}86, WG72, WG74a, WG74b, Wig76g]. Space [Ber77, Ble86, BC62, Cox61, KW87c, KW88c, KW88b, KW89c, KW90d, KNS04, PW67, SW58, Sch85, Tor87, Wig50a, Ara95, Bac89, BZ74, BVV84, BL72, Fon94, GI03, HL90, KPT86, KW87a, KW88a, KW89b, Lag84, MS06, PW68, Rno74, SV86, dBDW^{+}84]. Space-dependent [KNS04]. Space-time [KW89c, SV86]. spacelike [SW67]. Spacer [WO58]. spaces [APW02, APW03]. spacing [Wig57c, Wig65g]. spacings [Wig51g, Wig65j]. Spain [GHM^{+}87]. Spatiotemporal [JW87]. Spatiotemporal-Frequency [JW87]. Speak [BAA^{+}75]. Speaks [CH68, Mar69, Wig68i]. specchio [Wig79s]. Special [CV80, OW78b, Tal68, Wig67q, ZDF^{+}84, HS83, OW78a, Rno74, Bie69, NKW88]. specific [PČH02]. Spectra [Wig31b, vNW28a, vNW28b, vNW28c, NW70, Por65, Wig44, Wig59g, Fan60, Gut60, Sac59]. Spectral [AF95, Bai93, BY05, Rin92, Arn94, GS04b, GS04a]. spectral-collocation [Arn94]. Spectroscopy [Wig37b]. Spectrum [WB36, JD90]. speculates [MGM62]. Speech [Wig78a, Wig68j]. speeches [Nob72]. spektra [WW28]. Spektren [vNW28a, vNW28b, vNW28c]. Spheres [CJW55, WCJS55, WCJS56, APW03]. spherical [Dra86, MS06, Poi99]. Spin [Tem07, vNW28a, vNW28b, vNW28c, Bec72, DGW34, Dra86, FR66, Kot07, Lan87, LR83, OW84, Pos00, Pos01, RNM04, Wig58d, Wig66g, Wig79g]. spin- [OW84]. spin-weighted [Dra86]. Spinerhaltungssatzes [DGW34]. spinning [NW70]. spinors [HKS86]. Spirit [Gib19]. Split [NAKS04]. Splitting [AR96, Arn94]. sponsored [Zic78]. spontaneous [BH13, DK70a]. springs [LSV06]. St [Wig74c]. Stäbchensolen [SW23]. stability [CMS77, Fol78, PI02]. stable [Wig67s]. standardization [LLBC85]. star [GS04b, GS04a]. start [Ban04, Wei04]. start-up [Ban04, Wei04]. State [Rom00, SIM93, Wig67t, Ano65, GI03, HB90, Mei64, PČH02, SM06, Wig38b]. state-specific [PČH02]. statement [Wig67u, Wig69w, Wig79h, Wig79f, Wig82b]. statements [Wig77d]. States [NW49, Wig76d, BW95, CW71, LLW^{+}09, TKT^{+}05, Wig66h, Wig72d, Cob72, Dan95]. stati [Wig83l]. stationary [ALZ00]. Statistical [CW71, IÖRZ67, Por65, SP74, Wig59f, DK70a, DK70b, DK71, DK73a, DK73b, DK75, DK78, Wig25, Wig51g, Wig65j, Wig69s, wYDK74]. Statistics [AF95, AK00, FKS97, BLS03, DK04, GS04b, GS04a, Mar06a, OW84, RNM04, Wig29c]. Statistische [Wig25]. statisztikája [Wig29c]. Stephen [Wig73q, Wig95c]. sticks [SW23]. Still [Wig79c, Wig88e]. Stochastic [BT76, NAKS04, App87, dlPC79]. Stockholm [Wig78a, Wig68j]. stockholmi [Wig68j]. Stoops [Wig64f]. Storage [RY03]. Störungsrechnung [Ahl70]. Story [Bel73, MP01, TW79b, Wig82l]. Strahlung [WW30a]. strangeness [Vig86, Woo80]. strategic [CW74]. strategy [Wig76g]. Streuung [Wig33b]. strike [Wig70d]. strongly [Fre93]. Structural [PI02, Hal81, Hal83]. Structure [FW37, Hug89, SD96, Wig36c, Wig37c, Wu72, Wu73, BL72, EW58, EGW67, EGW69, LB73, MW24, WW28, Wig69y, Won76, De 58, Kin58, Set58, Bec58, Edw60, Wig49e, Wig49b, Wig58a]. Structures [Wig66a, Wig27a, Wig27b, Wig64-28]. Struktur [WW28]. Stuart [Wig67u]. student [Wig73r]. Studies [Ble86, CHS97, Löw64, Sch85, Tor87, BVV84, CS51, dBDW^{+}84]. Study [Ber77, NAH^{+}89, Zic78, DK70a, SNDS14, TKT^{+}05, Wig75d, Gre69, PČH02]. Stupochenko [Wig69d]. SU3lib [DLD^{+}21]. Subcommittee [Wig69w]. Subgroup [Wig71c, GW72b]. Subgroups [Wig73j, Wig73i]. subject [Wig71b]. Subnuclear [Zic78, Zic83a, Zic83b, Zic85, Zic88]. succès [Wig73f]. success [Wig73d, Wig73f, Wig82e]. successful [Wig70d]. such [Ano19]. Sufficient [BY88, MFAB85]. suggests [Ano19]. sulfur [MW24]. Sum [TW52, CM68, Ros98, Ros99, Wig31a]. Summary [PW64a, PW64b, Wig65k, Wig69u, Wig71l, Wig73o, Wig75f, Wig77g]. Summensatzes [Wig31a]. Sundberg [Wig96j]. Super [Abe92, RB92, DMMT93, MM92]. Super-Resolution [RB92]. super-rotation [DMMT93, MM92]. superalgebra [ZY88]. Superbomb [Wig76c]. Superconducting [Tes86]. superconductivity [SM05]. superconductor [SM06]. Supermultiplet [HP69]. Supermultiplets [FR13d]. superposition [Ell92]. Superselection [WWW70, HKW68]. Sure [BY88]. surest [WU75]. surface [KLW09]. surprising [Wig64-27]. surrender [Wig83b]. Survey [Ano65]. Survival [Wig69x, Ada70, Ros70, Wei69a, Wei69b, Wig70b]. survive [Wig86e]. Surviving [WH76]. suspected [Ano19]. sustained [And73]. Sustaining [COWY57]. Symbol [LY09, CD79, Hol73a]. Symbolic [Sha92]. Symbols [JF16, Lai94, Ear81, Jan68, Max10, Ros98, Ros99]. Symmetric [Wig58h, Bo09, HB90, JD90, LR78, Wig20, Wig59f, Wig64m, Wig65h]. Symmetrical [Maj37]. Symmetrie [Wig85c]. Symmetries [FR13d, GM80, RL81, WK72, Wig70k, Wig78l, ZDF^{+}84, Wig79w, Wig79q, Wig83f, GHM^{+}87, Gut68, Mar67, Wig67-28, Wig72n, MFG^{+}83, Cho68, Fla68, Kir73, Pai67a, Sch67]. symmetrischen [Wig20]. Symmetry [Bar64, BSW36, Ne’04, Ros82, Sch96, Stu03, Wig37b, WF41, Wig60h, Wig64s, Wig64t, Wig65p, Wig65q, Wig65t, Wig68f, Wig72k, Wig82h, BH13, BSW64, Hol73a, Wig54h, Wig76o, Wig78d, Wig80d, Wig84b, Wig84c, ZWE12, dDG83, Vig64, Vig66, Wig65r, Wig65s, Wig76b, Wig85c]. Symposia [Cut63, How62]. Symposium [BH71, BW61, Gau94, Meh73, Sal69, Sec62, Ano54a, Bug77, Dar47, LW78, MR76, Stu79, Woo80, Wig96d]. Syntheses [WMW95, Wig97c, Wig01b, Rom00]. System [NWEC61, WY60, Wig63c, Ano08, CHH05, CPCF99, DK75, DK78, GP86, LSV06]. Systeme [Wig20]. Systems [HMW01, NW49, OW77, OW78b, ACW98, DK70b, DK73a, DK73b, Duf71, FB88, KPT86, LZ94, LSV08, Mic85, OW78a, OW84, PS97, Ram00, SC83, Wig20, Wig29c, Wig64m, Wig77a, ZWE12, Wig72d]. szakvélemény [Frö93, Lad93]. Szanton [Bon92, WS92, Wig02c, WS03]. szeptemberében [Wig73q]. szerepérol [Wig92d]. szerint [Wig29c]. szerkezete [EGW69, Wig69y]. Szilard [Wig92e, Wig96i, Fra05, Hor97, Lan93b, Lan99, Wig69a, Wig92e]. szimmetria [Wig76o]. szimmetria-elvek [Wig76o]. Szimmetriák [Wig72n]. szimmetriákig [Wig95c]. színképek [NW70]. szorzatát [Wig71n].

T [HMW05, Wig66a, Wig77b]. table [MFG^{+}83]. Tacit [Jha11]. tagja [Wig02b]. Talk [GW72a, Wig69c, Wig71b, Wig72g]. talks [Wig70e]. Tallahassee [FR66, LW78, Wig87b]. Talman [Bie69]. Talmi [Wig63a]. tanárai [KW01]. tanuló [Wig73r]. Tasks [Wig73j, Wig73i]. Taylor [L.88, Stu88]. Teacher [Stu05]. Teacher/Eugene [Stu05]. Teachers [Stu05, KW01]. teaching [HKT77, Jan14]. Technique [Rin92]. Techniques [Ano54a, Gre86, Sha73]. technology [Bug77, Wig77i]. Teil [Wig26a, Wig26b]. teljes [Wag98]. Teller [Wig76c, BH89, Hor97, Lan93b, Lan99, MF69, Wig62b, Wig69f, Wig96b]. Temperature [CJW55]. Tennessee [ZHH60]. Tensor [MS00, BL72, Kib76, LB73, Mez77, Won71]. Tensorial [Tem07, Ell92]. tensors [HB90, RR96]. Teoria [Maj32, Maj37]. term [Wig27a, Wig27b]. Terme [Wig26a, Wig26b]. természettörvények [Wig65u]. természettörvényekrõl [Wig95c]. természettudományokban [Wig64-27]. Terms [LY09, HKS86, Wig26a, Wig26b]. Termstrukturen [Wig27a, Wig27b]. termszerkezetekre [Wig64-28]. Terrible [Duf46]. Test [Fin60]. testimony [Pla24, Pla25]. Testing [Ore58, Wig58e]. Texas [WNY^{+}84, Her79]. texture [JC90]. Their [IW53, Par65, WNY^{+}84, HS83, HB90, Hor97, Lan93a, LB73, Scu07b, Wig64x, Wig66i, Wig80d, Wig83f]. them [Wig76a]. Theodore [Hor97]. Theorem [Bar64, Gin63, HW57b, Agr80, BH13, BG03, Ell92, HW57a, Jan84, MS00, Mez77, Mol99, Pos00, Pos01, Sch96, BG02, LM63, WvN54]. Theorems [Wig67q, ZWE12]. Theoretic [Bie69, Tal68, Tem07, AW75]. Theoretical [Ano54b, Ano68, BW48, L.88, Meh73, Sal69, Wig46b, Wig68e, Zac84, DGW84, GW63, KR78, Wig79x, Wig53a]. Theorie [Wig27a, Wig27b, Wig51c, Wig62h]. Theories [Bri65b, FR13c, Eck33, Por65, Szm98, Szm99, Wig50e, Wig67l]. Theorists [Wig73j, Wig73i]. Theory [Ano78c, BSW36, BSW64, BM64, Coh59, CW41, Dir27, DW64, Fan60, Fer32, Gla59, Gut60, Hal01, HW39, HW01, Kou59, LB96, Löw64, MH98, OG07, PE77, RB92, Sac59, Stu04, Tau61a, Tau61b, Tau63b, TW52, Wen98, Wig31b, WB35, WCT39, Wig39c, Wig49d, Wig51f, Wig54b, WvN54, Wig64f, Wig64g, Wig86b, BT76, BV65, BW61, CvV82, Cha01, CW71, Fon94, Gre86, Hol73c, Hon76, Hoo73, HMW05, Kol60, LW78, LWA74, Loe75a, Loe71, Loe75b, Maj32, Maj37, MM73, Mei64, MS83, NW71, NKW88, PMHW07, Ram00, SW72, Sch96, WW58, WW30b, WZ83, Wig26a, Wig26b, Wig27a, Wig27b, Wig29b, Wig33d, Wig35b, Wig35c, WE39, Wig44, WE47b, Wig51c, Wig54e, Wig54i, WS55, Wig57g, Wig59g, Wig61f, Wig62g, Wig64-28, Wig64y, Wig65l, Wig65o, Wig69s, Wig71h, Wig73c]. theory [Wig82h, Wig86f, Wig88d, Wig88e, Wigxx, Zic84, Tau63c, Wig62h, Bel73, Cut63, Fes60, How62, Wig63a, Wig66a, Hei74, Wig66a]. there [Ano19, Wig64-30, Wig64-31, Wig88d]. thermal [Wig64-28]. Thermodynamic [Wig32a, Wig83a]. thesis [F^{+}76]. thing [Ano19]. things [Wig88c]. Third [Ano58, BH71, Zic88]. Thirteenth [Vil85]. Thirty [Vig86, Wig80e, Wig96l]. Thomas [AC09, Nau73]. Thompson [Wig62a]. those [CLM75]. thought [Wig72m]. Thoughts [Wig62i, Wig48g]. Threat [Duf46]. Three [And73, Ano08, BZ74, RV11, Rno74, Frö93, Lad93]. three-dimensional [RV11]. Thresholds [Wig48f, Wig32b]. tightening [Wig31a]. Tildon [Wig77b]. Time [Ble86, CM80, SW58, Sch85, SEC83, Tor87, VC01, Wig68g, BVV84, Fon94, HKW68, KW89c, KW90d, Rob93, SV86, Wig32c, Wig64z, Wig64q, Wig72c, dBDW^{+}84]. time-dependent [Rob93]. time-energy [Wig72c]. Time-Frequency [CM80, VC01]. time-reversal [HKW68]. times [Wig71k]. Titanic [Jon10]. Today [Vig68, WFS^{+}69]. Tokyo [Ano54b]. Told [Bon92, TW79b, WS92, Wig02c, WS03]. Tomographic [Ler90]. Tomography [BGA07]. Too [Ger75, Pal88]. Tool [CM80]. Topics [Tau62]. tops [FR03]. Total [CJW55]. Tóth [Wig92e, Wig92e]. Towering [Vog95]. trace [MM73]. trace-class [MM73]. Transcendentals [Wig52d]. Transcribed [Wig95c, Wig88d]. Transform [Col94, KS82, Zie88]. transformation [Ber06, KW89a, KW90a, Ras89, VGB87]. Transformations [QM71, HKS86, KP90, KW90b, KW90c, SW67, Vas89, Wig90a]. transition [CHDK08, Wig33c, Wig38b]. transitions [Hal81, Hal83, SM05]. Translation [Wig71i]. Transmission [Cob72]. Transport [AC09, KKF89, MH94, JC10, JLC14, RRC05]. Trapani [Zic78, Zic83a, Zic83b, Zic85]. treasury [FF91]. treatment [Dra86]. treatments [JC10]. Treist [Wig68e]. trends [LW78]. Tribute [Muk95, SW89, Wig78k]. Tributes [Stu88, Tay87, L.88]. Trieste [DGW84, Meh73, Sal69, Wig73q]. Triesztben [Wig73q]. triple [Ros98, Ros99]. True [MP01]. truncation [SAZ90]. Trustees [Ano30]. Tube [Wig61c]. Tübingen [KR78]. tud [Wig88c]. tudomány [Wig64-29, WÁ68]. tudománynak [Wig64-30, Wig64-31]. tudományos [GW03, Wig72n]. ‘Tudtam [Wig92f]. tulajdonságának [NW70]. túlságosan [Pal88]. Tunneling [But88, JCT11, NL69]. turbulent [TW79b].

TV [Wig73q, Wig73q]. TV-interjú [Wig73q]. Twentieth [Wig63f, Har06b]. Twenty [Wol67, Zic85, Zic88, Wig62d]. Twenty-first [Zic85]. Twenty-Five [Wol67]. twenty-third [Zic88]. twisted [CCT08]. Two [Fin60, OG07, Wig60d, Wig61a, Wig64u, Wig64v, Wig67j, Ano19, CV80, HKS86, KL03, Nar88, Pal82, PS02, RRC05, WWCS42, Wig64r, Wig65n, Wig66j, KPT86, Wig26b]. two-component [HKS86]. Two-Dimensional [Wig60d, CV80, RRC05, KPT86]. two-electron [PS02]. Two-Electron-Density [OG07]. Type [IW54, Kla91, PI02].

U [Wig56d, Wig82b, BGL67]. U.S [Wig69w]. U.S. [SBR^{+}86]. Überschreiten [Wig32b]. új [Wig29c]. újabb [Wig29b]. Ulehla [Wig65c]. ülésén [Wig88d]. ultimate [Hor97, Wig84f]. Umsetzungen [PW28]. Umwandlung [Wig33c]. Uncertainty [LLW^{+}09, Yan10, Wig72c]. Underdetermination [Gel14]. undermine [WG72]. Understanding [WNY^{+}84, BH13, Zic78]. unglaubliche [Hed93]. Unification [EW80]. unified [Dra86]. Uniform [LY09]. Unimodular [SW67]. Union [Wig62j, WG74a, WG74b]. uniqueness [OW81a]. Unit [KN94, Wig61b, Won71]. Unitarity [LM63]. Unitary [BMR65, Dir45, GP86, Kas21, Wig39b, Wig60h, Wig64e, Wig79a, BH13, BL72, Ell92, FB88, Kot05, Kot07, LB73, Mol99, Wig89a]. United [Dan95, Wig76d, Cob72]. Unity [Ano75, Ano77b, Ano78b, Ano79b, Ano83a, Gre69, Rub82, Vil85, Wig75e, Wig75f, Wig76m, Wig80c, Zic83b]. Universal [GS04b, GS04a, FS82]. universality [SM05]. Universe [Wig60d]. University [Bec58, Ble86, Edw60, Hen89, Hoo73, Kol60, MR76, Pai67a, Sch85, Wei69a, Wig64f, Wig64g, Wig79j, WNY^{+}84, Wig59b, Wig74h]. unknown [Wig35a, Wig65r, Wig65s, Wig69v]. ünnepi [Wig88d]. Unravelling [FR13e]. unreal [Wig79s]. Unreasonable [Gel14, HNP09, Isl22, Nic12, Omn11, Wig60j, Wig60k, Wig60l, Wig91, Jan14, Wig67k, Wig85d, Wig90b, Wig92b, Vig68]. Unusual [vNW29b, MF69]. upon [Wig66b]. Uranium [CJSW55a, CJW55, CJSW55b, CW58, CW61, WM58, Wig92a]. Ursache [PW28]. USA [FR66, WNY^{+}84, Zac84]. USCDC [Wig74c]. Use [HVW65, HVW82, LR78, Wig84f]. user [AD73, AD84]. uses [Wig79u]. Using [JW87, JF16, KNS04, GRT91, JC90, RRC05, VC01]. Utah [Wig74h]. ütközések [Wig64y]. Üzenet [Wig95d].

V [L.88, Wig53a, Wig61e, Wig69d, Vig66, Vig68, BB65, DK73b, Tau63b]. Valentine [Wig78k]. validity [Wig82i, Wig84f]. value [Wig80d]. valued [Bo09]. Values [WW84, Ano75, Ano77b, Ano78b, Ano79b, Ano83a, Dar47, Vil85, Ano48]. Van-e [Wig64-30, Wig64-31]. Vancouver [Gau94]. Varenna [d’E71]. várható [WÁ68]. Variable [Cla71b, Wig71j, SS05]. Variables [Wig70g, Wig76i, Wig76j]. Variance [NAKS04]. Variational [WN90]. városházán [Wig68j]. ve [Ano19]. Veblen [Veb60]. Vector [HB90].

Vectors [Wig55a, Wig57b]. véges [Wig71m]. Velocity [OW77, OW78b, AW75, ALZ00, OW78a, SP74, Wig77a]. Verhalten [vNW29a]. verification [DGW34]. Vernichtungschlag [Wig70d]. Verschärfung [Wig31a]. Version [Gou02, Gou03]. Versus [Duf46, WW95]. veszélyek [WÁ68]. Vey [DP01]. VI [Tau63c, Wig95b, wYDK74]. Via [Col94, RB92, KPT86, MR98, Pal82, Tem07]. vibration [PW28]. Video [Wig88d]. Videofelvételröl [Wig88d]. vie [DMS84]. view [WW60, Wig67d]. vigasztaló [Wig66k]. Vigier [Ble86, Sch85, Tor87]. Vignettes [Ano92, Sza92]. Vii [Sch85, DK75]. VIII [DK78, Wig98]. világformáló [Har06a]. Ville [VC01, Zie88]. Vincenzo [CGPT21]. vingt [Wig62d]. Violations [Wig65t]. Visionary [Mar95]. visszaemlékezés [Wig02a]. Visszaemlékezéseim [Wig73s]. Vistas [LS76]. Visualization [KKF89]. vita [Wig79p]. voction [WPG77]. Vol [Wig66a, Wig81a, Wig93, Wig95b, Wig96m, Wig97a, Wig97b, Wig98, Cut63, How62, Wig56c, Wig57h, Wig74a]. Volume [Rom00, FW52, Fre79, CHS97, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. vonatkozóan [Wig64-28]. Vortrag [Wig64n]. vote [Wig72l]. vs [Luo04, RI03, Wig69j, Wig69k]. vulnerability [CW74].

W [Wig63a]. Walk [Web73]. wall [LSV08]. Walsh [Wig92f]. Walter [CH68, Sch67]. War [Wig67o, BWD76b, Wig67y, Wig68h, Wig76d, WH76, Wig80b]. Warner [Ano52]. Wars [Stu18]. Warsaw [MR76]. Wary [Fin60]. Washington [Ano77b, Vil85, W^{+}75, Wig82e]. Water [Wig43, WOYW56a]. Watermarking [TC10]. Wave [BW48, BSW36, WBBE30, Wig47c, Wig48b, BT76, BSW64, OK78, Wig47b, Wig74g]. wave-kinetic [BT76]. wave-mechanical [OK78]. Waves [KW87c, KW87a, Wig69d]. Way [Duf46, WU75]. ways [Wig76a]. weakly [Wig63d]. weakness [Wig87a]. Weapons [Ore58, Wig58e, Wig62c]. Weighing [Wig76n]. Weighted [Jan81, Dra86]. Weinberg [Coh59, Gla59, Kol60, Kou59, Wig66d, Wig69t, Wig96j]. Weird [Str11]. Weisner [Wig64c]. Weisskopf [Wig53a, App87, DK70b, DK71, DK73a, DK73b, DK75, DK78, Dav74, wYDK74]. Weizmann [Zic78]. Wellengleichungen [Wig47c, Wig48b]. Wellenmechanik [WBBE30]. Wentzel [TCT16]. Werner [Wig76h]. Wesley [Vig86]. West [SW59]. Western [Hoo73]. Weyl [Maj05a, BBDP08, Ber06, DV97, DP01, Maj05b, MMZ^{+}05, Poi99, RSdG99]. Where [Lan93b, Lan99]. which [TW79b, Wig65i]. While [Wig61h, Wig67g]. White [Wig65m]. Whittlesey [Duf46]. Who [Duf46, Wig68i, Dem15, Har06b, Mar06b, Dem15, Mar69, CH68]. Whyte [Wig65d]. Wick [Ler90]. Widths [Wig49c, Wig51g, Wig65j]. Wiesskopf [DK70a]. Wightman [Rom00]. Wigner [Ada70, Ahl70, Bec58, BB93, Bie69, Coh59, CH68, DV97, Edw60, Fan60, Fes60, Fla68, Gut60, Gut68, Hei74, Hen89, Kir73, Kol60, L.88, Maj05a, Mar67, Mar69, Pai67a, Pai67b, Par65, Pos01, Rom00, Ros70, Sac59, Sch85, Sch88, Sei95, SVW98, SVW00, Set58, Stu88, Szm99, Tel95, Wei69a, Wei97, Wig72o, Wig73q, Abe92, Agr80, Ahl70, AD73, AD84, AC09, APW02, APW03, AF95, AK00, AN09, Ano48, Ano52, Ano58, Ano62, Ano83b, Ano92, Ano02, Ano04, Ano06, Ano08, App87, Ara95, AN92, Arn94, AR96, ALZ00, ACW98, Bac89, BY88, Bai93, BY05, BB65, BH13, BG02, BG03, Bal19, Ban09, Bo09, Bar64, BvdM83, BVV84, BBDP08, Beh91, Beh94, Bel73, BLS03, Ber06, Ber77, BB92, Bie61, BGL67, Ble86, Bon92, Bri65b, BH89, BC62]. Wigner [BW95, BMR65, BM64, BJH68, Bru85, Bru87, BGA07, CMS77, CHDK08, CCT08, CM66, CCB72, CLM75, CM68, CHH05, CD79, Cha01, CIMS08, Cho68, CV80, CPCF99, CM80, Col94, Cox61, Cri67, CUZ01, Cut63, Dah95, DK04, DMMT93, DK70a, DK70b, DK71, DK73a, DK73b, DK75, DK78, Dav74, De 58, DFS20, DB08, DP01, DP02, Die21, DMS84, DA73, Dra86, DLD^{+}21, Ear81, Eck33, Ell92, FP03, FS82, FR13b, FR13c, FR13d, FM03, Fle00, Flü46, Fol78, Fol80, Fon94, For97, Fre93, Frö93, FR03, FKS97, Gel14, GI03, GBO04, Gin63, Gla59, GS04b, GS04a, Goe01, Got71, Gou02, Gou03, Gou81, Gou86a, Gou86b, GP86, GLB92, Gou92, Gra74, GRT91, HS01, Hal81, Hal83, Hal01, HKS86, Har99, HH04b, HS95, HP69, HS83, HB90]. Wigner [HL88, HL90, Hod76, Hol73b, HST74, Hon76, Hor97, How62, HW01, HMW01, HMW02, HMW05, Hud78, IÖRZ67, Isl22, ISRK12, Jan68, Jan81, Jan84, JKM03, JC90, Jha11, JC10, JCT11, JLC14, JF16, JD90, Jor80, Joy01, JBR02, KL03, KPT86, KP90, Kas21, Ker05, Ker85, KN94, Kin58, Kla91, Kli74, KS82, KNS04, Kot05, Kot07, Kou59, KW01, KS07, KLW09, Kue61, KPM97, Lad93, Lag84, Lai94, Lan87, LZ94, Lan93b, Lan99, LWA74, LM89, LB96, Lee22, LNPC92, Ler90, LLW^{+}09, LPN11, Lie90, LO10, LSV06, LSV08, LY09, LM63, LR83, Löw64, LLBC85, Luo04, Mac04, Mae91, MH94, Maj05b, Maj93, Mar03, Mar06a, MMP94, MPPS02, MPPS03, MS00, Mar95, MH98, McD17, MS06, Mez77, MM92, MR98, MDH94, Mol99, MYPL90, Mos62, ML91, MFAB85]. Wigner [MO20, Muk95, MACS04, MMZ^{+}05, Nag88, Nar88, NL69, NAKS04, Ne’04, Nic12, Nou99, OG07, OW92, Olt74a, Omn11, PFG06, PI02, Pal82, PS97, Pal88, PD98, PMHW07, PL75, PČH02, Pla24, Pla25, Poi99, Pos00, PE77, PS02, Pul06, Put64, RNM04, RI03, RI05, Ram00, Ram89, RY03, Ras89, RRC05, RV10, RV11, Rin92, Rob93, RB92, RL81, Rom00, Ros98, Ros99, RR96, RSdG99, RdGS01, SW89, SS05, SSP^{+}91, SD96, Sch96, Scu07b, SIM93, Sec62, Seg03, SNDS14, Sha92, SCS99, SS79, SEC83, SC83, SM05, SM06, SK05, Sta10, Ste25, SAZ90, Ste91, Str79, Stu05, Stu03, Stu04, Sza92, Szm98, TP01, Tal68, TKT^{+}05, TEEBH03, Tem07, TN84, Tis03, TKC^{+}08, TCT16, TC10, VGB87, Vas89, VC01, Vog95, Vos93, Wag81]. Wigner [Wag98, Wal73, WW95, Wei69b, Wei02, Wen98, Wes21, Wig86a, Wig95a, Wig66c, Wig67-28, Wig67-27, WF68, Wig69l, Wig69w, Wig70k, Wig71d, Wig72e, Wig72f, Wig72m, Wig72n, Wig72o, Wig73q, Wig73b, Wig73a, Wig74c, Wig74i, Wig75d, Wig76f, Wig78l, Wig78k, Wig83j, Wig88c, Wig92c, WWM98, WS92, Wig93, Wig95b, Wig96m, Wig96j, Wig97a, Wig97b, Wig98, Wig01b, Wig02c, WS03, WI98, Win76, WN90, Wol67, Won76, Won79, WB05, Woo95, Wu72, Yan10, wYDK74, ZY88, Zha03, ZWE12, Zie88, Zsa76, dSFA05, dBDW^{+}84, dDG83, FB88, Sch67, Tor87]. Wigner-Based [Wen98]. Wigner-Distribution [RB92]. Wigner-function [MMP94]. Wigner-like [SM06, WI98]. Wigner-Measure [MPPS02, MPPS03]. Wignerrol [Frö93, Lad93]. Wigner’s [Cla71a]. Wiley [Wig53a, Wig64f]. Wilhelm [Wig67q]. Will [WG72, Ano30]. winner [Wol67]. Winners [Ano79a]. Wires [AF95]. Wissen [Wig76l]. Without [TWB73a, TWB73b, HKW68, LM89]. Wolfgang [Wig73p]. Work [PE77, WB35, WG96, GW57, GW03, Wig85b]. Working [San95]. Works [Rom00, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, Wig92c, WWM98, Wig93, Wig95b, Wig96m, Wig97a, Wig97b, Wig98, Wig01b]. World [BAB^{+}82, Duf46, GR63, KPS78, Stu18, Wig49b, Wig71k, Ano78b, Ano83a, CGPT21, FF91, Har06a, Jah81, Mar06b, MW46, MW07, PPG^{+}19, Wig49e, Wig79s, Wig82g]. world-shaping [Har06a]. Worlds [Ste25, Wig64r, Wig65n]. would [Wig83l]. WWII [San95].

x [Duf46, Hen89, Wei69a, Wig66f]. X-10 [Wig66f]. XI [Cut63, How62]. XIIIth [Zac84]. XIIth [DGW84]. xviii [Hen89]. XVIIIth [WNY^{+}84].

Yanase [LLW^{+}09, Yan10, GI03, Luo04]. Ye. [Wig69d]. Year [FR13e, Ano15]. Years [Mla98, Rog13, Wig96l, Wol67, Hoi01, TW79b, Vig86, Wig62d, Wig76o, Wig78d, Wig80b, Wig80e]. Yield [Kas21]. Yiftah [Wig61e]. York [Bug77, Duf46, Gre86, Wig53a, Wig64f, Wig65c, Wig69s, Ano75, Wig64c, Wig76c]. Yorker [BW61]. young [Frö93, Lad93, Wig95d]. Yukawa [FR13c, Hal83, PI02]. Yukawa-type [PI02].

Z [Wig65c, Ano15]. zakony [Vig64, Vig65]. Zbl [Wig85d]. Zeilinger [Lee22]. Zeitumkehr [Wig32c]. Zerfall [PW25, Wig25]. zero [Jor78, Wig56e]. zeros [Bru85, Bru87]. Ziele [Wig58i]. zone [CV80]. Zones [BSW36, BSW64]. Zur [vNW28a, vNW28b, vNW28c]. zweiatomigen [WW28].

References

Abe:1992:SWF

[Abe92] Sumiyoshi Abe. Super Wigner function. Journal of Mathematical Physics, 33(5):1690–1694, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

            //jmp.aip.org/resource/1/jmapaq/v33/i5/p1690_s1.

Amaldi:2012:ALF

[ABES12] Edoardo Amaldi, S. (Saverio) Braccini, Antonio Ereditato, and Paola Scampoli, editors. The adventurous life of Friedrich Georg Houtermans, physicist (1903–1966). Springerbriefs in Physics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-32854-7 (paperback), 3-642-32855-5 (e-book). ISSN 2191-5423. xiv + 152 pp. LCCN Q124.6-127.2; QC16.H688 A43 2012. URL http://www.sif.it/

attivita/saggiatore/recensioni/e_amaldi.

Alamo:2009:PTT

[AC09] Nieves Alamo and Carlos Criado. Parallel transport and Thomas– Wigner rotation. American Mathematical Monthly, 116(5):439– 446, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/
40391119.

Allison:1942:RCE

[ACF^{+}42] Samuel K. Allison, C. M. Cooper, Enrico Fermi, Eugene P. Wigner, and Leo Szilard. Report of the committee for the examination of the Moore–Leverett design of a He-cooled plant. Report CE-324, US Atomic Energy Commission, Washington, DC, USA, 1942. Included are four pages that seem to have been written by Enrico Fermi, chairman of the committee. Undated: year chosen according to report number.

Atakishiyev:1998:WDF

[ACW98] Natig M. Atakishiyev, Sergey M. Chumakov, and Kurt Bernardo Wolf. Wigner distribution function for finite systems. Journal of Mathematical Physics, 39(12):6247–6261, December 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Akiyama:1973:UGF

[AD73] Yoshimi Akiyama and J. P. Draayer. A user’s guide to Fortran programs for Wigner and Racah coefficients of SU_{3}. Computer Physics Communications, 5(6):405–406, June 1973. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465573900775.

Akiyama:1984:UGF

[AD84] Yoshimi Akiyama and J. P. Draayer. A user’s guide to Fortran programs for Wigner and Racah coefficients of SU(3). Computer Physics Communications, 35(1–3):C–194–??, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S001046558482439X.

Adair:1970:BRP

[Ada70] Robert K. Adair. Book review: E. P. Wigner, Survival and the Bomb: Methods of Civil Defense. American Journal of Physics, 38(11):1367, November 1970. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.

org/link/ajpias/v38/i11/p1367/s1.

Altland:1995:SSM

[AF95] Alexander Altland and Dirk Fuchs. Spectral statistics of mesoscopic wires: Crossover from Wigner–Dyson to Poisson regime. Physical Review Letters, 74(??):4269–??, May 22, 1995. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-

0145. URL http://link.aps.org/doi/10.1103/PhysRevLett.
74.4269.

Agrawala:1980:WET

[Agr80] Vishnu K. Agrawala. Wigner–Eckart theorem for an arbitrary group or Lie algebra. Journal of Mathematical Physics, 21(7): 1562–1565, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i7/p1562_s1.

Ahlrichs:1970:KBW

[Ahl70] Reinhart Ahlrichs. Die Konvergenz der Brillouin–Wigner Störungsrechnung. (German) [Convergence of the Brillouin– Wigner perturbation calculation]. International Journal of Quantum Chemistry, 4(2):149–172, March 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Altland:2000:WDS

[AK00] Alexander Altland and Alex Kamenev. Wigner–Dyson statistics from the Keldysh σ-model. Physical Review Letters, 85(??):5615– ??, December 25, 2000. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.

aps.org/doi/10.1103/PhysRevLett.85.5615.

Arnold:2000:DVS

[ALZ00] Anton Arnold, Horst Lange, and Paul F. Zweifel. A discretevelocity, stationary Wigner equation. Journal of Mathematical Physics, 41(11):7167–7180, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arnold:1992:NAD

[AN92] Anton Arnold and Francis Nier. Numerical analysis of the deterministic particle method applied to the Wigner equation. Mathematics of Computation, 58(198):645–669, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.jstor.org/stable/2153207.

Ammari:2009:MFL

[AN09] Z. Ammari and F. Nier. Mean field limit for bosons and propagation of Wigner measures. Journal of Mathematical Physics, 50(4): 042107, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i4/p042107_s1.

Anderson:1973:TQA

[And73] Herbert Anderson. Three questions about the sustained nuclear chain reaction. The University of Chicago Magazine, 65(5):3– 7, March/April 1973. ISSN 0041-9508. URL http://library.

ucsd.edu/dc/object/bb3960604v.

Anonymous:1930:PTM

[Ano30] Anonymous. Princeton Trustees meet: will concentrate departments — Berlin lecturers named. New York Times, ??(??):8, January 10, 1930. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://

search.proquest.com/hnpnewyorktimes/docview/99086947/
fulltextPDF/13EDE5EACA87A3ADEAE/1. This article appears to

be the first mention of John von Neumann in this newspaper, noting his appointment, with Eugene Wigner, as lecturers in mathematical physics.

Anonymous:1948:BAK

[Ano48] Anonymous. Book announcement: K. K. Darrow, Physical Science and Human Values, with a foreword by E. P. Wigner. Physics Today, 1(5):37, May 1948. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v1/i1/p37/s3.

Anonymous:1952:NMW

[Ano52] Anonymous. News-makers: Warner, Wigner, and Fisk appointed to AEC Committee. Chemical & Engineering News Archive, 30 (41):4303–4307, 1952. CODEN ???? ISSN ???? URL http://

pubs.acs.org/doi/abs/10.1021/cen-v030n041.p4303.

Anonymous:1954:NRT

[Ano54a] Anonymous, editor. New research Techniques in physics: symposium, July 15–29, 1952. Academia Brasileira de Ciencias, Rio de Janeiro, 1954. LCCN ????

Anonymous:1954:PIC

[Ano54b] Anonymous, editor. Proceedings of the International Conference on Theoretical Physics, Kyoto and Tokyo, September 1953. Science Council of Japan, Ueno Park, Tokyo, Japan, 1954. LCCN ????

Anonymous:1958:PWG

[Ano58] Anonymous. Princeton’s Wigner gets third Enrico Fermi Award. The Science News-Letter, 74(24):372, December 13, 1958. CO-DEN SNLEAI. ISSN 0096-4018 (print), 2326-1285 (electronic).

URL http://www.jstor.org/stable/3939817.

Anonymous:1961:LPK

[Ano61] Anonymous, editor. The Logic of Personal Knowledge: Essays presented to Michael Polányi on his seventieth birthday, 11th March 1961. Free Press, New York, NY, USA, 1961. xi + 247 pp. LCCN Q171 .L855 1961.

Anonymous:1962:EPW

[Ano62] Anonymous. To Eugene Paul Wigner on his sixtieth birthday. Reviews of Modern Physics, 34(4):587–591, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
34.587;   http://rmp.aps.org/abstract/RMP/v34/i4/p587_
1.

Anonymous:1963:ANN

[Ano63] Anonymous. Awards: Nobelmen & Nobelwoman. Time Magazine, ??(??):??, November 15, 1963. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic). News of the award of the 1963 Nobel Prize in Physics to Eugene Wigner, Maria Goeppert Mayer and J. Hans D. Jensen.

Anonymous:1965:SLE

[Ano65] Anonymous, editor. Survey (London, England: 1961): The state of Soviet science. MIT Press, Cambridge, MA, USA, 1965. 209 pp. LCCN Q127.R9 S95.

Anonymous:1968:LPE

[Ano68] Anonymous, editor. From a life of physics: evening lectures at the International Centre for Theoretical Physics. International Atomic Energy Agency, Vienna, Austria, 1968. 79 pp. LCCN ???? A special supplement of the IAEA Bulletin.

Anonymous:1975:CSA

[Ano75] Anonymous, editor. The centrality of science and absolute values: proceedings of the fourth International Conference on the Unity of the Sciences, November 27–30, 1975, New York. International Cultural Foundation, New York, NY, USA, 1975. ISBN 0-89226-003-3, 0-89226-004-1 (paperback). LCCN Q175.4 .I57 1975.

Anonymous:1977:HAH

[Ano77a] Anonymous, editor. Hellenike Anthropistike Hetaireira. ????, ????, 1977. ISBN ???? ???? pp. LCCN ????

Anonymous:1977:SAV

[Ano77b] Anonymous, editor. The search for absolute values: harmony among the sciences: proceedings of the Fifth International Conference on the Unity of the Sciences, November 26–28, 1976, Washington, DC. International Cultural Foundation Press, New York, NY, USA, 1977. ISBN 0-89226-005-X, 0-89226-006-8 (paperback). LCCN Q174 .I565 1976.

Anonymous:1978:PAP

[Ano78a] Anonymous, editor. Proceedings of the Atoms for Peace Awards, 1957–1969: a memorial to Henry Ford and Edsel Ford. MIT Press, Cambridge, MA, USA, 1978. ISBN 0-262-11068-7. LCCN QC792.7 .P76.

Anonymous:1978:SAV

[Ano78b] Anonymous, editor. The search for absolute values in a changing world: proceedings of the Sixth International Conference on the Unity of the Sciences, November 25–27, 1977, San Francisco. International Cultural Foundation Press, New York, NY, USA, 1978. ISBN 0-89226-007-6, 0-89226-008-4 (paperback). LCCN BD232 .I57 1977.

Anonymous:1978:TMQ

[Ano78c] Anonymous, editor. Theory of Measurement in Quantum Mechanics. Physical Society of Japan, ????, Japan, 1978. ISBN ???? LCCN ????

Anonymous:1979:MNP

[Ano79a] Anonymous, editor. 29th Meeting of Nobel Prize Winners in Lindau, June 25–29, 1979. ????, ????, 1979. ISBN ???? LCCN ????

Anonymous:1979:REE

[Ano79b] Anonymous, editor. The re-evaluation of existing values and the search for absolute values: proceedings of the seventh International Conference on the Unity of the Sciences, November 24– 26, 1978, Boston, Massachusetts.: Unity of the sciences. International Cultural Foundation Press, New York, NY, USA, 1979. ISBN 0-89226-010-6 (paperback) (set), 0-89226-009-2 (set). LCCN Q175.4 .I57 1978.

Anonymous:1983:AVC

[Ano83a] Anonymous, editor. Absolute values and the creation of the new world: proceedings of the Eleventh International Conference on the Unity of the Sciences, November 25–28, 1982, Philadelphia, Pennsylvania. International Cultural Foundation Press, New York, NY, USA, 1983. ISBN 0-89226-020-3 (set), 0-89226-021-1 (paperback : set). LCCN T14.5 .I58 1982.

Anonymous:1983:PEW

[Ano83b] Anonymous. Photo of Eugene Wigner. Foundations of Physics, 13(1):ii, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889405.

Anonymous:1992:VEW

[Ano92] Anonymous. Vignettes: Eugene Wigner. Science, 258(5089): 1821, December 11, 1992. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/258/5089/1821.full.pdf.

Anonymous:19xx:AEL

[Anoxxa] Anonymous. The birth of modern physics. Los Alamos National Laboratory Web document, 19xx. URL http://www.lanl.gov/

history/road/birthofmodernphysics.shtml. Section The Ein-

stein Letter discusses the Einstein–Roosevelt letter of 2 August 1939, and says “drafted mostly by Szilard.”.

Anonymous:19xx:BMP

[Anoxxb] Anonymous. The birth of modern physics. Los Alamos National Laboratory Web document, 19xx. URL http://www.lanl.gov/

history/road/birthofmodernphysics.shtml. Section The Ein-

stein Letter discusses the Einstein–Roosevelt letter of 2 August 1939, and says “drafted mostly by Szilard.”.

Anonymous:2002:EW

[Ano02] Anonymous. On Eugene Wigner (1902-1995). ScienceWeek, 6 (46):??, November 15, 2002. ISSN 1529-1472.

Anonymous:2004:EWA

[Ano04] Anonymous. Eugene Wigner (1902–1995) and the atomic bomb. ScienceWeek, ??(??):??, ???? 2004. ISSN 1529-1472. URL http:

//scienceweek.com/2004/rmps-38.htm.

Anonymous:2006:ODW

[Ano06] Anonymous. One-dimensional Wigner crystal. Science, 313 (5784):144, July 14, 2006. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/313/5784/144.5.full.pdf.

Anonymous:2008:WCQ

[Ano08] Anonymous. Wigner crystallization in a quasi-three-dimensional electronic system. Nature Physics, 4(12):936–939, October 5, 2008. CODEN NPAHAX. ISSN 1745-2473 (print), 1745-2481

(electronic). URL http://www.nature.com/nphys/journal/v4/
n12/full/nphys1094.html.

Anonymous:2015:SZM

[Ano15] Anonymous. Sandia’s Z Machine helps solve Saturn’s 2-billion-year age gap. Scientific Computing, ??(??):??, June 29, 2015. CO-DEN SCHRCU. ISSN 1930-5753 (print), 1930-6156 (electronic).

URL http://www.scientificcomputing.com/news/2015/06/
sandias-z-machine-helps-solve-saturns-2-billion-year-

age-gap. From the article: “Experiments at Sandia s Z machine may help solve that problem when they verified an 80-year-old untested proposition that molecular hydrogen, normally an insulator, becomes metallic if squeezed by enough pressure. At that point, a lattice of hydrogen molecules would break up into individual hydrogen atoms, releasing free-floating electrons that could carry a current, physicists Eugene Wigner and Hilliard Huntington [WH35] predicted in 1935.”.

Anonymous:2019:QES

[Ano19] Anonymous. A quantum experiment suggests there’s no such thing as objective reality: Physicists have long suspected that quantum mechanics allows two observers to experience different, conflicting realities. Now they ve performed the first experiment that proves it. Technology Review (M.I.T.), ?? (??):??, March 12, 2019. CODEN TEREAU. ISSN 0040-

1692. URL https://www.technologyreview.com/s/613092/
a-quantum-experiment-suggests-theres-no-such-thing-as-
objective-reality/.

Applebaum:1987:SWW

[App87] David Applebaum. The stochastic Wigner–Weisskopf atom. Journal of Mathematical Physics, 28(8):1858–1863, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i8/p1858_s1.

Alonso:2002:WFC

[APW02] Miguel Angel Alonso, George S. Pogosyan, and Kurt Bernardo Wolf. Wigner functions for curved spaces. I. On hyperboloids. Journal of Mathematical Physics, 43(12):5857–5871, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alonso:2003:WFC

[APW03] Miguel Angel Alonso, George S. Pogosyan, and Kurt Bernardo Wolf. Wigner functions for curved spaces. II. On spheres. Journal of Mathematical Physics, 44(4):1472–1489, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arnold:1996:OSM

[AR96] Anton Arnold and Christian Ringhofer. An operator splitting method for the Wigner–Poisson problem. SIAM Journal on Numerical Analysis, 33(4):1622–1643, August 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (elec-

tronic). URL http://epubs.siam.org/sam-bin/dbq/article/
23882;   http://www.jstor.org/stable/2158320.

Arai:1995:SES

[Ara95] Takahiro Arai. Some extensions of semiclassical limit h̄ → 0 for Wigner functions on phase space. Journal of Mathematical Physics, 36(2):622–630, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arnold:1994:MSC

[Arn94] Anton Arnold. A mixed spectral-collocation and operator splitting method for the Wigner–Poisson equation. In Gautschi [Gau94], pages 249–253. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Ahmad:1975:ITD

[AW75] S. M. W. Ahmad and E. P. Wigner. Invariant theoretic derivation of the connection between momentum and velocity. Il Nuovo Cimento A (11), 28(1):1–11, ???? 1975. CODEN NIFAAM.

ISSN ???? URL http://www.springerlink.com/content/
tq2h218t45302k37.

Bethe:1975:SSN

[BAA^{+}75] Hans A. Bethe, Luis Alvarez, Peter Auer, William O. Baker, John Bardeen, Robert F. Bacher, Felix Bloch, Norris E. Bradbury, Harold Brown, Richard H. Chamberlain, Cyril L. Comar, Arthur Kantrowitz, Ralph E. Lapp, Joshua Lederberg, Willard F. Libby, Franklin A. Long, Edwin M. McMillan, Kenneth S. Pitzer, Edward M. Purcell, I. I. Rabi, Norman Rasmussen, Roger Revelle, Glenn T. Seaborg, Frederick Seitz, Edward Teller, James A. Van Allen, Warren Weaver, Alvin Weinberg, Victor F. Weisskopf, Edward Wenk, Jr., Eugene Wigner, and Richard Wilson. 32 scientists speak out: “no alternative to nuclear power”. Bulletin of the Atomic Scientists, 31(3):4–5, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Wei75] about the incorrect inclusion of Victor Weisskopf in the authors. See comments [Gir75, Smi75, Sau75, Gul75].

Bardis:1982:OWP

[BAB^{+}82] P. D. Bardis, M. Alonso, R. Berger, C. S. Blanco, D. Charles, J. M. Chaves, M. Dufrenne, W. B. Lewis, G. Radnitzky, J. Silverman, and E. Wigner. Obstacles to world peace. International Social Science Review, 57(2):101–105, ???? 1982. ISSN 0278-2308.

Bacry:1989:NLS

[Bac89] Henri Bacry. The notions of localizability and space: From Eugene Wigner to Alain Connes. Nuclear Physics B, Proceedings Supplements, 6(1):222–230, March 1989. CODEN NPBSE7. ISSN 0920-5632 (print), 1873-3832 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0920563289904428.

Bai:1993:CREa

[Bai93] Z. D. Bai. Convergence rate of expected spectral distributions of large random matrices. Part I. Wigner matrices. Annals of Probability, 21(2):625–648, April 1993. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989261;
http://www.jstor.org/stable/2244669.

Ballentine:2019:MW

[Bal19] Leslie E. Ballentine. A meeting with Wigner. Foundations of Physics, 49(8):783–785, August 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Bankoff:2004:NHR

[Ban04] S. George Bankoff. Notes on Hanford reactor start-up. American Journal of Physics, 57(4):17–18, April 2004. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http:

//scitation.aip.org/content/aip/magazine/physicstoday/
article/57/4/10.1063/1.1752408. See reply [Wei04].

Bangu:2009:WPM

[Ban09] Sorin Bangu. Wigner’s puzzle for mathematical naturalism. International Studies in the Philosophy of Science, 23(3):245– 263, 2009. CODEN ???? ISSN 0269-8595 (print), 1469-9281

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/02698590903196015.

Bargmann:1964:NWT

[Bar64] V. Bargmann. Note on Wigner’s theorem on symmetry operations. Journal of Mathematical Physics, 5(7):862–868, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v5/i7/p862_s1.

Baird:1965:RSL

[BB65] G. E. Baird and L. C. Biedenharn. On the representations of the semisimple Lie groups. V. Some explicit Wigner operators for SU_{3}. Journal of Mathematical Physics, 6(12):1847–1854, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v6/i12/p1847_s1.

Bertrand:1992:CAW

[BB92] J. Bertrand and P. Bertrand. A class of affine Wigner functions with extended covariance properties. Journal of Mathematical Physics, 33(7):2515–2527, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v33/i7/p2515_s1. See er-

ratum [BB93].

Bertrand:1993:ECA

[BB93] J. Bertrand and P. Bertrand. Erratum: “A class of affine Wigner functions with extended covariance properties [J. Math. Phys. 33, 2515 (1992)]. Journal of Mathematical Physics, 34(2):885, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v34/i2/p885_s1. See [BB92].

Bastos:2008:WWF

[BBDP08] Catarina Bastos, Orfeu Bertolami, Nuno Costa Dias, and João Nuno Prata. Weyl–Wigner formulation of noncommutative quantum mechanics. Journal of Mathematical Physics, 49(7): 072101, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i7/p072101_s1.

Brittin:1962:WDF

[BC62] Wesley E. Brittin and Willard R. Chappell. The Wigner distribution function and second quantization in phase space. Reviews of Modern Physics, 34(4):620–627, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
34.620;   http://rmp.aps.org/abstract/RMP/v34/i4/p620_
1.

Bethe:1989:LP

[BDH^{+}89] Hans A. (Hans Albrecht) Bethe, P. A. M. Dirac, W. Heisenberg, E. P. Wigner, O. Klein, L. D. Landau, and E. M. Lifshitz, editors. From a life of physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1989. ISBN 9971-5-0937-7. ix + 92 pp. LCCN QC71 .F74 1989. URL http://

adsabs.harvard.edu/abs/1989liph.book.....B.

Becker:1958:BRB

[Bec58] Richard L. Becker. Book review: Nuclear structure, by Leonard Eisenbud and Eugene P. Wigner. 128 pages, 6 × 9.25 in. Princeton, Princeton University Press, 1958. Price $4.00. Journal of The Franklin Institute, 266(4):311–312, October 1958. CO-DEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0016003258904800.

Beckers:1972:HCS

[Bec72] J. Beckers. Helicity and canonical spin operators of the Poincaré algebra. Journal of Mathematical Physics, 13(6):915–918, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v13/i6/p915_s1.

Behncke:1991:ACH

[Beh91] Horst Behncke. Absolute continuity of Hamiltonians with von Neumann–Wigner potentials. Proceedings of the American Mathematical Society, 111(2):373–384, February 1991. CODEN PAM-YAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http:

//www.jstor.org/stable/2048326.

Behncke:1994:FHW

[Beh94] H. Behncke. The m-function for Hamiltonians with Wigner–von Neumann potentials. Journal of Mathematical Physics, 35(4): 1445–1462, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i4/p1445_s1.

Bell:1973:BRD

[Bel73] John S. Bell. Book review: The Dirac story Aspects of Quantum Theory, by Abdus Salam and E. P. Wigner. Science, 180 (4089):943–944, June 1, 1973. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1736041.

Berry:1977:SCM

[Ber77] M. V. Berry. Semi-classical mechanics in phase space: a study of Wigner’s function. Philosophical Transactions of the Royal Society of London. Series A, Mathematical and Physical Sciences, 287 (1343):237–271, October 4, 1977. CODEN ???? ISSN ???? URL

http://www.jstor.org/stable/74926.

Bergeron:2006:WWT

[Ber06] H. Bergeron. The Wigner–Weyl transformation and the quantum path integral. Journal of Mathematical Physics, 47(4): 042101, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i4/p042101_s1.

Bakic:2002:WTH

[BG02] Damir Bakić and Boris Guljaš. Wigner’s Theorem in Hilbert C^{∗}- modules over C^{∗}-algebras of compact operators. Proceedings of the American Mathematical Society, 130(8):2343–2349, August 2002. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826

(electronic). URL http://www.jstor.org/stable/2699471.

Bakic:2003:WTC

[BG03] Damir Bakic and Boris Guljas. Wigner’s theorem in a class of Hilbert C^{∗}-modules. Journal of Mathematical Physics, 44 (5):2186–2191, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Butucea:2007:MAE

[BGA07] Cristina Butucea, Mădălin Guţă, and Luis Artiles. Minimax and adaptive estimation of the Wigner function in quantum homodyne tomography with noisy data. Annals of Statistics, 35(2):465–494, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://www.jstor.org/stable/25463565.

Biedenharn:1967:CDW

[BGL67] L. C. Biedenharn, A. Giovannini, and J. D. Louck. Canonical definition of Wigner coefficients in U_{n}. Journal of Mathematical Physics, 8(4):691–700, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v8/i4/p691_s1.

Bromley:1970:FP

[BH70] D. Allan (David Allan) Bromley and Vernon W. Hughes, editors. Facets of physics. Academic Press, New York, USA, 1970. LCCN QC780 .F3.

Barschall:1971:PPN

[BH71] H. H. (Henry Herman) Barschall and W. (Willy) Haeberli, editors. Polarization phenomena in nuclear reactions: proceedings of the Third International Symposium, Madison 1970. University of Wisconsin Press, Madison, WI, USA, 1971. ISBN 0-299-05890-5. LCCN QC794 .I545 1970.

Brink:1989:IET

[BH89] Jean R. Brink and Roland Haden. Interviews with Edward Teller and Eugene P. Wigner. Annals of the History of Computing, 11(3):177–178, July/September 1989. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1989/pdf/a3177.pdf;   http://www.computer.org/
annals/an1989/a3177abs.htm.

Baker:2013:HSS

[BH13] David John Baker and Hans Halvorson. How is spontaneous symmetry breaking possible? Understanding Wigner’s theorem in light of unitary inequivalence. Studies in History and Philosophy of Modern Physics, 44(4):464–469, November 2013. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000798.

Badash:1986:NFR

[BHT86] Lawrence Badash, Elizabeth Hodes, and Adolph Tiddens. Nuclear fission: Reaction to the discovery in 1939. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 130(2):196–231, June 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://

www.jstor.org/stable/987181.

Biedenharn:1961:WCG

[Bie61] L. C. Biedenharn. Wigner coefficients for the R_{4} group and some applications. Journal of Mathematical Physics, 2(3):433– 441, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i3/p433_s1.

Biedenharn:1969:BRE

[Bie69] L. C. Biedenharn. Book review: Eugene P. Wigner, Jr. and James D. Talman, Jr. Special Functions. A Group Theoretic Approach.

American Journal of Physics, 37(10):1073–??, October 1969. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/ajpias/v37/i10/p1073/s2.

Broyles:1968:LEW

[BJH68] Arthur A. Broyles, Stanley L. Jaki, and William Cornelius Hall. Letters to the Editor: On the Wigner–Feld exchange. Bulletin of the Atomic Scientists, 24(5):45–46, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel67, WF68].

Biedenharn:1972:SCTb

[BL72] L. C. Biedenharn and J. D. Louck. On the structure of the canonical tensor operators in the unitary groups. II. The tensor operators in U(3) characterized by maximal null space. Journal of Mathematical Physics, 13(12):1985–2001, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i12/p1985_s1.

Bleyer:1986:BRA

[Ble86] Ulrich Bleyer. Book review: Asim O. Barut, Alwyn Van der Merwe and Jean-Pierre Vigier (eds.): Quantum, Space and Time — The Quest Continues. Studies and Essays in Honour of Louis de Broglie, Paul Dirac and Eugene Wigner. Cambridge University Press, Cambridge 1984, ISBN 0-521-31911-0, Preis: £25. Astronomische Nachrichten, 307(3), 1986. CODEN ASNAAN. ISSN 0004-6337 (print), 1521-3994 (electronic).

Benet:2003:WDS

[BLS03] L. Benet, F. Leyvraz, and T. H. Seligman. Wigner–Dyson statistics for a class of integrable models. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 68(??):045201, October 21, 2003. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRevE.68.045201.

Brown:1964:DEB

[BM64] W. Byers Brown and William J. Meath. On the differential equations of Brillouin–Wigner perturbation theory. Proceedings of the National Academy of Sciences of the United States of America, 52(1):65–67, July 15, 1964. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.jstor.org/

stable/72081.

Brody:1965:RRW

[BMR65] T. A. Brody, M. Moshinsky, and I. Renero. Recursion relations for the Wigner coefficients of unitary groups. Journal of Mathematical Physics, 6(10):1540–1546, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i10/p1540_s1.

Bar-on:2009:DWF

[Bo09] T. Bar-on. Discrete Wigner function by symmetric informationally complete positive operator valued measure. Journal of Mathematical Physics, 50(7):072106, July 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i7/
p072106_s1.

Bondi:1992:RWE

[Bon92] H. Bondi. The recollections of Eugene Wigner as told to Andrew Szanton. Nature, 360(6402):374–375, November 26, 1992. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Breit:1972:BRB

[Bre72] G. Breit. Book review: ???? American Scientist, 60(2): 246, March 1972. CODEN AMSCAC. ISSN 0003-0996 (print),

1545-2786 (electronic). URL http://www.jstor.org/stable/
27843047.

Brink:1965:NF

[Bri65a] David Maurice Brink. Nuclear Forces, volume 354 of Commonwealth and international library: Selected readings in physics. Pergamon Press, New York, NY, USA, 1965. ISBN 0-08-011034-7. viii + 232 pp. LCCN QC173 .B8513 1965.

Brink:1965:THW

[Bri65b] David Maurice Brink. The theories of Heisenberg, Wigner and Majorana. In Nuclear Forces [Bri65a], chapter 2, pages 13–25. ISBN 0-08-011034-7. LCCN QC173 .B8513 1965.

Brudno:1985:NZWa

[Bru85] Simcha Brudno. Nontrivial zeros of the Wigner (3-j ) and Racah (6-j ) coefficients. I. Linear solutions. Journal of Mathematical Physics, 26(3):434–435, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v26/i3/p434_s1.

Brudno:1987:NZW

[Bru87] Simcha Brudno. Nontrivial zeros of the Wigner (3-j ) and Racah (6-j ) coefficients. II. Some nonlinear solutions. Journal of Mathematical Physics, 28(1):124–127, January 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i1/p124_
s1.

Bouckaert:1936:TBZ

[BSW36] L. P. Bouckaert, R. Smoluchowski, and E. Wigner. Theory of Brillouin zones and symmetry properties of wave functions in crystals. Physical Review (2), 50(1):58–67, July 1, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.50.58.

Bouckaert:1964:TBZ

[BSW64] L. P. Bouckaert, R. Smoluchowski, and E. Wigner. Theory of Brillouin zones and symmetry properties of wave functions in crystals. In Meijer [Mei64], pages 1–10. LCCN ????

Besieris:1973:KEQ

[BT73] Ioannis M. Besieris and Frederick D. Tappert. Kinetic equations for the quantized motion of a particle in a randomly perturbed potential field. Journal of Mathematical Physics, 14(12):1829–1836, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i12/p1829_s1.

Besieris:1976:SWK

[BT76] Ioannis M. Besieris and Frederick D. Tappert. Stochastic wavekinetic theory in the Liouville approximation. Journal of Mathematical Physics, 17(5):734–743, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i5/p734_s1.

Broyles:1973:GH

[BTW73] A. A. Broyles, E. Teller, and E. Wigner. A glass house. Survive, 6(5):1–??, ???? 1973. CODEN ???? ISSN 0039-6354.

Bugliarello:1977:STM

[Bug77] George Bugliarello, editor. Science, technology, and modern society: inaugural symposium and lectures following the inauguration of George Bugliarello as first president of the Polytechnic Institute of New York, March 13–14, 1975. Polytechnic Press, Brooklyn, NY, USA, 1977. LCCN Q175.4 .S38.

Buttiker:1988:CST

[But88] M. Buttiker. Coherent and sequential tunneling in series barriers. IBM Journal of Research and Development, 32(1):63–75, January 1988. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Biedenharn:1965:QTA

[BV65] L. C. Biedenharn and Hendrik Van Dam, editors. Quantum theory of angular momentum: a collection of reprints and original papers. Perspectives in physics. Academic Press, New York, USA, 1965. vii + 332 pp. LCCN QC174.1 .B46.

Barut:1983:EPW

[BvdM83] Asim O. Barut and Alwyn van der Merwe. To Eugene Paul Wigner on his eightieth birthday. Foundations of Physics, 13 (1):3–5, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889406.

Barut:1984:QST

[BVV84] A. O. (Asim Orhan) Barut, Alwyn Van der Merwe, and Jean-Pierre Vigier, editors. Quantum, space, and time–the quest continues: studies and essays in honour of Louis de Broglie, Paul Dirac, and Eugene Wigner. Cambridge monographs on physics. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-31911-0 (paperback). vii + 659 pp. LCCN QC174.125 .Q376 1984.

Breit:1935:NME

[BW35] G. Breit and E. Wigner. Note on Majorana’s exchange energy. Physical Review (2), 48(11):918–919, December 1, 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.48.918.2.

Breit:1936:CSN

[BW36] G. Breit and E. Wigner. Capture of slow neutrons. Physical Review (2), 49(7):519–531, April 1, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

link.aps.org/doi/10.1103/PhysRev.49.519.

Breit:1937:DL

[BW37] G. Breit and E. Wigner. The disintegration of Li^{8}. Physical Review (2), 51(7):593, April 1, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRev.51.593.2.

Breit:1938:SRN

[BW38] G. Breit and E. Wigner. The saturation requirements for nuclear forces. Physical Review (2), 53(12):998–1003, June 15, 1938. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://link.aps.org/doi/10.1103/PhysRev.53.998.

Bohr:1939:MNF

[BW39] Niels Bohr and John Archibald Wheeler. The mechanism of nuclear fission. Physical Review (2), 56(5):426–450, September 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v56/
i5/p426_1.

Bargmann:1948:GTD

[BW48] V. Bargmann and E. P. Wigner. Group theoretical discussion of relativistic wave equations. Proceedings of the National Academy of Sciences of the United States of America, 34(5):211–223, May 15, 1948. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490

(electronic). URL http://www.jstor.org/stable/88422.

Birkhoff:1961:NRT

[BW61] Garrett Birkhoff and Eugene P. (Eugene Paul) Wigner, editors. Nuclear reactor theory: [proceedings of the 11th Symposium in Applied Mathematics of the American Mathematical Society held at the Hotel New Yorker, April 23–25, 1959], volume 11 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1961. ISSN 0160-7634. LCCN QC787.N8 S9 1959.

Boyer:1973:DIC

[BW73] Charles P. Boyer and Kurt Bernardo Wolf. Deformations of inhomogeneous classical Lie algebras to the algebras of the linear groups. Journal of Mathematical Physics, 14(12):1853–1859, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i12/p1853_s1.

Brocker:1995:MSP

[BW95] T. Bröcker and R. F. Werner. Mixed states with positive Wigner functions. Journal of Mathematical Physics, 36(1):62–75, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Broyles:1976:CDN

[BWD76a] Arthur A. Broyles, Eugene P. Wigner, and Sidney D. Drell. Civil defence: The new debate. Survival: Global Politics and Strategy, 18(5):217–224, April 1976. CODEN ???? ISSN 0039-6338 (print), 1468-2699 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00396337608441639.

Broyles:1976:CDL

[BWD76b] Arthur A. Broyles, Eugene P. Wigner, and Sidney D. Drell. Civil defense in limited war — a debate. Physics Today, 29(4):44– 47, 50, 52–53, 55–57, April 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v29/i4/p44/s1. In favor: Broyles and

Wigner; Opposed: Drell.

Bai:1988:NSC

[BY88] Z. D. Bai and Y. Q. Yin. Necessary and sufficient conditions for almost sure convergence of the largest eigenvalue of a Wigner matrix. Annals of Probability, 16(4):1729–1741, October 1988. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991594;
http://www.jstor.org/stable/2243989.

Bai:2005:CSE

[BY05] Z. D. Bai and J. Yao. On the convergence of the spectral empirical process of Wigner matrices. Bernoulli: official journal of the Bernoulli Society for Mathematical Statistics and Probability, 11 (6):1059–1092, December 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://www.jstor.org/

stable/25464779.

Balazs:1974:SFS

[BZ74] N. L. Balazs and G. G. Zipfel, Jr. The semiclassical fermion μ-space density in three dimensions. Journal of Mathematical Physics, 15(12):2086–2089, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i12/p2086_s1.

Chacon:1972:EMF

[CCB72] E. Chacón, Mikael Ciftan, and L. C. Biedenharn. On the evaluation of the multiplicity-free Wigner coefficients of U(n). Journal of Mathematical Physics, 13(5):577–590, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i5/
p577_s1.

Castro:2008:WOT

[CCT08] P. G. Castro, B. Chakraborty, and F. Toppan. Wigner oscillators, twisted Hopf algebras, and second quantization. Journal of Mathematical Physics, 49(8):082106, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i8/
p082106_s1.

Chatterjee:1979:CWS

[CD79] R. Chatterjee and J. M. Dixon. Comment on the Wigner 9-j symbol. Journal of Mathematical Physics, 20(8):1622–1623, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v20/i8/p1622_s1.

Castanos:1978:GFC

[CFM78] O. Castaños, A. Frank, and M. Moshinsky. The gradient formula for the O(5) ⊆ O(3) chain of groups. Journal of Mathematical Physics, 19(8):1781–1789, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v19/i8/p1781_s1.

Clark:1978:OEF

[CFW78] John W. Clark, Michael W. Friedlander, and Eugene P. Wigner. Obituary: Eugene Feenberg. Physics Today, 31(3):81–??, March 1978. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v31/i3/
p81/s1.

Calosi:2021:EAS

[CGPT21] Claudio Calosi, Pierluigi Graziani, Davide Pietrini, and Gino Tarozzi, editors. Experience, abstraction and the scientific image of the world: festschrift for Vincenzo Fano. Epistemologia. FrancoAngeli, Milano, Italy, 2021. ISBN 88-351-2139-6. 402 pp. LCCN Q175 .E95 2021.

Cooling:1968:BRB

[CH68] B. Franklin Cooling, III and Jack B. Hilliard. Book review: Who Speaks for Civil Defense?, by Eugene P. Wigner and Walter Cronkite. Military Affairs, 32(2):85, October 1968. CODEN ????

ISSN ???? URL http://www.jstor.org/stable/1985549.

Chayut:2001:PGE

[Cha01] Michael Chayut. From the periphery: the genesis of Eugene P. Wigner’s application of group theory to quantum mechanics. Foundations of Chemistry, 3(1):55–78, ???? 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A:1011431408763.

Camjayi:2008:CCW

[CHDK08] A. Camjayi, K. Haule, V. Dobrosavljević, and G. Kotliar. Coulomb correlations and the Wigner–Mott transition. Nature Physics, 4(12):932–935, October 26, 2008. CODEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (electronic). URL http:

//www.nature.com/nphys/journal/v4/n12/full/nphys1106.
html.

Chae:2005:RRW

[CHH05] Myeongju Chae, Seung-Yeal Ha, and Hyungjin Huh. Remarks on the repulsive Wigner–Poisson system. Journal of Mathematical Physics, 46(11):112103, November 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/
i11/p112103_s1.

Choppin:1968:BRB

[Cho68] Gregory R. Choppin. Book review: Symmetries and reflections, by Eugene P. Wigner. Journal of Chemical Education, 45(5):352, 1968. CODEN JCEDA8. ISSN 0021-9584 (print),

1938-1328 (electronic). URL http://pubs.acs.org/doi/abs/
10.1021/ed045p352.1.

Cohen:1997:EMQ

[CHS97] Robert S. Cohen, Michael Horne, and John Stachel, editors. Experimental Metaphysics: Quantum Mechanical Studies for Abner Shimony. Volume One, volume 193 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-4452-9 (volume 1), 0-7923-4454-5 (set). ISSN 0068-0346. LCCN

Q174 .B67 vol. 193. URL http://www.lehmanns.ch/shop/
naturwissenschaften/422840-9780792344520-experimental-
metaphysics;   http://www.loc.gov/catdir/enhancements/
fy0822/97002399-d.html;   http://www.loc.gov/catdir/enhancements/
fy0822/97002399-t.html.

Chmielinski:2008:PWE

[CIMS08] Jacek Chmieliński, Dijana Ilisević, Mohammad Sal Moslehian, and Ghadir Sadeghi. Perturbation of the Wigner equation in inner product C^{∗}-modules. Journal of Mathematical Physics, 49(3): 033519, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i3/p033519_s1.

Creutz:1955:EGR

[CJSW55a] E. Creutz, H. Jupnik, T. Snyder, and E. P. Wigner. Effect of geometry on resonance absorption of neutrons by uranium. Journal of Applied Physics, 26(3):271–275, March 1955. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-

8850. URL http://link.aip.org/link/japiau/v26/i3/p271/
s1.

Creutz:1955:RMR

[CJSW55b] E. Creutz, H. Jupnik, T. Snyder, and E. P. Wigner. Review of the measurements of the resonance absorption of neutrons by uranium in bulk. Journal of Applied Physics, 26:257–259, 1955. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-

8850. URL http://link.aip.org/link/japiau/v26/i3/p257/
s1.

Creutz:1955:ETT

[CJW55] E. Creutz, H. Jupnik, and E. P. Wigner. Effect of temperature on total resonance absorption of neutrons by spheres of uranium oxide. Journal of Applied Physics, 26(3):276–279, March 1955. CO-DEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic),

1520-8850. URL http://link.aip.org/link/japiau/v26/i3/
p276/s1.

Clauser:1971:RDW

[Cla71a] John F. Clauser. Reply to Dr. Wigner’s objections. American Journal of Physics, 39(9):1098, September 1971. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://ajp.aapt.org/resource/1/ajpias/v39/i9/p1098_s1.

See [Wig71j, Cla71b].

Clauser:1971:NIH

[Cla71b] John F. Clauser. Von Neumann’s informal hidden-variable argument. American Journal of Physics, 39(9):1095–1096, September 1971. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v39/

i9/p1095_s1. See [Wig71j, Cla71a].

Chacon:1975:ECW

[CLM75] E. Chacón, D. Levi, and M. Moshinsky. Equivalence of a class of Wigner coefficients of SU(1, 1) with those of SU(2). Journal of Mathematical Physics, 16(9):1876–1879, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i9/p1876_s1.

Chacon:1976:LAS

[CLM76] E. Chacón, D. Levi, and M. Moshinsky. Lie algebras in the Schrödinger picture and radial matrix elements. Journal of Mathematical Physics, 17(10):1919–1929, October 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i10/p1919_s1.

Caswell:1966:FPC

[CM66] Randall Smith Caswell and Leonard C. Maximon. Fortran programs for the calculation of Wigner 3j, 6j, and 9j coefficients for angular momenta ≤ 80. Technical note 409, United States Government Printing Office, Washington, DC, USA, 1966. 65 pp.

Chadan:1968:DNS

[CM68] K. Chadan and A. Montes. Derivation of nonrelativistic sum rules from the causality condition of Wigner and Van Kampen. Journal of Mathematical Physics, 9(11):1898–1914, November 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/
i11/p1898_s1.

Claasen:1980:WDT

[CM80] T. Claasen and W. Mecklenbrauker. The Wigner distribution— a tool for time-frequency signal analysis, I–III. Phillips Journal of Research, 35:217–250, 276–300, 372–389, 1980. CODEN ???? ISSN ????

Calinon:1977:SPE

[CMS77] R. Calinon, D. Merlini, and R. R. Sari. On the h-stability property and energy of Wigner lattices in the one component classical plasma: a numerical analysis. Computer Physics Communications, 13(1):1–8, May 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465577900200.

Coben:1972:SET

[Cob72] Stanley Coben. The scientific establishment and the transmission of quantum mechanics to the United States, 1919–32. The American Historical Review, 76(2):442–466, April 1972. ISSN 0002-8762 (print), 1937-5239 (electronic). URL http://www.jstor.org/

stable/1858707.

Cohen:1959:BRA

[Coh59] E. Richard Cohen. Book review: Alvin M. Weinberg and Eugene P. Wigner, The Physical Theory of Neutron Chain Reactors. Physics Today, 12(3):34, March 1959. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v12/i3/p34/s1.

Colin:1994:SED

[Col94] Thierry Colin. Smoothing effects for dispersive equations via a generalized Wigner transform. SIAM Journal on Mathematical Analysis, 25(6):1622–1641, November 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/24322.

Compton:1956:AQP

[Com56] Arthur Holly Compton. Atomic quest, a personal narrative. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1956. xix + 370 pp. LCCN QC773.A1 C65.

Conatser:1972:CLD

[Con72] C. W. Conatser. Contractions of the low-dimensional real Lie algebras. Journal of Mathematical Physics, 13(2):196–203, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v13/i2/p196_s1.

Creutz:1959:NR

[COW^{+}59] Edward C. Creutz, Leo A. Ohlinger, Alvin M. Weinberg, Eugene P. Wigner, and Gale J. Young. Neutronic reactor. US Patent 2,910,418., October 27, 1959. URL http://www.google. com/patents/US2910418. US Patent Application filed Jan 23, 1945.

Creutz:1961:NR

[COW^{+}61] Edward C. Creutz, Leo A. Ohlinger, Alvin M. Weinberg, Eugene P. Wigner, and Gale J. Young. Nuclear reactors. Canadian Patent 619064., April 25, 1961. URL http://brevets-patents.

ic.gc.ca/opic-cipo/cpd/eng/patent/619064/summary.html.

Creutz:1957:MSN

[COWY57] Edward C. Creutz, Leo A. Ohlinger, Eugene P. Wigner, and Gale J. Young. Means for sustaining a nuclear fission. US Patent 2,796,398., June 18, 1957. URL http://www.google.com/ patents/US2796398. US Patent Application filed Oct 8, 1946.

Coxeter:1961:WPR

[Cox61] H. S. M. Coxeter. On Wigner’s problem of reflected light signals in de Sitter space. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 261(1307):435–442, May 23, 1961. CODEN PRLAAZ. ISSN 0080-4630. URL http:

//www.jstor.org/stable/2414187.

Ciccotti:1999:WAS

[CPCF99] Giovanni Ciccotti, Carlo Pierleoni, Fabrizio Capuani, and Vladimir S. Filinov. Wigner approach to the semiclassical dynamics of a quantum many-body system: the dynamic scattering function of ^{4}He. Computer Physics Communications, 121–122:452–459, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465599003811.

Critchfield:1967:BWE

[Cri67] Charles L. Critchfield. Bargmann–Wigner equations. Journal of Mathematical Physics, 8(5):1037–1039, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i5/
p1037_s1.

Coryell:1951:RSF

[CS51] Charles D. (Charles Du Bois) Coryell and Nathan Sugarman, editors. Radiochemical studies: the fission products, volume 9 of National nuclear energy series. Manhattan Project technical section. Division IV: Plutonium Project record. McGraw-Hill, New York, NY, USA, 1951. xxx + 2086 (three volumes) pp. LCCN QD601 .C65. With special editorial assistance from R.A. Brightsen and others.

Cuthill:1963:BRB

[Cut63] Elizabeth Cuthill. Book review: Proceedings of Symposia in Applied Mathematics, vol. XI, Nuclear Reactor Theory, by G. Birkhoff and E. P. Wigner. Mathematics of Computation, 17(84): 475–476, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.jstor.org/

stable/2004032.

Curtright:2001:GAW

[CUZ01] Thomas Curtright, Tsuneo Uematsu, and Cosmas Zachos. Generating all Wigner functions. Journal of Mathematical Physics, 42(6):2396–2415, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chow:1980:SPT

[CV80] H. C. Chow and S. H. Vosko. Special points for two-dimensional Brillouin zone or Wigner–Seitz cell integrations. Canadian Journal of Physics = Journal canadien de physique, 58(4):497–503, 1980. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045

(electronic). URL http://www.nrcresearchpress.com/doi/
abs/10.1139/p80-070.

Charbonneau:1982:LRT

[CvV82] M. Charbonneau, K. M. van Vliet, and P. Vasilopoulos. Linear response theory revisited. III. one-body response formulas and generalized Boltzmann equations. Journal of Mathematical Physics, 23(2):318–336, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v23/i2/p318_s1.

Critchfield:1941:AIB

[CW41] Charles L. Critchfield and Eugene P. Wigner. The antisymmetrical interaction in beta-decay theory. Physical Review (2), 60 (5):412–413, September 1, 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.60.412.

Creutz:1958:SUC

[CW58] Edward C. Creutz and Eugene P. Wigner. Separating uranium containing solids. US Patent 2,833,618., May 6, 1958. URL http:

//www.google.com/patents/US2833618. US Patent Application

filed May 28, 1945.

Creutz:1961:SUC

[CW61] Edward C. Creutz and Eugene P. Wigner. Separation of uranium containing solids. Canadian Patent 616181., March 14,

1961. URL http://brevets-patents.ic.gc.ca/opic-cipo/
cpd/eng/patent/616181/summary.html.

Cleary:1971:SGT

[CW71] J. G. Cleary and B. G. Wybourne. Statistical group theory and the distribution of angular momentum states. Journal of Mathematical Physics, 12(1):45–52, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i1/p45_s1.

Chester:1974:PVN

[CW74] C. V. Chester and E. P. Wigner. Population vulnerability: The neglected issue in arms limitation and the strategic balance. Orbis, 18(??):763–769, ???? 1974. CODEN ???? ISSN ????

Draayer:1973:WRC

[DA73] J. P. Draayer and Yoshimi Akiyama. Wigner and Racah coefficients for SU_{3}. Journal of Mathematical Physics, 14(12):1904– 1912, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i12/p1904_s1.

Dahl:1995:ODA

[Dah95] Jens Peder Dahl. On the oscillatory decay of atomic Wigner functions. Journal of Physical Chemistry, 99(9):2727–2731, 1995. CO-DEN JPCHAX. ISSN 0022-3654 (print), 1541-5740 (electronic).

URL http://pubs.acs.org/doi/abs/10.1021/j100009a032.

Dannen:1995:PPU

[Dan95] Gene Dannen. A petition to the President of the United States: July 17, 1945. Web site., 1995. Contains text of the petition signed by Leo Szilard, Eugene Wigner, and 68 others to avoid use of the atomic bomb on Japan.

Darrow:1947:PSH

[Dar47] K. K. Darrow, editor. Physical science and human values, a symposium. Princeton University Press, Princeton, NJ, USA, 1947. LCCN Q171 .P947 1946a. Foreword by E. P. Wigner.

Davies:1974:DMW

[Dav74] E. B. Davies. Dynamics of a multilevel Wigner–Weisskopf atom. Journal of Mathematical Physics, 15(12):2036–2041, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v15/i12/p2036_s1.

Deshpande:2008:ODW

[DB08] Vikram V. Deshpande and Marc Bockrath. The one-dimensional Wigner crystal in carbon nanotubes. Nature Physics, 4(4):314– 318, March 9, 2008. CODEN NPAHAX. ISSN 1745-2473 (print),

1745-2481 (electronic). URL http://www.nature.com/nphys/
journal/v4/n4/full/nphys895.html.

deBroglie:1984:QST

[dBDW^{+}84] Louis de Broglie, P. A. M. (Paul Adrien Maurice) Dirac, Eugene Paul Wigner, A. O. (Asim Orhan) Barut, Alwyn Van der Merwe, and Jean-Pierre Vigier, editors. Quantum, space, and time — the quest continues: studies and essays in honour of Louis de Broglie, Paul Dirac, and Eugene Wigner. Cambridge monographs on physics. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-31911-0 (paperback). vii + 659 pp. LCCN QC174.125 .Q376 1984.

doAmaral:1983:HFB

[dDG83] A. F. Furtado do Amaral, F. A. Doria, and M. Gleiser. Higgs fields as Bargmann–Wigner fields and classical symmetry breaking. Journal of Mathematical Physics, 24(7):1888–1890, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v24/i7/p1888_s1.

DeBenedetti:1958:BRB

[De 58] S. De Benedetti. Book review: Nuclear Structure, by Leonard Eisenbud and Eugene P. Wigner. American Scientist, 46(4): 340A, ???? 1958. CODEN AMSCAC. ISSN 0003-0996 (print),

1545-2786 (electronic). URL http://www.jstor.org/stable/
27827220.

dEspagnat:1971:FQM

[d’E71] Bernard d’Espagnat, editor. Foundations of quantum mechanics: Proceedings of the International School of Physics ‘Enrico Fermi’, course 49, Varenna on Lake Como, 29 June–11 July 1970. Its Proceedings, course 49. Academic Press, New York, USA, 1971. ISBN 0-12-368849-3. xiv + 480 pp. LCCN QC174.1 .V36 1971.

Demortier:2015:QEQ

[Dem15] Guy Demortier. Qui est qui dans ^{≪}the letter to Roosevelt ^{≫}?. (French) [Who is who in “the letter to Roosevelt”?]. Revue des Questions Scientifiques, 186(1):125–152, ???? 2015. CO-DEN RQSCAN. ISSN 0035-2160. URL https://www.unamur.

be/sciences/philosoc/revueqs/textes-en-ligne/rqs_186_
1_demortier.

Dollard:1977:PIS

[DF77] John D. Dollard and Charles N. Friedman. Product integrals and the Schrödinger equation. Journal of Mathematical Physics, 18 (8):1598–1607, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v18/i8/p1598_s1.

DeBrota:2020:ROF

[DFS20] John B. DeBrota, Christopher A. Fuchs, and Rüdiger Schack. Respecting one’s fellow: QBism’s analysis of Wigner’s friend. Foundations of Physics, 50(12):1859–1874, December 2020. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL https://link.springer.com/article/10.1007/s10701-
020-00369-x.

Dopel:1934:EPS

[DGW34] R. Döpel, K. Gailer, and E. P. Wigner. Über die experimentelle Prüfung des Spinerhaltungssatzes. (German) [On the experimental verification of the spin conservation law]. Physikalische Zeitschrift, 35(8):336–337, ???? 1934. CODEN PHZTAO. ISSN 0369-982X.

Denardo:1984:GTM

[DGW84] G. Denardo, G. Ghirardi, and T. Weber, editors. Group theoretical methods in physics: proceedings of the XIIth International Colloquium held at the International Centre for Theoretical Physics, Trieste, Italy, September 5–11, 1983, volume 201 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 3-540-13335-6, 0-387-13335-6. LCCN ????

Daniels:1949:OER

[DHW49] Farrington Daniels, M. King Hubbert, and Eugene P. Wigner. Our energy resources. Physics Today, 2(4):18–??, April 1949. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v2/i4/p18/s1.

Dieks:2021:WF

[Die21] Dennis Dieks. Wigner’s friend. In Calosi et al. [CGPT21], pages 41–54. ISBN 88-351-2139-6. LCCN Q175 .E95 2021.

Dirac:1927:QTE

[Dir27] P. A. M. Dirac. The quantum theory of the emission and absorption of radiation. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 114(767):243–265, March 1, 1927. CODEN PRLAAZ. ISSN 0080-4630. URL http:

//www.jstor.org/stable/94746. Russian translation in [?, Vol.

II, pp. 285ff].

Dirac:1945:URL

[Dir45] P. A. M. Dirac. Unitary representations of the Lorentz group. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 183(994):284–295, February 22, 1945. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/97721.

Davidson:1970:RQSa

[DK70a] Russell Davidson and John J. Kozak. On the relaxation to quantum-statistical equilibrium of the Wigner–Wiesskopf atom in a one-dimensional radiation field. I. A study of spontaneous emission. Journal of Mathematical Physics, 11(1):189–202, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v11/i1/p189_s1.

Davidson:1970:RQSb

[DK70b] Russell Davidson and John J. Kozak. Relaxation to quantum statistical equilibrium of the Wigner–Weisskopf atom in a onedimensional radiation field. II. Finite systems. Journal of Mathematical Physics, 11(4):1420–1436, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i4/p1420_s1.

Davidson:1971:RQS

[DK71] Russell Davidson and John J. Kozak. Relaxation to quantumstatistical equilibrium of the Wigner–Weisskopf atom in a onedimensional radiation field. III. The quantum-mechanical solution. Journal of Mathematical Physics, 12(6):903–917, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v12/i6/p903_s1.

Davidson:1973:RQSa

[DK73a] Russell Davidson and John J. Kozak. On the relaxation to quantum-statistical equilibrium of the Wigner–Weisskopf atom in a one-dimensional radiation field. IV. Exact solution for finite systems. Journal of Mathematical Physics, 14(3):414–421, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v14/i3/p414_s1.

Davidson:1973:RQSb

[DK73b] Russell Davidson and John J. Kozak. On the relaxation to quantum-statistical equilibrium of the Wigner–Weisskopf atom in a one-dimensional radiation-field. V. Exact solution for infinite systems. Journal of Mathematical Physics, 14(4):423–431, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v14/i4/p423_s1.

Davidson:1975:RQS

[DK75] Russell Davidson and John J. Kozak. On the relaxation to quantum-statistical equilibrium of the Wigner–Weisskopf atom in a one-dimensional radiation field. VII. Emission in a finite system in the presence of an extra photon. Journal of Mathematical Physics, 16(5):1013–1022, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v16/i5/p1013_s1.

Davidson:1978:RQS

[DK78] Russell Davidson and John J. Kozak. On the relaxation to quantum statistical equilibrium of the Wigner–Weisskopf atom in a one-dimensional radiation field. VIII. Emission in an infinite system in the presence of an extra photon. Journal of Mathematical Physics, 19(5):1074–1086, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v19/i5/p1074_s1.

Datta:2004:RMA

[DK04] Nilanjana Datta and Hervé Kunz. A random matrix approach to the crossover of energy-level statistics from Wigner to Poisson. Journal of Mathematical Physics, 45(3):870–886, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v45/i3/p870_s1.

Dytrych:2021:SCL

[DLD^{+}21] Tomás Dytrych, Daniel Langr, Jerry P. Draayer, Kristina D. Launey, and Daniel Gazda. SU3lib: a C++ library for accurate computation of Wigner and Racah coefficients of SU(3). Computer Physics Communications, 269(??):Article 108137, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002496.

delaPena:1979:QHO

[dlPC79] L. de la Peña and A. M. Cetto. The quantum harmonic oscillator revisited: a new look from stochastic electrodynamics. Journal of Mathematical Physics, 20(3):469–483, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i3/
p469_s1.

Daumens:1993:SRR

[DMMT93] M. Daumens, P. Minnaert, M. Mozrzymas, and S. Toshev. The super-rotation Racah–Wigner calculus revisited. Journal of Mathematical Physics, 34(6):2475–2507, June 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i6/
p2475_s1.

Doncel:1984:IEP

[DMS84] M. G. Doncel, L. L. Michel, and J. Six. Interview de Eugene P. Wigner sur sa vie scientifique. (French) [Interview with Eugene P. Wigner on his scientific life]. Archives internationale d’historie des sciences, 34(112):177–217, ???? 1984. CODEN ???? ISSN 0003-9810.

Dias:2001:GWW

[DP01] Nuno Costa Dias and João Nuno Prata. Generalized Weyl– Wigner map and Vey quantum mechanics. Journal of Mathematical Physics, 42(12):5565–5579, December 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dias:2002:WFB

[DP02] Nuno Costa Dias and João Nuno Prata. Wigner functions with boundaries. Journal of Mathematical Physics, 43(10):4602–4627, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dray:1986:UTW

[Dra86] Tevian Dray. A unified treatment of Wigner D functions, spinweighted spherical harmonics, and monopole harmonics. Journal of Mathematical Physics, 27(3):781–792, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i3/
p781_s1.

daSilva:2005:BCN

[dSFA05] Antônio J. R. da Silva, A. Fazzio, and Alex Antonelli. Bundling up carbon nanotubes through Wigner defects. Nano Letters, 5 (6):1045–1049, 2005. CODEN NALEFD. ISSN ???? URL http:

//pubs.acs.org/doi/abs/10.1021/nl050457c.

Duffus:1946:ABV

[Duf46] R. L. Duffus. Atom bomb versus ‘human race’: The scientists who fashioned it remind us of its terrible threat: Review of One World or None. A Report to the Public on the Full Meaning of the Atomic Bomb. Edited by Dexer Masters and Kathlerine Way. Foreword by Niels Bohr. Introduction by Arthur H. Compton. x + 79 pp. New York: Whittlesey House. $1. New York Times, ??(??): BR1–BR2, March 17, 1946. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/docview/107455627.

Dufty:1971:GME

[Duf71] James W. Dufty. Generalized master equations for inhomogeneous systems. Journal of Mathematical Physics, 12(2):272–280, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v12/i2/p272_s1.

Dereli:1997:CWW

[DV97] T. Dereli and A. Vercin. W_{∞}-covariance of the Weyl–Wigner– Groenewold–Moyal quantization. Journal of Mathematical Physics, 38(11):5515–5530, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Duke:1964:AMC

[DW64] C. B. Duke and E. P. Wigner. Approximation method in collision theory based on R-matrix theory. Reviews of Modern Physics, 36(2):584–589, April 1, 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.36.584;   http://
rmp.aps.org/abstract/RMP/v36/i2/p584_1.

Earp:1981:WSP

[Ear81] C. Earp. Wigner 9j symbols and the product group. Journal of Mathematical Physics, 22(11):2357–2359, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2357_s1.

Eckart:1933:CNT

[Eck33] Carl Eckart. A comparison of the nuclear theories of Heisenberg and Wigner. I. Physical Review (2), 44(2):109, July 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v44/i2/p109_1.

Eckart:1973:CEP

[Eck73] Carl Eckart. Carl Eckart papers, 1921–1973 (bulk 1935–1970). US Library of Congress archival manuscript material (collection).,

1973. URL http://hdl.loc.gov/loc.mss/eadmss.ms001014.

Edwards:1960:BRB

[Edw60] S. F. Edwards. Book review: Nuclear structure, by Leonard Eisenbud and Eugene P. Wigner. 128 pages, 6 × 9.25 in. Princeton, Princeton University Press, 1958. Price $4.00. Journal of Nuclear Energy. Part A. Reactor Science, 12(4):179–180, August 1960. CODEN JNEAA8. ISSN 0368-3265 (print), 1878-2116

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0368326560900979.

Eisenbud:1967:GPN

[EGW67] L. Eisenbud, G. T. Garvey, and E. P. Wigner. General principles of nuclear structure. In Edward Uhler Condon and Hugh Odishaw, editors, Handbook of physics, chapter 1 (part 9), pages 9–54. Mc-Graw-Hill, New York, NY, USA, second edition, 1967. ISBN 0-07-012403-5. LCCN ????

Eisenbud:1969:ASH

[EGW69] L. Eisenbud, G. T. Garvey, and E. P. Wigner. Az atommag szerkezete. (Hungarian) [The structure of the nucleus]. Akad. K., Budapest, Hungary, 1969. 144 pp. LCCN ????

Einstein:1939:AEL

[Ein39] Albert Einstein. Albert Einstein’s letter to President Franklin Delano Roosevelt. World-Wide Web document, August 2, 1939.

URL http://hypertextbook.com/eworld/einstein.shtml;
http://www.anl.gov/Science_and_Technology/History/Anniversary_
Frontiers/aetofdr.html;   http://www.lanl.gov/history/
road/pdf/Einstein.pdf.

Ellzey:1992:NIT

[Ell92] M. L. Ellzey, Jr. Normalized irreducible tensorial matrices and the Wigner–Eckart theorem for unitary groups: a superposition Hamiltonian constructed from octahedral NITM. International Journal of Quantum Chemistry, 41(5):653–665, March 5, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Elvee:1982:MN

[Elv82] Richard Q. (Richard Quentin) Elvee, editor. Mind in nature. Harper and Row, San Francisco, CA, USA, 1982. ISBN 0-06-250285-9. xvi + 157 pp. LCCN BD161 .N6 1981.

Eisenbud:1941:IFI

[EW41] L. Eisenbud and E. P. Wigner. Invariant forms of interaction between nuclear particles. Proceedings of the National Academy of Sciences of the United States of America, 27(6):281–289, June 15, 1941. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490

(electronic). URL http://www.jstor.org/stable/87333.

Eisenbud:1958:NS

[EW58] Leonard Eisenbud and Eugene Paul Wigner. Nuclear structure, volume 8 of Investigations in physics. Princeton University Press, Princeton, NJ, USA, 1958. viii + 128 pp. LCCN QC173 .E36.

Eisenbud:1961:EKG

[EW61] Leonard Eisenbud and Eugene Paul Wigner. Einführung in die Kernphysik. (German) [Introduction to Nuclear Physics], volume 16 of B. I.-Hochschultaschenbücher. Bibliographisches Institut, Mannheim, Germany, 1961. vi + 145 pp. LCCN QC173 .E365.

Edsall:1980:UC

[EW80] John T. Edsall and E. P. Wigner. Unification church. Physics Today, 33(10):15–16, October 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v33/i10/p15/s1.

Flato:1976:QMD

[F^{+}76] M. (Moshé) Flato et al., editors. Quantum mechanics, determinism, causality, and particles: an international collection of contributions in honor of Louis de Broglie on the occasion of the jubilee of his celebrated thesis, volume 1 of Mathematical physics and applied mathematics. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1976. ISBN 90-277-0623-9. x + 252 pp. LCCN QC174.125 .Q36.

Fano:1960:BRP

[Fan60] U. Fano. Book review: E. P. Wigner, Group Theory and Its Application to the Quantum Mechanics of Atomic Spectra. American Journal of Physics, 28(4):408, April 1960. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

link.aip.org/link/ajpias/v28/i4/p408/s2.

Fawcett:1988:SOU

[FB88] R. J. B. Fawcett and A. J. Bracken. Simple orthogonal and unitary compact quantum systems and the İnönü-Wigner contraction. Journal of Mathematical Physics, 29(7):1521–1528, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v29/i7/p1521_s1.

Fermi:1941:NP

[FBR^{+}41] Enrico Fermi, Gregory Breit, I. I. Rabi, Eugene P. Wigner, and John H. van Vleck, editors. Nuclear physics. University Press of Pennsylvania, Philadelphia, PA, USA, 1941. v + 68 pp.

Feld:1967:DDE

[Fel67] B. T. Feld. The decision to deploy: An editorial opinion. Bulletin of the Atomic Scientists, 23(10):18–19, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [WF68, BJH68].

Fermi:1932:QTR

[Fer32] Enrico Fermi. Quantum theory of radiation. Reviews of Modern Physics, 4(1):87–132, January 1, 1932. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
4.87;   http://rmp.aps.org/abstract/RMP/v4/i1/p87_1.

Fermi:1968:III

[Fer68] Laura Fermi. Illustrious immigrants; the intellectual migration from Europe, 1930/41. University of Chicago Press, Chicago, IL, USA and London, UK, 1968. xi + 440 pp. LCCN E184.A1 F47.

Fermi:1971:III

[Fer71] Laura Fermi. Illustrious immigrants; the intellectual migration from Europe, 1930/41. University of Chicago Press, Chicago, IL, USA and London, UK, second edition, 1971. ISBN 0-226-24376-1, 0-226-24378-8 (paperback). xi + 431 pp. LCCN E184.A1 F47 1971.

Feshbach:1960:BRE

[Fes60] Herman Feshbach. Book review: Eugene P. Wigner, Group Theory. Physics Today, 13(1):62, January 1960. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/phtoad/v13/i1/p62/s1.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Finney:1960:TSW

[Fin60] John W. Finney. Two scientists wary on policing of atomic-test ban: Disclose concern at capital ceremony honoring four. New York Times, ??(??):3–??, May 19, 1960. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/docview/114981445. Report on the Atoms for

Peace Awards with shared $150,000 prize presented in Washington, DC, on 18 May 1960 to Leo Szilard, Alving M. Weinberg, Eugene Wigner, and Walter H. Zinn.

Fyodorov:1997:AHR

[FKS97] Yan V. Fyodorov, Boris A. Khoruzhenko, and Hans-Jürgen Sommers. Almost Hermitian random matrices: Crossover from Wigner–Dyson to Ginibre eigenvalue statistics. Physical Review Letters, 79(??):557–??, July 28, 1997. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http:

//link.aps.org/doi/10.1103/PhysRevLett.79.557.

Flatte:1968:BBR

[Fla68] Stanley M. Flatté. Books: Symmetries and Reflections, by Eugene P. Wigner; The Relevance of Physics, by Stanley L. Jaki. Bulletin of the Atomic Scientists, 24(3):28–30, March 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fleming:2000:RSM

[Fle00] Gordon N. Fleming. Reeh–Schlieder meets Newton–Wigner. Philosophy of Science, 67(3S):S495–S515, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188690. See comments [Hal01].

Flugge:1946:DBW

[Flü46] Siegfried Flügge. On the derivation of the Breit–Wigner formula. Zeitschrift für Naturforschung, 1(??):121–??, ???? 1946. CODEN ZNTFA2. ISSN 0372-9516.

Filippas:2003:SWF

[FM03] Stathis Filippas and George N. Makrakis. Semiclassical Wigner function and geometrical optics. Multiscale Modeling & Simulation, 1(4):674–710, 2003. CODEN MMSUBT. ISSN 1540-3459 (print), 1540-3467 (electronic). URL http://epubs.siam.org/

sam-bin/dbq/article/40979.

Foldy:1978:ESW

[Fol78] Leslie L. Foldy. Electrostatic stability of Wigner and Wigner– Dyson lattices. Physical Review B: Solid State, 17(??):4889–??, June 15, 1978. CODEN PLRBAQ. ISSN 0556-2805. URL http:

//link.aps.org/doi/10.1103/PhysRevB.17.4889.

Foldy:1980:NIW

[Fol80] L. L. Foldy. New instabilities in Wigner–Dyson lattices. Physical Review B: Condensed Matter and Materials Physics, 22(??):4992– ??, November 15, 1980. CODEN PRBMDO. ISSN 1098-0121.

URL http://link.aps.org/doi/10.1103/PhysRevB.22.4992.

Fonarev:1994:WFQ

[Fon94] Oleg A. Fonarev. Wigner function and quantum kinetic theory in curved space–time and external fields. Journal of Mathematical Physics, 35(5):2105–2129, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v35/i5/p2105_s1.

Forrest:1997:CSW

[For97] Peter Forrest. Common sense and a “Wigner–Dirac” approach to quantum mechanics. The Monist, 80(1):131–159, ???? 1997. CODEN ???? ISSN 0026-9662 (print), 2153-3601 (electronic).

URL http://www.jstor.org/stable/27903515.

Fannes:2003:PWM

[FP03] Mark Fannes and Dénes Petz. Perturbation of Wigner matrices and a conjecture. Proceedings of the American Mathematical Society, 131(7):1981–1988, July 2003. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http://

www.jstor.org/stable/1194011.

Fox:1966:ISN

[FR66] John D. (John David) Fox and Donald Robson, editors. Isobaric spin in nuclear physics; proceedings [17–19 March 1966: Tallahassee, FL, USA]. Academic Press, New York, USA, 1966. LCCN QC721 .C68 1966.

Fujiie:2003:BWF

[FR03] Setsuro Fujiié and Thierry Ramond. Breit–Wigner formula at barrier tops. Journal of Mathematical Physics, 44(5):1971–1983, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:2013:DPN

[FR13a] Bernard Fernandez and Georges Ripka. Do the protons and neutrons form shells as electrons do in the atom? In Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896– 1956 [FR13e], pages 271–278. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). LCCN QC773 .F47 2013.

Fernandez:2013:EPW

[FR13b] Bernard Fernandez and Georges Ripka. Eugene P. Wigner. In Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956 [FR13e], page 270. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). LCCN QC773 .F47 2013.

Fernandez:2013:PTY

[FR13c] Bernard Fernandez and Georges Ripka. A proliferation of theories: Yukawa, Breit and Wigner, Bohr. In Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956 [FR13e], page 331. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). LCCN QC773 .F47 2013.

Fernandez:2013:SSW

[FR13d] Bernard Fernandez and Georges Ripka. The symmetries and supermultiplets of Wigner and Feenberg. In Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956 [FR13e], page 434. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). LCCN QC773 .F47 2013.

Fernandez:2013:UMA

[FR13e] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). xviii + 522 pp. LCCN QC773 .F47 2013.

Frank:2005:ERG

[Fra05] Tibor Frank. Ever ready to go: The multiple exiles of Leo Szilard. Physics in Perspective (PIP), 7(2):204–252, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://adsabs.harvard.edu/abs/2005PhP...
..7..204F;   http://www.springerlink.com/content/1422-
6960/.

French:1979:ECVa

[Fre79] A. P. (Anthony Philip) French, editor. Einstein: a centenary volume. Heinemann for the International Commission of Physics Education, London, UK, 1979. ISBN 0-435-58200-3. xx + 332 pp. LCCN QC16.E5 E37 1979b.

Fredrickson:1993:NWS

[Fre93] Glenn H. Fredrickson. Noncircular Wigner–Seitz cells in strongly segregated block copolymers. Macromolecules (Washington, DC, USA), 26(16):4351–4355, 1993. CODEN MAMOBX. ISSN 0024-9297 (print), 1520-5835 (electronic). URL http://pubs.acs.

org/doi/abs/10.1021/ma00068a043.

FreireJunior:2015:QDR

[Fre15] Olival Freire Junior. The Quantum Dissidents: Rebuilding the Foundations of Quantum Mechanics (1950–1990). Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-662-44661-8, 3-662-44662-6 (e-book). xvi + 356 pp. LCCN QC173.98 .F74 2015.

Fronsdal:1979:IPQ

[Fro79] Christian Fronsdal. Invariant ^{∗}-product quantization of the onedimensional Kepler problem. Journal of Mathematical Physics, 20(11):2226–2232, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v20/i11/p2226_s1.

Frohlich:1993:HHS

[Frö93] Fröhlich Pál. Három hazai szakvélemény a fiatal wignerről és neumannról. (Hungarian) [Three young domestic opinions (??) and Wigner and von Neumann]. Fizikai Szemle (Budapest), 43 (5):197–201, 1993. CODEN FISZA6. ISSN ????

Fernandez:1982:LIW

[FS82] Ariel Fernández and Oktay Sinanoğlu. The lifting of an Inönü– Wigner contraction at the level of universal coverings. Journal of Mathematical Physics, 23(12):2234–2235, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2234_s1.

Farkas:1936:CRE

[FW36] L. Farkas and E. Wigner. Calculation of the rate of elementary reactions of light and heavy hydrogen. Transactions of the Faraday Society, 32:708–723, 1936. CODEN TFSOA4. ISSN 0014-7672.

Feenberg:1937:SNB

[FW37] E. Feenberg and E. Wigner. On the structure of the nuclei between helium and oxygen. Physical Review (2), 51(2):95–106, January 15, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.51.95.

Farkas:1952:FMV

[FW52] Adalbert Farkas and Eugene Paul Wigner, editors. L. Farkas memorial volume, volume 1 of Research Council of Israel. Special publication. Research Council of Israel, Jerusalem, Israel, 1952. 309 pp. LCCN QD455 .F3.

Freedman:1973:BMB

[FW73] S. Freedman and E. Wigner. On Bub’s misunderstanding of Bell’s locality argument. Foundations of Physics, 3(4):457–458, December 1973. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://www.springerlink.com/
content/r17r1242148r3870.

Fermi:1945:BPD

[FWN^{+}45] Enrico Fermi, Eugene P. Wigner, L. W. Nordheim, Alvin M. Weinberg, H. Soodak, H. S. Brown, Miles C. Leverett, F. Daniels, G. Young, and Glenn T. Seaborg. Breeder pile discussion. Report, US Atomic Energy Commission, Washington, DC, USA, June 19–20, 1945.

Gardiner:1988:QNQ

[Gar88] C. W. Gardiner. Quantum noise and quantum Langevin equations. IBM Journal of Research and Development, 32(1):127–136, January 1988. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Gautschi:1994:MCH

[Gau94] Walter Gautschi, editor. Mathematics of computation, 1943– 1993: a half-century of computational mathematics: Mathematics of Computation 50th Anniversary Symposium, August 9–13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Gill:2004:WIK

[GBO04] Peter M. W. Gill, Nicholas A. Besley, and Darragh P. O’Neill. Wigner intracule for the Kellner helium-like ions. International Journal of Quantum Chemistry, 100(2):166–171, ???? 2004. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gelfert:2014:AIU

[Gel14] Axel Gelfert. Applicability, indispensability, and underdetermination: Puzzling over Wigner’s ‘unreasonable effectiveness of mathematics’. Science & Education (Springer), 23(5):997–1009, May 2014. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Gerrish:1975:LTM

[Ger75] Laina Gerrish. Letters: Too much. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

GarciaDoncel:1987:SPP

[GHM^{+}87] Manuel García Doncel, Armin Hermann, Louis Michel, et al., editors. Symmetries in physics (1600–1980): proceedings of the 1st International Meeting on the History of Scientific Ideas held at Sant Feliu de Guíxols, Catalonia, Spain, September 20–26, 1983. Seminari d’História de les Ciéncies, Universitat Autónoma de Barcelona, Bellaterra, Barcelona, Spain, 1987. ISBN 84-7488-148-8. LCCN QC174.17.S9 I57 1983.

Gibilisco:2003:WYI

[GI03] Paolo Gibilisco and Tommaso Isola. Wigner–Yanase information on quantum state space: The geometric approach. Journal of Mathematical Physics, 44(9):3752–3762, September 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gibson:2019:SIH

[Gib19] Susannah Gibson. The Spirit of Inquiry: How One Extraordinary Society Shaped Modern Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2019. ISBN 0-19-883337-7 (hardcover). xxi + 377 pp. LCCN Q41.C194 G537 2019.

Ginibre:1963:WET

[Gin63] Jean Ginibre. Wigner–Eckart theorem and simple Lie groups. Journal of Mathematical Physics, 4(5):720–726, May 1963. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i5/p720_s1.

Girdner:1975:LNA

[Gir75] William Girdner. Letters: No alternative? Bulletin of the Atomic Scientists, 31(6):4–5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Golub:2023:HPF

[GL23] Robert Golub and Steve Keith Lamoreaux. The Historical and Physical Foundations of Quantum Mechanics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2023. ISBN 0-19-186123-5, 0-19-255536-7, 0-19-882218-9 (hardcover), 0-19-882219-7 (paperback). xiii + 747 pp. LCCN QC174.12 .G65 2023.

Glasstone:1959:BRB

[Gla59] Samuel Glasstone. Book review: The Physical Theory of Neutron Chain Reactors, by Alvin M. Weinberg and Eugene P. Wigner. American Scientist, 47(1):56A, ???? 1959. CODEN AMSCAC.

ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://

www.jstor.org/stable/27827259.

Gould:1992:MEW

[GLB92] M. D. Gould, J. Links, and A. J. Bracken. Matrix elements and Wigner coefficients for U_{q}[gl(n)]. Journal of Mathematical Physics, 33(3):1008–1022, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v33/i3/p1008_s1.

Gruber:1980:SS

[GM80] Bruno Gruber and Richard S. Millman, editors. Symmetries in science. Plenum Press, New York, NY, USA; London, UK, 1980. ISBN 0-306-40541-5. LCCN Q172.5.S95 S92.

GoeppertMayer:1949:CSN

[Goe49] Maria Goeppert Mayer. On closed shells in nuclei. II. Physical Review (2), 75(12):1969–1970, June 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:

//link.aps.org/doi/10.1103/PhysRev.75.1969;   http://
prola.aps.org/abstract/PR/v75/i12/p1969_1;   http://www.
nobelprize.org/nobel_prizes/physics/laureates/1963/.

GoeppertMayer:2001:WEP

[Goe01] Maria Goeppert Mayer. Wigner, Eugene Paul. In Hoiberg [Hoi01], page 464. ISBN 81-88237-00-0. LCCN AS911.N9 A15 2001.

Gottlieb:1971:WCG

[Got71] H. P. W. Gottlieb. Wigner coefficients for the group E(2). Journal of Mathematical Physics, 12(1):39–44, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i1/
p39_s1.

Gould:1981:WCS

[Gou81] M. D. Gould. Wigner coefficients for a semisimple Lie group and the matrix elements of the O(n) generators. Journal of Mathematical Physics, 22(11):2376–2388, November 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i11/p2376_s1.

Gould:1986:MFWa

[Gou86a] M. D. Gould. Multiplicity-free Wigner coefficients for semisimple Lie groups. I. The U(n) pattern calculus. Journal of Mathematical Physics, 27(8):1944–1963, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v27/i8/p1944_s1.

Gould:1986:MFWb

[Gou86b] M. D. Gould. Multiplicity-free Wigner coefficients for semisimple Lie groups. II. A pattern calculus for O(n). Journal of Mathematical Physics, 27(8):1964–1979, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i8/p1964_s1.

Gould:1992:RWC

[Gou92] M. D. Gould. Reduced Wigner coefficients for U_{q}[gl(n)]. Journal of Mathematical Physics, 33(3):1023–1031, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i3/
p1023_s1.

Goudon:2002:ASV

[Gou02] Thierry Goudon. Analysis of a semidiscrete version of the Wigner equation. SIAM Journal on Numerical Analysis, 40(6):2007–2025, December 2002. CODEN SJNAAM. ISSN 0036-1429 (print),

1095-7170 (electronic). URL http://epubs.siam.org/sam-bin/
dbq/article/38836.

Goudon:2003:ASV

[Gou03] Thierry Goudon. Analysis of a semidiscrete version of the Wigner equation. SIAM Journal on Numerical Analysis, 40(6):2007–2025, ???? 2003. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170

(electronic). URL http://www.jstor.org/stable/4100981.

Gould:1986:UGAa

[GP86] M. D. Gould and J. Paldus. Unitary group approach to general system partitioning. I. Calculation of U(n = n_{1} + n_{2}): U(n_{1}) × u(n_{2}) reduced matrix elements and reduced Wigner coefficients. International Journal of Quantum Chemistry, 30(3):327–363, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Grodzins:1963:AAS

[GR63] Morton Grodzins and Eugene I. Rabinowitch, editors. The Atomic Age: Scientists in National and World Affairs. Articles from the Bulletin of the Atomic Scientists 1945–1962. Basic Books, New York, NY, USA, 1963. xviii + 616 pp. LCCN D842 .B78. With the assistance of Harvey Flaumenhaft and Lois Gradner.

Graffi:1974:RPB

[Gra74] S. Graffi. Regular perturbations, Brillouin–Wigner expansion, and continued fractions. Journal of Mathematical Physics, 15(5): 521–523, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i5/p521_s1.

Grene:1969:AKP

[Gre69] Marjorie Grene, editor. The anatomy of knowledge: papers presented to the Study Group on Foundations of Cultural Unity, Bowdoin College 1965 and 1966. Routledge & Kegan Paul, London, UK, 1969. xii + 367 pp. LCCN ????

Greenberger:1986:NTI

[Gre86] Daniel M. Greenberger, editor. New techniques and ideas in quantum measurement theory: [papers presented at a conference entitled … held by the New York Academy of Sciences, on January 21–24, 1986 in New York City]. The New York Academy of Sciences, New York, NY, USA, 1986. ISBN 0-89766-356-X. LCCN ????

Griffin:1991:REA

[GRT91] C. Griffin, P. Rao, and F. Taylor. Roundoff error analysis of the discrete Wigner distribution using fixed-point arithmetic. IEEE Transactions on Signal Processing, 39(9):2096–2098, September 1991. CODEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic).

Gnutzmann:2004:USSb

[GS04a] Sven Gnutzmann and Burkhard Seif. Universal spectral statistics in Wigner–Dyson, chiral, and Andreev star graphs. I. Construction and numerical results. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 69(??):056219, May 25, 2004. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRevE.69.056219.

Gnutzmann:2004:USSa

[GS04b] Sven Gnutzmann and Burkhard Seif. Universal spectral statistics in Wigner–Dyson, chiral, and Andreev star graphs. II. Semiclassical approach. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 69(??):056220, May 25, 2004. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRevE.
69.056220.

Gulbransen:1975:LSE

[Gul75] Earl A. Gulbransen. Letters: Not safe enough. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Guth:1960:BRB

[Gut60] Eugene Guth. Book review: Group Theory. And Its Application to the Quantum Mechanics of Atomic Spectra, by Eugene P. Wigner and J. J. Griffin. American Scientist, 48(2):144A, ???? 1960. CO-DEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

URL http://www.jstor.org/stable/27827524.

Guth:1968:BRE

[Gut68] Eugene Guth. Book review: Eugene Wigner, Symmetries and Reflections, Scientific Essays. Physics Today, 21(5):95, May 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v21/i5/
p95/s1.

Goldstine:1957:SWJ

[GW57] Herman H. Goldstine and Eugene P. Wigner. The scientific work of John von Neumann. Science, 125(3250):683–684, April 12, 1957. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/125/
3250/683.full.pdf.

Greenberg:1963:GTM

[GW63] O. W. Greenberg and E. P. Wigner. Group theoretical methods in elementary particle physics. Physics Today, 16(4):62–65, April 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v16/i4/
p62/s1.

Gailar:1971:SCD

[GW71] J. Gailar and E. Wigner. Is the Soviet civil defense program really better than ours? US Congressional Record, Proc. and Debates of the 92nd Congress, First Session, vol. 117, January 28, 1971.

Goldrich:1972:OCM

[GW72a] F. Goldrich and Eugene P. Wigner. On the observability of the collision matrix. (talk). In J. R. Klauder, editor, Magic Without Magic, pages 147–160. W. H. Freeman and Company, New York, NY, USA, 1972.

Goldrich:1972:CAI

[GW72b] Fredric E. Goldrich and Eugene P. Wigner. Condition that all irreducible representations of a compact Lie group, if restricted to a subgroup, contain no representation more than once. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):432–438, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldstine:2003:NJT

[GW03] Herman H. Goldstine and Wigner Jenő. Neumann János tudományos munkássága. (Hungarian) [Scientific work of John von Neumann]. Természet Világa [Natural World], 134(3):17–19, ???? 2003. CODEN ???? ISSN ????

Hall:1981:EST

[Hal81] George L. Hall. Electrostatic structural transitions in a Gaussian Wigner solid. Journal of Mathematical Physics, 22(12):2996–3001, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i12/p2996_s1.

Hall:1983:EST

[Hal83] George L. Hall. Electrostatic structural transitions in a Yukawa– Wigner solid. Journal of Mathematical Physics, 24(1):209–214, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i1/p209_s1.

Halvorson:2001:RSD

[Hal01] Hans Halvorson. Reeh–Schlieder defeats Newton–Wigner: On alternative localization schemes in relativistic quantum field theory.

Philosophy of Science, 68(1):111–133, March 2001. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3081027. See [Fle00].

Hargittai:1999:YPD

[Har99] I. Hargittai. ‘You are Pleasantly Disagreeable’: Eugene P. Wigner remembers. The Chemical Intelligencer, 5(3):50–52, ???? 1999. CODEN CHEIFK. ISSN 0947-0662. This article is based on a video recording on March 4, 1986, by Clarence Larson.

Hargittai:2006:OVM

[Har06a] István Hargittai. Az öt világformáló marslakó. (Hungarian) [The five world-shaping Martians]. Vince, Budapest, Hungary, 2006. ISBN 963-9552-77-1. 397 + 32 pp. LCCN QC15 .H27155 2006.

Hargittai:2006:MSF

[Har06b] István Hargittai. The Martians of Science: five physicists who changed the twentieth century. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-517845-9 (hardcover). xxiv + 313 + 32 pp. LCCN QC15 .H27 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0636/2005029427-d.
html;   http://www.loc.gov/catdir/enhancements/fy0724/
2005029427-b.html;   http://www.loc.gov/catdir/toc/ecip061/
2005029427.html.

Hecht:1990:VCS

[HB90] K. T. Hecht and L. C. Biedenharn. Vector coherent state constructions of U(3) symmetric tensors and their SU(3) ⊃ SU(2) × U(1) Wigner coefficients. Journal of Mathematical Physics, 31(12): 2781–2796, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v31/i12/p2781_s1.

Hedrich:1993:GUE

[Hed93] Reiner Hedrich. Die nicht ganz so unglaubliche Effizienz der Mathematik in den Naturwissenschaften. (German) [The not-so-incredible effectiveness of mathematics in science]. Philosophia Naturalis, 30(1):106–125, 1993. CODEN ???? ISSN 0031-8027.

Heilbron:1974:BRA

[Hei74] J. L. Heilbron. Book review: Abdus Salam and Eugene Wigner, Aspects of Quantum Theory. Isis, 65(2):292–293, June 1974. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229412.

Hendry:1989:BRB

[Hen89] John Hendry. Book review: Behram N. Kurşunoğlu & Eugene P. Wigner (eds). Reminiscences about a Great Physicist: Paul Adrien Maurice Dirac. Cambridge: Cambridge University Press, 1987. Pp. xviii + 297. ISBN 0-521-34013-6. £30.00. C.W. Kilmister (ed.). Schrödinger: Centenary Celebration of a Polymath. Cambridge: Cambridge University Press, 1987. Pp. x + 253. ISBN 0-521-34017-9. £30.00. British Journal for the History of Science, 22(1):87–88, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4026687.

Herbig:1976:KDA

[Her76] Jost Herbig. Kettenreaktion: das Drama der Atomphysiker. (German) [Nuclear reaction: the drama of the nuclear physicist]. Carl Hanser, München, Germany, 1976. ISBN ???? 514 + 8 pp. LCCN ???? Neumann János (1903–1957), Teller Ede (1908– 2003), Szilárd Leó (1898-1964), Wigner Jenő (1902–1995).

Herschman:1979:APN

[Her79] Arthur Herschman, editor. Abstracts of papers of the 145th national meeting, 3–8 January 1979, Houston, Texas, volume 79-2 of AAAS publication. American Association for the Advancement of Science, Washington, DC, USA, 1979. ISBN ???? LCCN ????

Hargittai:2004:CSI

[HH04a] Magdolna Hargittai and István Hargittai, editors. Candid science IV: conversations with famous physicists. Imperial College Press, London, UK, 2004. ISBN 1-86094-414-0, 1-86094-416-7 (paperback). xvi + 711 pp. LCCN QC15 .H295 2004. URL http://

www.worldscibooks.com/physics/p304.html.

Hargittai:2004:EPW

[HH04b] Magdolna Hargittai and István Hargittai. Eugene P. Wigner. In Candid science IV: conversations with famous physicists [HH04a], pages 1–19. ISBN 1-86094-414-0, 1-86094-416-7 (paperback).

LCCN QC15 .H295 2004. URL http://www.worldscibooks.
com/physics/p304.html.

Han:1986:EPW

[HKS86] D. Han, Y. S. Kim, and D. Son. Eulerian parametrization of Wigner’s little groups and gauge transformations in terms of rotations in two-component spinors. Journal of Mathematical Physics, 27(9):2228–2235, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v27/i9/p2228_s1.

Hook:1977:ETS

[HKT77] Sidney Hook, Paul Kurtz, and Miro Todorovich, editors. The Ethics of teaching and scientific research. Prometheus Books, Buffalo, NY, USA, 1977. ISBN 0-87975-068-5. xiii + 212 pp. LCCN LB2331.4 .E86. US$11.95.

Hegerfeldt:1968:PFS

[HKW68] G. C. Hegerfeldt, K. Kraus, and E. P. Wigner. Proof of the fermion superselection rule without the assumption of timereversal invariance. Journal of Mathematical Physics, 9(12): 2029–2031, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v9/i12/p2029_s1;   http:
//link.aip.org/link/jmapaq/v9/i12/p2029/s1.

Heidenreich:1988:QCI

[HL88] W. F. Heidenreich and M. Lorente. Quantization of conformally invariant Bargmann–Wigner equations with gauge freedom. Journal of Mathematical Physics, 29(7):1698–1704, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v29/i7/p1698_s1.

Heidenreich:1990:BWE

[HL90] W. F. Heidenreich and M. Lorente. Bargmann–Wigner equations in de Sitter space. Journal of Mathematical Physics, 31(4):939– 947, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i4/p939_s1.

Hubac:2001:MBW

[HMW01] I. Hubač, P. Mach, and S. Wilson. Multireference Brillouin– Wigner methods for many-body systems. Advances in Quantum Chemistry, 39:225–240, 2001. CODEN AQCHA9. ISSN

0065-3276. URL http://www.sciencedirect.com/science/
article/pii/S0065327605390149.

Hubac:2002:PCM

[HMW02] I. Hubać, P. Mach, and S. Wilson. A posteriori corrections to multireference limited configuration interaction based on a Brillouin– Wigner perturbative analysis. International Journal of Quantum Chemistry, 89(4):198–207, ???? 2002. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hubac:2005:MBW

[HMW05] I. Hubač, P. Mach, and S. Wilson. Multireference Brillouin– Wigner coupled cluster (MR-BWCC) theory applied to the H8 model: Comparison with CCSD(T) theory. International Journal of Quantum Chemistry, 104(4):387–396, ???? 2005. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Halevy:2009:UED

[HNP09] Alon Halevy, Peter Norvig, and Fernando Pereira. The unreasonable effectiveness of data. IEEE Intelligent Systems, 24(2): 8–12, March/April 2009. ISSN 1541-1672 (print), 1941-1294

(electronic). URL https://static.googleusercontent.com/
media/research.google.com/en//pubs/archive/35179.pdf.

Hodgson:1976:WCN

[Hod76] P. E. Hodgson. Wigner cusps in nuclear reactions. Nature, 259 (5542):364–365, February 5, 1976. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v259/n5542/full/259364a0.html.

Hoiberg:2001:YNL

[Hoi01] Dale Hoiberg, editor. 100 years with Nobel laureates. Encyclopaedia Britannica (India) and I.K. International, New Delhi, India, 2001. ISBN 81-88237-00-0. xiii + 1098 pp. LCCN AS911.N9 A15 2001.

Holman:1973:ARS

[Hol73a] Wayne J. Holman III. The application of Regge symmetry to the 9-j symbol. Journal of Mathematical Physics, 14(3):330– 336, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i3/p330_s1.

Holman:1973:WC

[Hol73b] Wayne J. Holman III. On the Wigner coefficients of Sp(4) and SO(5). Journal of Mathematical Physics, 14(4):440–447, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v14/i4/p440_s1.

Holman:1973:RT

[Hol73c] Wayne J. Holman III. The representation theory of SU(4) and Sp(4). Journal of Mathematical Physics, 14(4):448–455, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v14/i4/p448_s1.

Hongoh:1976:WCT

[Hon76] M. Hongoh. Wigner coefficients for the theory of five-dimensional quasispin. Journal of Mathematical Physics, 17(4):535–541, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v17/i4/p535_s1.

Hooker:1973:CRF

[Hoo73] C. A. (Clifford Alan) Hooker, editor. Contemporary research in the foundations and philosophy of quantum theory; proceedings of a conference held at the University of Western Ontario, London, Canada, volume 2 of The University of Western Ontario series in philosophy of science. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1973. ISBN 90-277-0271-3. LCCN QC174.12 .C66.

Horvath:1997:TKE

[Hor97] T. Horvath. Theodore Kármán, [Eugene] Paul Wigner, John [von] Neumann, Leo Szilard, Edward Teller and their ideas of ultimate reality and meaning. Ultimate Reality and Meaning, 20(2–3):123– 146, June/September 1997. ISSN 0709-549X.

Hillery:1984:DFP

[HOSW84] M. Hillery, R. F. O’Connell, M. O. Scully, and E. P. Wigner. Distribution functions in physics: fundamentals. Physics Reports, 106 (3):121–167, 1984. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270 (electronic).

Howlett:1962:BRB

[How62] J. Howlett. Book review: Proceedings of Symposia in Applied Mathematics, vol. XI, Nuclear Reactor Theory, by G. Birkhoff and E. P. Wigner. Mathematical Gazette, 46(357):248–250, October 1962. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL http://www.jstor.org/stable/3614048.

Hecht:1969:WSS

[HP69] K. T. Hecht and Sing Chin Pang. On the Wigner supermultiplet scheme. Journal of Mathematical Physics, 10(9):1571–1616, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i9/p1571_s1.

Hecht:1983:SSW

[HS83] K. T. Hecht and Y. Suzuki. Some special SU(3) ⊇ R(3) Wigner coefficients and their application. Journal of Mathematical Physics, 24(4):785–792, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i4/p785_s1.

Healy:1995:ICC

[HS95] D. M. Healy, Jr. and F. E. Schroeck, Jr. On informational completeness of covariant localization observables and Wigner coefficients. Journal of Mathematical Physics, 36(1):453–507, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hahn:2001:WDA

[HS01] Stefan L. Hahn and Kajetana M. Snopek. Wigner distributions and ambiguity functions in image analysis. Lecture Notes in Computer Science, 2124:537–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/link/service/series/0558/bibs/
2124/21240537.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2124/21240537.pdf.

Hongoh:1974:EWC

[HST74] M. Hongoh, R. T. Sharp, and D. E. Tilley. Explicit O(5) Wigner coefficients. Journal of Mathematical Physics, 15(6):782–788, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v15/i6/p782_s1.

Huddleston:1978:IWC

[Hud78] P. L. Huddleston. Inönü-Wigner contractions of the real fourdimensional Lie algebras. Journal of Mathematical Physics, 19 (8):1645–1649, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v19/i8/p1645_s1.

Hughes:1989:SIQ

[Hug89] R. I. G. Hughes. The Structure and Interpretation of Quantum Mechanics. Harvard University Press, Cambridge, MA, USA, 1989. ISBN 0-674-84391-6 (hardcover), 0-674-84392-4 (paperback). xiii + 369 pp. LCCN QC174.12 .H82 1989.

Houtappel:1965:CBU

[HVW65] R. M. F. Houtappel, H. Van Dam, and E. P. Wigner. The conceptual basis and use of the geometric invariance principles. Reviews of Modern Physics, 37(4):595–632, October 1, 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
37.595;   http://rmp.aps.org/abstract/RMP/v37/i4/p595_

1. Reprinted in [HVW82].

Houtappel:1982:CBU

[HVW82] R. M. F. Houtappel, H. Van Dam, and E. P. Wigner. The conceptual basis and use of the geometric invariance principles. In Rosen [Ros82], page ?? ISBN ???? LCCN ???? Reprint of [HVW65].

Hirschfelder:1935:SRC

[HW35] J. O. Hirschfelder and E. Wigner. Separation of rotational coordinates from the Schrödinger equation for N particles. Proceedings of the National Academy of Sciences of the United States of America, 21(2):113–119, February 15, 1935. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://

www.jstor.org/stable/86865.

Hirschfelder:1939:SQM

[HW39] J. O. Hirschfelder and E. Wigner. Some quantum-mechanical considerations in the theory of reactions involving an activation energy. Journal of Chemical Physics, 7(??):616–628, ???? 1939. CO-DEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

URL http://link.aip.org/link/jcpsa6/v7/i8/p616/s1.

Hewitt:1957:TM

[HW57a] Edwin Hewitt and Eugene P. Wigner. On a theorem of Magnus. Proceedings of the American Mathematical Society, 8:740– 744, 1957. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Hewitt:1957:TMI

[HW57b] Edwin Hewitt and Eugene P. Wigner. On a theorem of Magnus. Proceedings of the American Mathematical Society, 8(4): 740–744, August 1957. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http://www.jstor.org/

stable/2033291.

Haaland:1974:DCA

[HW74] C. M. Haaland and E. Wigner. Defense of cities by antiballistic missiles. Bulletin of the Operations Research Society of America, 22(Supplement 1):??, Spring 1974. CODEN ???? ISSN ????

Haaland:1977:DCA

[HW77] C. M. Haaland and E. P. Wigner. Defense of cities by antiballistic missiles. SIAM Review, 19(2):279–296, April 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://www.jstor.org/stable/2029503.

Hubac:2001:GBW

[HW01] I. Hubac and S. Wilson. On the generalized Brillouin–Wigner perturbation theory and the many-body problem. Advances in Quantum Chemistry, 39:209–223, 2001. CODEN AQCHA9. ISSN

0065-3276. URL http://www.sciencedirect.com/science/
article/pii/S0065327605390137.

Hinshelwood:1938:GD

[HWWJ^{+}38] C. N. Hinshelwood, E. Wigner, W. F. K. Wynne-Jones, V. K. La Mer, E. A. Guggenheim, M. C. Evans, J. A. Christiansen, M. G. Evans, A. Wassermann, A. R. Ubbelohde, M. Polányi, R. H. Fowler, E. A. Moelwyn-Hughes, L. P. Hammett, C. F. Goodeve, and E. Rabinowitch. General discussion. Transactions of the Faraday Society, 34:70–81, 1938. CODEN TFSOA4. ISSN 0014-7672.

Imre:1967:WMQ

[IÖRZ67] Kaya Imre, Ercüment Özizmir, Marcos Rosenbaum, and P. F. Zweifel. Wigner method in quantum statistical mechanics. Journal of Mathematical Physics, 8(5):1097–1108, May 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/
i5/p1097_s1.

Islami:2022:UEM

[Isl22] Arezoo Islami. The unreasonable effectiveness of mathematics: From Hamming to Wigner and back again. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00592-8. See [?].

Iyakutti:2012:WCQ

[ISRK12] Kombiah Iyakutti, Velappa Jayaraman Surya, Ratnavelu Rajeswarapalanichamy, and Yoshiyuki Kawazoe. Wigner crystallization of quadratically dispersing electrons in graphene. International Journal of Quantum Chemistry, 112(6):1725–1736, March 15, 2012. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Inonu:1952:RGG

[IW52] E. Inönü and E. P. Wigner. Representations of the Galilei group. Il Nuovo Cimento (9), 9(8):705–718, ???? 1952. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://

www.springerlink.com/content/vw86571351k88751.

Inonu:1953:CGT

[IW53] E. Inonu and E. P. Wigner. On the contraction of groups and their representations. Proceedings of the National Academy of Sciences of the United States of America, 39(6):510–524, June 15, 1953. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490

(electronic). URL http://www.jstor.org/stable/88703.

Inonu:1954:PTC

[IW54] E. Inonu and E. P. Wigner. On a particular type of convergence to a singular matrix. Proceedings of the National Academy of Sciences of the United States of America, 40(2):119–121, February 15, 1954. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490

(electronic). URL http://www.jstor.org/stable/89064.

Jahn:1981:RCP

[Jah81] Robert G. Jahn, editor. The role of consciousness in the physical world, volume 57 of AAAS selected symposium. Westview Press, Boulder, CO, USA, 1981. ISBN 0-89158-955-4. xiii + 136 pp. LCCN QC30 .R64.

Jang:1968:RAW

[Jan68] S. Jang. Relationships among the Wigner 9-j symbols. Journal of Mathematical Physics, 9(3):397–402, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i3/
p397_s1.

Janssen:1981:PWW

[Jan81] A. J. E. M. Janssen. Positivity of weighted Wigner distributions. SIAM Journal on Mathematical Analysis, 12(5):752–758, September 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Janssen:1984:NHT

[Jan84] A. J. E. M. Janssen. A note on Hudson’s theorem about functions with nonnegative Wigner distributions. SIAM Journal on Mathematical Analysis, 15(1):170–176, January 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jankvist:2014:HTM

[Jan14] Uffe Thomas Jankvist. A historical teaching module on ‘the unreasonable effectiveness of mathematics’: Boolean algebra and Shannon circuits. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(2):120–133, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17498430.2014.874869.

Joyce:2002:RWC

[JBR02] W. P. Joyce, P. H. Butler, and H. J. Ross. The Racah–Wigner category. Canadian Journal of Physics = Journal canadien de physique, 80(6):613–632, 2002. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic). URL http://www.

nrcresearchpress.com/doi/abs/10.1139/p02-020.

Jau:1990:STU

[JC90] J. Y. Jau and R. T. Chin. Shape from texture using the Wigner distribution. Computer Vision, Graphics, and Image Processing, 52(2):248–263, November 1990. CODEN CVGPDB. ISSN 0734-189x (print), 1557-895x (electronic).

Jiang:2010:EBT

[JC10] Haiyan Jiang and Wei Cai. Effect of boundary treatments on quantum transport current in the Green’s function and Wigner distribution methods for a nano-scale DG–MOSFET. Journal of Computational Physics, 229(12):4461–4475, June 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999110000811.

Jiang:2011:AFI

[JCT11] Haiyan Jiang, Wei Cai, and Raphael Tsu. Accuracy of the Frensley inflow boundary condition for Wigner equations in simulating resonant tunneling diodes. Journal of Computational Physics, 230(5):2031–2044, March 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999110006662.

Jones:1990:SRE

[JD90] Raymund C. Jones and Gurjeet S. Dhesi. Some remarks on the eigenvalue spectrum of a large symmetric Wigner randomsign matrix. Canadian Journal of Physics = Journal canadien de physique, 68(11):1304–1311, 1990. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic). URL http://www.

nrcresearchpress.com/doi/abs/10.1139/p90-187.

Johansson:2016:FAE

[JF16] H. T. Johansson and C. Forssén. Fast and accurate evaluation of Wigner 3j, 6j, and 9j symbols using prime factorization and multiword integer arithmetic. SIAM Journal on Scientific Computing, 38(1):A376–A384, ???? 2016. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Jha:2011:WPE

[Jha11] Stefania Jha. Wigner’s “Polanyian” epistemology and the measurement problem: The Wigner–Polanyi dialog on tacit knowledge. Physics in Perspective (PIP), 13(3):329–358, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-010-0050-5.

Janszky:2003:WCI

[JKM03] Jozsef Janszky, Young S. Kim, and Margarita A. Man’ko. Wigner Centennial issue. Journal of Optics B: Quantum and Semiclassical Optics, 5(3):S219–S220, 2003. CODEN JOBOFD. ISSN 1464-4266 (print), 1741-3575 (electronic). URL http://iopscience.

iop.org/1464-4266/5/3/003.

Jiang:2014:DAI

[JLC14] Haiyan Jiang, Tiao Lu, and Wei Cai. A device adaptive inflow boundary condition for Wigner equations of quantum transport. Journal of Computational Physics, 258(??):773–786, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007572.

Jones:2010:THR

[Jon10] Derry W. Jones. Titanic Hungarian refugee physicists. Contemporary Physics, 51(3):267–271, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jordan:1978:SPN

[Jor78] Thomas F. Jordan. Simple proof of no position operator for quanta with zero mass and nonzero helicity. Journal of Mathematical Physics, 19(6):1382–1385, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i6/p1382_s1.

Jordan:1980:SDN

[Jor80] Thomas F. Jordan. Simple derivation of the Newton–Wigner position operator. Journal of Mathematical Physics, 21(8):2028– 2032, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i8/p2028_s1.

Joyce:2001:DPR

[Joy01] William P. Joyce. Diagram projection rules for recoupling diagrams in the Racah–Wigner category. Journal of Mathematical Physics, 42(3):1346–1363, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jordan:1934:AGQ

[JvNW34] P. Jordan, J. von Neumann, and E. Wigner. On an algebraic generalization of the quantum mechanical formalism. Annals of Mathematics (2), 35(1):29–64, January 1934. CODEN AN-MAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL

http://www.jstor.org/stable/1968117. Reprinted in [Tau61b,

Paper 21].

Jordan:1928:PAG

[JW28] P. Jordan and E. Wigner. Über das Paulische Äquivalenzverbot. (German) [On Pauli’s equivalence prohibition]. Zeitschrift für Physik, 47(9–10):631–651, ???? 1928. CODEN ZEPYAA.

ISSN ???? URL http://www.springerlink.com/content/
hx1t32272451437h.

Jacobson:1987:DOF

[JW87] Lowell Jacobson and Harry Wechsler. Derivation of optical flow using a spatiotemporal-frequency approach. Computer Vision, Graphics, and Image Processing, 38(1):29–65, April 1987. CO-DEN CVGPDB. ISSN 0734-189x (print), 1557-895x (electronic).

Jauch:1967:SCC

[JWY67] J. M. Jauch, E. P. Wigner, and M. M. Yanase. Some comments concerning measurements in quantum mechanics. Il Nuovo Cimento B, 48(1):144–151, ???? 1967. CODEN NCIBAW.

ISSN ???? URL http://www.springerlink.com/content/
071173w138r47g34.

Kerr:1984:SCI

[K^{+}84] Donald M. Kerr et al., editors. Science, computers, and the information onslaught: a collection of essays. Academic Press, New York, USA, 1984. ISBN 0-12-404970-2. xiii + 276 pp. LCCN Q223 .S24 1984.

Kastner:2021:UID

[Kas21] R. E. Kastner. Unitary interactions do not yield outcomes: Attempting to model “Wigner’s friend”. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00492-3.

Khan:1971:I

[KAUZ71] M. N. Khan and Arif-Uz-Zaman. Invariants of πN → πB_{s}. Journal of Mathematical Physics, 12(7):1059–1066, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i7/p1059_s1.

Kerimov:1985:WCG

[Ker85] G. A. Kerimov. On the Wigner coefficients of the generalized Lorentz groups in the parabolic basis. Journal of Mathematical Physics, 26(8):1885–1888, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v26/i8/p1885_s1.

Keraani:2005:WMD

[Ker05] Sahbi Keraani. Wigner measures dynamics in a Coulomb potential. Journal of Mathematical Physics, 46(6):063512, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v46/i6/p063512_s1.

Kuhn:1967:SHQ

[KHFA67] Thomas S. Kuhn, John L. Heilbron, Paul Forman, and Lini Allen. Sources for history of quantum physics: an inventory and report, volume 68 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1967. ix + 176 pp. LCCN QC174.1 .S66.

URL http://www.amphilsoc.org/guides/ahqp/;   http://
www.amphilsoc.org/guides/ahqp/s-t.htm#schrodinger.

Kibler:1976:ITM

[Kib76] Maurice R. Kibler. Irreducible tensor method for a chain SU_{2} ⊃ ⋯ ⊃ G′′ ⊃ G′ ⊃ G and molecular physics. Journal of Mathematical Physics, 17(6):855–858, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i6/p855_s1.

King:1958:BRB

[Kin58] R. W. King. Book review: Nuclear Structure, by Leonard Eisenbud and Eugene P. Wigner. Science, 128(3321):409, August 22, 1958. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.jstor.org/stable/1755061.

Kirschenmann:1973:BRB

[Kir73] Peter Kirschenmann. Book review: Symmetries and reflections: Scientific essays: E. P. Wigner Cambridge, Mass.: M.I.T. Press, 1970. Studies in History and Philosophy of Science Part A, 4(2): 193–207, August 1973. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0039368173900046. See [Wig67-28,

Wig70k, Wig78l].

Kluksdahl:1989:RVS

[KKF89] N. C. Kluksdahl, A. M. Kriman, and D. K. Ferry. The role of visualization in the simulation of quantum electronic transport in semiconductors. Computer, 22(8):60–66, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kammerer:2003:WMC

[KL03] Clotilde Fermanian Kammerer and Caroline Lasser. Wigner measures and codimension two crossings. Journal of Mathematical Physics, 44(2):507–527, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klaus:1991:ABJ

[Kla91] Martin Klaus. Asymptotic behavior of Jost functions near resonance points for Wigner–von Neumann type potentials. Journal of Mathematical Physics, 32(1):163–174, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i1/
p163_s1.

Klink:1974:EBW

[Kli74] William H. Klink. Expansions in Breit–Wigner amplitudes and biorthogonal functions. Journal of Mathematical Physics, 15(5): 565–569, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i5/p565_s1.

Kube:2009:MCS

[KLW09] Susanna Kube, Caroline Lasser, and Marcus Weber. Monte Carlo sampling of Wigner functions and surface hopping quantum dynamics. Journal of Computational Physics, 228(6):1947– 1962, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108006104.

Kettle:1994:WSU

[KN94] Sidney F. A. Kettle and Lars J. Norrby. The Wigner–Seitz unit cell. Journal of Chemical Education, 71(12):1003, 1994. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic). URL

http://pubs.acs.org/doi/abs/10.1021/ed071p1003.

Kim:2019:NPE

[KN19] Y. S. Kim and Marilyn E. Noz. New perspectives on Einstein’s E = mc^{2}. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2019. ISBN 981-323-770-8 (hardcover), 981-323-771-6 (e-book). xi + 192 pp. LCCN QC173.59.S65 K56 2018.

Kosina:2004:SSD

[KNS04] H. Kosina, M. Nedjalkov, and S. Selberherr. Solution of the space-dependent Wigner equation using a particle model. Monte Carlo Methods and Applications, 10(3–4):359–368, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2004.10.issue-3-4/mcma.2004.10.3-4.359/mcma.2004.10.3-
4.359.xml.

Kolas:1960:BRB

[Kol60] H. Kolas. Book review: The physical theory of neutron chain reactors, Alvin M. Weinberg and Eugene P. Wigner, University of Chicago Press. 800 pp., $15.00. Journal of Nuclear Energy. Part A. Reactor Science, 12(4):179, August 1960. CO-DEN JNEAA8. ISSN 0368-3265 (print), 1878-2116 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0368326560900955.

Kolman:1973:LAA

[Kol73] Bernard Kolman, editor. Lie algebras: applications and computational methods; papers. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1973. 155 pp. LCCN QA252.3 .C66 1972. Reprinted from the September, 1973, issue of the SIAM Journal on Applied Mathematics, volume 25, number 2.

Kota:2005:WRA

[Kot05] V. K. B. Kota. SU(N) Wigner–Racah algebra for the matrix of second moments of embedded Gaussian unitary ensemble of random matrices. Journal of Mathematical Physics, 46(3): 033514, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i3/p033514_s1.

Kota:2007:WRA

[Kot07] V. K. B. Kota. U(2Ω) ⊃ U(Ω) ⊃ ⊗SU(2) Wigner–Racah algebra for embedded Gaussian unitary ensemble of random matrices with spin. Journal of Mathematical Physics, 48(5):053304, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v48/i5/p053304_s1.

Kouts:1959:BRB

[Kou59] Herbert Kouts. Book review: The Physical Theory of Neutron Chain Reactors, by Alvin M. Weinberg and Eugene P. Wigner. Science, 129(3356):1135, April 24, 1959. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1756793.

Kasperkovitz:1990:STI

[KP90] P. Kasperkovitz and M. K. Peev. Similarity transformations of irreducible corepresentations in Wigner canonical form. Journal of Mathematical Physics, 31(6):1304–1309, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i6/
p1304_s1.

Kurtsiefer:1997:MWF

[KPM97] Ch. Kurtsiefer, T. Pfau, and J. Mlynek. Measurement of the Wigner function of an ensemble of helium atoms. Nature, 386 (6621):150–153, March 13, 1997. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v386/n6621/pdf/386150a0.pdf.

Kadiroglu:1978:NEA

[KPS78] Osman Kemal Kadiroğlu, Arnold Perlmutter, and Linda F. Scott, editors. Nuclear energy and alternatives: proceedings of the International Scientific Forum on an Acceptable Nuclear Energy Future of the World, November 7–11, 1977. Ballinger Publishing Company, Cambridge, MA, USA, 1978. ISBN 0-88410-081-2. LCCN TK9006 .I6 1977.

Kamupingene:1986:WQS

[KPT86] A. H. Kamupingene, T. D. Palev, and S. P. Tsaneva. Wigner quantum systems. Two particles interacting via a harmonic potential. I. Two-dimensional space. Journal of Mathematical Physics, 27(8):2067–2075, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v27/i8/p2067_s1.

Kramer:1978:GTM

[KR78] Peter Kramer and Alfred Rieckers, editors. Group theoretical methods in physics: sixth international colloquium, Tübingen, 1977, volume 79 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. ISBN 0-387-08848-2. xviii + 546 pp. LCCN QC20.7.G76 I57 1977.

Knight:1982:WTS

[KS82] B. W. Knight and L. Sirovich. The Wigner transform and some exact properties of linear operators. SIAM Journal on Applied Mathematics, 42(2):378–389, April 1982. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

www.jstor.org/stable/2101219.

Kozlov:2007:WFD

[KS07] V. V. Kozlov and O. G. Smolyanov. Wigner function and diffusion in a collision-free medium of quantum particles. Theory of Probability and its Applications, 51(1):168–181, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuenzli:1961:LER

[Kue61] Alfred E. Kuenzli. Letter to the Editor: Response to Wigner. Bulletin of the Atomic Scientists, 17(5–6):256, May/June 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wig61h].

Kearny:1977:SCD

[KW77] Cresson H. Kearny and Eugene P. Wigner. Soviet civil defense. Science, 195(4275):243, January 21, 1977. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.jstor.org/stable/1743238;   http://www.sciencemag.
org/content/195/4275/243.1.full.pdf.

Kim:1987:CPSb

[KW87a] Y. S. Kim and E. P. Wigner. Covariant phase-space representation for localized light waves. Physical Review A (Atomic, Molecular, and Optical Physics), 36(??):1293–1297, August 1, 1987. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-

4519. URL http://link.aps.org/doi/10.1103/PhysRevA.36.
1293.

Kim:1987:CGM

[KW87b] Y. S. Kim and E. P. Wigner. Cylindrical group and massless particles. Journal of Mathematical Physics, 28(5):1175–1179, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v28/i5/p1175_s1;   http://link.aip.org/link/
jmapaq/v28/i5/p1175/s1.

Kim:1987:CPSa

[KW87c] Y. S. Kim and Eugene P. Wigner. Covariant phase space representation for localized light waves. Technical Report MdDP-PP-87-197, University of Maryland, College Park, MD, USA, 1987. 18 pp.

Kursunoglu:1987:RAG

[KW87d] Behram Kurşunoğlu and Eugene Paul Wigner, editors. Reminiscences about a great physicist: Paul Adrien Maurice Dirac. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34013-6. xviii + 297 pp. LCCN QC16.D57 R46 1987. URL http:/

/adsabs.harvard.edu/abs/1987ragp.book.....D;   http://
www.loc.gov/catdir/description/cam023/86033409.html;
http://www.loc.gov/catdir/toc/cam028/86033409.html.

Kim:1988:CPSb

[KW88a] Y. S. Kim and E. P. Wigner. Covariant phase-space representation for harmonic oscillators. Physical Review A (Atomic, Molecular, and Optical Physics), 38(3):1159–1167, August 1, 1988. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-

4519. URL http://link.aps.org/doi/10.1103/PhysRevA.38.
1159.

Kim:1988:CPSc

[KW88b] Y. S. Kim and Eugene P. Wigner. Covariant phase space representation and overlapping distribution functions. Report MdDP-PP-89-092, University of Maryland, College Park, MD, USA, December 1988. 26 pp.

Kim:1988:CPSa

[KW88c] Y. S. Kim and Eugene P. Wigner. Covariant phase space reprresentation for harmonic oscillators. Report MdDP-PP-88-211, University of Maryland, College Park, MD, USA, April 1988. 32 pp.

Kim:1989:CTQ

[KW89a] Y. S. Kim and E. P. Wigner. Canonical transformation in quantum mechanics. Report MdDP-PP-89-165, University of Maryland, College Park, MD, USA, April 1989. ???? pp.

Kim:1989:CPS

[KW89b] Y. S. Kim and E. P. Wigner. Covariant phase-space representation and overlapping distribution functions. Physical Review A (Atomic, Molecular, and Optical Physics), 39(6):2829–2834, March 15, 1989. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://link.aps.org/

doi/10.1103/PhysRevA.39.2829.

Kim:1989:STG

[KW89c] Y. S. Kim and E. P. Wigner. Space-time geometry of relativistic particles. Report MdDP-PP-90-014, University of Maryland, College Park, MD, USA, July 1989. 25 pp. Published in [KW90d].

Kim:1990:CTQ

[KW90a] Y. S. Kim and E. P. Wigner. Canonical transformation in quantum mechanics. American Journal of Physics, 58(5):439–??, May 1990. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://link.aip.org/link/ajpias/v58/i5/
p439/s1.

Kim:1990:ELTa

[KW90b] Y. S. Kim and E. P. Wigner. Entropy and Lorentz transformations. Report MdDP-PP-90-237, University of Maryland, College Park, MD, USA, May 1990. 11 pp.

Kim:1990:ELTb

[KW90c] Y. S. Kim and E. P. Wigner. Entropy and Lorentz transformations. Physics Letters A, 147(7):343–347, 1990. CODEN PY-LAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

Kim:1990:STG

[KW90d] Y. S. Kim and E. P. Wigner. Space–time geometry of relativistic particles. Journal of Mathematical Physics, 31(1): 55–60, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v31/i1/p55_s1;   http://
link.aip.org/link/jmapaq/v31/i1/p55/s1.

Kovacs:2001:WJT

[KW01] László Kovács and Eugene Paul Wigner. Wigner Jenő és tanárai. (Hungarian) [Eugene Wigner’s teachers], volume 7 of Habilitationes savarienses. Savaria University Press, Szombathely, Hungary, 2001. ISBN 963-8275-90-1. 63 + 4 pp. LCCN QC15 .K68 2001.

L:1988:BRS

[L.88] K. L. Book review: Some other books of interest: Tributes to Paul Dirac, by J. G. Taylor; Reminiscences about a Great Physicist, by Behram N. Kurşunoğlu and Eugene P. Wigner; Instruments and Experiences, by R. V. Jones; Kelvin’s Baltimore Lectures and Modern Theoretical Physics, by Robert Kargon and Peter Achinstein. Science, 241(4870):1239–1240, September 2, 1988. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.jstor.org/stable/1702737.

Ladanyi:1993:HHS

[Lad93] Ladányi Andor. Három hazai szakvélemény a fiatal Wignerről és Neumannról. (Hungarian) [Three young domestic opinions (??) and Wigner and von Neumann]. Hiány, 4(6-7):56–60, 1993. CO-DEN ???? ISSN ????

Lagolnitzer:1984:WPS

[Lag84] D. Lagolnitzer. Wigner’s phase-space density. Nature, 310(5979): 635, August 23, 1984. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v310/n5979/pdf/310635b0.pdf.

Lai:1994:CAF

[Lai94] Shan-Tao Lai. Computation of algebraic formulas for Wigner 3-j, 6-j, and 9-j symbols by Maple. International Journal of Quantum Chemistry, 52(3):593–607, October 15, 1994. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lanouette:1983:DM

[Lan83] William Lanouette. Dream machine. Atlantic Monthly, ??(??):35– 52, 85–86, April 1983. ISSN 1072-7825 (print), 2151-9463 (electronic). A 20-page cover story about the history of the nuclear breeder reactor, a power plant designed to make more fuel than it consumes, including Szilard’s role both devising and naming the nuclear “chain reaction” in 1933 and the “breeder” in 1943. Letters and author’s replies in June, pages 6-7; July, pages 6 & 8; and August.

Landau:1987:EWD

[Lan87] Basil V. Landau. Extensions of Wigner’s distribution to particles with spin 1/2. Journal of Mathematical Physics, 28(7):1508–1511, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v28/i7/p1508_s1.

Lanouette:1993:HGT

[Lan93a] William Lanouette. Hungarian geniuses and their bomb. Seminar on Science and Social Responsibility, about Szilard, Teller, von Neumann, and Wigner. Georgetown University, Washington, DC, USA., May 10, 1993.

Lanouette:1993:WFM

[Lan93b] William Lanouette. Where the four Martians of the Manhattan Project landed: The national laboratory connections of Teller, Wigner, von Neumann, and Szilard. Talk at the American Physical Society, Washington, DC, USA., April 13, 1993.

Lanouette:1999:WFM

[Lan99] William Lanouette. Where the four Martians of the Manhattan Project landed: The National Laboratory connections of Teller, Wigner, von Neumann, and Szilard. A Talk at a Conference on Laboratory History and Sociology, State University of New York at Stony Brook & Brookhaven National Laboratory, June 10, 1999.

Lanouette:2011:FHM

[Lan11] William Lanouette. The Four Hungarians of the Manhattan Project. Washington Magyar Club. Described the roles of Leo Szilard, Eugene Wigner, John von Neumann, and Edward Teller in the Manhattan Project., January 15, 2011.

Lapp:1960:LEG

[Lap60] Ralph E. Lapp. Letter to the Editor: The GAC Report. Bulletin of the Atomic Scientists, 16(5):193, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wig60c].

Louck:1971:ISR

[LB71] J. D. Louck and L. C. Biedenharn. Identity satisfied by the Racah coefficients of u(n). Journal of Mathematical Physics, 12(2):173– 177, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i2/p173_s1.

Louck:1973:SCT

[LB73] J. D. Louck and L. C. Biedenharn. On the structure of the canonical tensor operators in the unitary groups. III. Further developments of the boson polynomials and their implications. Journal of Mathematical Physics, 14(10):1336–1357, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i10/p1336_s1.

Lee:1996:WWM

[LB96] James B. Lee and Arthur A. Broyles. What Wigner meant to signal theory, what ‘particle’ meant to Wigner. Physics Today, 49(9):11, 13, September 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v49/i9/p11_s1.

Leverett:1942:DHC

[LCM^{+}42] M. C. Leverett, C. M. Cooper, T. V. Moore, E. P. Wigner, E. S. Steinbach, E. Fermi, J. A. Wheeler, S. K. Allison, and Leo Szilard. Discussion of helium cooled power plant. Report CS-267, US Atomic Energy Commission, Washington, DC, USA, September 16, 1942.

Leegwater:2022:WGH

[Lee22] Gijs Leegwater. When Greenberger, Horne and Zeilinger meet Wigner’s friend. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s10701-022-00586-6.

Lerner:1990:WWF

[Ler90] Nicolas Lerner. Wick–Wigner functions and tomographic methods. SIAM Journal on Mathematical Analysis, 21(4):1083–1092, July 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lieb:1990:IBR

[Lie90] Elliott H. Lieb. Integral bounds for radar ambiguity functions and Wigner distributions. Journal of Mathematical Physics, 31 (3):594–599, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i3/p594_s1.

Litvin:1972:DDP

[Lit72] Daniel B. Litvin. On the decomposition of direct products of irreducible representations. Journal of Mathematical Physics, 13 (9):1386–1389, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v13/i9/p1386_s1.

Lawson:2004:DHM

[LK04] C. George Lawson and Carolyn Krause. Documenting history: Minutes of the New Piles Committee meetings. Nuclear News, 47(12):36–38, November 2004. CODEN NUNWA8. ISSN 0029-

5574. URL http://www.ans.org/pubs/magazines/nn/docs/
2004-11-3.pdf.

Lulek:1985:RWA

[LLBC85] B. Lulek, T. Lulek, J. Biel, and R. Chatterjee. Racah–Wigner approach to standardization of permutation representations for finite groups. Canadian Journal of Physics = Journal canadien de physique, 63(8):1065–1073, 1985. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic). URL http://www.

nrcresearchpress.com/doi/abs/10.1139/p85-174.

Li:2009:URM

[LLW^{+}09] D. Li, X. Li, F. Wang, H. Huang, X. Li, and L. C. Kwek. Uncertainty relation of mixed states by means of Wigner–Yanase–Dyson information. Physical Review A (Atomic, Molecular, and Optical Physics), 79(??):052106, May 6, 2009. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http:

//link.aps.org/doi/10.1103/PhysRevA.79.052106.

Lomont:1963:WUA

[LM63] J. S. Lomont and P. Mendelson. The Wigner Unitarity– Antiunitarity Theorem. Annals of Mathematics, 78(3):548–559, November 1963. CODEN ANMAAH. ISSN 0003-486X (print),

1939-8980 (electronic). URL http://www.jstor.org/stable/
1970540.

Lea:1989:QMW

[LM89] M. J. Lea and N. H. March. Quantum-mechanical Wigner electron crystallization with and without magnetic fields. International Journal of Quantum Chemistry, 36(S23):717–729, April 1– 8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lemke:1992:WAC

[LNPC92] Jürgen Lemke, Yuval Ne’eman, and Jose Pecina-Cruz. Wigner analysis and Casimir operators of \overline{SA}(4, R). Journal of Mathematical Physics, 33(8):2656–2659, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i8/p2656_s1.

Lieb:2010:LMW

[LO10] Elliott H. Lieb and Yaron Ostrover. Localization of multidimensional Wigner distributions. Journal of Mathematical Physics, 51(10):102101, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v51/i10/p102101_s1.

Lockwood:1976:CPC

[Loc76] Loren A. Lockwood. Canonical parameters of the 3j coefficient. Journal of Mathematical Physics, 17(9):1671–1672, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v17/i9/p1671_s1.

Loebl:1971:GTA

[Loe71] Ernest M. (Ernest Moshe) Loebl, editor. Group theory and its applications, volume 2. Academic Press, New York, USA, 1971. ISBN 0-12-455152-1. 312 pp. LCCN ????

Loebl:1968:GTA

[Loe75a] Ernest M. (Ernest Moshe) Loebl, editor. Group theory and its applications, volume 1. Academic Press, New York, USA, 1968– 1975. ISBN 0-12-455150-5. xxvii + 696 pp. LCCN ????

Loebl:1975:GTA

[Loe75b] Ernest M. (Ernest Moshe) Loebl, editor. Group theory and its applications, volume 3. Academic Press, New York, USA, 1975. ISBN 0-12-455153-X (vol. 3). xvi + 480 pp. LCCN ????

Lowdin:1964:SPTa

[Löw64] Per-Olov Löwdin. Studies in perturbation theory. II. Generalization of the Brillouin–Wigner formalism. Journal of Molecular Spectroscopy, 13(3):326–331, July 1964. CODEN JMOSA3. ISSN 0022-2852 (print), 1096-083x (electronic).

Li:2011:CWM

[LPN11] Huaqing Li, Jens Aage Poulsen, and Gunnar Nyman. The classical Wigner method with an effective quantum force: Application to the collinear H + H_{2} reaction. Journal of Physical Chemistry A, 115(25):7338–7345, 2011. CODEN JPCAFH. ISSN ???? URL

http://pubs.acs.org/doi/abs/10.1021/jp200886v.

Larsen:1978:USG

[LR78] M. L. Larsen and W. W. Repko. The use of the symmetric group in the construction of multispinor Lagrangians. Journal of Mathematical Physics, 19(5):930–934, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i5/p930_s1.

Lorente:1983:LBW

[LR83] M. Lorente and M. A. Rodríguez. A Lagrangian of Bargmann– Wigner equations for massive particles of spin 2. Journal of Mathematical Physics, 24(12):2823–2827, December 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i12/p2823_s1.

Margenau:1976:VPR

[LS76] Ervin Laszlo and Emily B. Sellon, editors. Vistas in physical reality: a festschrift for Henry Margenau. Plenum Press, New York, NY, USA; London, UK, 1976. ISBN 0-306-30884-3. xii + 228 pp. LCCN Q175.3 .V57.

Lievens:2006:HOC

[LSV06] S. Lievens, N. I. Stoilova, and J. Van der Jeugt. Harmonic oscillators coupled by springs: Discrete solutions as a Wigner quantum system. Journal of Mathematical Physics, 47(11):113504, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v47/i11/p113504_s1.

Lievens:2008:HOC

[LSV08] S. Lievens, N. I. Stoilova, and J. Van der Jeugt. Harmonic oscillator chains as Wigner quantum systems: Periodic and fixed wall boundary conditions in }↕(1|n) solutions. Journal of Mathematical Physics, 49(7):073502, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v49/i7/p073502_s1.

Lane:1955:GRI

[LTW55] A. M. Lane, R. G. Thomas, and E. P. Wigner. Giant resonance interpretation of the nucleon–nucleus interaction. Physical Review (2), 98(3):693–701, May 1, 1955. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.98.693.

Luo:2004:WYS

[Luo04] Shunlong Luo. Wigner–Yanase skew information vs. quantum Fisher information. Proceedings of the American Mathematical Society, 132(3):885–890, March 2004. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http://www.

jstor.org/stable/1194711.

Ladenburg:1934:ANP

[LW34] Rudolf Ladenburg and Eugene Wigner. Award of the Nobel Prizes in Physics to Professors Heisenberg, Schrödinger and Dirac. The Scientific Monthly, 38(1):86–91, January 1934. CODEN SC-MOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL

http://www.jstor.org/stable/15534.

Lannutti:1978:CTT

[LW78] J. E. Lannutti and P. K. Williams, editors. Current trends in the theory of fields: (Tallahassee 1978): a symposium in honor of P.

A. M. Dirac, volume 48 of AIP conference proceedings. American Institute of Physics, Woodbury, NY, USA, 1978. ISBN 0-88318-147-9. LCCN QC793.3.F5 C87.

Laughlin:1974:BWP

[LWA74] C. Laughlin, M. R. Woodward, and A. T. Amos. Brillouin–Wigner perturbation theory and the generalized eigenvalue equation. International Journal of Quantum Chemistry, 8(4):491–498, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Littlejohn:2009:USA

[LY09] Robert G. Littlejohn and Liang Yu. Uniform semiclassical approximation for the Wigner 6j-symbol in terms of rotation matrices? Journal of Physical Chemistry A, 113(52):14904–14922, 2009. CODEN JPCAFH. ISSN 1089-5639 (print), 1520-5215

(electronic). URL http://pubs.acs.org/doi/abs/10.1021/

jp905056y. PMID: 19817389.

Lange:1994:DWP

[LZ94] Horst Lange and P. F. Zweifel. Dissipation in Wigner–Poisson systems. Journal of Mathematical Physics, 35(4):1513–1521, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v35/i4/p1513_s1.

Macphail:1959:PCM

[Mac59] M. S. Macphail, editor. Proceedings of the 4th Canadian Mathematical Congress, Banff, 1957. University of Toronto Press, Toronto, ON, Canada, 1959. LCCN ????

Macia:2004:WMD

[Mac04] Fabricio Macià. Wigner measures in the discrete setting: Highfrequency analysis of sampling and reconstruction operators. SIAM Journal on Mathematical Analysis, 36(2):347–383, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.siam.org/sam-bin/dbq/article/
43152.

Mukunda:2004:WDQ

[MACS04] N. Mukunda, Arvind, S. Chaturvedi, and R. Simon. Wigner distributions and quantum mechanics on Lie groups: The case of the regular representation. Journal of Mathematical Physics, 45 (1):114–148, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i1/p114_s1.

Maekawa:1991:WRC

[Mae91] Takayoshi Maekawa. On the Wigner and the Racah coefficients of su_{q}(2) and su_{q}(1, 1). Journal of Mathematical Physics, 32(10): 2598–2604, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i10/p2598_s1.

Majorana:1932:TRP

[Maj32] Ettore Majorana. Teoria relativistica di particelle con momento intrinseco arbitrario. (Italian) [Relativistic theory of particles with arbitrary intrinsic momentum]. Il Nuovo Cimento (8), 9(10):335– 344, December 1932. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). See comments [?]. That paper reports that both the neutron [?, ?] and positron [?] had already been discovered when Majorana’s article was published, but that news had not yet reached Majorana when he was preparing this paper.

Majorana:1937:TSD

[Maj37] Ettore Majorana. Teoria simmetrica dell elettrone e del positrone. (Italian) [Symmetrical theory of the electron and the positron]. Il Nuovo Cimento (8), 14(4):171–184, April 1937. CODEN NU-CIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL

http://en.wikipedia.org/wiki/Majorana_mass;   http://
nl.wikipedia.org/wiki/Majorana-deeltje. In this paper, Ma-

jorana predicted the existence of a new type of particle, now called a Majorana fermion, which is its own antiparticle, and whose existence may have finally been confirmed by experiment seventy years later [?, ?, ?]. See also comments in [?]. Esposito [?] reports about this paper “With amazing farsightedness Majorana suggested that the neutrino, which had just been postulated by Wolfgang Pauli and Fermi to explain puzzling features of radioactive beta decay, could be such a particle. This would make the neutrino unique among the elementary particles and, moreover, enable it to have mass. Today many experiments are still devoted to detect these peculiar properties, which include the phenomenon of neutrino oscillations: we have not yet succeeded to find a definite answer to Majorana’s proposal.”.

Majumdar:1993:IWC

[Maj93] Parthasarathi Majumdar. Inönü-Wigner contraction of Kac– Moody algebras. Journal of Mathematical Physics, 34(5):2059– 2065, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i5/p2059_s1.

Maj:2005:PNR

[Maj05a] Omar Maj. Publisher’s note: “The relationship between the Wigner–Weyl kinetic formalism and the complex geometrical optics method”, [J. Math. Phys. 46, 083510 (2005)]. Journal of Mathematical Physics, 46(8):089901, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i8/

p089901_s1. See [Maj05b].

Maj:2005:RBW

[Maj05b] Omar Maj. The relationship between the Wigner–Weyl kinetic formalism and the complex geometrical optics method. Journal of Mathematical Physics, 46(8):083510, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i8/

p083510_s1. See publisher’s note [Maj05a].

Margenau:1967:BRE

[Mar67] Henry Margenau. Book review: Eugene Wigner, Symmetries and Reflections, Scientific Essays. American Journal of Physics, 35 (12):1169, December 1967. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.org/

link/ajpias/v35/i12/p1169/s1.

Marton:1969:BRE

[Mar69] L. Marton. Book review: Eugene P. Wigner, Who Speaks for Civil Defense? Physics Today, 22(3):76–??, March 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v22/i3/p76/s1.

Marx:1995:VSW

[Mar95] G. Marx. A visionary scientist, Eugene P. Wigner 1902–1995. Hungarian Quarterly, 36(137):51–61, Spring 1995. ISSN 0028-

Marx:1996:MMG

[Mar96] George Marx. The myth of the Martians and the golden age of Hungarian science. Science & Education (Springer), 5(3):225– 234, July 1996. CODEN SCEDE9. ISSN 0926-7220 (print),

1573-1901 (electronic). URL http://adsabs.harvard.edu/abs/
1996Sc%26Ed...5..225M.

March:2003:CIC

[Mar03] N. H. March. Classic ionic crystals and quantal Wigner electron solids: Role of electron correlation. International Journal of Quantum Chemistry, 92(1):11–21, March 5, 2003. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

March:2006:QSC

[Mar06a] N. H. March. Quantum statistics of charged particles and fingerprints of Wigner crystallization in D dimensions. International Journal of Quantum Chemistry, 106(15):3032–3042, ???? 2006. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Marton:2006:GEN

[Mar06b] Kati Marton. The great escape: nine Jews who fled Hitler and changed the world. Simon and Schuster, New York, NY, USA, 2006. ISBN 0-7432-6115-1 (hardcover), 0-7432-6116-X (paperback). 271 + 16 pp. LCCN DS135.H93 A153 2006. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0665/2006049162-
d.html;   http://catdir.loc.gov/catdir/enhancements/fy0668/
2006049162-s.html;   http://catdir.loc.gov/catdir/enhancements/
fy0668/2006049162-t.html;   http://catdir.loc.gov/catdir/
enhancements/fy0912/2006049162-b.html.

Maximon:2010:S

[Max10] Leonard C. Maximon. 3j, 6j, 9j symbols. In Olver et al. [OLBC10], chapter 34, pages 757–766. ISBN 0-521-19225-0. LCCN QA331 .N57 2010. US$99.00. URL http://dlmf.nist.

gov/;   http://www.cambridge.org/9780521140638.

McDonnell:2017:WPP

[McD17] Jane McDonnell. Wigner’s puzzle and the Pythagorean heuristic. Synthese, 194(8):2931–2948, August 2017. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:

//link.springer.com/article/10.1007/s11229-016-1080-6;

 http://link.springer.com/content/pdf/10.1007/s11229-
016-1080-6.pdf.

McDayter:1967:GBB

[MD67] Walt McDayter and Norman Drew. The giants: The bomb builders. Denver Post, ??(??):??, February 3, 1967. URL http:

//library.ucsd.edu/dc/object/bb0103915g. This is a reason-

ably accurate 83-frame comic strip on the history of the building of the atomic bomb, with Leo Szilard as the central figure of the story.

Moller:1994:WMD

[MDH94] Klaus B. Moller, Jens Peder Dahl, and Niels E. Henriksen. Wigner method dynamics in the interaction picture. Journal of Physical Chemistry, 98(13):3272–3279, 1994. CODEN JPCHAX. ISSN 0022-3654 (print), 1541-5740 (electronic). URL http://pubs.

acs.org/doi/abs/10.1021/j100064a005.

Mehra:1973:PCN

[Meh73] Jagdish Mehra, editor. The physicist’s conception of nature: Symposium on the Development of the Physicist’s Conception of Nature in the 20th century. Held at the International Centre for Theoretical Physics, Miramare, Trieste, Italy, 18–25 September 1972. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1973. ISBN 90-277-0345-0, 90-277-2536-5. LCCN QC173.96 .S95 1972. URL

http://www.springer.com/us/book/9789027703453.

Meier:1961:LER

[Mei61] Richard L. Meier. Letter to the Editor: Is reactor development economical? Bulletin of the Atomic Scientists, 17(2):40, 81, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [WW60].

Meijer:1964:GTS

[Mei64] Paul H. Meijer, editor. Group theory and solid state physics: a selection of papers, volume 7 of International science review series. Gordon and Breach, New York, NY, USA, 1964. ???? pp. LCCN ????

Mezincescu:1977:WET

[Mez77] L. Mezincescu. Wigner–Eckart theorem for tensor operators of graded Lie algebras. Journal of Mathematical Physics, 18(3):453– 455, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i3/p453_s1.

Mark:1969:PMU

[MF69] Hans Mark and Sidney Fernbach, editors. Properties of matter under unusual conditions: in honor of Edward Teller’s 60th birthday. Interscience Publishers, New York, NY, USA, 1969. ISBN 0-470-56990-5. ix + 389 pp. LCCN Q171 .P98.

Mourgues:1985:NSC

[MFAB85] G. Mourgues, M. R. Feix, J. C. Andrieux, and P. Bertrand. Not necessary but sufficient condition for the positivity of generalized Wigner functions. Journal of Mathematical Physics, 26(10):2554– 2555, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i10/p2554_s1.

Michel:1992:RTE

[MFG^{+}83] M. Michel, V. L. Fitch, F. Gursey, A. Pais, R. U. Sexl, et al. Round table on the evolution of symmetries. In Sant Feliu de Guixols 1983, Proceedings, Symmetries in physics (1600–1980). ????, ????, 1983.

Mayne:1962:SSA

[MGM62] Alan J. Mayne, Irving John Good, and John Maynard Smith, editors. The scientist speculates: an anthology of partly-baked ideas. Heinemann, London, UK, 1962. xvii + 413 pp. LCCN ????

Mains:1994:ARF

[MH94] R. K. Mains and G. I. Haddad. An accurate re-formulation of the Wigner function method for quantum transport modeling. Journal of Computational Physics, 112(1):149–161, May 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710886.

Masik:1998:MBW

[MH98] J. Mášik and I. Hubač. Multireference Brillouin–Wigner coupledcluster theory: Single-root approach. Advances in Quantum Chemistry, 31:75–104, 1998. CODEN AQCHA9. ISSN 0065-3276.

URL http://www.sciencedirect.com/science/article/pii/
S0065327608601847.

Mickens:1985:MAP

[Mic85] Ronald E. Mickens, editor. Mathematical analysis of physical systems. Van Nostrand Reinhold Co., New York, NY, USA, 1985. ISBN 0-442-26077-6. x + 357 pp. LCCN QC20 .M36 1985.

Moshinsky:1991:RIM

[ML91] M. Moshinsky and G. López Laurrabaquio. Relativistic interactions by means of boundary conditions: The Breit–Wigner formula. Journal of Mathematical Physics, 32(12):3519–3528, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i12/p3519_s1.

Mladjenovic:1998:DYN

[Mla98] Milorad Mladjenović. The Defining Years in Nuclear Physics, 1932–1960s. IOP Publishing, Bristol, UK, 1998. ISBN 0-7503-0472-3 (hardcover). xx + 441 pp. LCCN QC773 .M54 1998.

Martin:1973:TCO

[MM73] Ph. Martin and B. Misra. On trace-class operators of scattering theory, and the asymptotic behavior of scattering cross section at high energy. Journal of Mathematical Physics, 14(8):997– 1005, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i8/p997_s1.

Minnaert:1992:RWC

[MM92] Pierre Minnaert and Marek Mozrzymas. Racah–Wigner calculus for the super-rotation algebra. I. Journal of Mathematical Physics, 33(5):1582–1593, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v33/i5/p1582_s1.

Markowich:1994:WFA

[MMP94] P. A. Markowich, N. J. Mauser, and F. Poupaud. A Wignerfunction approach to (semi)classical limits: electrons in a periodic potential. Journal of Mathematical Physics, 35(3):1066– 1094, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i3/p1066_s1.

Mukunda:2005:WWI

[MMZ^{+}05] N. Mukunda, G. Marmo, A. Zampini, S. Chaturvedi, and R. Simon. Wigner–Weyl isomorphism for quantum mechanics on Lie groups. Journal of Mathematical Physics, 46(1):012106, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v46/i1/p012106_s1.

Mucino:2020:WCF

[MO20] R. Muciño and E. Okon. Wigner’s convoluted friends. Studies in History and Philosophy of Modern Physics, 72(??):87–90, November 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219820300976.

Molnar:1999:GWU

[Mol99] Lajos Molnár. A generalization of Wigner’s unitary–antiunitary theorem to Hilbert modules. Journal of Mathematical Physics, 40 (11):5544–5554, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moshinsky:1962:WCG

[Mos62] Marcos Moshinsky. Wigner coefficients for the SU_{3} group and some applications. Reviews of Modern Physics, 34(4): 813–828, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.34.813;   http://
rmp.aps.org/abstract/RMP/v34/i4/p813_1.

Medawar:2001:HGT

[MP01] J. S. Medawar and David Pyke. Hitler’s Gift: the True Story of the Scientists Expelled by the Nazi Regime. Arcade Publishing, New York, NY, USA, 2001. ISBN 1-55970-564-7. xx + 268 pp. LCCN Q141 .M385 2001. Foreword by Max Perutz.

Markowich:2002:WMA

[MPPS02] Peter A. Markowich, Paola Pietra, Carsten Pohl, and Hans Peter Stimming. A Wigner-measure analysis of the Dufort–Frankel scheme for the Schrödinger equation. SIAM Journal on Numerical Analysis, 40(4):1281–1310, August 2002. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/38173.

Markowich:2003:WMA

[MPPS03] Peter A. Markowich, Paola Pietra, Carsten Pohl, and Hans Peter Stimming. A Wigner-measure analysis of the Dufort–Frankel scheme for the Schrödinger equation. SIAM Journal on Numerical Analysis, 40(4):1281–1310, ???? 2003. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

www.jstor.org/stable/4101006.

MacDowell:1972:RPG

[MR72] W. W. MacDowell and Ralph Roskies. Reduction of the Poincaré group with respect to the Lorentz group. Journal of Mathematical Physics, 13(10):1585–1592, October 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v13/i10/p1585_s1. See errata [MR73].

MacDowell:1973:ERP

[MR73] S. W. MacDowell and Ralph Roskies. Errata: “Reduction of the Poincaré group with respect to the Lorentz group”. Journal of Mathematical Physics, 14(12):2018, December 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i12/p2018_s2. See [MR72].

Maurin:1976:MPP

[MR76] Krzysztof Maurin and Ryszard Rñaczka, editors. Mathematical physics and physical mathematics: proceedings of the international symposium organized by the Mathematical Institute of the Polish Academy of Sciences, the Institute for Nuclear Research, and University of Warsaw, held in Warsaw, 25–30 March 1974, volume 2 of Mathematical physics and applied mathematics. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1976. ISBN 90-277-0537-2. LCCN QC19.2 .M37.

Mitra:1998:SLF

[MR98] A. N. Mitra and R. Ramanathan. On simulating Liouvillian flow from quantum mechanics via Wigner functions. Journal of Mathematical Physics, 39(9):4492–4498, September 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meystre:1983:QOE

[MS83] Pierre Meystre and Marlan O. (Marlan Orvil) Scully, editors. Quantum optics, experimental gravity, and measurement theory, volume 94 of NATO advanced science institutes series. Series B, Physics. Plenum Press, New York, NY, USA; London, UK, 1983. ISBN 0-306-41354-X. LCCN QC446.15 .N37 1981. Proceedings of the NATO Advanced Study Institute on Quantum Optics and Experimental General Relativity, held August 16–29, 1981, in Bad Windsheim, Federal Republic of Germany.

Marotta:2000:TOW

[MS00] V. Marotta and A. Sciarrino. Tensor operators and Wigner– Eckart theorem for U_{q→0}(sl(2)). Journal of Mathematical Physics, 41(8):5735–5744, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McKeon:2006:BWE

[MS06] D. G. C. McKeon and T. N. Sherry. The Bargmann–Wigner equations in spherical space. Canadian Journal of Physics = Journal canadien de physique, 84(1):37–52, 2006. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic). URL http://

www.nrcresearchpress.com/doi/abs/10.1139/p06-011.

Mukunda:1995:WEP

[Muk95] N. Mukunda. Eugene Paul Wigner — a tribute. Current Science, 69(4):375–385, August 25, 1995. CODEN CUSCAM. ISSN 0011-3891.

Mark:1924:GRS

[MW24] H. Mark and E. Wigner. Die Gitterstruktur des rhombischen Schwefels. (German) [The grid structure of rhombic sulfur]. Z. phys. Chem., 111(??):398–414, ???? 1924. CODEN ???? ISSN ????

Margenau:1940:MMO

[MW40] H. Margenau and E. Wigner. Magnetic moments of odd nuclei. Physical Review (2), 58(2):103–110, July 15, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.58.103.

Masters:1946:OWN

[MW46] Dexter Masters and Katharine Way, editors. One world or none: a report to the public on the full meaning of the atomic bomb.

Whittlesey House, McGraw-Hill Book Co., Inc., New York, NY, USA, 1946. x + 79 pp. LCCN UF767 .M3 1946a. Foreword by Niels Bohr. Introduction by Arthur H. Compton. See also reprint [MW07].

Margenau:1962:CPP

[MW62] H. Margenau and E. P. Wigner. Comments on Professor Putnam’s comments. Philosophy of Science, 29(3):292–293, July 1962. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186048.

Margenau:1964:RPP

[MW64] Henry Margenau and Eugene P. Wigner. Reply to Professor Putnam. Philosophy of Science, 31(1):7–9, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186741.

Murphey:1975:LAN

[MW75] Walter Murphey and Eugene P. Wigner. Letters: Armed neutrality. Bulletin of the Atomic Scientists, 31(6):2, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morris:1979:NRO

[MW79] Peter A. Morris and E. P. Wigner. Nuclear reactor operation. Science, 205(4402):148, July 13, 1979. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.

org/stable/1748912.

Masters:2007:OWN

[MW07] Dexter Masters and Katharine Way, editors. One world or none: a report to the public on the full meaning of the atomic bomb. New Press, New York, NY, USA, 2007. ISBN 1-59558-227-4 (hardcover). xx + 220 pp. LCCN UG1282.A8 O54 2007. URL http:/

/thenewpress.com/index.php?option=com_title&task=view_
title&metaproductid=1703;   http://www.loc.gov/catdir/
toc/ecip0718/2007020838.htm. Foreword by Niels Bohr. Intro-

duction by Arthur H. Compton. Reprint of [MW46].

Miller:1966:DBH

[MWW66] Carl F. Miller, Sidney G. Winter, Jr., and Eugene P. Wigner. Dialogue: Between Harbor participants and S/C Scientific Advisory Board. Scientist and Citizen, 8(4–5):17–37, March 1966. CODEN

???? ISSN 2155-1278. URL http://www.tandfonline.com/doi/
abs/10.1080/21551278.1966.10114741.

Morrison:1990:ADM

[MYPL90] Robert C. Morrison, Weitao Yang, Robert G. Parr, and Chengteh Lee. Approximate density matrices and Wigner distribution functions from density, kinetic energy density, and idempotency constraints. International Journal of Quantum Chemistry, 38(6):819– 830, December 1990. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nagy:1988:WJK

[Nag88] Nagy Károly. Wigner Jenő köszöntése. (Hungarian) [Eugene Wigner greeting]. Fizikai Szemle (Budapest), 38(5):161, May 1988. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540

(electronic). URL http://www.kfki.hu/fszemle/archivum/
fsz8805/tart8805.html.

Nagy:1989:JNC

[NAH^{+}89] D. Nagy, W. Aspray, P. Horvath, E. Teller, N. Vonneuman, and E. P. Wigner. John von Neumann — a case-study of scientific creativity — discussion. Annals of the History of Computing, 11 (3):165–169, ???? 1989. CODEN AHCOE5. ISSN 0164-1239.

Nedjalkov:2004:OSM

[NAKS04] M. Nedjalkov, E. Atanassov, H. Kosina, and S. Selberherr. Operator-split method for variance reduction in stochastic solutions of the Wigner equation. Monte Carlo Methods and Applications, 10(3–4):461–468, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http:/

/www.degruyter.com/view/j/mcma.2004.10.issue-3-4/mcma.
2004.10.3-4.461/mcma.2004.10.3-4.461.xml.

Narcowich:1988:CCT

[Nar88] Francis J. Narcowich. Conditions for the convolution of two Wigner distributions to be itself a Wigner distribution. Journal of Mathematical Physics, 29(9):2036–2041, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i9/p2036_s1.

Nauenberg:1973:HDE

[Nau73] Michael Nauenberg. High density expansions in the Thomas– Fermi approximation. Journal of Mathematical Physics, 14(4):

537–539, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i4/p537_s1.

Neeman:2004:SMN

[Ne’04] Y. Ne’eman. Symmetry and “magic” numbers or from the Pythagoreans to Eugene Wigner. Acta Physica Hungarica New Series: Heavy Ion Physics, 19(3–4):241–246, ???? 2004. CODEN APHPFM. ISSN 1219-7580 (print), 1588-2675 (electronic).

Nicholson:2012:PWU

[Nic12] Jason Scott Nicholson. A perspective on Wigner’s “unreasonable effectiveness of mathematics”. Notices of the American Mathematical Society, 59(1):38–42, January 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/201201/rtx120100038p.pdf.

Noz:1988:SRQ

[NKW88] Marilyn E. Noz, Y. S. Kim, and Eugene Paul Wigner, editors. Special Relativity and quantum theory: a collection of papers on the Poincaré group. Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 90-277-2799-6. xiii + 504 pp. LCCN QC174.46 .S64 1988.

Nash:1969:SCW

[NL69] J. C. Nash and W. G. Laidlaw. Some comments on the Wigner tunneling formula. Canadian Journal of Chemistry == Journal canadien de chimie, 47(8):1441–1443, 1969. CODEN CJCHAG. ISSN 0008-4042 (print), 1480-3291 (electronic). URL http://

www.nrcresearchpress.com/doi/abs/10.1139/v69-239.

Nobel:1972:NLI

[Nob72] Nobel Foundation, editor. Nobel lectures, including presentation speeches and Laureates’ biographies, 1963–1970. Elsevier, Amsterdam, The Netherlands, 1972. ISBN 0-444-40993-9. xi + 349 pp. LCCN QC71.N735.

Nouri:1999:DWD

[Nou99] Saeid Nouri. Determination of Wigner distribution function for the d-dimensional Coulomb problem. Journal of Mathematical Physics, 40(3):1294–1299, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Newton:1949:LSE

[NW49] T. D. Newton and E. P. Wigner. Localized states for elementary systems. Reviews of Modern Physics, 21(3):400–406, July 1, 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.400;   http://rmp.aps.org/abstract/

RMP/v21/i3/p400_1. Special issue in honor of Albert Einstein on his 70th birthday.

Nordheim:1961:MCR

[NW61] Lothar W. Nordheim and Eugene P. Wigner. Means for controlling reactions. US Patent 2,990,355., June 27, 1961. URL http:

//www.google.com/patents/US2990355. US Patent Application

filed Nov 2, 1945.

Neumann:1970:SNT

[NW70] Neumann János and Wigner Jenő. A színképek néhány tulajdonságának értelmezésér?l a pörg? elektron kvantummechanikája alapján. (Hungarian) [The spectra of some properties: interpretation by quantum mechanics of the spinning electron]. Magyar Fizikai Folyóirat, 18(6):567–582, ???? 1970. CODEN MGFFAC. ISSN 0025-0104.

Narcovich:1971:EMT

[NW71] E. Narcovich and E. Wigner. Extension of the R-matrix theory: Abstract. Bulletin of the American Physical Society, ??(??):623– ??, April 1971. CODEN BAPSA6. ISSN 0003-0503.

Nichols:1972:MRD

[NWA72] Rodney W. Nichols, Eugene P. Wigner, and Robert K. Adair. Military research and development. Science, 175(4020):356–357, January 28, 1972. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
1733555.

Newson:1961:OCS

[NWEC61] Henry W. Newson, Eugene P. Wigner, Elbert P. Epler, and Thomas E. Cole. Overall control system for high flux pile. US Patent 2,985,574., May 23, 1961. URL http://www.google.com/ patents/US2985574. US Patent Application filed May 25, 1953.

Ozay:2025:CCS

[ÖAÖ25] S. Özay, S. Akdemir, and E. Öztekin. The coupling coefficients with six parameters and the generalized hypergeometric functions. Computer Physics Communications, 315(??):??, October 2025. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465525001584.

ONeill:2007:SCH

[OG07] Darragh P. O’Neill and Peter M. W. Gill. Self-Consistent Hartree– Fock–Wigner Calculations: a Two-Electron-Density Functional Theory, volume 958 of ACS Symposium Series, chapter 4, pages 27–35. American Chemical Society, Washington, DC, USA, 2007.

URL http://pubs.acs.org/doi/abs/10.1021/bk-2007-0958.
ch003.

Ohnuki:1978:WMR

[OK78] Y. Ohnuki and S. Kamefuchi. On the wave-mechanical representation of a Bose-like oscillator. Journal of Mathematical Physics, 19 (1):67–78, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i1/p67_s1.

Olver:2010:NHM

[OLBC10] Frank W. J. Olver, Daniel W. Lozier, Ronald F. Boisvert, and Charles W. Clark, editors. NIST Handbook of Mathematical Functions. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19225-0. xv + 951 pp. LCCN QA331 .N57

2010. US$99.00. URL http://dlmf.nist.gov/;   http://
www.cambridge.org/9780521140638.

Oltmans:1974:IEW

[Olt74a] Willem L. Oltmans. Interview with Eugene Wigner. In On growth [Olt74b], pages 349–355. ISBN 0-399-11233-2. LCCN HC59

.O563. URL http://catalog.hathitrust.org/api/volumes/
oclc/979019.html.

Oltmans:1974:G

[Olt74b] Willem L. Oltmans, editor. On growth. Capricorn Books, New York, 1974. ISBN 0-399-11233-2. xii + 493 pp. LCCN HC59

.O563. URL http://catalog.hathitrust.org/api/volumes/
oclc/979019.html.

Omnes:2011:WUE

[Omn11] Roland Omnès. Wigner’s “Unreasonable Effectiveness of Mathematics”, revisited. Foundations of Physics, 41(11):1729– 1739, November 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-011-9587-7.

Orear:1958:DNW

[Ore58] Jay Orear. Detection of nuclear weapons testing. Bulletin of the Atomic Scientists, 14(3):98–101, March 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [?, Wig58e].

OConnell:1977:RBM

[OW77] R. F. O’Connell and Eugene P. Wigner. On the relation between momentum and velocity for elementary systems. Physics Letters A, A61(6):353–354, June 13, 1977. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

OConnell:1978:POSb

[OW78a] R. F. O’Connell and E. P. Wigner. Position operators for systems exhibiting the special relativistic relation between momentum and velocity. Physics Letters A, 67(5–6):319–321, September 4, 1978. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0375960178903171.

OConnell:1978:POSa

[OW78b] R. F. O’Connell and Eugene P. Wigner. Position operators for systems exhibiting the special relativistic relation between momentum and velocity. Technical Report Print-78-0772, Louisiana State, ????, 1978.

OConnell:1981:QMD

[OW81a] R. F. O’Connell and E. P. Wigner. Quantum-mechanical distribution functions: Conditions for uniqueness. Physics Letters A, 83(4):145–148, May 25, 1981. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0375960181908707.

OConnell:1981:SPN

[OW81b] R. F. O’Connell and E. P. Wigner. Some properties of a nonnegative quantum-mechanical distribution function. Physics Letters A, 85(3):121–126, September 21, 1981. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0375960181908811.

OConnell:1984:MBF

[OW84] R. F. O’Connell and E. P. Wigner. Manifestations of Bose and Fermi statistics on the quantum distribution function for systems of spin-0 and spin-1/2 particles. Physical Review A (Atomic, Molecular, and Optical Physics), 30(5):2613–2618, November 1, 1984. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://link.aps.org/doi/

10.1103/PhysRevA.30.2613.

Ohnuki:1992:SAO

[OW92] Yoshio Ohnuki and Shuji Watanabe. Self-adjointness of the operators in Wigner’s commutation relations. Journal of Mathematical Physics, 33(11):3653–3665, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i11/p3653_s1.

Ohlinger:1958:NR

[OWWY58] Leo A. Ohlinger, Eugene P. Wigner, Alvin M. Weinberg, and Gale J. Young. Neutronic reactor. US Patent 2,850,447., Septem-

ber 2, 1958. URL http://www.google.com/patents/US2850447.

US Patent Application filed Nov 2, 1945.

Ohlinger:1959:NR

[OWWY59] Leo A. Ohlinger, Eugene P. Wigner, Alvin M. Weinberg, and Gale J. Young. Neutronic reactor. US Patent 2,890,158., June 9,

1959. URL http://www.google.com/patents/US2890158. US

Patent Application filed Dec 19, 1944.

Ohlinger:1956:R

[OWYW56] Leo A. Ohlinger, Eugene P. Wigner, Gale J. Young, and Alvin M. Weinberg. Reactor. US Patent 2,743,225., April 24, 1956. URL

http://www.google.com/patents/US2743225. US Patent Ap-

plication filed Aug 27, 1946.

Ohlinger:1961:NRF

[OYWW61] Leo A. Ohlinger, Gale J. Young, Eugene P. Wigner, and Alvin M. Weinberg. Nuclear reactor fuel charging/discharging device.

Canadian Patent 613039., January 24, 1961. URL http://

brevets-patents.ic.gc.ca/opic-cipo/cpd/eng/patent/613039/
summary.html.

Pais:1967:BRB

[Pai67a] A. Pais. Book review: Symmetries and Reflections. Scientific Essays. Eugene P. Wigner. Indiana University Press, Bloomington, 1967. 288 pp., illus. $7.50. Science, 157(3791):911–912, August 25, 1967. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/157/
3791/911.full.pdf.

Pais:1967:BRE

[Pai67b] A. Pais. Book review: Eugene P. Wigner, Physics, Life, and the Mind. Science, 157(3791):911–912, August 25, 1967. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.jstor.org/stable/1722102.

Pais:2000:GSP

[Pai00] Abraham Pais. The Genius of Science: a Portrait Gallery. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850614-7 (hardcover). 356 pp. LCCN Q141 .P29 2000.

URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
;   http://www.loc.gov/catdir/toc/fy02/99046603.html.

Palev:1982:WAQ

[Pal82] Tchavdar D. Palev. Wigner approach to quantization. Noncanonical quantization of two particles interacting via a harmonic potential. Journal of Mathematical Physics, 23(10):1778–1784, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v23/i10/p1778_s1.

Pallo:1988:FRC

[Pal88] Palló Gábor. A fizika érdekessége csökkent, mert túlságosan nagyra nött: Budapesti beszélgetés Wigner Jenövel. (Hungarian) [The interesting physics reduced because they are too big : Budapest conversation with Eugene Wigner]. Magyar Tudomány [Hungarian Science], 33(12):961–966, ???? 1988. CODEN ???? ISSN ????

Park:1965:BRW

[Par65] David Park. Book review: E. Wigner (editor), Dispersion Relations and Their Connection with Causality. American Journal of Physics, 33(5):418, May 1965. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.

aip.org/link/ajpias/v33/i5/p418/s2.

Pittner:2002:FRS

[PČH02] Jiří Pittner, Petr Čársky, and Ivan Hubač. Four- and 8-reference state-specific Brillouin–Wigner coupled-cluster method: Study of the singlet oxygen. International Journal of Quantum Chemistry, 90(3):1031–1037, ???? 2002. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pan:1998:CGRb

[PD98] Feng Pan and J. P. Draayer. Complementary group resolution of the SU (n) outer multiplicity problem. II. recoupling approach for SU (3) ⊃ U (2) reduced Wigner coefficients. Journal of Mathematical Physics, 39(10):5642–5662, October 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Primas:1977:CRW

[PE77] Hans Primas and Michael Esfeld. A critical review of Wigner’s work on the conceptual foundations of quantum theory. Not published because of the extravagant length of the review., September 1977. URL http://unil.academia.

edu/MichaelEsfeld/Papers/1012982/A_Critical_Review_of_
Wigners_Work_on_the_Conceptual_Foundations_of_Quantum_
Theory.

Pagaran:2006:MPC

[PFG06] J. Pagaran, S. Fritzsche, and G. Gaigalas. Maple procedures for the coupling of angular momenta. IX. Wigner D-functions and rotation matrices. Computer Physics Communications, 174(8):616– 630, April 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550600004X.

Palmer:1976:MCD

[PFW^{+}76] Thomas Y. Palmer, Carl E. Friedberg, Eugene P. Wigner, Arthur A. Broyles, Walter Murphey, and Sidney D. Drell. More on civil defense. Physics Today, 29(12):11, 13, 15, December 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v29/i12/p11/s1.

Palanichamy:2002:SSA

[PI02] R. Rajeswara Palanichamy and K. Iyakutti. Structural stability analysis of Wigner crystal with Gaussian and Yukawatype positive background. International Journal of Quantum Chemistry, 86(5):478–486, ???? 2002. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://

www3.interscience.wiley.com/cgi-bin/abstract/86511023/
START;   http://www3.interscience.wiley.com/cgi-bin/fulltext/
86511023/FILE?TPL=ftx_start;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=86511023&PLACEBO=IE.pdf.

Perry:1975:DBW

[PL75] W. L. Perry and C. D. Luning. On the density of the Breit– Wigner functions. Journal of Mathematical Physics, 16(8):1569– 1572, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i8/p1569_s1.

Placek:2024:TSW

[Pla24] Tomasz Placek. On testimony in scenarios with Wigner and Friend. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-

024-04754-1. See correction [Pla25].

Placek:2025:CTS

[Pla25] Tomasz Placek. Correction to: On testimony in scenarios with Wigner and Friend. Synthese, 205(2):??, February 2025. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-

024-04849-9. See [Pla24].

Papp:2007:MBB

[PMHW07] P. Papp, P. Mach, I. Hubač, and S. Wilson. Many-body Brillouin– Wigner second-order perturbation theory: a robust and efficient approach to the multireference correlation problem. International Journal of Quantum Chemistry, 107(14):2622–2631, ????

2007. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Poirier:1999:WWC

[Poi99] Bill Poirier. Wigner–Weyl correspondence and semiclassical quantization in spherical coordinates. Journal of Mathematical Physics, 40(12):6302–6318, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Porter:1965:STS

[Por65] Charles E. Porter, editor. Statistical theories of spectra: fluctuations: a collection of reprints and original papers, with an introductory review. Perspectives in physics. Academic Press, New York, USA, 1965. xv + 576 pp. LCCN QC173 .P572.

Poschadel:2000:NFS

[Pos00] Norbert Poschadel. A note on families of spin observables and a generalization of Wigner’s theorem in C^{2}. Journal of Mathematical Physics, 41(11):7832–7838, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [Pos01].

Poschadel:2001:ENF

[Pos01] Norbert Poschadel. Erratum: “A note on families of spin observables and a generalization of Wigner’s theorem in C^{2}”, [J. Math. Phys. 41, 7832 (2000)]. Journal of Mathematical Physics, 42(2): 978, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [Pos00].

Proietti:2019:ERO

[PPG^{+}19] Massimiliano Proietti, Alexander Pickston, Francesco Graffitti, Peter Barrow, Dmytro Kundys, Cyril Branciard, Martin Ringbauer, and Alessandro Fedrizzi. Experimental rejection of observer-independence in the quantum world. arXiv.org, ??(??):

1–10, February 13, 2019. URL https://arxiv.org/abs/1902.
05080.

Palev:1997:MBW

[PS97] T. D. Palev and N. I. Stoilova. Many-body Wigner quantum systems. Journal of Mathematical Physics, 38(5):2506–2523, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Puente:2002:WLD

[PS02] Antonio Puente and Llorenç Serra. Wigner localization and dynamics in two-electron semiconductor rings. International Journal of Quantum Chemistry, 86(1):27–34, ???? 2002. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/85008604/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext/85008604/FILE?TPL=ftx_start;
 http://www3.interscience.wiley.com/cgi-bin/fulltext?
ID=85008604&PLACEBO=IE.pdf.

Pulvirenti:2006:SEW

[Pul06] M. Pulvirenti. Semiclassical expansion of Wigner functions. Journal of Mathematical Physics, 47(5):052103, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i5/
p052103_s1.

Putnam:1964:CCC

[Put64] Hilary Putnam. Comments on comments on comments: A reply to Margenau and Wigner. Philosophy of Science, 31(1):1–6, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186740.

Polanyi:1925:BZM

[PW25] M. Polányi and E. Wigner. Bildung und Zerfall von Molekülen. (German) [Formation and decay of molecules]. Zeitschrift für Physik, 33(1):429–434, ???? 1925. CODEN ZEPYAA.

ISSN ???? URL http://www.springerlink.com/content/
r7316281773443v2.

Polanyi:1928:IEU

[PW28] M. Polányi and E. Wigner. Über die Interferenz von Eigenschwingungen als Ursache von Energieschwankungen und chemischer Umsetzungen. (German) [On the interference of vibration as the cause of energy fluctuations and chemical conversions]. Z. phys. Chem., Abt. A. Haber-Band, 139(??):439–452, ???? 1928. CODEN ???? ISSN ????

Pelzer:1932:GAR

[PW32] H. Pelzer and E. Wigner. Über die Geschwindigkeitskonstante von Austausch-Reaktionen (German) [On the rate constant of exchange reactions]. Zeitschrift für Physikalische Chemie, Abteilung B: Chemie der Elementarprozesse, Aufbau der Materie, 15(??): 445–453, ???? 1932. CODEN ZPCBAL. ISSN ????

Park:1964:CDPa

[PW64a] R. Park and E. Wigner. Civil defense: Project Harbor summary report. Report 1237, National Academy of Sciences of the United States of America, Washington, DC, USA, 1964.

Park:1964:CDPb

[PW64b] R. Park and E. Wigner. Civil defense: Project Harbor summary report. Report, United States Atomic Energy Commission, Oak Ridge, TN, USA, 1964.

Philips:1967:SSP

[PW67] T. O. Philips and Eugene Paul Wigner. De Sitter space and positive energy. ????, ??(??):??, 1967. CODEN ???? ISSN ????

Philips:1968:SSP

[PW68] T. O. Philips and Eugene P. Wigner. De Sitter space and positive energy. In Loebl [Loe75a], pages 631–676. ISBN 0-12-455150-5. LCCN ????

Quesne:1971:CTM

[QM71] C. Quesne and M. Moshinsky. Canonical transformations and matrix elements. Journal of Mathematical Physics, 12(8):1780– 1783, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i8/p1780_s1.

Quesne:1976:SEA

[Que76] C. Quesne. SU(2) × SU(2) scalars in the enveloping algebra of SU(4). Journal of Mathematical Physics, 17(8):1452–1467, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v17/i8/p1452_s1.

Ramond:1989:BRB

[Ram89] Pierre Ramond. Book review: Reminiscences about a Great Physicist: Paul Adrien Maurice Dirac, by Behram N. Kurşunoğlu and Eugene P. Wigner. American Scientist, 77(1):71–72, January 1989. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786

(electronic). URL http://www.jstor.org/stable/27855557.

Ramacher:2000:MLE

[Ram00] Pablo Ramacher. Modular localization of elementary systems in the theory of Wigner. Journal of Mathematical Physics, 41(9): 6079–6089, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rashid:1989:SPR

[Ras89] M. A. Rashid. A simple proof of the result that the Wigner transformation is of finite order. Journal of Mathematical Physics, 30 (9):1999–2000, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v30/i9/p1999_s1.

Rodenburg:1992:TSR

[RB92] J. M. Rodenburg and R. H. T. Bates. The theory of superresolution electron microscopy via Wigner-distribution deconvolution. Philosophical Transactions: Physical Sciences and Engineering, 339(1655):521–553, June 15, 1992. CODEN ???? ISSN

???? URL http://www.jstor.org/stable/53996.

Rowe:2001:ALW

[RdGS01] D. J. Rowe, H. de Guise, and B. C. Sanders. Asymptotic limits of SU(2) and SU(3) Wigner functions. Journal of Mathematical Physics, 42(5):2315–2342, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rajeswarapalanichamy:2003:FVN

[RI03] R. Rajeswarapalanichamy and K. Iyakutti. Ferromagnetic vs. nonmagnetic phases of 2-D Wigner electron crystal. International Journal of Quantum Chemistry, 92(3):337–343, April 5, 2003. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rajeswarapalanichamy:2005:APD

[RI05] R. Rajeswarapalanichamy and K. Iyakutti. Antiferromagnetic phase of a 2-D Wigner crystal. International Journal of Quantum Chemistry, 102(1):112–117, ???? 2005. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ringhofer:1992:SCT

[Rin92] Christian Ringhofer. A spectral collocation technique for the solution of the Wigner–Poisson problem. SIAM Journal on Numerical Analysis, 29(3):679–700, June 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

www.jstor.org/stable/2158273.

Rodriguez:1981:BWE

[RL81] M. A. Rodríguez and M. Lorente. Bargmann–Wigner equations: Symmetries of multispinors and equations of motion. Journal of Mathematical Physics, 22(6):1283–1288, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i6/
p1283_s1.

Rabson:2004:CPW

[RNM04] D. A. Rabson, B. N. Narozhny, and A. J. Millis. Crossover from Poisson to Wigner–Dyson level statistics in spin chains with integrability breaking. Physical Review B: Condensed Matter and Materials Physics, 69(??):054403, February 9, 2004. CODEN PRB-

MDO. ISSN 1098-0121. URL http://link.aps.org/doi/10.
1103/PhysRevB.69.054403.

Rno:1974:CGC

[Rno74] Jung Sik Rno. Clebsch–Gordan coefficients and special functions related to the Euclidean group in three space. Journal of Mathematical Physics, 15(12):2042–2047, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i12/p2042_s1.

Robinson:1993:SMT

[Rob93] Sam L. Robinson. Semiclassical mechanics for time-dependent Wigner functions. Journal of Mathematical Physics, 34(6):2185– 2205, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i6/p2185_s1.

Rogers:2013:NDY

[Rog13] J. D. Rogers. The neutron’s discovery — 80 years on. Physics Procedia, 43:1–9, 2013. CODEN PPHRCK. ISSN 1875-3892. URL

http://adsabs.harvard.edu/abs/2013PhPro..43....1R.

Romer:1991:EMP

[Rom91] Robert H. Romer. Editorial: Memorable papers from the American Journal of Physics, 1933–1990. American Journal of Physics, 59(3):201–207, March 1991. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Roman:2000:BRE

[Rom00] Paul Roman. Book review: Eugene Paul Wigner: The Collected Works of Eugene Paul Wigner. Arthur S. Wightman: The Scientific Papers. Volume 3, Part 1: Particles and Fields. The Scientific Papers. Volume 3, Part 2: Foundations of Quantum Mechanics. The Scientific Papers. Volume 4. Part 1: Physical Chemistry The Scientific Papers. Volume 4. Part 2: Solid State Physics. Jagdish Mehra: Historical, Philosophical, and Socio-Political Papers. Volume 6: Philosophical Reflections and Syntheses. Isis, 91(1):190–192, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237627.

Ross:1970:BRP

[Ros70] Marc Ross. Book review: E. P. Wigner, Survival and the Bomb: Methods of Civil Defense. Physics Today, 23(11):51, November 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v23/i11/p51/s1.

Rosen:1982:SPS

[Ros82] Joe Rosen, editor. Symmetry in physics: selected reprints. American Association of Physics Teachers, Stony Brook, NY, USA, 1982. ISBN ???? 153 pp. LCCN ????

Rosengren:1998:TSF

[Ros98] Hjalmar Rosengren. On the triple sum formula for Wigner 9jsymbols. Journal of Mathematical Physics, 39(12):6730–6744, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rosengren:1999:APT

[Ros99] Hjalmar Rosengren. Another proof of the triple sum formula for Wigner 9j-symbols. Journal of Mathematical Physics, 40(12): 6689–6691, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rowe:1996:RWA

[RR96] D. J. Rowe and J. Repka. The Racah–Wigner algebra and coherent tensors. Journal of Mathematical Physics, 37(5):2498–2509, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Recine:2005:NST

[RRC05] Greg Recine, Bernard Rosen, and Hong-Liang Cui. Numerical simulation of two-dimensional electron transport in cylindrical nanostructures using Wigner function methods. Journal of Computational Physics, 209(2):421–447, November 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999105001361.

Rowe:1999:RWG

[RSdG99] D. J. Rowe, B. C. Sanders, and H. de Guise. Representations of the Weyl group and Wigner functions for SU (3). Journal of Mathematical Physics, 40(7):3604–3615, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rubenstein:1982:MHR

[Rub82] Richard L. Rubenstein, editor. Modernization: the humanist response to its promise and problems: selected readings from the proceedings of the International Conferences on the Unity of the Sciences. Paragon House, Washington, DC, USA, 1982. ISBN 0-89226-015-7. LCCN CB478 .I57 1982. US$24.95.

Regniers:2010:WQS

[RV10] G. Regniers and J. Van der Jeugt. Wigner quantization of some one-dimensional Hamiltonians. Journal of Mathematical Physics, 51(12):123515, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v51/i12/p123515_s1.

Regniers:2011:AMD

[RV11] G. Regniers and J. Van der Jeugt. Angular momentum decomposition of the three-dimensional Wigner harmonic oscillator. Journal of Mathematical Physics, 52(11):113503, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v52/i11/p113503_s1.

Redlich:1954:DME

[RW54] Martin G. Redlich and Eugene P. Wigner. A β-decay matrix element for a deformed core model. Physical Review (2), 95(1):

122–126, July 1, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRev.95.122.

Rasch:2003:ESS

[RY03] J. Rasch and A. C. H. Yu. Efficient storage scheme for precalculated Wigner 3-j, 6-j and Gaunt coefficients. SIAM Journal on Scientific Computing, 25(4):1416–1428, July 2003. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL

http://epubs.siam.org/sam-bin/dbq/article/42293.

Sachs:1959:BRB

[Sac59] Robert G. Sachs. Book review: Group Theory. And Its Application to the Quantum Mechanics of Atomic Spectra, by Eugene P. Wigner and J. J. Griffin. Science, 130(3382):1106–1107, October 23, 1959. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1758126.

Salam:1969:CPT

[Sal69] Abdus Salam, editor. Contemporary physics: Trieste Symposium 1968: proceedings of the International Symposium on Contemporary Physics organized by and held at the International Centre for Theoretical Physics, Trieste, from 7 to 28 June 1968. International Atomic Energy Agency, Vienna, Austria, 1969. LCCN ???? Two volumes.

Sanger:1995:WBO

[San95] Stephen L. Sanger. Working on the Bomb: an Oral History of WWII Hanford. Portland State University, Continuing Education Press, Portland, OR, USA, 1995. ISBN 0-87678-115-6 (paperback). xiii + 264 pp. LCCN QC773.3.U5 S265 1995. US$18.00. Edited by Craig Wollner.

Sauberman:1975:LCL

[Sau75] Nat. H. Sauberman. Letters: Credibility lost. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Stefanou:1990:ENM

[SAZ90] N. Stefanou, H. Akai, and R. Zeller. An efficient numerical method to calculate shape truncation functions for Wigner–Seitz atomic polyhedra. Computer Physics Communications, 60(2):231–238, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090009P.

Seitz:1986:RCA

[SBR^{+}86] Frederick Seitz, Hans Bethe, Dixy Lee Ray, Miro M. Todorovich, Robert K. Adair, Bernard L. Cohen, Thomas J. Connolly, Herbert Goldstein, Herbert Kouts, Leon Lidofski, John McCarthy, Eugene Wigner, and Edwin L. Zebroski. Review of the Chernobyl accident, and comparison of Soviet with U.S. reactors, by a Select Panel of Scientists and Engineers for Secure Energy (SE_{2} ). Report, Select Panel for Post-Chernobyl Safety Review, 570 Seventh Avenue, Room 1007, New York, NY 10018, USA, August 29,

1986. 64–71 pp. URL http://www.energytruth.com/Strike_
Back_Or_Die.pdf.

Soto:1983:WWF

[SC83] Francisco Soto and Pierre Claverie. When is the Wigner function of multidimensional systems nonnegative? Journal of Mathematical Physics, 24(1):97–100, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i1/p97_s1.

Schwinger:1958:SPQ

[Sch58] Julian Schwinger, editor. Selected Papers on Quantum Electrodynamics. Dover books on engineering and engineering physics. Dover, New York, NY, USA, 1958. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35.

Schlegel:1967:BRB

[Sch67] Richard Schlegel. Book review: Symmetries and Reflections: Scientific Essays of Eugene P. Wigner edited by Walter J. Moore; Michael Scriven. Philosophy of Science, 34(4):383–386, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186127.

Schopf:1985:BRB

[Sch85] H.-G. Schöpf. Book review: Barut, A. O. /van der Merwe, A. /Vigier, J.-P. (eds.), Quantum, Space and Time — The Quest Continues. Studies and Essays in Honour of Louis de Broglie, Paul Dirac and Eugene Wigner. Cambridge et al., Cambridge University Press 1984. Vii, 662 S., £25.00 A P/ b. US $49.50. ISBN 0-521-31911-0. (Cambridge Monographs on Physics). Zeitschrift für Angewandte Mathematik und Mechanik, 65(10), 1985. CODEN ZAMMAX. ISSN 0044-2267 (print),

1521-4001 (electronic). URL http://onlinelibrary.wiley.
com/doi/10.1002/zamm.19850651014/abstract.

Schweber:1988:BRB

[Sch88] S. S. Schweber. Book review: Behram N. Kurşunoğlu and Eugene P. Wigner: Reminiscences about a Great Physicist: Paul Adrien Maurice Dirac. Isis, 79(2):356–357, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233680.

Schreckenberg:1996:SHQ

[Sch96] S. Schreckenberg. Symmetry and history quantum theory: an analog of Wigner’s theorem. Journal of Mathematical Physics, 37 (12):6086–6105, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schwinger:2003:SPQ

[Sch03] Julian Schwinger. Selected Papers on Quantum Electrodynamics. Dover, New York, NY, USA, 2003. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35. Reprint of [Sch58] with ISBN.

Simon:1999:CCF

[SCS99] R. Simon, S. Chaturvedi, and V. Srinivasan. Congruences and canonical forms for a positive matrix: Application to the Schweinler–Wigner extremum principle. Journal of Mathematical Physics, 40(7):3632–3642, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scully:2007:DQP

[Scu07a] Robert J. Scully. The demon and the quantum: from the Pythagorean mystics to Maxwell’s demon and quantum mystery. Wiley-VCH, Weinheim, Germany, 2007. ISBN 3-527-40688-3 (paperback). viii + 271 pp. LCCN QC174.12 .S43 2007. with endnotes by Marlan O. (Marlan Orvil) Scully.

Scully:2007:WFQ

[Scu07b] Robert J. Scully. From Wigner’s friend to quantum eraser: Of Wigner’s friends and their amnesia. In The demon and the quantum: from the Pythagorean mystics to Maxwell’s demon and quantum mystery [Scu07a], chapter 9, pages 121–136. ISBN 3-527-40688-3 (paperback). LCCN QC174.12 .S43 2007. with endnotes by Marlan O. (Marlan Orvil) Scully.

Schmider:1996:NSE

[SD96] Hartmut Schmider and Jens Peder Dahl. Nodal structure of the electronic Wigner function of many-electron atoms and molecules. International Journal of Quantum Chemistry, 60(1):439–452, October 5, 1996. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-

461X (electronic). URL http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=60607.

Secrest:1962:BRB

[Sec62] E. Leigh Secrest. Book review: Proceedings of the Eleventh Symposium in Applied Mathematics of the American Mathematical Society, edited by Birkhoff and Wigner. Mathematics Magazine, 35(5):307–308, November 1962. CODEN MAMGA8. ISSN 0025-

570X. URL http://www.jstor.org/stable/2688212.

Soto-Eguibar:1983:TEW

[SEC83] Francisco Soto-Eguibar and Pierre Claverie. Time evolution of the Wigner function. Journal of Mathematical Physics, 24(5):1104– 1109, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i5/p1104_s1.

Segre:1985:HPR

[Seg85] Emilio Segrè. Historical perspective: Refugee scientists and nuclear energy. Annals of the New York Academy of Sciences, 452(1):xv–xix, 1985. CODEN ANYAA9. ISBN 0-89766-298-9, 0-89766-299-7 (paperback). ISSN 0077-8923 (print), 1749-6632 (electronic). Sixth International Conference on Collective Phenomena: reports from the Moscow Refusnik Seminar / edited by Inga Fischer-Hjalmars and Joel L. Lebowitz. Contributions from the Moscow Refusnik Seminar and from two International Conferences on Collective Phenomena, one held in Stockholm, Sweden, 1–2 December 1983, and the other in Tel Aviv, Israel, 31 May–1 June 1984.

Segev:2003:FGR

[Seg03] Bilha Segev. Fermi’s golden rule in the Wigner representation. Journal of Optics B: Quantum and Semiclassical Optics, 5(3):S219–S220, 2003. CODEN JOBOFD. ISSN 1464-4266 (print), 1741-3575 (electronic). URL http://iopscience.iop.

org/1464-4266/5/3/373.

Seitz:1995:OEW

[Sei95] Frederick Seitz. Obituary: Eugene Wigner (1902–1995). Nature, 373(6512):288, January 26, 1995. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v373/n6512/pdf/373288a0.pdf.

Seth:1958:BRL

[Set58] Kamal K. Seth. Book review: Leonard Eisenbud and Eugene P. Wigner, Nuclear Structure. Physics Today, 11(12):54, December 1958. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v11/i12/p54/s2.

Schulten:1975:ERE

[SG75] Klaus Schulten and Roy G. Gordon. Exact recursive evaluation of 3j- and 6j-coefficients for quantum-mechanical coupling of angular momenta. Journal of Mathematical Physics, 16(10):1961– 1970, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i10/p1961_s1.

Shapiro:1973:POT

[Sha73] Joseph Y. Shapiro. Projection operator techniques for compact groups. Journal of Mathematical Physics, 14(9):1262–1270, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i9/p1262_s1.

Sharan:1992:SCW

[Sha92] Pankaj Sharan. Symbolic computation of Wigner–Kirkwood expansion to O(h̄8). Computer Physics Communications, 69 (2–3):235–242, March/April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290163S.

Sebasdiyar:1993:EPB

[SIM93] I. Sebasdiyar, K. Iyakutti, and M. Mahendran. Effect of positive background on the ground-state energy of a Wigner lattice. International Journal of Quantum Chemistry, 47(3):177–183, August 5, 1993. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stanek:2005:WFR

[SK05] Jerzy Stanek and Jerzy Konarski. Wigner function of the rotating Morse oscillator. International Journal of Quantum Chemistry, 103(1):10–18, ???? 2005. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Squire:2005:QPT

[SM05] Richard H. Squire and Norman H. March. Quantum phase transitions and superconductivity: Proposal for the phase diagram of alkali-doped fullerides and by universality HTSC including Wigner solids. International Journal of Quantum Chemistry, 105(6):883–897, ???? 2005. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Squire:2006:CIS

[SM06] Richard H. Squire and Norman H. March. Changing an insulating state with Wigner-like correlations into a superconductor with applications to doped fullerides. International Journal of Quantum Chemistry, 106(15):3343–3363, ???? 2006. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simos:2007:CMS

[SM07] Theodore E. Simos and George Maroulis, editors. Computation in Modern Science and Engineering: Proceedings of the International Conference on Computational Methods in Science and Engineering 2007 (ICCMSE 2007), Corfu, Greece, 25–30 September 2007, volume 2A. American Institute of Physics, Woodbury, NY, USA, 2007. ISBN 0-7354-0476-3 (set), 0-7354-0477-1 (vol. 1), 0-7354-0478-X (vol. 2). LCCN Q183.9 2007. Two volumes.

Smith:1975:LWN

[Smi75] Anthony Wayne Smith. Letters: We need restraint. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Sellier:2014:BSW

[SNDS14] Jean Michel Sellier, Mihail Nedjalkov, Ivan Dimov, and Siegfried Selberherr. A benchmark study of the Wigner Monte Carlo method. Monte Carlo Methods and Applications, 20(1):43–??, March 2014. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/

j/mcma.2014.20.issue-1/mcma-2013-0018/mcma-2013-0018.
xml.

Srinivas:1974:SMV

[SP74] M. D. Srinivas and Y. S. Prahalad. Statistical mechanics for velocity dependent interactions. Journal of Mathematical Physics, 15 (1):67–69, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i1/p67_s1.

So:1979:WC

[SS79] S. I. So and D. Strottman. Wigner coefficients for SU(6) ⊇ SU(3) ⊗ SU(2). Journal of Mathematical Physics, 20(1):153– 176, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i1/p153_s1.

Santos:2005:WMD

[SS05] J. P. Santos and L. O. Silva. Wigner–Moyal description of free variable mass Klein–Gordon fields. Journal of Mathematical Physics, 46(10):102901, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v46/i10/p102901_s1.

Schleich:1991:JSW

[SSP^{+}91] Wolfgang P. Schleich, Georg Süssmann, John Philpott, Malvin C. Teich, and Bahaa E. A. Saleh. A jump shot at the Wigner distribution. Physics Today, October 1, 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2810308.

Stanek:2010:WFR

[Sta10] Jerzy Stanek. The Wigner function of the rotating Kratzer oscillator mapped onto the Morse oscillator. International Journal of Quantum Chemistry, 110(9):1615–1621, August 5, 2010. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Steinruck:1991:ODW

[Ste91] H. Steinrück. The one-dimensional Wigner–Poisson problem and its relation to the Schrödinger–Poisson problem. SIAM Journal on Mathematical Analysis, 22(4):957–972, July 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Steane:2025:EWF

[Ste25] Andrew Steane. The extended Wigner’s friend, many-and single– worlds and reasoning from observation. Foundations of Physics, 55(2):??, April 2025. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-025-00831-8.

Strottman:1979:EWC

[Str79] D. Strottman. Evaluation of SU(6) ⊇ SU(3) ⊗ SU(2) Wigner coefficients. Journal of Mathematical Physics, 20(8):1643–1647, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v20/i8/p1643_s1.

Strickland:2011:WSC

[Str11] Jeffrey Strickland. Weird scientists — the creators of quantum physics. Lulu.com, ????, 2011. ISBN 1-257-97624-9. LCCN ????

Stuewer:1979:NPR

[Stu79] Roger H. Stuewer, editor. Nuclear physics in retrospect: proceedings of a symposium on the 1930s. University of Minnesota Press, Minneapolis, MN, USA, 1979. ISBN 0-8166-0869-5. LCCN QC773 .S95 1977.

Stuewer:1988:BRB

[Stu88] Roger H. Stuewer. Book review: Behram N. Kurşunoğlu, Eugene P. Wigner, and John G. Taylor, Reminiscences about a Great Physicist: Paul Adrien Maurice Dirac and Tributes to Paul Dirac. Physics Today, 41(2):84, February 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/phtoad/v41/i2/p84/s2.

Stumpf:2003:SPP

[Stu03] H. Stumpf. Symmetry properties of photon eigenstates of generalized de Broglie–Bargmann–Wigner equations. Annales de la Fondation Louis de Broglie, 28(1):65–76, 2003. ISSN 0182-4295 (print), 2108-6397 (electronic).

Stumpf:2004:GBB

[Stu04] H. Stumpf. Generalized de Broglie–Bargmann–Wigner equations, a modern formulation of de Broglie’s fusion theory. Annales de la Fondation Louis de Broglie, 29(??):785–806, 2004. CODEN ????

ISSN 0182-4295 (print), 2108-6397 (electronic). URL http://

aflb.ensmp.fr/AFLB-295/aflb295m307.htm.

Stuewer:2005:GHL

[Stu05] Roger H. Stuewer. George de Hevesy/Loránd Eötvös: Scientist– teacher/Eugene P. Wigner and his Hungarian teachers. Physics in Perspective (PIP), 7(4):498–499, December 2005. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Stuewer:2018:AIN

[Stu18] Roger H. Stuewer. The Age of Innocence: Nuclear Physics Between the First and Second World Wars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-186658-X, 0-19-882787-3 (hardback), 0-19-256290-8 (e-book). xv + 484 pp. LCCN QC773 .S78 2018.

Saleker:1986:QLM

[SV86] G. Saleker and E. Vigner. Quantum limitations of the measurement of space-time distances. In Einstein collection, 1982–1983 (Russian), pages 285–301. Nauka, Moscow, Russia, 1986. Translated from the English by Yu. A. Danilov.

Seitz:1998:EPW

[SVW98] Frederick Seitz, Erich Vogt, and Alvin M. Weinberg. Eugene Paul Wigner. 17 November 1902–1 January 1995. Biographical memoirs — National Academy of Sciences of the United States of America, 74:1–26, 1998. CODEN BMNSAC. ISSN

???? URL http://books.nap.edu/html/biomems/ewigner.

pdf. Reprinted in [SVW00].

Seitz:2000:EPW

[SVW00] Frederick Seitz, Erich Vogt, and Alvin M. Weinberg. Eugene Paul Wigner. 17 November 1902–1 January 1995. Biographical Memoirs of Fellows of the Royal Society, 46(2): 577–592, November 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL http://rsbm.

royalsocietypublishing.org/content/46/577.abstract;
http://www.jstor.org/stable/770417;   https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0102. Reprint of [SVW98].

Szegvari:1923:EES

[SW23] A. Szegvari and E. Wigner. Über elektrische Erscheinungen bei Stäbchensolen. (German) [On electrical phenomena in brine sticks]. Kolloid-Zeitschrift & Zeitschrift für Polymere, 33(4):218– 222, ???? 1923. CODEN CPMSB6. ISSN 0303-402X (print),

1435-1536 (electronic). URL http://www.springerlink.com/
content/h7166xnv1uj81w18.

Seitz:1956:GCD

[SW56a] F. Seitz and E. Wigner. On the Geneva Conference: a dissenting opinion. Bulletin of the Atomic Scientists, 12(1):23–24, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seitz:1956:ERS

[SW56b] Frederick Seitz and Eugene P. Wigner. The effects of radiation on solids. Scientific American, 195(2):76–86, August 1956. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/
v195/n2/pdf/scientificamerican0856-76.pdf.

Salecker:1958:QLM

[SW58] H. Salecker and E. P. Wigner. Quantum limitations of the measurement of space–time distances. Physical Review (2), 109(2): 571–577, January 15, 1958. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.109.571.

Seitz:1959:GPA

[SW59] Frederick Seitz and Eugene P. Wigner. Geneva, 1958. pure and applied nuclear physics in East and West. Bulletin of the Atomic Scientists, 15(3):127–131, March 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Soodak:1961:FNR

[SW61] Harry Soodak and Eugene P. Wigner. Fast neutron reactor. US Patent 2,993,850., July 25, 1961. URL http://www.google.com/ patents/US2993850. US Patent Application filed Jun 14, 1948.

Saenz:1967:UMH

[SW67] A. W. Sáenz and E. P. Wigner. Unimodular matrices homomorphic to Lorentz transformations in n ≥ 2 spacelike dimensions. Zeitschrift für Naturforschung, 22a(??):1293–1299, ???? 1967. CODEN ZNTFA2. ISSN 0372-9516. Jordan Festschrift.

Schweinler:1970:OM

[SW70] H. C. Schweinler and E. P. Wigner. Orthogonalization methods. Journal of Mathematical Physics, 11(5):1693–1694, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v11/i5/p1693_s1;   http://link.aip.org/link/
jmapaq/v11/i5/p1693/s1.

Salam:1972:AQT

[SW72] Abdus Salam and Eugene Paul Wigner, editors. Aspects of quantum theory. Cambridge University Press, Cambridge, UK, 1972. ISBN 0-521-08600-0. xvi + 268 pp. LCCN QC174.1

.A85 1972. URL http://hooke.lib.cam.ac.uk/cgi-bin/
bib_seek.cgi?cat=ul&bib=1733506;   http://www.loc.gov/
catdir/enhancements/fy1001/72075298-d.html;   http://
www.loc.gov/catdir/enhancements/fy1001/72075298-t.html.

Salam:1989:TEP

[SW89] Abdus Salam and Victor F. Weisskopf. Tribute to Eugene Paul Wigner. Nuclear Physics B, Proceedings Supplements, 6(1):3– 6, March 1989. CODEN NPBSE7. ISSN 0920-5632 (print),

1873-3832 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0920563289904003.

Szilard:1959:JFE

[SWC59] Leo Szilard, Eugene P. Wigner, and Edward C. Creutz. Jacketed fuel elements for graphite moderated reactors. US Patent 2,886,503., May 12, 1959. URL http://www.google.com/ patents/US2886503. US Patent Application filed Feb 20, 1946, serial number 649,080.

Szanton:1992:VEW

[Sza92] A. Szanton. Vignettes — Eugene Wigner. Science, 258(5089): 1821, December 11, 1992. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Szmytkowski:1998:OFW

[Szm98] Radosław Szmytkowski. Operator formulation of Wigner’s Rmatrix theories for the Schrödinger and Dirac equations. Journal of Mathematical Physics, 39(10):5231–5252, October 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [Szm99].

Szmytkowski:1999:EOF

[Szm99] Radosław Szmytkowski. Erratum: “Operator formulation of Wigner’s R-matrix theories for the Schrödinger and Dirac equations”, [J. Math. Phys. 39, 5231 (1998)]. Journal of Mathematical Physics, 40(8):4181, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [Szm98].

Talman:1968:SFG

[Tal68] James D. Talman. Special Functions: a Group Theoretic Approach Based on Lectures by Eugene P. Wigner. The mathematical physics monograph series. W. A. Benjamin, Inc., New York, NY, USA, 1968. xii + 260 pp. LCCN ???? With an introduction by Eugene P. Wigner.

Taub:1961:JNCa

[Tau61a] A. H. Taub, editor. John von Neumann: Collected Works: Volume I: Logic, Theory of Sets and Quantum Mechanics. Pergamon Press, New York, NY, USA, 1961. x + 654 pp. LCCN ???? See also volumes II–VI [Tau61b, Tau63a, Tau62, Tau63b, Tau63c].

Taub:1961:JNCb

[Tau61b] A. H. Taub, editor. John von Neumann: Collected Works. Volume II: Operators, Ergodic Theory and Almost Periodic Functions in a Group. Pergamon Press, New York, NY, USA, 1961. x + 568 pp. LCCN ???? See also volumes I, III–VI [Tau61a, Tau63a, Tau62, Tau63b, Tau63c].

Taub:1962:JNC

[Tau62] A. H. Taub, editor. John von Neumann: Collected Works. Volume IV: Continuous Geometry and Other Topics. Pergamon Press, New York, NY, USA, 1962. x + 516 pp. LCCN ???? See also volumes I–III, V–VI [Tau61a, Tau61b, Tau63a, Tau63b, Tau63c].

Taub:1961:JNCc

[Tau63a] A. H. Taub, editor. John von Neumann: Collected Works. Volume III: Rings of Operators. Pergamon Press, New York, NY, USA, 1961–1963. ix + 574 pp. LCCN ???? See also volumes I–II, IV–VI [Tau61a, Tau61b, Tau62, Tau63b, Tau63c].

Taub:1963:JNCa

[Tau63b] A. H. Taub, editor. John von Neumann: Collected Works. Volume V: Design of Computers, Theory of Automata and Numerical Analysis. Pergamon Press, New York, NY, USA, 1963. ix + 784 pp. LCCN ???? See also volumes I–IV, VI [Tau61a, Tau61b, Tau63a, Tau62, Tau63c].

Taub:1963:JNCb

[Tau63c] A. H. Taub, editor. John von Neumann: Collected Works. Volume VI: Theory of Games, Astrophysics, Hydrodynamics and Meteorology. Pergamon Press, New York, NY, USA, 1963. x + 538 pp. LCCN ???? See also volumes I–V [Tau61a, Tau61b, Tau63a, Tau62, Tau63b].

Taylor:1987:TPD

[Tay87] John Gerald Taylor, editor. Tributes to Paul Dirac. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987. US$10.00. URL http://www.loc.gov/

catdir/enhancements/fy0745/87153334-d.html. Based on the

papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Tuan:2010:AWB

[TC10] Do Van Tuan and Ui-Pil Chong. Audio watermarking based on advanced Wigner distribution and important frequency peaks. The International Journal of High Performance Computing Applications, 24(2):154–163, May 2010. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://hpc.sagepub.

com/content/24/2/154.full.pdf+html.

Tosiek:2016:WKB

[TCT16] J. Tosiek, R. Cordero, and F. J. Turrubiates. The Wentzel– Kramers–Brillouin approximation method applied to the Wigner function. Journal of Mathematical Physics, 57(6):062103, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Telling:2003:WDB

[TEEBH03] Rob H. Telling, Chris P. Ewels, Ahlam A. El-Barbary, and Malcolm I. Heggie. Wigner defects bridge the graphite gap. Nature Materials, 2(5):333–337, April 13, 2003. CODEN NMAACR. ISSN 1476-1122 (print), 1476-4660 (electronic). URL http://

www.nature.com/nmat/journal/v2/n5/full/nmat876.html.

Teller:1995:EPW

[Tel95] E. Teller. Eugene Paul Wigner, 1902–1995. Physics World, 8(2): 62, February 1995. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

Temme:2007:LRW

[Tem07] F. P. Temme. Limitations of Racah–Wigner algebra in S_{n}- invariant theoretic modelling of NMR spin dynamics via dual tensorial sets. In Simos and Maroulis [SM07], pages 264–268. ISBN 0-7354-0476-3 (set), 0-7354-0477-1 (vol. 1), 0-7354-0478-X (vol. 2). LCCN Q183.9 2007. URL http://proceedings.

aip.org/getpdf/servlet/GetPDFServlet?filetype=pdf&amp;
 id=APCPCS000963000002000264000001&amp; idtype=cvips.

Two volumes.

Tesche:1986:SCR

[Tes86] C. D. Tesche. Schrödinger’s cat: A realization in superconducting devices. Annals of the New York Academy of Sciences, 480(??): 36–50, December 30, 1986. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic).

Tisza:1998:REM

[Tis98] Laszlo Tisza. The reasonable effectiveness of mathematics in the natural sciences. In Cohen et al. [CHS97], pages 213–238. ISBN 0-7923-4452-9 (volume 1), 0-7923-4454-5 (set). ISSN 0068-0346. LCCN Q174 .B67 vol. 193.

URL http://www.lehmanns.ch/shop/naturwissenschaften/
422840-9780792344520-experimental-metaphysics;   http:
//www.loc.gov/catdir/enhancements/fy0822/97002399-d.html;
 http://www.loc.gov/catdir/enhancements/fy0822/97002399-
t.html.

Tisza:2003:REW

[Tis03] Laszlo Tisza. Remembering Eugene Wigner and pondering his legacy. Europhysics News, 34(2):58–61, March/April 2003. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (elec-

tronic). URL http://www.europhysicsnews.org/articles/
epn/abs/2003/02/epn03205/epn03205.html.

Toscano:2008:HWR

[TKC^{+}08] Fabricio Toscano, Anatole Kenfack, Andre R. R. Carvalho, Jan M. Rost, and Alfredo M. Ozorio de Almeida. Husimi–Wigner representation of chaotic eigenstates. Proceedings: Mathematical, Physical and Engineering Sciences, 464(2094):1503–1524, June 8, 2008. CODEN ???? ISSN ???? URL http://www.jstor.org/

stable/20209482.

Teberekidis:2005:GSB

[TKT^{+}05] Vasilios I. Teberekidis, Ioannis S. K. Kerkines, Constantinos A. Tsipis, Petr Čársky, and Aristides Mavridis. Ground states of BeC and MgC: a comparative multireference Brillouin–Wigner coupled cluster and configuration interaction study. International Journal of Quantum Chemistry, 102(5):762–774, ???? 2005. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tennyson:1984:RPD

[TN84] Jonathan Tennyson and Cliff J. Noble. RESON — a program for the detection and fitting of Breit–Wigner resonances. Computer Physics Communications, 33(4):421–424, October 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465584901474.

Torretti:1987:BRA

[Tor87] Roberto Torretti. Book review: Asim O. Barut, Alwyn van der Merwe, and Jean-Pierre Vigier, editors, Quantum, Space, and Time — The Quest Continues, Studies and Essays in Honour of Louis de Broglie, Paul Dirac, and Eugene Wigner. Noûs, 21 (3):442–444, September 1987. CODEN ???? ISSN 0029-4624 (print), 1468-0068 (electronic). URL http://www.jstor.org/

stable/2215196.

Takagaki:2001:PWD

[TP01] Y. Takagaki and K. H. Ploog. Poisson and Wigner–Dyson distributions of conductance fluctuations in quantum cavities. Physical Review B: Condensed Matter and Materials Physics, 64(??): 245336, December 10, 2001. CODEN PRBMDO. ISSN 1098-

0121. URL http://link.aps.org/doi/10.1103/PhysRevB.64.
245336.

True:1951:PS

[Tru51] Webster Prentiss True, editor. Panorama of science. Smithsonian Series Publishers, New York, NY, USA, 1951. ???? pp. LCCN Q9 .P285.

Teichmann:1952:SRD

[TW52] T. Teichmann and E. P. Wigner. Sum rules in the dispersion theory of nuclear reactions. Physical Review (2), 87(1):123–135, July 1, 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.87.123.

Teichmann:1953:EFE

[TW53] T. Teichmann and E. P. Wigner. Electromagnetic field expansions in loss-free cavities excited through holes. Journal of Applied Physics, 24:262–267, 1953. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. URL http://link.

aip.org/link/japiau/v24/i3/p262/s1.

Teller:1979:ANB

[TW79a] Edward Teller and Eugene P. Wigner. America needs a better civil defense program. New York Times, ??(??):??, July 22, 1979. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1979:EHE

[TW79b] Edward Teller and Eugene Paul Wigner. Energy from Heaven and Earth: in which a story is told about energy from its origins 15,000,000,000 years ago to its present adolescence — turbulent, hopeful, beset by problems, and in need of help. W. H. Freeman and Company, New York, NY, USA, 1979. ISBN 0-7167-1063-3, 0-7167-1064-1 (paperback). xiv + 322 pp. LCCN TJ163.2 .T4.

Teller:1973:CDWa

[TWB73a] E. Teller, E. Wigner, and A. Broyles. Without civil defense we are in a glass house. Technical report, Washington Report of the American Security Council, Washington, DC, USA, May 1973. ???? pp.

Teller:1973:CDWb

[TWB73b] E. Teller, E. Wigner, and A. Broyles. Without civil defense we are in a glass house. Survive, 6(5):1–2, ???? 1973. CODEN ???? ISSN 0039-6354.

Ufford:1942:CDF

[UW42] C. W. Ufford and E. P. Wigner. On the calculation of the distribution function. Physical Review (2), 61(7–8):524–527, April 1, 1942. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.61.524.

VanDam:1994:BRB

[Van94] Hendrik Van Dam. Book review: ???? American Scientist, 82 (2):183–184, ???? 1994. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.jstor.org/

stable/29775160.

Vassiliadis:1989:WLG

[Vas89] Dimitris V. Vassiliadis. Wigner’s little group and decomposition of Lorentz transformations. Journal of Mathematical Physics, 30 (9):2177–2180, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v30/i9/p2177_s1.

Vasudevan:2001:TFA

[VC01] Kris Vasudevan and Frederick A. Cook. Time-frequency analysis of deep crustal reflection seismic data using Wigner–Ville distributions. Canadian Journal of Earth Sciences = Journal canadien des sciences de la terre, 38(7):1027–1035, 2001. CODEN CJE-SAP. ISSN 0008-4077 (print), 1480-3313 (electronic). URL http:

//www.nrcresearchpress.com/doi/abs/10.1139/e01-003.

Veblen:1960:OVP

[Veb60] Oswald Veblen. Oswald Veblen papers, 1881–1960 (bulk 1920– 1960). US Library of Congress archival manuscript mate-

rial (collection)., 1960. URL http://hdl.loc.gov/loc.mss/
eadmss.ms011049;   http://hdl.loc.gov/loc.mss/eadmss.
ms011049.3.

Varilly:1987:WTF

[VGB87] Joseph C. Várilly and José M. Gracia-Bondía. The Wigner transformation is of finite order. Journal of Mathematical Physics, 28 (10):2390–2392, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v28/i10/p2390_s1.

Vigner:1958:RIK

[Vig58] E. Vigner. Reljativistskaja invariantnost’ i kvantovye javlenija (Russian) [Relativistic invariance and quantum phenomena]. Uspekhi Fizicheskikh Nauk, 67(6):257–281, June 1958. CODEN UF-NAAG. ISSN ???? Russian translation of [Wig57f].

Vigner:1964:SZS

[Vig64] E. Vigner. Simmetrija i zakony sohranenija. (Russian) [Symmetry and conservation laws]. Uspekhi Fizicheskikh Nauk, 83(8): 729–739, August 1964. CODEN UFNAAG. ISSN ???? URL

http://ufn.ru/ru/articles/1964/8/f/. Russian translation of

[Wig64s].

Vigner:1965:SZP

[Vig65] E. Vigner. Sobytija, zakony prirody i principy invariantnosti. (Russian) [The events of the laws of nature, and invariance principles]. Uspekhi Fizicheskikh Nauk, 85(4):727–736, April 1965. CO-

DEN UFNAAG. ISSN ???? URL http://ufn.ru/ru/articles/

1965/4/g/. Russian translation of [Wig64p].

Vigner:1966:NSV

[Vig66] E. Vigner. Narushenie simmetrii v fizike. (Russian) [Symmetry breaking in physics]. Uspekhi Fizicheskikh Nauk, 89(7):453–466, July 1966. CODEN UFNAAG. ISSN ???? URL http://ufn. ru/ru/articles/1966/7/e/. Russian translation of [Wig65t].

Vigner:1968:NJM

[Vig68] E. Vigner. Nepostizhimaja jeffektivnost’ matematiki v estestvennyh naukah (Fizika Nashih Dnej). (Russian) [Unreasonable effectiveness of mathematics in the natural sciences (Physics Today)]. Uspekhi Fizicheskikh Nauk, 94(3):535–546, March 1968. CODEN

UFNAAG. ISSN ???? URL http://ufn.ru/ru/articles/1968/

3/f/. Russian translation of [Wig60j].

Vigner:1986:TYK

[Vig86] E. Vigner. Thirty years of knowing Einstein [Some strangeness in the proportion (Princeton, N.J., 1979), 461–472, Addison-Wesley, Reading, Mass., 1980]. In Einstein collection, 1982–1983 (Russian), pages 149–169. Nauka, Moscow, Russia, 1986. Translated from the English by Yu. A. Danilov.

Villee:1985:ICU

[Vil85] Claude Alvin Villee, editor. Absolute values and the new cultural revolution: selected proceedings of the Thirteenth International Conference on the Unity of the Sciences: September 2–5, 1984, Washington, DC, Unity of knowledge series. International Cultural Foundation, New York, NY, USA, 1985. ISBN 0-89226-040-8. LCCN AC5 .I696 1984.

Villars:1986:PSC

[Vil86] C. N. Villars. The paradox of Schrödinger’s cat. Physics Education, 21(4):232–237, July 1986. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic).

vonNeumann:1996:PJN

[vN96] John von Neumann. Papers of John von Neumann, 1912–1996 (bulk 1935–1957). US Library of Congress archival manuscript material (collection)., 1996. 11,660 items. 34 containers plus 1 vault container. 13.4 linear feet. Manuscript number MSS44180. Correspondence, memoranda, journals, speeches, article and book drafts, notes, charts, graphs, patent, biographical material, family papers, printed materials, newspaper clippings, photographs, and other materials pertaining primarily to von Neumann’s career as professor of mathematics at the Institute for Advanced Study including his directorship of the Electronic Computer Project; adviser and commissioner on the U.S. Atomic Energy Commission; scientific consultant to government and private concerns, including the Los Alamos Scientific Laboratory, Los Alamos, New Mexico, and the U.S. Army Ballistic Research Laboratory, Aberdeen, Maryland; and author of works on ballistic research, computers, continuous geometries, logic, operator theory, quantum mechanics, and the theory of games. Includes evaluations of his work written after his death by colleagues including Herman Heine Goldstine, Paul R. Halmos, and Abraham H. Taub. Of special interest are an Albert Einstein letter and report on theoretical physics (1937). Also includes a small amount of material pertaining to Eva and Peter Aldor. Correspondents include Eva Aldor, Frank Aydelotte, Hans Albrecht Bethe, Garrett Birkhoff, S. Chandrasekhar, George Bernard Dantzig, P. A. M. Dirac, Carl Eckart, Enrico Fermi, Abraham Flexner, George Gamow, Kurt Gödel, Herman Heine Goldstine, Werner Heisenberg, L. van Hove, Cuthbert Corwin Hurd, Pascual Jordan, R. H. Kent, George B. Kistiakowsky, Oskar Morgenstern, J. Robert Oppenheimer, Rudolf Ortvay, Wolfgang Pauli, Marshall H. Stone, Lewis L. Strauss, Abraham Haskel Taub, Edward Teller, Stanislaw M. Ulam, Oswald Veblen, Klara Dan Von Neumann, Warren Weaver, Hermann Weyl, Norbert Wiener, and Eugene Paul Wigner. Gift, Marina Von Neumann Whitman, 1974–1975. Gift, Nicholas A. Vonneuman, 1993.

vonNeumann:1928:EEEa

[vNW28a] John von Neumann and Eugene Wigner. Zur Erklärung einiger Eigenschaften der Spektren aus der Quantenmechanik des Drehelektrons. (German) [On the explanation of some of the eigenvalue spectra from the quantum mechanics of electron spin]. Zeitschrift für Physik, 47(3–4):203–220, ???? 1928. CODEN

ZEPYAA. ISSN ???? URL http://www.springerlink.com/

content/n814671704213174. Reprinted in [Tau61a, Paper 18].

vonNeumann:1928:EEEb

[vNW28b] John von Neumann and Eugene Wigner. Zur Erklärung einiger Eigenschaften der Spektren aus der Quantenmechanik des Drehelektrons. (German) [On the explanation of some of the eigenvalue spectra from the quantum mechanics of electron spin]. Zeitschrift für Physik, 49(1–2):73–94, ???? 1928. CODEN

ZEPYAA. ISSN ???? URL http://www.springerlink.com/

content/q26q292617797417. Reprinted in [Tau61a, Paper 19].

vonNeumann:1928:EEEc

[vNW28c] John von Neumann and Eugene Wigner. Zur Erklärung einiger Eigenschaften der Spektren aus der Quantenmechanik des Drehelektrons. (German) [On the explanation of some of the eigenvalue spectra from the quantum mechanics of electron spin]. Zeitschrift für Physik, 51(11–12):845–888, ???? 1928. CODEN

ZEPYAA. ISSN ???? URL http://www.springerlink.com/

content/j7802882351u2312. Reprinted in [Tau61a, Paper 20].

vonNeumann:1929:VEA

[vNW29a] John von Neumann and Eugene Wigner. Über das Verhalten von Eigenwerten bei adiabatischen Prozessen. (German) [On the behavior of the eigenvalues of adiabatic processes]. Physikalische Zeitschrift, 30(15):467–470, ???? 1929. CODEN PHZTAO. ISSN ???? Reprinted in [Tau61a, Paper 24].

vonNeumann:1929:MDE

[vNW29b] John von Neumann and Eugene Wigner. Über merkwürdige diskrete Eigenwerte. (German) [On unusual discrete eigenvalues]. Physikalische Zeitschrift, 30(15):465–467, ???? 1929. CODEN PHZTAO. ISSN 0369-982X. Reprinted in [Tau61a, Paper 23].

vonNeumann:1940:MAP

[vNW40] John von Neumann and Eugene P. Wigner. Minimally almost periodic groups. Annals of Mathematics (2), 41(4):746–750, October 1940. CODEN ANMAAH. ISSN 0003-486X (print),

1939-8980 (electronic). URL http://www.jstor.org/stable/

1968853. Reprinted in [Tau62, Paper 21].

Vogt:1995:WEP

[Vog95] E. Vogt. Eugene Paul Wigner — a towering figure of modern physics. Physics Today, 48(12):40–44, December 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

vonNeumann:1996:JNP

[von96] John von Neumann. John von Neumann papers, 1912–1996 (bulk 1935–1957). US Library of Congress archival manuscript material (collection)., 1996. URL http://hdl.loc.gov/loc.mss/

eadmss.ms996003;   http://hdl.loc.gov/loc.mss/eadmss.
ms996003.3.

Voss:1993:RWE

[Vos93] D. Voss. The recollections of Eugene P. Wigner. New York Times Book Review, ??(??):25, January 24, 1993. ISSN 0028-7806.

VanDam:1965:CRM

[VW65] H. Van Dam and E. P. Wigner. Classical relativistic mechanics of interacting point particles. Physical Review (2), 138(6B): B1576–B1582, June 21, 1965. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.138.B1576.

VanDam:1966:IAC

[VW66] H. Van Dam and E. P. Wigner. Instantaneous and asymptotic conservation laws for classical relativistic mechanics of interacting point particles. Physical Review (2), 142(4):838–843, February 25, 1966. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.142.838.

Wigner:1975:RDS

[W^{+}75] Eugene P. Wigner et al. Roundtable discussion of the significance and accomplishments of the conference, Washington, DC, March 7, 1975. In ????, editor, [unknown], volume 425, pages 964–965. National Bureau of Standards, Gaithersburg, MD, USA, 1975. ISBN ???? LCCN ????

Wigner:1968:TNK

[WÁ68] Wigner Jenő and Ákos Károly. A tudomány növekedése — kedvező kilátások és várható veszélyek. (Hungarian) [The growth of science and a positive outlook: expected hazards]. Magyar Tudomány [Hungarian Science], 75(5):304–318, ???? 1968. CODEN ???? ISSN ????

Wigner:1972:LEM

[WA72] Eugene P. Wigner and Robert K. Adair. Letter to the Editor: Military research and development. Science, 175(4020):356–357, January 28, 1972. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/content/
175/4020/356.2.full.pdf.

Wagner:1981:EPW

[Wag81] Francis S. Wagner. Eugene P. Wigner: an architect of the Atomic Age: highlights of a career with a comprehensive bibliography, volume 1 of Rákóczi Foundation bio-bibliographies. Rákóczi Foundation, Toronto, ON, Canada, 1981. ISBN 0-919545-00-9 (paperback). 80 pp. LCCN QC16.W52 W33. With the research assistance of Christina Maria T. Wagner-Jones. Epilogue by Edward Teller.

Wagner:1998:WJP

[Wag98] Wagner Ferenc. Wigner Jenő Pál: az atomkor egyik megalapítója: egy karrier fénypontjai, teljes bibliográfiával. (Hungarian) [Eugene Paul Wigner: one of the founders of the nuclear age: highlights of a career full bibliography], volume 4 of Studia physica Savariensia. BDTF, Szombathely, Hungary, 1998. ISBN 963-9017-36-1. ISSN 1219-2678. 78 pp. LCCN ???? With the assistance of Christina Maria T. Wagner-Jones. Epilog by Edward Teller.

Walsh:1973:CEW

[Wal73] John Walsh. A conversation with Eugene Wigner. Science, 181(4099):527–533, August 10, 1973. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1736495;   http://www.sciencemag.org/
content/181/4099/527.full.pdf.

Wattenberg:1993:BNA

[Wat93] Albert Wattenberg. The birth of the nuclear age. Physics Today, 46(1):44–51, January 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://physicstoday.

scitation.org/doi/pdf/10.1063/1.881378.

Wigner:1935:TWF

[WB35] E. Wigner and J. Bardeen. Theory of the work functions of monovalent metals. Physical Review (2), 48(1):84–87, July 1, 1935. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://link.aps.org/doi/10.1103/PhysRev.48.84.

Wigner:1936:RSL

[WB36] E. Wigner and G. Breit. The β-ray spectrum of Li^{8}. Physical Review (2), 50(12):1191–1911, December 15, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.50.1191.3.

Wigner:1977:WHD

[WB77] Eugene P. Wigner and A. A. Broyles. We heartily disagree. Journal of Civil Defense, 10(??):4–8, ???? 1977. CODEN ???? ISSN 0740-5537.

Wood:2005:BIW

[WB05] J. G. Wood and A. J. Bracken. Bounds on integrals of the Wigner function: The hyperbolic case. Journal of Mathematical Physics, 46(4):042103, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i4/p042103_s1.

Wigner:1930:BBL

[WBBE30] E. Wigner, J. Bartels, A. Betz, and S. Erk. Besprechungen: de Broglie, Louis, Einführung in die Wellenmechanik (German) [Book reviews: de Broglie, Louis, Introduction to Wave Mechanics]. Naturwissenschaften, 18(14):307–310, ???? 1930. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://www.springerlink.com/content/m5071k1380584921/.

Wigner:1960:CNR

[WC60] Eugene P. Wigner and Edward C. Creutz. Cooled neutronic reactor. US Patent 2,928,781., March 15, 1960. URL http://www.

google.com/patents/US2928781. US Patent Application filed

Jun 12, 1952.

Wigner:1955:RAN

[WCJS55] E. P. Wigner, E. Creutz, H. Jupnik, and T. Snyder. Resonance absorption of neutrons by spheres. Journal of Applied Physics, 26(3):260–270, March 1, 1955. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. URL http:

//link.aip.org/link/japiau/v26/i3/p260/s1. See erratum

[WCJS56].

Wigner:1956:ERA

[WCJS56] E. P. Wigner, E. Creutz, H. Jupnik, and T. Snyder. Errata: Resonance absorption of neutrons by spheres. Journal of Applied Physics, 27:839–??, 1956. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. URL http://link.

aip.org/link/japiau/v27/i7/p839/s1. See [WCJS55].

Wigner:1939:EPF

[WCT39] Eugene P. Wigner, Charles L. Critchfield, and Edward Teller. The electron–positron field theory of nuclear forces. Physical Review (2), 56(6):530–539, September 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v56/i6/p530_1.

Wigner:1937:RCR

[WE37] Eugene Wigner and Henry Eyring. On the rate of chemical reactions. The Scientific Monthly, 44(6):564–567, June 1937. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL

http://www.jstor.org/stable/16206.

Wigner:1939:SFD

[WE39] E. Wigner and L. Eisenbud. On the saturation of forces derived from meson theory: a abstract. Physical Review (2), 56(2):214, July 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v56/i2/p206_1.

Wigner:1947:HAM

[WE47a] E. P. Wigner and L. Eisenbud. Higher angular momenta and long range interaction in resonance reactions. Physical Review (2), 72(1):29–41, July 1, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.72.29.

Wigner:1947:MFR

[WE47b] E. P. Wigner and L. Eisenbud. A mathematical foundation of the resonance theory. MDDC 744. United States Atomic Energy Commission, Oak Ridge, TN, USA, 1947. 4 pp.

Weber:1973:RWS

[Web73] Robert L. Weber, editor. A Random Walk In Science. The Institute of Physics, Bristol and London, 1973. ISBN 0-85498-027-X. xvii + 206 pp. LCCN Q167.W42. Compiled by R. L. Weber, edited by E. Mendoza, with a foreword by William Cooper.

Weinberg:1969:BRB

[Wei69a] Steven Weinberg. Book review: Survival and the Bomb. Methods of Civil Defense. Eugene P. Wigner, Ed., Indiana University Press, Bloomington, 1969. x + 310 pp., illus. $7.50. Science, 166(3909):1131–1132, November 28, 1969. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.sciencemag.org/content/166/3909/1131.full.pdf.

Weinberg:1969:BRP

[Wei69b] Steven Weinberg. Book review: Protection: Survival and the Bomb. Methods of Civil Defense, by Eugene P. Wigner. Science, 166(3909):1131–1132, November 28, 1969. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.jstor.org/stable/1727463.

Weisskopf:1975:LCE

[Wei75] Victor Weisskopf. Letters: A clerical error. Bulletin of the Atomic Scientists, 31(4):3, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reports incorrect inclusion of Weisskopf’s name in [BAA^{+}75].

Weinberg:1997:EPW

[Wei97] Alvin M. Weinberg. Eugene Paul Wigner (17 November 1902–1 January 1995). Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 141(3): 376–381, September 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor.org/

stable/987207.

Weinberg:2002:EWN

[Wei02] A. M. Weinberg. Eugene Wigner, nuclear engineer. Physics Today, 55(10):42–46, October 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Weinberg:2004:NHR

[Wei04] Alvin Weinberg. Notes on Hanford reactor start-up. American Journal of Physics, 57(4):18–19, April 2004. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http:

//scitation.aip.org/content/aip/magazine/physicstoday/
article/57/4/10.1063/1.4796463. Reply to [Wei04].

Wenzel:1998:EEB

[Wen98] W. Wenzel. Excitation energies in Brillouin–Wigner-based multireference perturbation theory. International Journal of Quantum Chemistry, 70(4–5):613–622, ???? 1998. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://

www3.interscience.wiley.com/cgi-bin/abstract?ID=75039;
 http://www3.interscience.wiley.com/cgi-bin/fulltext?
ID=75039&PLACEBO=IE.pdf.

Weststeijn:2021:WFR

[Wes21] Nikki Weststeijn. Wigner’s friend and relational quantum mechanics: a reply to Laudisa. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00487-0.

Wigner:1941:SPN

[WF41] Eugene P. Wigner and Eugene Feenberg. Symmetry properties of nuclear levels. Reports on Progress in Physics, 8:274–317, 1941. CODEN RPPHAG. ISSN 0034-4885 (print), 1361-6633

(electronic). URL http://iopscience.iop.org/0034-4885/8/
1/313.

Wigner:1968:CCW

[WF68] Eugene P. Wigner and Bernard T. Feld. Current comments: Wigner and Feld. Bulletin of the Atomic Scientists, 24(2):29, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel67, BJH68].

Wald:1969:TAD

[WFS^{+}69] George Wald, George W. Frimpter, Stephen E. Szasz, Harry L. Berman, Paul B. Porter, Theodor D. Sterling, Eugene P. Wigner, and W. Allen Wallis. Today’s army: Drafted or professional? Science, 163(3874):1396–1400, March 28, 1969. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.jstor.org/stable/1726119.

Wigner:1972:WSC

[WG72] Eugene P. Wigner and J. Gailar. Will Soviet civil defense undermine SALT? Human Events, 32(??):9–??, July 8, 1972. CODEN ???? ISSN ????

Wigner:1974:CDSa

[WG74a] Eugene P. Wigner and J. S. Gailar. Civil defense in the Soviet Union. Survive, 7(6):2–??, ???? 1974. CODEN ???? ISSN 0039-6354.

Wigner:1974:CDSb

[WG74b] Eugene P. Wigner and J. S. Gailar. Civil defense in the Soviet Union. Foresight, 1(3):9–11, ???? 1974. CODEN ???? ISSN ????

Wigner:1974:REP

[WG74c] Eugene P. Wigner and J. S. Gailar. Russian evacuation plans — the fears they create. Survive, 7(5):4–5, ???? 1974. CODEN ???? ISSN 0039-6354.

Wigner:1996:SWJ

[WG96] E. P. Wigner and H. H. Goldstine. The scientific work of John von Neumann. In Wigner [Wig96m], pages 123–126. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1935:PMM

[WH35] E. Wigner and H. B. Huntington. On the possibility of a metallic modification of hydrogen. Journal of Chemical Physics, 3 (??):764–770, ???? 1935. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://link.aip.org/

link/jcpsa6/v3/i12/p764/s1.

Wigner:1976:SNW

[WH76] Eugene P. Wigner and C. M. Haaland. Surviving a nuclear war. National Review, 28(5):1005–??, ???? 1976. CODEN ???? ISSN 0028-0038.

Wigner:1977:MPM

[WH77a] E. P. Wigner and R. A. Hodgkin. Michael Polányi. 12 March 1891–22 February 1976. Biographical Memoirs of Fellows of the Royal Society, 23(2):413–448, November 1977. CODEN BM-FRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL

http://www.jstor.org/stable/769621.

Wigner:1977:CDN

[WH77b] Eugene P. Wigner and Carsten M. Haaland. Civil defense and nuclear blackmail. Science, 196(4292):912–913, May 20, 1977. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1744075;
 http://www.sciencemag.org/content/196/4292/912.full.
pdf.

Wheeler:2009:MF

[Whe09] John A. Wheeler. Mechanism of fission. Physics Today, 62(4): 35–38, April 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Whitaker:1996:EBQ

[Whi96] Andrew Whitaker. Einstein, Bohr, and the quantum dilemma. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-48220-8 (hardcover), 0-521-48428-6 (paperback). xvii + 349 pp.

LCCN QC174.12 .W48 1996. URL ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;   http://www.cambridge.org/
uk/catalogue/catalogue.asp?isbn=0521484286;   http://
www.loc.gov/catdir/description/cam027/95018270.html;
http://www.loc.gov/catdir/toc/cam021/95018270.html.

Wilson:1998:NWL

[WI98] Leslie C. Wilson and Stanislav Ivanov. A new Wigner-like correlation-energy functional from coordinate scaling requirements. International Journal of Quantum Chemistry, 69(4):523– 532, ???? 1998. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-

461X (electronic). URL http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=30025;   http://www3.interscience.

wiley.com/cgi-bin/fulltext?ID=30025&PLACEBO=IE.pdf. Spe-

cial Issue: Symposium on Density Functional Theory and Applications (Part II of II). Issue Edited by Sam Trickey, Weitao Yang, Mel Levy.

Wiener:1953:RPS

[Wie53] Philip P. (Philip Paul) Wiener, editor. Readings in philosophy of science; introduction to the foundations and cultural aspects of the sciences. Scribner, New York, NY, USA, 1953. ix + 645 pp. LCCN Q175 .W6519.

Wigner:1920:EES

[Wig20] Eugene P. Wigner. Über die elastischen Eigenschwingungen symmetrischen Systeme. (German) [On elastic oscillations of symmetric systems]. Nachrichten der Ges. der Wissenschaften zu Göttingen. Mathematisch-Physikalische Klasse, ??(??):133–146, ???? 1920. CODEN ???? ISSN ????

Wigner:1925:BZM

[Wig25] Eugen Paul Wigner. Bildung und Zerfall von Molekülen, Statistische Mechanik und Reaktionsgeschwindigkeit. (German) [Formation and decay of molecules, statistical mechanics and reaction rate]. Dr.-Ing., Technische Universität, Berlin, Germany, 1925.

Wigner:1926:KTNa

[Wig26a] E. Wigner. Über nicht kombinierende Terme in der neueren Quantentheorie: Erster Teil. (German) [On noncombining terms in modern quantum theory: Part One]. Zeitschrift für Physik, 40 (7):492–500, ???? 1926. CODEN ZEPYAA. ISSN ???? URL

http://www.springerlink.com/content/t72k522376w51532.

Wigner:1926:KTNb

[Wig26b] E. Wigner. Über nicht kombinierende Terme in der neueren Quantentheorie. II. Teil. (German) [On noncombining terms in modern quantum theory: Part Two]. Zeitschrift für Physik, 40(11–12): 883–892, ???? 1926. CODEN ZEPYAA. ISSN ???? URL http:

//www.springerlink.com/content/u867j2643g460g78.

Wigner:1927:EFSa

[Wig27a] E. Wigner. Einige Folgerungen aus der Schrödingerschen Theorie für die Termstrukturen. (German) [Some implications of the Schrödinger theory of term structures]. Zeitschrift für Physik, 43 (9–10):624–652, ???? 1927. CODEN ZEPYAA. ISSN ???? URL

http://www.springerlink.com/content/w781433358348155.

Wigner:1927:EFSb

[Wig27b] E. Wigner. Einige Folgerungen aus der Schrödingerschen Theorie für die Termstrukturen. (German) [Some implications of the Schrödinger theory of term structures]. Zeitschrift für Physik, 45 (7–8):601–602, ???? 1927. CODEN ZEPYAA. ISSN ???? URL

http://www.springerlink.com/content/h34u54l439122380.

Wigner:1927:EQG

[Wig27c] E. Wigner. Über die Erhaltungssätze in der Quantenmechanik. (German) [On conservation in quantum mechanics]. Nachrichten der Ges. der Wissenschaften zu Göttingen. Mathematisch-Physikalische Klasse, ??(??):375–381, ???? 1927. CODEN ???? ISSN ????

Wigner:1929:BEN

[Wig29a] E. Wigner. Eine Bemerkung zu Einsteins neuer Formulierung des allgemeinen Relativitätsprinzips. (German) [A note on Einstein’s new formulation of the Principle of General Relativity]. Zeitschrift für Physik, 53(7–8):592–596, ???? 1929. CODEN

ZEPYAA. ISSN ???? URL http://www.springerlink.com/
content/u1lm233463601018.

Wigner:1929:KKJ

[Wig29b] Wigner Jenő. A kémiai kötés újabb elmélete. (Hungarian) [The new theory of the chemical bond]. Chem. Rundschau, 6(3):24–27, ???? 1929. CODEN ???? ISSN ????

Wigner:1929:ORS

[Wig29c] Wigner Jenő. Összetett rendszerek statisztikája az új quantummechanika szerint. (Hungarian) [Complex systems in the new statistics of quantum mechanics]. Matematikai és Természettudomanyi Értes’itő, 46(??):576–583, ???? 1929. CODEN MNAUAD. ISSN ????

Wigner:1931:VSG

[Wig31a] E. Wigner. Über eine Verschärfung des Summensatzes. (German) [On a tightening of the sum set]. Physikalische Z., 32(11):450–453, ???? 1931. CODEN ???? ISSN ????

Wigner:1931:GIA

[Wig31b] Eugen Wigner. Gruppentheorie und ihre Anwendung auf die Quantenmechanik der Atomspektren. (German) [Group Theory and its Application to the Quantum Mechanics of Atomic Spectra], volume 85 of Die Wissenschaft; Sammlung von Einzeldarstellungen . . .. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1931. ISBN 3-528-08353-0. viii + 332 pp. LCCN QA171 .W65.

Wigner:1932:QCT

[Wig32a] E. Wigner. On the quantum correction for thermodynamic equilibrium. Physical Review (2), 40(5):749–759, June 1, 1932. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://link.aps.org/doi/10.1103/PhysRev.40.749.

Wigner:1932:UPC

[Wig32b] E. P. Wigner. Über das Überschreiten von Potentialschwellen bei chemischen Reaktionen. (German) [On the crossing of thresholds in chemical potential reactions]. Zeitschrift für Physikalische Chemie, Abteilung B: Chemie der Elementarprozesse, Aufbau der Materie, 19(??):203–216, ???? 1932. CODEN ZPCBAL. ISSN ????

Wigner:1932:OZQ

[Wig32c] E. P. Wigner. Über die Operation der Zeitumkehr in der Quantenmechanik. (German) [On the action of time reversal in quantum mechanics]. Nachrichten der Ges. Wiss. zu Göttingen. Mathematisch-Physikalische Klasse, ??(??):546–559, ???? 1932. CODEN ???? ISSN ????

Wigner:1933:MDH

[Wig33a] E. Wigner. On the mass defect of helium. Physical Review (2), 43 (4):252–257, February 15, 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.43.252. Reprinted in [Wig65a].

Wigner:1933:SNP

[Wig33b] E. Wigner. Über die Streuung von Neutronen an Protonen. (German) [On the scattering of neutrons by protons]. Zeitschrift für Physik, 83(3–4):253–258, ???? 1933. CODEN ZEPYAA.

ISSN ???? URL http://www.springerlink.com/content/

g8q0342686060553. Reprinted in [Wig65b].

Wigner:1933:PUP

[Wig33c] E. P. Wigner. Über die paramagnetische Umwandlung von Para-Orthowasserstoff. (German) [On the paramagnetic transition of para-ortho hydrogen]. Zeitschrift für Physikalische Chemie, Abteilung B: Chemie der Elementarprozesse, Aufbau der Materie, 23(??):28–32, ???? 1933. CODEN ZPCBAL. ISSN ????

Wigner:1933:ANE

[Wig33d] Wigner Jenő. Adalékok a neutron elméletéhez. (Hungarian) [Contributions to the theory of neutron]. Matematikai és Természettudomanyi Értes’itő, 39(??):142–146, ???? 1933. CO-DEN MNAUAD. ISSN ????

Wigner:1934:IEM

[Wig34] E. Wigner. On the interaction of electrons in metals. Physical Review (2), 46(11):1002–1011, December 1, 1934. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.46.1002.

Wigner:1935:U

[Wig35a] E. P. Wigner. [unknown]. Bulletin of the American Mathematical Society, 41(??):306–??, ???? 1935. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Wigner:1935:MRS

[Wig35b] Wigner Jenő. On a modification of the Rayleigh–Schrödinger perturbation theory. Matematikai és Természettudomanyi Értes’itő, 53(??):477–482, ???? 1935. CODEN MNAUAD. ISSN ????

Wigner:1935:RSF

[Wig35c] Wigner Jenő. A Rayleigh-Schrödinger-féle perturbáció-elmélet módosításáról. (Hungarian) [The Rayleigh–Schrödinger perturbation theory’s modification]. Matematikai és Természettudomanyi Értes’itő, 53(??):474–475, ???? 1935. CODEN MNAUAD. ISSN ????

Wigner:1936:CRE

[Wig36a] E. Wigner. On the constant A in Richardson’s equation. Physical Review (2), 49(9):696–700, May 1, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRev.49.696.

Wigner:1936:SEF

[Wig36b] E. Wigner. On the saturation of exchange forces. Proceedings of the National Academy of Sciences of the United States of America, 22(11):662–666, November 15, 1936. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.

jstor.org/stable/86634.

Wigner:1936:SSB

[Wig36c] Eugen Wigner. On the structure of solid bodies. The Scientific Monthly, 42(1):40–46, January 1936. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://

www.jstor.org/stable/16056.

Wigner:1937:CRE

[Wig37a] E. Wigner. Calculation of the rate of elementary association reactions. Journal of Chemical Physics, 5(9):720–725, September 1, 1937. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690

(electronic). URL http://link.aip.org/link/jcpsa6/v5/i9/
p720/s1.

Wigner:1937:CSN

[Wig37b] E. Wigner. On the consequences of the symmetry of the nuclear Hamiltonian on the spectroscopy of nuclei. Physical Review (2), 51(2):106–119, January 15, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.51.106.

Wigner:1937:SNB

[Wig37c] E. Wigner. On the structure of nuclei beyond oxygen. Physical Review (2), 51(11):947–958, June 1, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

link.aps.org/doi/10.1103/PhysRev.51.947.

Wigner:1938:EEI

[Wig38a] E. Wigner. Effects of the electron interaction on the energy levels of electrons in metals. Transactions of the Faraday Society, 34: 678–685, 1938. CODEN TFSOA4. ISSN 0014-7672.

Wigner:1938:TSM

[Wig38b] E. Wigner. The transition state method. Transactions of the Faraday Society, 34:29–41, 1938. CODEN TFSOA4. ISSN 0014-

Wigner:1939:CCL

[Wig39a] E. P. Wigner. On coupling conditions in light nuclei and the lifetimes of β-radioactivities. Physical Review (2), 56(6):519– 527, September 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRev.56.519.

Wigner:1939:URI

[Wig39b] E. P. Wigner. On unitary representations of the inhomogeneous Lorentz group. Annals of Mathematics (2), 40(1):149–204, January 1939. CODEN ANMAAH. ISSN 0003-486X (print),

1939-8980 (electronic). URL http://www.jstor.org/stable/

1968551. Reprinted in [Wig89a].

Wigner:1939:SRT

[Wig39c] Eugene P. Wigner. Some remarks on the theory of reaction rates. Journal of Chemical Physics, 7(??):646–652, ???? 1939. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL

http://link.aip.org/link/jcpsa6/v7/i8/p646/s1.

Wigner:1941:NMB

[Wig41a] Eugene P. Wigner. Nuclear masses and binding energies. In Fermi et al. [FBR^{+}41], pages 27–38.

Wigner:1941:RCF

[Wig41b] Eugene P. Wigner. On representations of certain finite groups. American Journal of Mathematics, 63(1):57–63, January 1941. CODEN AJMAAN. ISSN 0002-9327 (print), 1080-6377 (elec-

tronic). URL http://www.jstor.org/stable/2371276.

Wigner:1943:RCW

[Wig43] Eugene P. Wigner. Radioactivity of the cooling water. Report CP-499, United States Department of Energy, Washington, DC, USA, March 1, 1943. 10 pp. URL http://www.osti.gov/

accomplishments/documents/fullText/ACC0143.pdf.

Wigner:1944:GIA

[Wig44] Eugen Wigner. Gruppentheorie und ihre Anwendung auf die Quantenmechanik der Atomspektren. (German) [Group theory and its application to the quantum mechanics of atomic spectra]. J. W. Edwards, Ann Arbor, MI, USA, 1944. viii + 332 pp. LCCN QA171 .W66.

Wigner:1946:RSC

[Wig46a] E. P. Wigner. Reaction and scattering cross-sections. Proceedings of the National Academy of Sciences of the United States of America, 32(12):302–306, December 15, 1946. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://

www.jstor.org/stable/87869.

Wigner:1946:TPM

[Wig46b] E. P. Wigner. Theoretical physics in the Metallurgical Laboratory of Chicago. Journal of Applied Physics, 17(??):857–836, ???? 1946. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (elec-

tronic), 1520-8850. URL http://link.aip.org/link/japiau/
v17/i11/p857/s1.

Wigner:1946:RR

[Wig46c] Eugene P. Wigner. Resonance reactions. Physical Review (2), 70(9–10):606–608, November 1, 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRev.70.606.

Wigner:1946:RRI

[Wig46d] Eugene P. Wigner. Resonance reactions. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 90(1):25–29, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://

www.jstor.org/stable/3301035.

Wigner:1946:RRA

[Wig46e] Eugene P. Wigner. Resonance reactions and anomalous scattering. Physical Review (2), 70(1–2):15–33, July 1, 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.70.15.

Wigner:1946:RAA

[Wig46f] Eugene P. Wigner. Roots of the atomic age. In Masters and Way [MW46], chapter 3, pages 11–15. LCCN UF767 .M3 1946a. Foreword by Niels Bohr. Introduction by Arthur H. Compton. See also reprint [MW07].

Wigner:1947:F

[Wig47a] E. P. Wigner. Foreword. In Darrow [Dar47], page ?? LCCN Q171 .P947 1946a. Foreword by E. P. Wigner.

Wigner:1947:RWE

[Wig47b] E. P. Wigner. Relativistic wave equations [abstract]. Science, 106(2761):506–??, November 28, 1947. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/106/2761/505.full.pdf.

Wigner:1947:RWG

[Wig47c] E. P. Wigner. Relativistische Wellengleichungen. (German) [Relativistic wave mechanics]. Zeitschrift für Physik, 124(7–12):665– 684, ???? 1947. CODEN ZEPYAA. ISSN ???? URL http://

www.springerlink.com/content/g6731t22263772u8.

Wigner:1948:AEc

[Wig48a] E. P. Wigner. Atomic energy. Research: a journal of science and its applications, 1(13):577–580, October 1948.

Wigner:1948:RWG

[Wig48b] E. P. Wigner. Relativistische Wellengleichungen. (German) [Relativistic wave equations]. Zeitschrift für Physik, 124:665–684, 1948. CODEN ZEPYAA. ISSN 0170-9739 (print), 1431-5858 (electronic).

Wigner:1948:AEa

[Wig48c] Eugene P. Wigner. Atomic energy. Science, 108(2811):517–521, November 12, 1948. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
1677610;   http://www.sciencemag.org/content/108/2811/
517.full.pdf.

Wigner:1948:AEb

[Wig48d] Eugene P. Wigner. Atomic energy. Pegasus, 12(5):??, ???? 1948. CODEN ???? ISSN ????

Wigner:1948:CAE

[Wig48e] Eugene P. Wigner. Commentary : Atomic energy. Research, 1 (??):577–??, ???? 1948. CODEN ???? ISSN ????

Wigner:1948:BCS

[Wig48f] Eugene P. Wigner. On the behavior of cross sections near thresholds. Physical Review (2), 73(9):1002–1009, May 1, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.73.1002.

Wigner:1948:STP

[Wig48g] Eugene P. Wigner. Second thoughts to the problem of nuclear energy. Science, 108(??):??, November ??, 1948. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Wigner:1949:AES

[Wig49a] E. P. Wigner. Albert Einstein seventy. Physics Today, 2(7):24–??, July 1949. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v2/i7/
p24/s1.

Wigner:1949:BRM

[Wig49b] E. P. Wigner. Book review: E. A. Milne, Kinematic relativity: a sequel to Relativity, Gravitation, and World Structure. Science, 110(2849):149–150, August 5, 1949. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.jstor.org/stable/1676545.

Wigner:1949:NRL

[Wig49c] E. P. Wigner. Nuclear reactions and level widths. American Journal of Physics, 17(3):99–109, March 1949. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

link.aip.org/link/ajpias/v17/i3/p99/s1.

Wigner:1949:IPT

[Wig49d] Eugene P. Wigner. Invariance in physical theory. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 93(7):521–526, December 30, 1949. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL http://www.jstor.org/stable/3143140. Reprinted in

[Wig82h].

Wigner:1949:RMB

[Wig49e] Eugene P. Wigner. Review of E. A. Milne: Kinematic relativity; a sequel to relativity, gravitation, and world structure. Science, 110(2849):149–150, August 5, 1949. CODEN SCIEAS. ISSN

???? URL http://www.sciencemag.org/content/110/2849/
149.3.extract.

Wigner:1950:SRI

[Wig50a] E. P. Wigner. Some remarks on the infinite de Sitter space. Proceedings of the National Academy of Sciences of the United States of America, 36(3):184–188, March 15, 1950. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://

www.jstor.org/stable/88185.

Wigner:1950:BRM

[Wig50b] Eugene P. Wigner. Book review: Max von Laue and Ralph E. Oesper, History of Physics. Physics Today, 3(11):32, November 1950. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v3/i11/
p32/s1.

Wigner:1950:DEM

[Wig50c] Eugene P. Wigner. Do the equations of motion determine the quantum mechanical commutation relations? Physical Review (2), 77(5):711–712, March 1, 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRev.77.711.

Wigner:1950:LS

[Wig50d] Eugene P. Wigner. The limits of science. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 94(5):422–427, October 19, 1950. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143609. Reprinted in Ger-

man in Comm. Friends of Student Body, Univ. of Heidelberg, 36, 7–??.

Wigner:1950:RMM

[Wig50e] Eugene P. Wigner. The role of mathematical methods in physical theories. Journal of The Franklin Institute, 250(6):477–479, December 1950. CODEN JFINAB. ISSN 0016-0032 (print),

1879-2693 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0016003250903103.

Wigner:1951:IOC

[Wig51a] E. P. Wigner. The International Oxford Conference on Nuclear Physics. Science, 113(2926):107–108, January 26, 1951. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Wigner:1951:IDA

[Wig51b] Eugene P. Wigner. Impact of the developments in atomic energy on the sciences. Bulletin of the Atomic Scientists, 7(3):66–69, 80, March 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wigner:1951:IPT

[Wig51c] Eugene P. Wigner. Invarianz in der physikalischen Theorie. (German) [Invariance in physical theory]. Physikalische Blätter, 7(10): 433–441, October 1951. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19510071001/abstract.

Wigner:1951:LS

[Wig51d] Eugene P. Wigner. The limits of science. In True [Tru51], pages 403–?? LCCN Q9 .P285.

Wigner:1951:NNI

[Wig51e] Eugene P. Wigner. News and notes: The International Oxford Conference on Nuclear Physics. Science, 113(2926):107–108, January 26, 1951. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/content/
113/2926/107.extract.

Wigner:1951:CAF

[Wig51f] Eugene P. Wigner. On a class of analytic functions from the quantum theory of collisions. Annals of Mathematics (2), 53 (1):36–67, January 1951. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL http://www.jstor.org/

stable/1969342.

Wigner:1951:SDW

[Wig51g] Eugene P. Wigner. On the statistical distribution of the widths and spacings of nuclear resonance levels. Mathematical proceedings of the Cambridge Philosophical Society, 47(4):790–798, ???? 1951. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Wigner:1952:MQO

[Wig52a] E. P. Wigner. Die Messung quantenmechanischer Operatoren. (German) [The measurement of quantum-mechanical operators]. Zeitschrift für Physik, 133(1–2):101–108, ???? 1952. CODEN ZEPYAA. ISSN 0170-9739 (print), 1431-5858 (electronic). URL

http://www.springerlink.com/content/ug64182071489377.

Wigner:1952:CBD

[Wig52b] E. P. Wigner. On the connection between the distribution of poles and residues for an R function and its invariant derivative. Annals of Mathematics (2), 55(1):7–18, January 1952. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL

http://www.jstor.org/stable/1969415.

Wigner:1952:LCH

[Wig52c] E. P. Wigner. On the law of conservation of heavy particles. Proceedings of the National Academy of Sciences of the United States of America, 38(5):449–451, May 15, 1952. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://

www.jstor.org/stable/88602.

Wigner:1952:SDP

[Wig52d] E. P. Wigner. Simplified derivation of the properties of elementary transcendentals. American Mathematical Monthly, 59(10): 669–683, December 1952. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2307544.

Wigner:1952:DMSa

[Wig52e] Eugene P. Wigner. Derivative matrix and scattering matrix. Revista Mexicana de Física, 1:81–90, 1952. CODEN RMXFAT. ISSN 0035-001X.

Wigner:1952:DMSb

[Wig52f] Eugene P. Wigner. Derivative matrix and scattering matrix. Revista Mexicana de Física, 1:91–101, 1952. CODEN RMXFAT. ISSN 0035-001X.

Wigner:1952:SMN

[Wig52g] Eugene P. Wigner. On the shell model for nuclei. In Farkas and Wigner [FW52], pages 45–61. LCCN QD455 .F3.

Wigner:1953:RBN

[Wig53a] E. Wigner. Review of Theoretical nuclear physics by John M. Blatt, V. F. Weisskopf. New York: Wiley, 1952. American Scientist, 41(2):310–314, April 1953. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.jstor.

org/stable/27826494.

Wigner:1953:LS

[Wig53b] Eugene P. Wigner. The limits of science. In Wiener [Wie53], pages 757–?? LCCN Q175 .W6519.

Wigner:1954:PMS

[Wig54a] E. P. Wigner. The problem of multiple scattering. Physical Review (2), 94(1):17–25, April 1, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.94.17.

Wigner:1954:ARS

[Wig54b] Eugene P. Wigner. Application of the Rayleigh–Schrödinger perturbation theory to the hydrogen atom. Physical Review (2), 94(1):77–78, April 1, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.94.77.

Wigner:1954:CLC

[Wig54c] Eugene P. Wigner. Conservation laws in classical and quantum physics. Progress of Theoretical Physics, 11:437–440, 1954. CO-DEN PTPKAV. ISSN 0033-068X (print), 1347-4081 (electronic).

Wigner:1954:DOR

[Wig54d] Eugene P. Wigner. Derivations of Onsager’s reciprocal relations. Journal of Chemical Physics, 22(11):1912–1915, November 1, 1954. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-

7690 (electronic). URL http://link.aip.org/link/jcpsa6/
v22/i11/p1912/s1.

Wigner:1954:ICT

[Wig54e] Eugene P. Wigner. Intermediate coupling theory for nuclear reactions: Abstract. In Anonymous, editor, Autumn Meeting, National Academy of Sciences, November 8–10, 1954, pages 1–?? Columbia University Press, New York, NY, USA, 1954. LCCN ????

Wigner:1954:IRC

[Wig54f] Eugene P. Wigner. The interpretation of Racah’s coefficients. In Anonymous [Ano54a], pages 279–?? LCCN ????

Wigner:1954:NBD

[Wig54g] Eugene P. Wigner. Note on the beta-decay. In Anonymous [Ano54b], pages 365–369. LCCN ????

Wigner:1954:KDL

[Wig54h] Eugene P. Wigner. On kinematic and dynamic laws of symmetry. In Anonymous [Ano54a], pages 199–200. LCCN ????

Wigner:1954:QTC

[Wig54i] Eugene P. Wigner. Qualitative theory of the cohesion in metals. In Anonymous [Ano54b], pages 649–663. LCCN ????

Wigner:1955:CVB

[Wig55a] Eugene P. Wigner. Characteristic vectors of bordered matrices with infinite dimensions. Annals of Mathematics (2), 62(3): 548–564, November 1955. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL http://www.jstor.org/

stable/1970079.

Wigner:1955:EF

[Wig55b] Eugene P. Wigner. Enrico Fermi, 1901–1954. In ????, editor, Yearbook of the American Philosophical Society, pages 435–439. American Philosophical Society, Philadelphia, PA, USA, 1955. ISSN 0065-9762. LCCN ???? URL http://www.aps-pub.com/.

Wigner:1955:LLE

[Wig55c] Eugene P. Wigner. Lower limit for the energy derivative of the scattering phase shift. Physical Review (2), 98(1):145–147, April 1, 1955. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.98.145.

Wigner:1955:DCN

[Wig55d] Eugene P. Wigner. On the development of the compound nucleus model. American Journal of Physics, 23(6):371–380, June 1955. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/ajpias/v23/i6/p371/s1.

Wigner:1956:RIQa

[Wig56a] E. P. Wigner. Relativistic invariance in quantum mechanics. Il Nuovo Cimento (10), 3(3):517–532, ???? 1956. CODEN NU-CIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL

http://www.springerlink.com/content/d738l3035285327g.

Wigner:1956:RIQc

[Wig56b] E. P. Wigner. Relativistic invariance in quantum mechanics. Il Nuovo Cimento (10), 4(2):690, ???? 1956. CODEN NUCIAD.

ISSN ???? URL http://www.springerlink.com/content/
f7413431135kw683.

Wigner:1956:BRB

[Wig56c] Eugene P. Wigner. Book review: Handbuch der Physik. vol. 1, Mathematical Methods by S. Flügge. Science, 124(3233):1214, December 14, 1956. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
1752855.

Wigner:1956:UCN

[Wig56d] Eugene P. Wigner. E. U. Condon: New Editor of Reviews of Modern Physics. Physics Today, 9(11):30–36, November 1956. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://link.aip.org/link/phtoad/v9/i11/p30/s1.

Wigner:1956:MPM

[Wig56e] Eugene P. Wigner. Must the photon mass be zero?: Discussion of paper by L. Bass and E. Schrödinger. Il Nuovo Cimento (10), Supplement 10(4):825–826, ???? 1956. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Wigner:1956:RIQb

[Wig56f] Eugene P. Wigner. Relativistic invariance of quantum-mechanical equations. Helvetica Physica Acta, Supplement 4:210–212, 1956. CODEN HPACAK. ISSN 0018-0238.

Wigner:1957:AMN

[Wig57a] Eugene P. Wigner. Apparatus for the measurement of neutron absorption. US Patent 2,781,307., February 12, 1957. URL http:

//www.google.com/patents/US2781307. US Patent Application

filed Mar 4, 1947.

Wigner:1957:CVB

[Wig57b] Eugene P. Wigner. Characteristics vectors of bordered matrices with infinite dimensions. II. Annals of Mathematics (2), 65 (2):203–207, March 1957. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL http://www.jstor.org/

stable/1969956.

Wigner:1957:DNR

[Wig57c] Eugene P. Wigner. Distribution of neutron resonance level spacing. In Anonymous, editor, International Conference on the Neutron Interactions with Nucleus, Columbia University, September 9–13 1957, pages 49–50. Columbia University, New York, NY, USA, 1957. LCCN ???? Technical Information Service Extension, Oak Ridge, Tennessee, Report CU-175 (TID-7547).

Wigner:1957:JN

[Wig57d] Eugene P. Wigner. John von Neumann, 1903–1957. In ????, editor, Yearbook of the American Philosophical Society, pages 149–153. American Philosophical Society, Philadelphia, PA, USA, 1957. ISSN 0065-9762. LCCN ????

Wigner:1957:NR

[Wig57e] Eugene P. Wigner. Neutronic reactor. US Patent 2,806,820.,

September 17, 1957. URL http://www.google.com/patents/

US2806820. US Patent Application filed Aug 18, 1947.

Wigner:1957:RIQ

[Wig57f] Eugene P. Wigner. Relativistic invariance and quantum phenomena. Reviews of Modern Physics, 29(3):255–268, July 1, 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.29.255;   http://rmp.aps.org/abstract/
RMP/v29/i3/p255_1.

Wigner:1957:RTR

[Wig57g] Eugene P. Wigner. Results and theory of resonance absorption. In Anonymous, editor, Conference on Neutron Physics by Time of Flight, held at Gatlinburg, Tennessee, November 1–2, 1956, pages 59–70. Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1957. LCCN ???? Report 2-30-9 (or ORNL-2309).

Wigner:1957:RBN

[Wig57h] Eugene P. Wigner. Review of Progress in nuclear energy, vol. 2. Reactors by Charpie, Hughes and Littler. Physics Today, 10(3): ??, March 1957. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Wigner:1957:RRC

[Wig57i] Eugene P. Wigner. Review of resonance capture by lumps. In Anonymous, editor, Proceedings of the Brookhaven Conference on resonance absorption of neutrons in nuclear reactors, pages 67– ?? Brookhaven National Laboratory, New York, NY, USA, 1957. LCCN ????

Wigner:1958:BRB

[Wig58a] Eugene P. Wigner. Book review: Proceedings of the Rehovoth Conference on Nuclear Structure. Science, 128(3317):195, July 25, 1958. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1754691.

Wigner:1958:CC

[Wig58b] Eugene P. Wigner. Certificate of correction. US Patent 2,860,093.,

November 11, 1958. URL http://www.google.com/patents/

US2860093. US Patent Application filed Nov 13, 1945.

Wigner:1958:FAN

[Wig58c] Eugene P. Wigner. Fuel assembly for a neutronic reactor. US Patent 2,832,732., April 29, 1958. URL http://www.google. com/patents/US2832732. US Patent Application filed Jun 8, 1953.

Wigner:1958:ISQ

[Wig58d] Eugene P. Wigner. Isotopic spin: a quantum number for nuclei. In W. O. (Winfred Oliver) Milligan, editor, Proceedings of the Robert A. Welch Foundation Conferences on Chemical Research. I, the structure of the nucleus, volume 1, pages 67–91. Robert A. Welch Foundation, Houston, TX, USA, 1958. LCCN QD1 R63 1957.

Wigner:1958:LED

[Wig58e] Eugene P. Wigner. Letter to the Editor: Detection of nuclear weapons testing. Bulletin of the Atomic Scientists, 14(6):233–234, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wigner:1958:NRb

[Wig58f] Eugene P. Wigner. Neutronic reactor. US Patent 2,831,806., April

22, 1958. URL http://www.google.com/patents/US2831806.

US Patent Application filed Oct 14, 1952.

Wigner:1958:NRc

[Wig58g] Eugene P. Wigner. Neutronic reactors. US Patent 2,856,339.,

October 14, 1958. URL http://www.google.com/patents/

US2856339. US Patent Application filed May 28, 1945.

Wigner:1958:DRC

[Wig58h] Eugene P. Wigner. On the distribution of the roots of certain symmetric matrices. Annals of Mathematics (2), 67(2):325– 327, March 1958. CODEN ANMAAH. ISSN 0003-486X (print),

1939-8980 (electronic). URL http://www.jstor.org/stable/
1970008.

Wigner:1958:ZPR

[Wig58i] Eugene P. Wigner. Ziele und Probleme von Reaktoren. (German) [Objectives and problems of reactors]. Acta Physica Austriaca, 11 (??):410–??, ???? 1958. CODEN APASAP. ISSN ????

Wigner:1959:CHC

[Wig59a] Eugene P. Wigner. Conducting hydrocarbon conversions in nuclear reactors. US Patent 2,905,610., September 22, 1959. URL

http://www.google.com/patents/US2905610. US Patent Ap-

plication filed Jun 22, 1956.

Wigner:1959:CAU

[Wig59b] Eugene P. Wigner. Convocation address, University of Alberta. In Macphail [Mac59], pages 11–?? LCCN ????

Wigner:1959:IOR

[Wig59c] Eugene P. Wigner. Interview with the Oak Ridger. The Oak Ridger, ??(??):??, December 14, 1959. CODEN ???? ISSN 0890-6009.

Wigner:1959:NIN

[Wig59d] Eugene P. Wigner. New ideas for nuclear reactors. Nuclear Science and Engineering, 6(??):420–??, ???? 1959. CODEN NSE-NAO. ISSN 0029-5639 (print), 1943-748X (electronic).

Wigner:1959:RS

[Wig59e] Eugene P. Wigner. Reactor shield. US Patent 2,874,307., February

17, 1959. URL http://www.google.com/patents/US2874307.

US Patent Application filed Jan 9, 1947.

Wigner:1959:SPR

[Wig59f] Eugene P. Wigner. Statistical properties of real symmetric matrices with many dimensions. In Macphail [Mac59], pages 174–184. LCCN ????

Wigner:1959:GTA

[Wig59g] Eugene Paul Wigner. Group theory and its application to the quantum mechanics of atomic spectra, volume 5 of Pure and applied physics. Academic Press, New York, USA, 1959. xi + 372 pp. LCCN QA171 .W653 1959. Translation by J. J. Griffin of the German original [Wig31b].

Wigner:1960:CHC

[Wig60a] Eugene P. Wigner. Conducting hydrocarbon conversions in nuclear reactors. Canadian Patent 601556., July 12,

1960. URL http://brevets-patents.ic.gc.ca/opic-cipo/
cpd/eng/patent/601556/summary.html.

Wigner:1960:CI

[Wig60b] Eugene P. Wigner. Conference on invariance. Physics Today, 13(3):78–79, March 1960. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v13/i3/p78/s1.

Wigner:1960:FCC

[Wig60c] Eugene P. Wigner. Fallout: Criticism of a criticism. Bulletin of the Atomic Scientists, 16(3):107–108, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lap60].

Wigner:1960:MCT

[Wig60d] Eugene P. Wigner. Measurement of the curvature in a twodimensional universe. Physical Review (2), 120(2):643, October 15, 1960. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.120.643.

Wigner:1960:NRa

[Wig60e] Eugene P. Wigner. Neutronic reactor. US Patent 2,954,335.,

September 27, 1960. URL http://www.google.com/patents/

US2954335. US Patent Application filed Feb 4, 1946.

Wigner:1960:NRb

[Wig60f] Eugene P. Wigner. Neutronic reactors. US Patent 2,961,392.,

November 22, 1960. URL http://www.google.com/patents/

US2961392. US Patent Application filed Aug 28, 1945.

Wigner:1960:NFA

[Wig60g] Eugene P. Wigner. Normal form of antiunitary operators. Journal of Mathematical Physics, 1(5):409–413, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/i5/
p409_s1;   http://link.aip.org/link/jmapaq/v1/i5/p409/
s1.

Wigner:1960:PDB

[Wig60h] Eugene P. Wigner. Phenomenological distinction between unitary and antiunitary symmetry operators. Journal of Mathematical Physics, 1(5):414–416, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v1/i5/p414_s1;   http:
//link.aip.org/link/jmapaq/v1/i5/p414/s1.

Wigner:1960:RC

[Wig60i] Eugene P. Wigner. Review of the conference. In Zucker et al. [ZHH60], pages 305–313. LCCN ????

Wigner:1960:UEMa

[Wig60j] Eugene P. Wigner. The unreasonable effectiveness of mathematics in the natural sciences. Communications on Pure and Applied Mathematics (New York), 13(1):1–14, February 1960. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic). Richard Courant Lecture in mathematical sciences delivered at New York University, May 11, 1959.

Wigner:1960:UEMb

[Wig60k] Eugene P. Wigner. The unreasonable effectiveness of mathematics in the natural sciences. In Anonymous, editor, Technion Yearbook, pages 73–?? American Technion Society, New York, NY, USA, 1960. LCCN ????

Wigner:1960:UEMc

[Wig60l] Eugene P. Wigner. The unreasonable effectiveness of mathematics in the natural sciences. Kayaku, 31(9):450–??, ???? 1960. CODEN ???? ISSN ????

Wigner:1961:GLP

[Wig61a] E. P. Wigner. Geometry of light paths between two material bodies. Journal of Mathematical Physics, 2(2):207– 211, February 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v2/i2/p207_s1;   http://
link.aip.org/link/jmapaq/v2/i2/p207/s1.

Wigner:1961:PES

[Wig61b] E. P. Wigner. The probability of the existence of a selfreproducing unit. In Anonymous [Ano61], pages 231–248. LCCN Q171 .L855 1961.

Wigner:1961:BSR

[Wig61c] Eugene P. Wigner. Bellows seal on reactor coolant tube. US Patent 2,975,115., March 14, 1961. URL http://www.google. com/patents/US2975115. US Patent Application filed May 14, 1946.

Wigner:1961:BRD

[Wig61d] Eugene P. Wigner. Book review: Daniel Blanc and Georges Ambrosino, Eléments de Physique nucléaire. Physics Today, 14(6): 52, June 1961. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v14/i6/p52/s1.

Wigner:1961:BRY

[Wig61e] Eugene P. Wigner. Book review: S. Yiftah and D. Okrent and P. A. Moldauer and J. V. Dunworth, Fast Reactor Cross Sections. Physics Today, 14(7):43, July 1961. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v14/i7/p43/s1.

Wigner:1961:MPN

[Wig61f] Eugene P. Wigner. Mathematical problems of nuclear reactor theory. In Proceedings of the Symposium on Applied Mathematics, Vol. XI, pages 89–104. American Mathematical Society, Providence, RI, USA, 1961.

Wigner:1961:MAP

[Wig61g] Eugene P. Wigner. Method and apparatus for producing power. US Patent 2,990,348., June 27, 1961. URL http://www.google. com/patents/US2990348. US Patent Application filed May 26,

Wigner:1961:REW

[Wig61h] Eugene P. Wigner. Recall the ends — while pondering means. Bulletin of the Atomic Scientists, 17(3):82–85, March 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wigner:1962:BBI

[Wig62a] Eugene Wigner. Books: The Inspiration of Science, by Sir George Thompson. Bulletin of the Atomic Scientists, 18(9):27– 28, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wigner:1962:FAA

[Wig62b] Eugene Wigner. Fermi Award: AEC honors Teller for contributions to nuclear science. Science, 138(3545):1087–1088, December 7, 1962. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
1709475;   http://www.sciencemag.org/content/138/3545/
1087.full.pdf.

Wigner:1962:BRB

[Wig62c] Eugene P. Wigner. Book review: American Scientists and Nuclear Weapons Policy, by Robert Gilpin. Bulletin of the Atomic Scientists, 18(8):29–30, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wigner:1962:VAP

[Wig62d] Eugene P. Wigner. Il y a vingt ans: Le premier réaction en chaine: Réflexions sur un anniversaire. (French) [Twenty years ago: The first chain reaction: Reflections on an anniversary]. Agence Internat. de l’Energie Atomic Bull., Special number:31–??, December 2, 1962. CODEN ???? ISSN ????

Wigner:1962:NR

[Wig62e] Eugene P. Wigner. Neutronic reactor. US Patent 3,070,529.,

December 25, 1962. URL http://www.google.com/patents/

US3070529. US Patent Application filed Mar 1, 1946.

Wigner:1962:RMB

[Wig62f] Eugene P. Wigner. Remarks on the mind-body question. In Mayne et al. [MGM62], pages 284–304. LCCN ????

Wigner:1962:RCT

[Wig62g] Eugene P. Wigner. Review of collision theory. In R. Stopper, editor, Transfert d’énergie dans les gaz: 12. conseil de chimie tenu à l’Université libre de Bruxelles du 5 au 10 novembre 1962, pages 211–239, 303–306, 308, 453–457, 515–516. Interscience Publishers, New York, NY, USA, 1962. LCCN ???? Discussion du Rapport de M. J. D. Morrison, page 453.

Wigner:1962:TQM

[Wig62h] Eugene P. Wigner. Theorie der quantenmechanischen Messung.. (German) [Theory of quantum-mechanical measurement]. In Ernst Brüche, editor, Physikertagung, Wien: Hauptvorträge von der gemeinsamen Jahrestagung 1961 des Verbandes Deutscher Physikalischer Gesellschaften und der Österreichischen Physikalischen Gesellschaft. (German) [Conference on Physics, Vienna. Keynotes of the 1961 joint annual meeting of the Association of German Physical Societies and the Austrian Physical societies], pages 1–8. Physik Verlag, Mosbach/Baden, West Germany, 1962. LCCN ????

Wigner:1962:TAC

[Wig62i] Eugene P. Wigner. Thoughts on the anniversary of CP-I. International Atomic Energy Agency Bulletin, Special number:31–??, December 2, 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic).

Wigner:1962:UGP

[Wig62j] Eugene P. Wigner. Union of German Physical Societies. Physics Today, 15(1):98–??, January 1962. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v15/i1/p98/s1.

Wigner:1963:BRA

[Wig63a] Eugene P. Wigner. Book review: Amos De-Shalit, Igal Talmi, and H. S. W. Massey, editors, Nuclear Shell Theory. Physics Today, 16(9):67, September 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v16/i9/p67/s1.

Wigner:1963:IQM

[Wig63b] Eugene P. Wigner. Invariant quantum mechanical equations of motion. In Richard Huntley Capps, editor, Theoretical Physics:

Lectures at the seminar on theoretical physics organized by the International Atomic Energy Agency, held at Trieste, from 16 July to 25 August 1962, pages 59–82. International Atomic Energy Agency, Vienna, Austria, 1963. LCCN ????

Wigner:1963:NRS

[Wig63c] Eugene P. Wigner. Neutronic reaction system. US Patent 3,102,851., September 3, 1963. URL http://www.google.com/ patents/US3102851. US Patent Application filed Mar 19, 1947.

Wigner:1963:WPM

[Wig63d] Eugene P. Wigner. On weakly positive matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):313–317, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wigner:1963:PM

[Wig63e] Eugene P. Wigner. The problem of measurement. American Journal of Physics, 31(1):6–15, January 1963. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.

aip.org/link/ajpias/v31/i1/p6/s1.

Wigner:1963:TBA

[Wig63f] Eugene P. Wigner. Twentieth birthday of the atomic age: Reflections on the atomic bomb. The New York Times Magazine, ?? (??):34–??, December 2, 1963. CODEN ???? ISSN ????

Wigner:1964:CMCa

[Wig64a] E. P. Wigner. Causality, R-matrix, and collision matrix. In Dispersion Relations and Their Connection with Causality (Proceedings of the International School of Physics “Enrico Fermi”, Course XXIX, Varenna, 1963), pages 40–67. Academic Press, New York, USA, 1964.

Wigner:1964:CDA

[Wig64b] E. P. Wigner. Civil defense: Address to the Mercer County, New Jersey Civil Defense and Disaster Control Organization, May 26, 1964. Journal of the New Jersey Chiropodists Society, 1(1):6–??, 1964. CODEN ???? ISSN ????

Wigner:1964:LAW

[Wig64c] E. P. Wigner. Letter on article by Weisner and York. Scientific American, 211(6):8, 10, 12, December 1964. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wigner:1964:RIP

[Wig64d] E. P. Wigner. The role of invariance principles in natural philosophy. In J. Rosen, editor, Dispersion Relations and Their Connection with Causality (Proceedings of the International School of Physics “Enrico Fermi”, Course XXIX, Varenna, 1963), pages ix–xvi. Academic Press, New York, USA, 1964.

Wigner:1964:URI

[Wig64e] E. P. Wigner. Unitary representations of the inhomogeneous Lorentz group including reflections. In Group theoretical concepts and methods in elementary particle physics (Lectures Istanbul Summer School Theoretical Physics, 1962), pages 37–80. Gordon and Breach, New York, NY, USA, 1964.

Wigner:1964:BRB

[Wig64f] Eugene P. Wigner. Book review: The Quantum Theory of Fields. Proceedings of the conference held at the University of Brussels, October 1961. Stoops, Brussels; Interscience, (Wiley), New York, 1963. 261 pp. Illus. $8. Science, 143(3605):462, January 31, 1964. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/143/3605/462.1.
full.pdf.

Wigner:1964:BRS

[Wig64g] Eugene P. Wigner. Book review: The Solvay Congresses: The Quantum Theory of Fields. Proceedings of the conference held at the University of Brussels, October 1961. Science, 143(3605):462, January 31, 1964. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
1712047.

Wigner:1964:CMCb

[Wig64h] Eugene P. Wigner. Causality, R-matrix and collision matrix. Technology Review (M.I.T.), 66(8):40–67, ???? 1964. CODEN TEREAU. ISSN 0040-1692.

Wigner:1964:CDWa

[Wig64i] Eugene P. Wigner. Civil defense — why and how. Nuclear News (ANS), ??(??):3–??, February 1964. CODEN ???? ISSN ????

Wigner:1964:CDWb

[Wig64j] Eugene P. Wigner. Civil defense — why and how. Civilforsvarsstyrelsen, 9(??):??, December 18, 1964. CODEN ???? ISSN ????

Wigner:1964:C

[Wig64k] Eugene P. Wigner. Cuba. Science, 145(3628):110, July 10, 1964. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/145/3628/110.1.
full.pdf.

Wigner:1964:RIN

[Wig64l] Eugene P. Wigner. Die Rolle von Invarianzprinzipien in der Naturphilosophie. (German) [The role of invariance principles in natural philosophy]. Physikalische Blätter, 20(9):393–406, September 1964. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19640200901/abstract.

Wigner:1964:EES

[Wig64m] Eugene P. Wigner. Elastic eigenvibrations of symmetric systems. In Meijer [Mei64], pages 247–264. LCCN ????

Wigner:1964:ENI

[Wig64n] Eugene P. Wigner. Ereignisse, Naturgesetze und Invarianzprinzipien: Nobel-Vortrag am 12. Dezember 1963. (German) [Events, laws of nature, and invariance principles: Nobel Lecture, 12 December 1963]. Angewandte Chemie, 76(17), September 7, 1964. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Wigner:1964:ELNa

[Wig64o] Eugene P. Wigner. Events, laws of nature, and invariance principles: Nobel lecture, December 12, 1963. In Anonymous, editor, Les Prix Nobel en 1963, pages 120–132. Nobel Foundation, Stockholm, Sweden, 1964. LCCN ????

Wigner:1964:ELNb

[Wig64p] Eugene P. Wigner. Events, laws of nature, and invariance principles: Nobel lecture, December 12, 1963. Science, 145 (3636):995–999, September 4, 1964. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1714068;   http://www.sciencemag.org/
content/145/3636/995.full.pdf.

Wigner:1964:OTR

[Wig64q] Eugene P. Wigner. The operation of time reversal in quantum mechanics. In Meijer [Mei64], pages 265–278. LCCN ????

Wigner:1964:STW

[Wig64r] Eugene P. Wigner. Science in two worlds. Survey, 51(??):22–23, April 1964. CODEN ???? ISSN ????

Wigner:1964:SCLa

[Wig64s] Eugene P. Wigner. Symmetry and conservation laws. Physics Today, 17(3):34–40, March 1964. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v17/i3/p34/s1.

Wigner:1964:SCLb

[Wig64t] Eugene P. Wigner. Symmetry and conservation laws. Proceedings of the National Academy of Sciences of the United States of America, 51(5):956–965, May 15, 1964. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://

www.jstor.org/stable/71915.

Wigner:1964:TKRa

[Wig64u] Eugene P. Wigner. Two kinds of reality. The Monist, 48(2):248– 264, April 1964. CODEN ???? ISSN 0026-9662 (print), 2153-3601

(electronic). URL http://www.jstor.org/stable/27901548.

Wigner:1964:TKRb

[Wig64v] Eugene P. Wigner. Two kinds of reality. In Edward D. Simmons, editor, Essays on knowledge and methodology, pages 41–?? Ken Cook, Milwaukee, WI, USA, 1964. LCCN ???? Papers prepared originally for presentation at workshops in philosophy, held at Marquette University between 1955 and 1964.

Wigner:1964:WCD

[Wig64w] Eugene P. Wigner. Why civil defense? Technology Review (M.I.T.), 66(8):21–23, June 1964. CODEN TEREAU. ISSN 0040-1692.

Wigner:1964:DRT

[Wig64x] Eugene Paul Wigner, editor. Dispersion relations and their connection with causality. Its Proceedings. Academic Press, New York, USA, 1964. xvi + 256 pp. LCCN QC174.45 .V3.

Wigner:1964:TTE

[Wig64y] Wigner Jenő. Áttekintés az ütközések elméletéről. (Hungarian) [Overview of the theory of collisions]. Fizikai Szemle (Budapest), 14(2):35–44, ???? 1964. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1964:IKH

[Wig64z] Wigner Jenő. Az időtükrözés a kvantummechanikában. (Hungarian) [Mirroring time in quantum mechanics]. Magyar Fizikai Folyóirat, 12(2):213–222, ???? 1964. CODEN MGFFAC. ISSN 0025-0104.

Wigner:1964:MMH

[Wig64-27] Wigner Jenő. A matematika meghökkentő hatékonysága a természettudományokban. (Hungarian) [The surprising effectiveness of mathematics in the natural sciences]. Fizikai Szemle (Budapest), 14(6):173–180, ???? 1964. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1964:SFE

[Wig64-28] Wigner Jenő. A schrödinger-féle elméletből a termszerkezetekre vonatkozóan adódó néhány következtetés. (Hungarian) [Schrödinger’s theory of the thermal structures: some conclusions]. Magyar Fizikai Folyóirat, 12(1):85–110, ???? 1964. CODEN MGFFAC. ISSN 0025-0104.

Wigner:1964:THH

[Wig64-29] Wigner Jenő. A tudomány határai. (Hungarian) [The boundaries of science]. Élet és Irodalom [Life and Literature], 10(40):6–7, ???? 1964. CODEN ???? ISSN 0424-8848 (print), 1588-0362 (electronic).

Wigner:1964:VTHa

[Wig64-30] Wigner Jenő. Van-e a tudománynak határa ?. (Hungarian) [Is there a limit to science?]. Fizikai Szemle (Budapest), 14(11):334– 339, ???? 1964. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1964:VTHb

[Wig64-31] Wigner Jenő. Van-e a tudománynak határa ?. (Hungarian) [Is there a limit to science?]. Élet és Irodalom [Life and Literature], VIII(40):6–7, ???? 1964. CODEN ???? ISSN 0424-8848 (print), 1588-0362 (electronic).

Wigner:1965:MDH

[Wig65a] E. Wigner. On the mass defect of helium. In Nuclear Forces [Bri65a], pages 170–181. ISBN 0-08-011034-7. LCCN QC173 .B8513 1965. Reprint of [Wig33a].

Wigner:1965:SNP

[Wig65b] E. Wigner. On the scattering of neutrons by protons. In Nuclear Forces [Bri65a], pages 182–188. ISBN 0-08-011034-7. LCCN QC173 .B8513 1965. Reprint of [Wig33b].

Wigner:1965:BRB

[Wig65c] Eugene P. Wigner. Book review: Optical model of the atomic nucleus, by I. Ulehla, L. Gomolcak and Z. Pluhar. 147 pages, diagrams, 6×9 in. Publishing House of the Czechoslovak Academy of Sciences and Academic Press, Inc., New York, 1965. Price, $7.75. Journal of The Franklin Institute, 280(4):356, October 1965. CO-DEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0016003265903157.

Wigner:1965:BRL

[Wig65d] Eugene P. Wigner. Book review: Lancelot Law Whyte, Internal Factors in Evolution. Physics Today, 18(10):62, October 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v18/i10/p62/s1.

Wigner:1965:CTE

[Wig65e] Eugene P. Wigner. Commentary: Toward an effective civil defense. Nuclear Applications, ??(??):287–??, August 1, 1965. CO-DEN ???? ISSN ????

Wigner:1965:DLR

[Wig65f] Eugene P. Wigner. Distribution laws for the roots of a random Hermitean matrix. In Porter [Por65], pages 446–?? LCCN ????

Wigner:1965:DNR

[Wig65g] Eugene P. Wigner. Distribution of neutron resonance level spacing. In Porter [Por65], pages 224–225. LCCN ????

Wigner:1965:DRC

[Wig65h] Eugene P. Wigner. On the distribution of the roots of certain symmetric matrices. In Porter [Por65], pages 226–227. LCCN ????

Wigner:1965:MWR

[Wig65i] Eugene P. Wigner. On the matrices which reduce the Kronecker products of representations of S. R. groups. In Biedenharn and Van Dam [BV65], pages 87–133. LCCN QC174.1 .B46.

Wigner:1965:SDW

[Wig65j] Eugene P. Wigner. On the statistical distribution of the widths and spacings of nuclear resonance levels. In Porter [Por65], pages 120–128. LCCN ????

Wigner:1965:RCC

[Wig65k] Eugene P. Wigner. Reply to S/C’s criticism of the Project Harbor summary. Scientist and Citizen, 7(9):1–5, August 1965. CODEN

???? ISSN 2155-1278. URL http://www.tandfonline.com/doi/
abs/10.1080/21551278.1965.9958623.

Wigner:1965:RTR

[Wig65l] Eugene P. Wigner. Results and theory of resonance absorption. In Porter [Por65], pages 200–211. LCCN ????

Wigner:1965:RBF

[Wig65m] Eugene P. Wigner. Review of Internal factors in evolution, by Lancelot Law White. Physics Today, 18(10):62, October 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://link.aip.org/link/phtoad/v18/i10/
p62/s1.

Wigner:1965:STW

[Wig65n] Eugene P. Wigner. Science in two worlds. In Anonymous [Ano65], pages 451–457. LCCN Q127.R9 S95.

Wigner:1965:SCS

[Wig65o] Eugene P. Wigner. Some consequences of the Schrödinger theory for eigenvalues. In Biedenharn and Van Dam [BV65], pages 87– 133. LCCN QC174.1 .B46.

Wigner:1965:SCLa

[Wig65p] Eugene P. Wigner. Symmetry and conservation laws. Scientific Endeavor, ??(??):140–??, ???? 1965. CODEN ???? ISSN ????

Wigner:1965:SCLb

[Wig65q] Eugene P. Wigner. Symmetry and conservation laws. In National Academy of Sciences Centennial Book, page ?? National Academy Press, Washington, DC, USA, 1965.

Wigner:1965:UGS

[Wig65r] Eugene P. Wigner. [unknown]. (German) [Symmetry and conservation laws]. Jenaer Rundschau, ??(??):251–??, May 1965. CO-DEN ???? ISSN ????

Wigner:1965:USS

[Wig65s] Eugene P. Wigner. [unknown]. (Serbian) [Symmetry and conservation laws]. Obzornik za Matematiko in Fiziko, 12(??):32–??, ???? 1965. CODEN ???? ISSN ????

Wigner:1965:VSP

[Wig65t] Eugene P. Wigner. Violations of symmetry in physics. Scientific American, 213(6):28–36, December 1965. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/
v213/n6/pdf/scientificamerican1265-28.pdf.

Wigner:1965:ETI

[Wig65u] Wigner Jenő. Események, természettörvények és invarianciaelvek: 1963. évi nobel-előadás. (Hungarian) [Events, laws of nature and invariance: 1963 annual Nobel Lecture]. Fizikai Szemle (Budapest), 15(1):1–6, ???? 1965. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1966:BRB

[Wig66a] Eugene P. Wigner. Book review: Theory of Groups in Classical & Quantum Physics, Vol. I: Mathematical Structures & the Foundations of Quantum Theory, by T. Kahan. American Scientist, 54(4):478A–479A, December 1966. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://

www.jstor.org/stable/27836649.

Wigner:1966:CRU

[Wig66b] Eugene P. Wigner. Citation and response upon receipt of honorary doctorate. Akademische Reden, 39(??):??, February 25, 1966. CODEN ???? ISSN ????

Wigner:1966:CDW

[Wig66c] Eugene P. Wigner. Civil defense: Wigner on Project Harbor. Bulletin of the Atomic Scientists, 22(2):21–22, February 1966. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wigner:1966:IAK

[Wig66d] Eugene P. Wigner. Introducing the 1966 Alfred Korzybski memorial lecturer (Alvin M. Weinberg). General Semantics Bull., 34 (??):8–??, ???? 1966. CODEN ???? ISSN ????

Wigner:1966:PEC

[Wig66e] Eugene P. Wigner. The possible effectiveness of civil defense. In Civil defense, volume 82, pages 33–51, 121–124. American Association for the Advancement of Science, ????, 1966.

Wigner:1966:RIX

[Wig66f] Eugene P. Wigner. Remarks at the inauguration of X-10 reactor as a national landmark, September 16, 1966. Oak Ridge Nat. Lab. News, 3(??):??, ???? 1966. CODEN ???? ISSN ????

Wigner:1966:RIS

[Wig66g] Eugene P. Wigner. Remarks: Isobaric spin in nuclear physics. In Fox and Robson [FR66], pages 437–466. LCCN QC721 .C68 1966.

Wigner:1966:RGR

[Wig66h] Eugene P. Wigner. Remarks on the giant resonance interpretation of analog states. In Fox and Robson [FR66], page ?? LCCN QC721 .C68 1966.

Wigner:1966:SNP

[Wig66i] Eugene P. Wigner. Science: The new particles and their radiations. Medical Opinion & Review, 1(6):40–??, ???? 1966. CODEN ???? ISSN ????

Wigner:1966:TRC

[Wig66j] Eugene P. Wigner. The two roles of civil defense. Princeton Engineer, 26(5):13–??, February 1966. CODEN ???? ISSN ????

Wigner:1966:KVE

[Wig66k] Wigner Jenő. Keserű és vigasztaló emlékezés. (Hungarian) [Bitter memories and comforting ones]. Új Európa (Múnchen) [New Europe], 5(??):17–??, ???? 1966. CODEN ???? ISSN ????

Wigner:1967:ELN

[Wig67a] Eugene Wigner. Events, laws of nature, and invariance principles: Nobel lecture, December 12, 1963. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67-28], pages 38–50. LCCN Q171 .W65.

Wigner:1967:JN

[Wig67b] Eugene Wigner. John von Neumann, 1903–1957. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67-28], page ?? LCCN Q171 .W65.

Wigner:1967:LS

[Wig67c] Eugene Wigner. The limits of science. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67-28], pages 211–221. LCCN Q171 .W65.

Wigner:1967:LRV

[Wig67d] Eugene Wigner. Longer range view of nuclear energy. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67-28], page ?? LCCN Q171 .W65.

Wigner:1967:DCN

[Wig67e] Eugene Wigner. On the development of the compound nucleus model. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67-28], pages 93–109. LCCN Q171 .W65.

Wigner:1967:PM

[Wig67f] Eugene Wigner. The problem of measurement. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67-28], pages 153–170. LCCN Q171 .W65.

Wigner:1967:REW

[Wig67g] Eugene Wigner. Recall the ends — while pondering the means. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67-28], page ?? LCCN Q171 .W65.

Wigner:1967:RIQ

[Wig67h] Eugene Wigner. Relativistic invariant and quantum phenomena. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67-28], page ?? LCCN Q171 .W65.

Wigner:1967:RMB

[Wig67i] Eugene Wigner. Remarks on the mind–body problem. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67-28], page ?? LCCN Q171 .W65.

Wigner:1967:TKR

[Wig67j] Eugene Wigner. Two kinds of reality. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67-28], page ?? LCCN Q171 .W65.

Wigner:1967:UEM

[Wig67k] Eugene Wigner. The unreasonable effectiveness of mathematics in the natural sciences. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67-28], page ?? LCCN Q171 .W65.

Wigner:1967:ARB

[Wig67l] Eugene P. Wigner. Abstract, review of the Theories of resonance reactions (Madison Conference). Bulletin of the American Physical Society, 12(8):1169–??, ???? 1967. CODEN BAPSA6. ISSN 0003-0503.

Wigner:1967:ACC

[Wig67m] Eugene P. Wigner. Address at the Chicago Commemoration of the first chain reaction. In Chicago Commemorative volume, December 2. 1967, page ?? ????, ????, 1967. LCCN ????

Wigner:1967:BRBb

[Wig67n] Eugene P. Wigner. Book review: Explaining Consciousness. Science, 156(3776):798–799, May 12, 1967. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1721997.

Wigner:1967:BRBa

[Wig67o] Eugene P. Wigner. Book review: Nuclear War & Nuclear Peace, by Y. Harkabi and Y. Shenkman. American Scientist, 55(2): 194A, ???? 1967. CODEN AMSCAC. ISSN 0003-0996 (print),

1545-2786 (electronic). URL http://www.jstor.org/stable/
27836843.

Wigner:1967:BRS

[Wig67p] Eugene P. Wigner. Book review: Salomon Bochner, The Role of Mathematics in the Rise of Science. Physics Today, 20 (11):93, November 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v20/i11/p93/s1.

Wigner:1967:BRW

[Wig67q] Eugene P. Wigner. Book review: Wilhelm Magnus, Fritz Oberhettinger, and R. P. Soni, Formulas and Theorems for the Special Functions of Mathematical Physics. Physics Today, 20(12):81, December 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v20/i12/p81/s1.

Wigner:1967:CDD

[Wig67r] Eugene P. Wigner. Civil defense and disarmament. Industrial Research, 9(1):1–??, ???? 1967. CODEN ???? ISSN ????

Wigner:1967:CDS

[Wig67s] Eugene P. Wigner. Civil defense for stable peace. Industrial Research, 9(??):49–50, July 1967. CODEN ???? ISSN ????

Wigner:1967:GSP

[Wig67t] Eugene P. Wigner. The growth of science — its promise and its dangers: E. P. Wigner’s J. F. Carlson lecture, presented April 13, 1964, at Iowa’s State College. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67-28], page ?? LCCN Q171 .W65.

Wigner:1967:LDN

[Wig67u] Eugene P. Wigner. Letter: Defense from nuclear attack: Remarks on Stuart Chase’s statement. New York Times, ??(??):??, February 27, 1967. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Wigner:1967:GEA

[Wig67v] Eugene P. Wigner. On a generalization of Euler’s angles. Medical Opinion & Review, 3(2):119–??, ???? 1967. CODEN ???? ISSN ????

Wigner:1967:RMP

[Wig67w] Eugene P. Wigner. Random matrices in physics. SIAM Review, 9(1):1–23, January 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://link.aip.org/

link/siread/v9/i1/p1/s1;   http://www.jstor.org/stable/
2027409.

Wigner:1967:RBM

[Wig67x] Eugene P. Wigner. Review of Molecules and men, by Francis Crick. Science, 156(??):798–799, May 12, 1967. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Wigner:1967:RBW

[Wig67y] Eugene P. Wigner. Review of Nuclear war and nuclear peace, by Y. Harkabi. American Scientist, 55(2):194A–??, ???? 1967. CO-DEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

Wigner:1967:RDR

[Wig67z] Eugene P. Wigner. The road to disarmament?: Rebuttal. Industrial Research, 9(??):80–??, July 1967. CODEN ???? ISSN ????

Wigner:1967:WRW

[Wig67-27] Eugene P. Wigner. What’s rotten? : Wigner’s letter. Medical Opinion & Review, 3(2):12–??, ???? 1967. CODEN ???? ISSN ????

Wigner:1967:SRS

[Wig67-28] Eugene Paul Wigner. Symmetries and reflections; scientific essays of Eugene P. Wigner. Indiana University Press, Bloomington, IN, USA, 1967. viii + 280 pp. LCCN Q171 .W65.

Wigner:1967:NJH

[Wig67-29] Wigner Jenő. Neumann János, 1903–1957. (Hungarian) [John von Neumann, 1903–1957]. Fizikai Szemle (Budapest), 17(8):227–229, ???? 1967. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1968:EQM

[Wig68a] E. P. Wigner. Epistemology of quantum mechanics. ????, ??(??): ??, ???? 1968. CODEN ???? ISSN ????

Wigner:1968:APL

[Wig68b] Eugene P. Wigner. Address: a physicist looks at the soul. In Anonymous, editor, Second conference on science for clergymen, August 5–16, 1968, Oak Ridge, Tennessee, pages 243–246. ????, ????, 1968. LCCN ???? Lecture 13.

Wigner:1968:GEA

[Wig68c] Eugene P. Wigner. On a generalization of Euler’s angles. In Loebl [Loe75a], pages 119–129. ISBN 0-12-455150-5. LCCN ????

Wigner:1968:RCD

[Wig68d] Eugene P. Wigner. Roadblocks to civil defense. Survive, 1(1): 1–??, May/June 1968. CODEN ???? ISSN 0039-6354.

Wigner:1968:SSL

[Wig68e] Eugene P. Wigner. The scientist and society: Lecture 4 at the 1968 Meeting at the International Centre for Theoretical Physics, Treist, Italy, June 1968. In Anonymous [Ano68], pages 49–56. LCCN ???? A special supplement of the IAEA Bulletin.

Wigner:1968:SPO

[Wig68f] Eugene P. Wigner. Symmetry principles in old and new physics. Bulletin of the American Mathematical Society, 74(5):793–815, 1968. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Wigner:1968:TER

[Wig68g] Eugene P. Wigner. Time and effort reports defended. Science, 159(3812):255, January 19, 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1722559;   http://www.sciencemag.org/
content/159/3812/255.1.full.pdf.

Wigner:1968:NWC

[Wig68h] Eugene Paul Wigner. Nuclear war and civil defense. In Who speaks for civil defense? [Wig68i], pages 13–27. LCCN UA926 .W45.

Wigner:1968:WSC

[Wig68i] Eugene Paul Wigner, editor. Who speaks for civil defense? Scribner, New York, NY, USA, 1968. 125 pp. LCCN UA926 .W45.

Wigner:1968:BSV

[Wig68j] Wigner Jenő. Beszéd a stockholmi városházán. (Hungarian) [Speech at the Stockholm City Hall]. Fizikai Szemle (Budapest), 18(7):224–??, ???? 1968. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1969:LSB

[Wig69a] E. P. Wigner. Leo Szilard 1898–1964: A biographical memoir. Biographical memoirs — National Academy of Sciences of the United States of America, 40:335–347, 1969. CODEN BMNSAC.

ISSN 0077-2933. URL http://books.nap.edu/html/biomems/
lszilard.pdf.

Wigner:1969:RICb

[Wig69b] E. P. Wigner. Relativistic interaction of classical particles. In Timm Gudehus, Geoffrey Kaiser, and Arnold Perlmutter, editors, Coral Gables conference on fundamental interactions at high energy: Center for theoretical studies, January 22–24, 1969, University of Miami, pages 344–361. Gordon and Breach, New York, NY, USA, 1969. LCCN QC794 .C583.

Wigner:1969:RICa

[Wig69c] E. P. Wigner. Relativistic interaction of classical particles. (talk). AIP Conference Proceedings, C6901221:344–361, 1969. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Wigner:1969:BRY

[Wig69d] Eugene Wigner. Book review: Ye. V. Stupochenko, S. A. Losev, and A. I. Osipov, Relaxation in Shock Waves. Physics Today, 22(5):91, May 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v22/i5/p91/s1.

Wigner:1969:ARB

[Wig69e] Eugene P. Wigner. Advance review of Brain, mind, and computers, by S. L. Jaki. Harder and Harder, 1969. Unknown publication location., 1969.

Wigner:1969:ABE

[Wig69f] Eugene P. Wigner. An appreciation on the 60th birthday of Edward Teller. In Mark and Fernbach [MF69], chapter 1, pages 1–6. ISBN 0-470-56990-5. LCCN Q171 .P98.

Wigner:1969:WM

[Wig69g] Eugene P. Wigner. Are we machines? Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 113(2):95–101, April 17, 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985959.

Wigner:1969:CDLa

[Wig69h] Eugene P. Wigner. Civil defense — Little Harbor report. Report TID-24690, Division of Technical Information, U.S. Atomic Energy Commission, ????, ???? 1969.

Wigner:1969:CDLb

[Wig69i] Eugene P. Wigner. Civil defense — Little Harbor report. EMO National Digest, 9(??):??, February 4, 1969. CODEN ???? ISSN 0012-7787.

Wigner:1969:DVRa

[Wig69j] Eugene P. Wigner. Defense vs. retaliation. Congressional Record, ??(??):S4465–??, May 1, 1969. CODEN ???? ISSN ????

Wigner:1969:DVRb

[Wig69k] Eugene P. Wigner. Defense vs. retaliation. Survive, 2(4):16–19, July/August 1969. CODEN ???? ISSN 0039-6354.

Wigner:1969:DFW

[Wig69l] Eugene P. Wigner. Djilas forgiven : Wigner’s letter. Princeton Alumni Weekly, 69(??):3–??, February 4, 1969. CODEN ???? ISSN 0149-9270.

Wigner:1969:EQMa

[Wig69m] Eugene P. Wigner. Epistemiology of quantum mechanics — its appraisal and demands. Psychological Issues, 6(2):22–??, ???? 1969. CODEN ???? ISSN ????

Wigner:1969:EQMb

[Wig69n] Eugene P. Wigner. Epistemiology of quantum mechanics — its appraisal and demands. In Salam [Sal69], pages 431–437. LCCN ???? Two volumes.

Wigner:1969:LE

[Wig69o] Eugene P. Wigner. Letters to the Editor. Science, 163(3874): 1397–1398, March 28, 1969. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/163/3874/1397.6.full.pdf.

Wigner:1969:MC

[Wig69p] Eugene P. Wigner. Manpower contradictions. Physics Today, 22(12):11–??, December 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v22/i12/p11/s2.

Wigner:1969:OCD

[Wig69q] Eugene P. Wigner. The objectives of civil defense. In Wigner [Wig69x], pages 3–23. ISBN 0-253-18588-2. LCCN UA926 .S85 1969.

Wigner:1969:SPP

[Wig69r] Eugene P. Wigner. On some problems of physics. Tennessee Alumnus, 49(5):35–??, ???? 1969. CODEN ???? ISSN ????

Wigner:1969:RBM

[Wig69s] Eugene P. Wigner. Review of Random matrices and the statistical theory of energy levels, New York: Academic Press, 1967. Quarterly of Applied Mathematics, 26(??):613–614, ???? 1969. CODEN QAMAAY. ISSN 0033-569X (print), 1552-4485 (electronic).

Wigner:1969:RBB

[Wig69t] Eugene P. Wigner. Review of Reflections on big science, by Alvin Weinberg. Nuclear Applications and Technology, 7(4):399– ??, ???? 1969. CODEN ???? ISSN ????

Wigner:1969:SC

[Wig69u] Eugene P. Wigner. Summary of the conference. In M. Harvey et al., editors, Proceedings. Comptes rendus. International Conference on Properties of Nuclear States. Conférence internationale sur les propriétés des états nucléaires. Montreal, Canada, August 25-30, 1969, pages 633–647. Press de l’université de Montréal, Montréal, QC, Canada, 1969. LCCN QC770 .I518 1969.

Wigner:1969:U

[Wig69v] Eugene P. Wigner. [unknown]. In Grene [Gre69], page ?? LCCN ????

Wigner:1969:WSB

[Wig69w] Eugene P. Wigner. Wigner’s statement before the Subcommittee on International Organization and Disarmament Affairs (Gore Committee) of the Committee on Foreign Relations, U.S. Senate, May 21, 1969.

Wigner:1969:SBM

[Wig69x] Eugene Paul Wigner, editor. Survival and the Bomb: Methods of Civil Defense. Indiana University Press, Bloomington, IN, USA, 1969. ISBN 0-253-18588-2. ix + 307 pp. LCCN UA926 .S85 1969.

Wigner:1969:EAS

[Wig69y] Wigner Jenő. Előszó az atommag szerkezete című műhöz. (Hungarian) [Preface to the nuclear plant’s structure]. Akadémiai k., Budapest, Hungary, 1969. ???? pp. LCCN ????

Wigner:1970:BSE

[Wig70a] Eugene P. Wigner. Blast shelter effectiveness and cost. EMO National Digest, 10(1):27–28, ???? 1970. CODEN ???? ISSN 0012-7787.

Wigner:1970:BRC

[Wig70b] Eugene P. Wigner. Book review: Civil defense: Survival and the Bomb — Methods of Civil Defense. Science, 167(3921):1076, 1079, February 20, 1970. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/ stable/1728662. Criticism of S. Weinberg’s review.

Wigner:1970:CD

[Wig70c] Eugene P. Wigner. Civil defense. Science, 167(3921):1076– 1079, February 20, 1970. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/167/3921/1076.2.full.pdf.

Wigner:1970:EGV

[Wig70d] Eugene P. Wigner. Der erfolgreiche Gegen- und Vernichtungschlag ein Märchen. (German) [The successful counter-strike and destruction of a myth]. Zivilschutz, 11(??):289–??, ???? 1970. CO-DEN ???? ISSN ????

Wigner:1970:DNB

[Wig70e] Eugene P. Wigner. Discussion notes : Behind the SALT talks. In Anonymous, editor, Fourth International Arms Control Symposium, Airlie House, Warrenton, Virginia, May 1–2, 1970. Fourth Section, pages 67–?? University of Pennsylvania, Philadelphia, PA, USA, 1970. LCCN ????

Wigner:1970:MAD

[Wig70f] Eugene P. Wigner. The myth of assured destruction. Survive, 3 (4):2–4, July/August 1970. CODEN ???? ISSN 0039-6354.

Wigner:1970:HVQ

[Wig70g] Eugene P. Wigner. On hidden variables and quantum mechanical probabilities. American Journal of Physics, 38(8):1005–1009, August 1970. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://link.aip.org/link/ajpias/v38/i8/
p1005/s1.

Wigner:1970:PEL

[Wig70h] Eugene P. Wigner. Physics and the explanation of life. Foundations of Physics, 1(1):35–45, March 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

www.springerlink.com/content/p0453n1p7835xt66.

Wigner:1970:SGC

[Wig70i] Eugene P. Wigner. Some general consequence of the short-range nature of nuclear forces. In Bromley and Hughes [BH70], pages 43–91. LCCN QC780 .F3.

Wigner:1970:WCD

[Wig70j] Eugene P. Wigner. Why civil defense? EMO National Digest, 11 (4):16–18, October/November 1970. CODEN ???? ISSN 0012-7787.

Wigner:1970:SRS

[Wig70k] Eugene Paul Wigner. Symmetries and reflections: scientific essays of Eugene P. Wigner. MIT Press, Cambridge, MA, USA, 1970. ISBN 0-262-73021-9. viii + 280 pp. LCCN Q171 .W65 1970.

Wigner:1971:QMD

[Wig71a] E. Wigner. Quantum mechanical distribution functions revisited. In W. Yourgrau and A. van der Merwe, editors, Perspectives in Quantum Theory, pages 25–36. MIT Press, Cambridge, MA, USA, 1971.

Wigner:1971:SOD

[Wig71b] E. Wigner. The subject of our discussions. (talk). In d’Espagnat [d’E71], pages 124–125. ISBN 0-12-368849-3. LCCN QC174.1 .V36 1971.

Wigner:1971:RIR

[Wig71c] E. P. Wigner. Restriction of irreducible representations of groups to a subgroup. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 322(1549):181–189, April 20, 1971. CODEN PRLAAZ. ISSN 0962-8444. URL http://

www.jstor.org/stable/77722.

Wigner:1971:IEP

[Wig71d] Eugene P. Wigner. Interview : Eugene Paul Wigner. New Engineer, 1(2):4–??, November 1971. CODEN ???? ISSN ????

Wigner:1971:CDN

[Wig71e] Eugene P. Wigner. Is civil defense necessary? Muse News, 3(1): 5–??, ???? 1971. CODEN ???? ISSN ????

Wigner:1971:LBA

[Wig71f] Eugene P. Wigner. Letter: The bomb again. Environment, ?? (??):51–??, November 1971. CODEN ???? ISSN ????

Wigner:1971:SPP

[Wig71g] Eugene P. Wigner. On some of physics’s problems. Main Currents in Modern Thought, 28(??):75–??, ???? 1971. CODEN ???? ISSN ????

Wigner:1971:PPQ

[Wig71h] Eugene P. Wigner. The philosophical problem — questions of physical theory. In d’Espagnat [d’E71], pages 122–124. ISBN 0-12-368849-3. LCCN QC174.1 .V36 1971.

Wigner:1971:PBS

[Wig71i] Eugene P. Wigner. Preface of Translation of Soviet civil defense handbook. US Congressional Record, June 16, 1971.

Wigner:1971:RNI

[Wig71j] Eugene P. Wigner. Rejoinder: [Von Neumann’s informal hiddenvariable argument]. American Journal of Physics, 39(9):1097– 1098, September 1971. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.org/

link/ajpias/v39/i9/p1097/s1. See [Cla71b, Cla71a].

Wigner:1971:REB

[Wig71k] Eugene P. Wigner. Review entitled Remembering Einstein’s life and times, of Einstein: The life and times, by Ronald Clark. World Publishers, 1971. Philadelphia Inquirer, ??(??):23–??, September 22, 1971. CODEN ???? ISSN ????

Wigner:1971:SC

[Wig71l] Eugene P. Wigner. Summary of the conference. In Barschall and Haeberli [BH71], pages 389–398. ISBN 0-299-05890-5. LCCN QC794 .I545 1970.

Wigner:1971:BVC

[Wig71m] Wigner Jenő. Bizonyos véges csoportok ábrázolásairól. (Hungarian) [Certain representations of finite groups]. Magyar Fizikai Folyóirat, 19(2):159–164, ???? 1971. CODEN MGFFAC. ISSN 0025-0104.

Wigner:1971:MAE

[Wig71n] Wigner Jenő. A mátrixokról, amelyek egyszerûen reducibilis csoportok ábrázolásainak kronecker-szorzatát kiredukálják. (Hungarian) [The matrix that may be provided simply reducible group representation of the Kronecker-multiplied ????]. Magyar Fizikai Folyóirat, 19(2):165–207, ???? 1971. CODEN MGFFAC. ISSN 0025-0104.

Wigner:1972:CS

[Wig72a] E. P. Wigner. Cloud seeding. Science, 177(4050):651, August 25, 1972. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/177/
4050/651.2.full.pdf.

Wigner:1972:ELN

[Wig72b] E. P. Wigner. Events, laws of nature, and invariance principles. In Nobel Foundation [Nob72], pages 6–17. ISBN 0-444-40993-9. LCCN QC71.N735.

Wigner:1972:TEU

[Wig72c] E. P. Wigner. On the time-energy uncertainty relation. In Salam and Wigner [SW72], pages 237–247. ISBN 0-521-08600-0.

LCCN QC174.1 .A85 1972. URL http://hooke.lib.cam.ac.uk/
cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http://www.
loc.gov/catdir/enhancements/fy1001/72075298-d.html;
http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Wigner:1972:ACM

[Wig72d] Eugene P. Wigner. Appendix C of M. L. Goldberger: Density of states in periodic systems. In Problems of theoretical physics, pages 396–416. Nauka, Moscow, Russia, 1972. ISBN ???? LCCN ????

Wigner:1972:EMW

[Wig72e] Eugene P. Wigner. Enough missiles?: Wigner’s reply to Richard English’s criticism of The bomb again. Environment, 14(3):54–??, ???? 1972. CODEN ???? ISSN ????

Wigner:1972:ERB

[Wig72f] Eugene P. Wigner. Ethics in relationship between science and society : Interview with Eugene P. Wigner. Impact of Science on Society, 22(??):283–288, ???? 1972. CODEN ???? ISSN ????

Wigner:1972:IT

[Wig72g] Eugene P. Wigner. Introductory talk. In J. B. Garg, editor, Statistical properties of nuclei: proceedings of the International Conference on Statistical Properties of Nuclei, held at Albany, New York, August 23–27, 1971. Plenum Press, New York, NY, USA; London, UK, 1972. ISBN 0-306-30576-3. LCCN ????

Wigner:1972:LPD

[Wig72h] Eugene P. Wigner. Letter: Premature demands. Phoenix (Swarthmore College), 93(1):??, September 19, 1972. CODEN ???? ISSN ????

Wigner:1972:OMG

[Wig72i] Eugene P. Wigner. Obituary: Maria Goeppert Mayer. Physics Today, 25(5):77–79, May 1972. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v25/i5/p77/s1.

Wigner:1972:PCM

[Wig72j] Eugene P. Wigner. The place of consciousness in modern physics. In ?. Outbridge and ?. Lazard, editors, Consciousness and reality. Avon Books, ????, 1972. ISBN ???? LCCN ????

Wigner:1972:SN

[Wig72k] Eugene P. Wigner. Symmetry in nature. In Anonymous, editor, [Program] / The Robert A. Welch Foundation Conferences on Chemical Research: XVI. Theoretical chemistry, November 20–22, 1972, Houston, Texas, pages 231–?? Robert A. Welch Foundation, Houston, TX, USA, 1972. LCCN ????

Wigner:1972:WPV

[Wig72l] Eugene P. Wigner. Why I plan to vote Republican. Town Topics (Princeton), ??(??):??, November 2, 1972. CODEN ???? ISSN ????

Wigner:1972:LAM

[Wig72m] Wigner Jenő. … amin én mint dilettáns filozófus gondolkodtam. (Hungarian) [???? I, as a dilettante philosopher thought: [Conversation with Wigner]]. Mérleg, 8(4):319–336, ???? 1972. CODEN ???? ISSN ????

Wigner:1972:SRW

[Wig72n] Wigner Jenő. Szimmetriák és reflexiók: Wigner Jenő tudományos esszéi. (Hungarian) [Symmetries and Reflections: Eugene Wigner scientific essays.]. Gondolat, Budapest, Hungary, 1972. ISBN ???? 355 pp. LCCN ????

Wigner:1972:WJL

[Wig72o] Wigner Jenő. Wigner Jenő levelei Ortvay Rudolfhoz [Közreadta Györgyi Géza]. (Hungarian) [Eugene Wigner leaves Rudolf Ortvay [Participation By George Géza]]. Fizikai Szemle (Budapest), 22 (2):45–58, ???? 1972. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1973:IEW

[Wig73a] Eugene Wigner. An interview with Eugene Wigner by I. Kardos. The New Hungarian Quarterly, 14(51):141–150, ???? 1973. CODEN ???? ISSN ????

Wigner:1973:CDA

[Wig73b] Eugene P. Wigner. Civil defense : Abstract of Wigner’s lecture. Technical report, University of Utah, Salt Lake City, UT, USA, March 9, 1973.

Wigner:1973:EPQ

[Wig73c] Eugene P. Wigner. Epistemological perspective on quantum theory. In Hooker [Hoo73], pages 369–385. ISBN 90-277-0271-3. LCCN QC174.12 .C66.

Wigner:1973:ISS

[Wig73d] Eugene P. Wigner. The impact of success on science. Cahiers de l’Association Lecomte de Nouÿ, 4(??):37–46, Spring 1973. CO-DEN ???? ISSN ????

Wigner:1973:IHM

[Wig73e] Eugene P. Wigner. Introduction : In honor of M. Moshinsky. Revista Mexicana de Física, 22(??):xi–??, ???? 1973. CODEN RMXFAT. ISSN ????

Wigner:1973:RDS

[Wig73f] Eugene P. Wigner. Le retentissement du succès sur la science. (French) [The impact of success on science]. Cahiers de l’Association Lecomte de Nouÿ, 4(??):4–14, Spring 1973. CODEN ???? ISSN ????

Wigner:1973:LFD

[Wig73g] Eugene P. Wigner. Letter: Falk debated. Princeton Alumni Weekly, ??(??):7–??, February 27, 1973. CODEN ???? ISSN 0149-9270.

Wigner:1973:SPP

[Wig73h] Eugene P. Wigner. On some physics’ problems. In ????, editor, Southeastern Section, American Physical Society, 14th Meeting. Wake Forest University, November 1973, pages 15–?? ????, ????, 1973. ISBN ???? LCCN ????

Wigner:1973:RDPb

[Wig73i] Eugene P. Wigner. Reduction of direct products and restriction of representation to subgroups: the everyday tasks of the quantum theorists. In Kolman [Kol73], pages 169–185. LCCN QA252.3 .C66 1972. Reprinted from the September, 1973, issue of the SIAM Journal on Applied Mathematics, volume 25, number 2.

Wigner:1973:RDPa

[Wig73j] Eugene P. Wigner. Reduction of direct products and restriction of representations to subgroups: The everyday tasks of the quantum theorists (in Lie algebras: Applications and computational methods). SIAM Journal on Applied Mathematics, 25(2):169–185, September 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

link.aip.org/link/smjmap/v25/i2/p169/s1;   http://www.
jstor.org/stable/2099938. Reprinted in [Wig73i].

Wigner:1973:REA

[Wig73k] Eugene P. Wigner. Relativistic equations admitting only positive energy. In ????, editor, Southeastern Section, American Physical Society, 14th Meeting. Wake Forest University, November 1973, pages 7–?? ????, ????, 1973. ISBN ???? LCCN ????

Wigner:1973:REQ

[Wig73l] Eugene P. Wigner. Relativistic equations in quantum mechanics. In Mehra [Meh73], pages 320–330. ISBN 90-277-0345-0, 90-277-2536-5. LCCN QC173.96 .S95 1972. URL http://www.springer.

com/us/book/9789027703453.

Wigner:1973:RCD

[Wig73m] Eugene P. Wigner. Roadblocks to civil defense. Survive, ??(??): 3–??, May/June 1973. CODEN ???? ISSN 0039-6354.

Wigner:1973:SSP

[Wig73n] Eugene P. Wigner. Science and society : Pamphlet. Van Leer Jerusalem Foundation and Technion, Haifa, Israel, 1973.

Wigner:1973:SCW

[Wig73o] Eugene P. Wigner. Summary of the course: What I have learned. In Zichichi [Zic73], pages 829–839. LCCN QC793 .I555 1972.

Wigner:1973:WP

[Wig73p] Eugene P. Wigner. Wolfgang Pauli. In Weber [Web73], page ?? ISBN 0-85498-027-X. LCCN Q167.W42. Compiled by R. L. Weber, edited by E. Mendoza, with a foreword by William Cooper.

Wigner:1973:BWJ

[Wig73q] Wigner Jenő. Beszélgetés Wigner Jenővel: TV-interjú, 1973. március: Az interjút kardos istván készítette 1972. szeptemberében, triesztben. (Hungarian) [Interview with Eugene Wigner: TV interview, March 1973: The interview with Stephen Kardos created September 1972, Trieste]. Valóság [Reality], 16(2):73–81, ???? 1973. CODEN ???? ISSN ????

Wigner:1973:HGJ

[Wig73r] Wigner Jenő. Hámos György: Egy jó tanuló a Fasori Gimnáziumból: Interjú. (Hungarian) [George Hamos: a good student in Fasori high school]. Élet és Irodalom [Life and Literature], 17(11):13–??, ???? 1973. CODEN ???? ISSN 0424-8848 (print), 1588-0362 (electronic).

Wigner:1973:VIH

[Wig73s] Wigner Jenő. Visszaemlékezéseim az iskolára. (Hungarian) [Recollections of school]. Fizikai Szemle (Budapest), 23(10):297–298, ???? 1973. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1974:BRB

[Wig74a] Eugene P. Wigner. Book review: Annual review of nuclear science, vol. 22. Nuclear Science and Engineering, 3(??):352–353, ???? 1974. CODEN NSENAO. ISSN ????

Wigner:1974:CR

[Wig74b] Eugene P. Wigner. Concluding remarks. In Symmetry properties of nuclei: Proceedings of the 15th Solvay Conference, September 28–October 3, 1970, pages 351–362. Gordon and Breach, New York, NY, USA, 1974. ISBN ???? LCCN ????

Wigner:1974:EWA

[Wig74c] Eugene P. Wigner. Excerpts from E. Wigner’s address to the USCDC Conference on October 11, 1973, St. Paul, Minnesota. Survive, 7(1):8–??, ???? 1974. CODEN ???? ISSN 0039-6354.

Wigner:1974:IIP

[Wig74d] Eugene P. Wigner. The intellectual influence of physics. Inside AIP, ??(4):??, February 13, 1974. CODEN ???? ISSN ????

Wigner:1974:SPP

[Wig74e] Eugene P. Wigner. On some physics’ problems: Abstract. Bulletin of the American Physical Society, 19(??):688–??, ???? 1974. CODEN BAPSA6. ISSN ????

Wigner:1974:PEL

[Wig74f] Eugene P. Wigner. Physics and the explanation of life. Boston Studies in the Philosophy of Science, 11(??):119–??, ???? 1974. CODEN ???? ISSN ????

Wigner:1974:RWE

[Wig74g] Eugene P. Wigner. Relativistic wave equations admitting only positive energy. Bulletin of the American Physical Society, 19 (??):683–??, ???? 1974. CODEN BAPSA6. ISSN 0003-0503.

Wigner:1974:RPK

[Wig74h] Eugene P. Wigner. The roles of primitive knowledge, of language and of mathematics in the physical sciences: Lecture at the University of Utah. Abstract., May 10, 1974.

Wigner:1974:WR

[Wig74i] Eugene P. Wigner. Wigner replies. Physics Today, 27(6):13–??, June 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v27/i6/p13/s1.

Wigner:1975:ERB

[Wig75a] Eugene P. Wigner. Ethics in the relationship between science and society. In H. Narasimhaiah, editor, Thoughts on science and society, page ?? Bangalore University Press, Bangalore, India, second edition, 1975. ISBN ???? LCCN ????

Wigner:1975:LQL

[Wig75b] Eugene P. Wigner. Letter: Question for Lewis. International Herald Tribune, ??(??):??, October 27, 1975. CODEN ???? ISSN ????

Wigner:1975:REW

[Wig75c] Eugene P. Wigner. Recollections and expectations: E. Wigner’s address at the inauguration of the Kramers Leerstoel, October 13, 1975. In [unknown], pages 1–9. Noord-Hollandsche Uitgers Maats, Amsterdam, The Netherlands, 1975.

Wigner:1975:SSS

[Wig75d] Eugene P. Wigner. Science should study emotions, says Wigner: Interview. The Oak Ridger, ??(??):1–??, May 14, 1975. CODEN ???? ISSN 0890-6009.

Wigner:1975:US

[Wig75e] Eugene P. Wigner. The unity of science. In Anonymous [Ano75], pages 25–33. ISBN 0-89226-003-3, 0-89226-004-1 (paperback). LCCN Q175.4 .I57 1975.

Wigner:1975:USC

[Wig75f] Eugene P. Wigner. The unity of science — closing summary. In Anonymous [Ano75], pages 1291–1296. ISBN 0-89226-003-3, 0-89226-004-1 (paperback). LCCN Q175.4 .I57 1975.

Wigner:1975:EKF

[Wig75g] Wigner Jenő. Emberi korba fogunk lépni. (Hungarian) [We are going to move on]. Magyar Hírek [Hungarian News], 27(3):11–??, ???? 1975. CODEN ???? ISSN ????

Wigner:1976:ONE

[Wig76a] E. P. Wigner. Our needs for energy and ways to satisfy them. Revista Interamericana, 6(??):485–??, ???? 1976. CODEN ???? ISSN ????

Wigner:1976:BRJ

[Wig76b] Eugene Wigner. Book review: Joe Rosen, Symmetry Discovered: Concepts and Applications in Nature and Science. American Scientist, 64(3):335–336, May 1976. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.jstor.

org/stable/27847274.

Wigner:1976:BRB

[Wig76c] Eugene P. Wigner. Book review: The Advisors: Oppenheimer, Teller, and the Superbomb, by Herbert York. American Scientist, 64(5):561, September 1976. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.jstor.org/

stable/27847469.

Wigner:1976:CPL

[Wig76d] Eugene P. Wigner. Civil preparedness and limited nuclear war: Hearing before the Joint Committee on Defense Production, Congress of the United States, Second Session, April 1976.

Wigner:1976:CBD

[Wig76e] Eugene P. Wigner. Comment on civil defense and the Dutch, by K. Muth. Journal of Civil Defense, 9(??):4–??, ???? 1976. CODEN ???? ISSN 0740-5537.

Wigner:1976:IQA

[Wig76f] Eugene P. Wigner. Iranian query answered : E. Wigner’s letter. Daily Reveille, ??(??):4–??, February 12, 1976. CODEN ???? ISSN ????

Wigner:1976:LSN

[Wig76g] Eugene P. Wigner. Letter: Soviet nuclear strategy. New York Times, ??(??):??, July 17, 1976. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Wigner:1976:OWK

[Wig76h] Eugene P. Wigner. Obituary: Werner K. Heisenberg. Physics Today, 29(4):86–87, April 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v29/i4/p86/s1.

Wigner:1976:HVQa

[Wig76i] Eugene P. Wigner. On hidden variables and quantum mechanical probabilities. In Maurin and Rñaczka [MR76], pages xviii + 504. ISBN 90-277-0537-2. LCCN QC19.2 .M37.

Wigner:1976:HVQb

[Wig76j] Eugene P. Wigner. On hidden variables and quantum mechanical probabilities. In Flato et al. [F^{+}76], page ?? ISBN 90-277-0623-9. LCCN QC174.125 .Q36.

Wigner:1976:SPP

[Wig76k] Eugene P. Wigner. On some physics’ problems. In Laszlo and Sellon [LS76], page ?? ISBN 0-306-30884-3. LCCN Q175.3 .V57.

Wigner:1976:PMW

[Wig76l] Eugene P. Wigner. Physik und menschliches Wissen. (German) [Physics and human knowledge]. In Werner Becker and Kurt Hübner, editors, Objektivität in den Natur- und Geisteswissenschaften. (German) [Objectivity in the natural sciences and humanities], page ?? Hoffmann und Campe, Hamburg, West Germany, 1976. ISBN 3-455-09210-1. LCCN BD220 .O24.

Wigner:1976:USE

[Wig76m] Eugene P. Wigner. The unity of science. Excerpts from E. Wigner’s address on frontiers of science at the University of Utah., March 9, 1976.

Wigner:1976:WEO

[Wig76n] Eugene P. Wigner. Weighing on energy options. Prism, 6(51): 485–??, ???? 1976. CODEN ???? ISSN ????

Wigner:1976:SET

[Wig76o] Wigner Jenő. A szimmetria-elvek ötven esztendeje. (Hungarian) [The principles of symmetry and fifty years ago]. Fizikai Szemle (Budapest), 26(10):361–367, ???? 1976. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1977:RBM

[Wig77a] E. P. Wigner. On the relation between momentum and velocity for elementary systems. Physics Letters A, 61(6):353–354, June 13, 1977. CODEN PYLAAG. ISSN 0375-9601 (print),

1873-2429 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0375960177903292.

Wigner:1977:AAA

[Wig77b] Eugene P. Wigner. Atoms, arms, and apathy: an interview by E. T. Tildon. Journal of Civil Defense, 12(??):24–27, ???? 1977. CODEN ???? ISSN 0740-5537.

Wigner:1977:FMI

[Wig77c] Eugene P. Wigner. Fisica — matematica. (Italian) [Physics — mathematics]. In Enciclopedia Italiana del 900, page ?? ????, ????, 1977.

Wigner:1977:ISS

[Wig77d] Eugene P. Wigner. Iranian statements shocking. Daily Reveille, 5(??):81–??, ???? 1977. CODEN ???? ISSN ????

Wigner:1977:MIM

[Wig77e] Eugene P. Wigner. Metodo. (Italian) [Method]. In Enciclopedia Italiana del 900, pages 1–24. ????, ????, 1977.

Wigner:1977:PS

[Wig77f] Eugene P. Wigner. The physical sciences. In Anonymous [Ano77b], pages 79–85. ISBN 0-89226-005-X, 0-89226-006-8 (paperback). LCCN Q174 .I565 1976.

Wigner:1977:PSC

[Wig77g] Eugene P. Wigner. The physical sciences: Committee chairman’s summary. In Anonymous [Ano77b], pages 983–988. ISBN 0-89226-005-X, 0-89226-006-8 (paperback). LCCN Q174 .I565 1976.

Wigner:1977:PRH

[Wig77h] Eugene P. Wigner. Physics and its relation to human knowledge. In Anonymous [Ano77a], pages 283–294. ISBN ???? LCCN ????

Wigner:1977:STA

[Wig77i] Eugene P. Wigner. Science and technology for affluence and satisfaction. In Bugliarello [Bug77], pages 44–50. LCCN Q175.4 .S38.

Wigner:1977:SPS

[Wig77j] Eugene P. Wigner. The scope and promise of science. In Hook et al. [HKT77], pages 131–133. ISBN 0-87975-068-5. LCCN LB2331.4 .E86. US$11.95.

Wigner:1978:CHS

[Wig78a] E. Wigner. City Hall Speech — Stockholm, 1963. In Symmetries and reflections: scientific essays of Eugene P. Wigner [Wig78l], pages 262–263. ISBN 0-313-20107-2. LCCN Q171 .W65 1978.

Wigner:1978:CRa

[Wig78b] Eugene P. Wigner. Concluding remarks. AIP Conference Proceedings, 48:175–177, 1978. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. URL http://link.

aip.org/link/apcpcs/v48/i1/p175/s1.

Wigner:1978:CRb

[Wig78c] Eugene P. Wigner. Concluding remarks. In Lannutti and Williams [LW78], pages 175–178. ISBN 0-88318-147-9. LCCN QC793.3.F5 C87.

Wigner:1978:FYS

[Wig78d] Eugene P. Wigner. Fifty years of symmetry operators. In Zichichi [Zic78], pages 879–892. ISBN 0-306-38183-4. LCCN QC793 .I555 1976.

Wigner:1978:FSW

[Wig78e] Eugene P. Wigner. The future of science — what one can hope for. Science, Technology and Humanities, 1(1):1–5, ???? 1978. CODEN ???? ISSN ????

Wigner:1978:NEA

[Wig78f] Eugene P. Wigner. Nuclear energy and alternatives. In Kadiroğlu et al. [KPS78], pages 22–23. ISBN 0-88410-081-2. LCCN TK9006 .I6 1977.

Wigner:1978:PM

[Wig78g] Eugene P. Wigner. The problem of measurement. In Anonymous [Ano78c], pages 123–132. ISBN ???? LCCN ????

Wigner:1978:PFL

[Wig78h] Eugene P. Wigner. The problems, future and limits for science. In Anonymous [Ano78b], pages 869–877. ISBN 0-89226-007-6, 0-89226-008-4 (paperback). LCCN BD232 .I57 1977.

Wigner:1978:RIK

[Wig78i] Eugene P. Wigner. Reply to Isa Khubas’ remarks. In Anonymous [Ano78b], pages 882–?? ISBN 0-89226-007-6, 0-89226-008-4 (paperback). LCCN BD232 .I57 1977.

Wigner:1978:RCJ

[Wig78j] Eugene P. Wigner. Response to citation by James R. Killians, Jr. In Anonymous [Ano78a], page 101. ISBN 0-262-11068-7. LCCN QC792.7 .P76.

Wigner:1978:WMT

[Wig78k] Eugene P. Wigner. The Wigner Medal: a tribute to Valentine Bargmann. In Kramer and Rieckers [KR78], page ?? ISBN 0-387-08848-2. LCCN QC20.7.G76 I57 1977.

Wigner:1978:SRS

[Wig78l] Eugene Paul Wigner. Symmetries and reflections: scientific essays of Eugene P. Wigner. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1978. ISBN 0-313-20107-2. viii + 280 pp. LCCN Q171 .W65 1978.

Wigner:1979:CCU

[Wig79a] E. P. Wigner. Citation classic — unitary representations of the inhomogeneous Lorentz group. Current Contents/Physical Chemical & Earth Sciences, ??(24):20, ???? 1979. ISSN 0163-2574.

Wigner:1979:NRO

[Wig79b] E. P. Wigner. Nuclear reactor operation. Science, 205(4402):148, July 13, 1979. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/content/
205/4402/148.2.full.pdf.

Wigner:1979:PCS

[Wig79c] E. R. Wigner. Is the principle of causality still applicable? Chemiker-Zeitung, 103(10):329, ???? 1979. ISSN 0009-2894.

Wigner:1979:EI

[Wig79d] Eugene Wigner. Einstein’s ideals. Nature, 282(5735):179–180, November 8, 1979. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v282/n5735/pdf/282179a0.pdf.

Wigner:1979:BCB

[Wig79e] Eugene P. Wigner. The basic conflict between the concepts of general relativity and of quantum mechanics. Bulletin of the American Physical Society, 24(4):633–??, ???? 1979. CODEN BAPSA6. ISSN 0003-0503.

Wigner:1979:COS

[Wig79f] Eugene P. Wigner. Chairman’s opening statement. In Anonymous [Ano79b], pages 13–?? ISBN 0-89226-010-6 (paperback) (set), 0-89226-009-2 (set). LCCN Q175.4 .I57 1978.

Wigner:1979:CIS

[Wig79g] Eugene P. Wigner. Come inventai lo spin isotopico. (Italian) [How I invented the isotopic spin]. Il Tempo della Cultura Moderna, ?? (??):??, November 10, 1979. CODEN ???? ISSN ????

Wigner:1979:CCC

[Wig79h] Eugene P. Wigner. Conference chaiman’s closing statement. In Anonymous [Ano79b], pages 1141–?? ISBN 0-89226-010-6 (paperback) (set), 0-89226-009-2 (set). LCCN Q175.4 .I57 1978.

Wigner:1979:EM

[Wig79i] Eugene P. Wigner. Einstein — a memoir. In French [Fre79], page ?? ISBN 0-435-58200-3. LCCN QC16.E5 E37 1979b.

Wigner:1979:EIR

[Wig79j] Eugene P. Wigner. Einstein’s ideals: Review of Albert Einstein, The human side: New glimpses from his archives, Selected and edited by Helen Dukas and Banesh Hoffman. Pp. 168. Princeton University, 1979. Nature, 282(5735):179–180, November 8, 1979. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v282/
n5735/pdf/282179a0.pdf.

Wigner:1979:EC

[Wig79k] Eugene P. Wigner. Existence of conciousness. In Anonymous [Ano79b], pages 135–143. ISBN 0-89226-010-6 (paperback) (set), 0-89226-009-2 (set). LCCN Q175.4 .I57 1978.

Wigner:1979:EAS

[Wig79l] Eugene P. Wigner. Extension of the area of science. In Herschman [Her79], page ?? ISBN ???? LCCN ????

Wigner:1979:FRE

[Wig79m] Eugene P. Wigner. A friend remembers Einstein. The News World, ??(??):1–??, March 9, 1979. CODEN ???? ISSN ????

Wigner:1979:GKG

[Wig79n] Eugene P. Wigner. Gilt das Kausalitätsprinzip?. (German) [Does the Principle of Causality apply?]. In Anonymous [Ano79a], page ?? ISBN ???? LCCN ????

Wigner:1979:GBI

[Wig79o] Eugene P. Wigner. I giorno di Berlino. (Italian) [The day in Berlin]. Il Tempo della Cultura Moderna, ??(??):??, October 20, 1979. CODEN ???? ISSN ????

Wigner:1979:VNF

[Wig79p] Eugene P. Wigner. La mia vita nella fisica. (Italian) [My life in the physics]. Il Tempo della Cultura Moderna, ??(??):??, September 29, 1979. CODEN ???? ISSN ????

Wigner:1979:SNL

[Wig79q] Eugene P. Wigner. La simmetrie nelle leggi del microcosmo. (Italian) [The symmetries in the laws of the microcosm]. Il Tempo della Cultura Moderna, ??(??):??, May 19, 1979. CODEN ???? ISSN ????

Wigner:1979:LNP

[Wig79r] Eugene P. Wigner. Letter: Nuclear plant safety. Morning Advocate (Baton Rouge, LA), ??(??):??, April 10, 1979. CODEN ???? ISSN ????

Wigner:1979:LMD

[Wig79s] Eugene P. Wigner. L’irreale mondo dietro lo specchio. (Italian) [The unreal world behind the mirror]. Il Tempo della Cultura Moderna, ??(??):??, June 23, 1979. CODEN ???? ISSN ????

Wigner:1979:MNS

[Wig79t] Eugene P. Wigner. Mutamenti nella scienza. (Italian) [Changes in science]. Il Tempo della Cultura Moderna, ??(??):??, June 23, 1979. CODEN ???? ISSN ????

Wigner:1979:NID

[Wig79u] Eugene P. Wigner. The neutron: The impact of its discovery and its uses. In Stuewer [Stu79], pages 159–178. ISBN 0-8166-0869-5. LCCN QC773 .S95 1977.

Wigner:1979:NR

[Wig79v] Eugene P. Wigner. Nuclear risks. The News World, ??(??):??, April 13, 1979. CODEN ???? ISSN ????

Wigner:1979:QSN

[Wig79w] Eugene P. Wigner. Quando le simmetrie non sono perfette. (Italian) [When symmetries are not perfect]. Il Tempo della Cultura Moderna, ??(??):??, June 16, 1979. CODEN ???? ISSN ????

Wigner:1979:CMK

[Wig79x] Wigner Jenő. Csoportelméleti módszer a kvantummechanikában. (Hungarian) [Group theoretical methods in quantum mechanics].

Akad. K., Budapest, Hungary, 1979. ISBN ???? 391 pp. LCCN ????

Wigner:1980:BJN

[Wig80a] Eugene P. Wigner. Biography of John von Neumann, Dec. 28, 1903–Feb. 8, 1957. In John A. Garraty, editor, Dictionary of American Biography, Supplement Six: 1956–1960, pages 655–656. Scribner’s Sons, New York, NY, USA, 1980. ISBN ???? LCCN E176 .D563.

Wigner:1980:ADG

[Wig80b] Eugene P. Wigner. Gli anni della guerra. Poi la pace. (Italian) [The years of war. Then peace]. Il Tempo della Cultura Moderna, ??(??):??, January 5, 1980. CODEN ???? ISSN ????

Wigner:1980:LAU

[Wig80c] Eugene P. Wigner. Letter: About the unity of the sciences. Physics Today, 33(10):15, October 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Wigner:1980:RVS

[Wig80d] Eugene P. Wigner. The role and value of symmetry principles and Einstein’s contribution to their recognition. In Gruber and Millman [GM80], pages 13–21. ISBN 0-306-40541-5. LCCN Q172.5.S95 S92.

Wigner:1980:TYK

[Wig80e] Eugene P. Wigner. Thirty years of knowing Einstein. In Woolf [Woo80], pages 461–468. ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Wigner:1980:WWL

[Wig80f] Eugene P. Wigner. What we have learned. In Zichichi [Zic83a], pages xi + 1104. ISBN 0-306-41036-2. LCCN QC793 .I555 1980.

Wigner:1981:BRS

[Wig81a] E. P. Wigner. Book review: Sir Rudolp Peierls, A Perspective of Physics, Vol. 2: Selections from 1977 Comments on Modern Physics. American Scientist, 69(1):88, January 1981. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL

http://www.jstor.org/stable/27850258.

Wigner:1981:EAS

[Wig81b] Eugene P. Wigner. The extension of the area of science. In Jahn [Jah81], pages 7–16. ISBN 0-89158-955-4. LCCN QC30 .R64.

Wigner:1982:SE

[Wig82a] E. Wigner. On science and its evolution. Journal de physique. Colloques, 43(C8):435–438, ???? 1982. CODEN JOPQAG. ISSN 0302-0738.

Wigner:1982:CCD

[Wig82b] Eugene P. Wigner. A case for civil defense: Prepared statement to the Committee on Foreign Relations, U. S. Senate. In Ninety-Seventh Congress, Second Session, March 16 and 31, 1982, pages 21–23. US Government Printing Office, Washington, DC, USA, 1982. ISBN ???? LCCN ????

Wigner:1982:CDC

[Wig82c] Eugene P. Wigner. Civil defense I and civil defense II. Progress in scientific culture, 7(3):181–182, ???? 1982. CODEN ???? ISSN ????

Wigner:1982:CDO

[Wig82d] Eugene P. Wigner. Civil defense: Our no. 1. requirement. Journal of Civil Defense, 15(2):22–??, ???? 1982. CODEN ???? ISSN 0740-5537.

Wigner:1982:EKC

[Wig82e] Eugene P. Wigner. Education — key to civil defense success: From an address to the American Civil Defense Association on October 9, 1981 in Washington. Journal of Civil Defense, 15(2): 18–20, ???? 1982. CODEN ???? ISSN 0740-5537.

Wigner:1982:EC

[Wig82f] Eugene P. Wigner. The existence of conciousness. In Rubenstein [Rub82], pages 279–285. ISBN 0-89226-015-7. LCCN CB478 .I57 1982. US$24.95.

Wigner:1982:FSN

[Wig82g] Eugene P. Wigner. Family, science, Nobel physicists gift to world. Press Republican, 3(??):??, May 15, 1982. CODEN ???? ISSN ????

Wigner:1982:IPT

[Wig82h] Eugene P. Wigner. Invariance in physical theory; symmetry and conservation laws; the role of invariance principles in natural philosophy. In Rosen [Ros82], pages 33–38. ISBN ???? LCCN ???? Reprint of [Wig49d].

Wigner:1982:LVP

[Wig82i] Eugene P. Wigner. The limitations of the validity of present-day physics. In Elvee [Elv82], page ?? ISBN 0-06-250285-9. LCCN BD161 .N6 1981.

Wigner:1982:PCS

[Wig82j] Eugene P. Wigner. Prima che si ascendesse il fuoco nucleare. (Italian) [Before they ascended the nuclear fire]. Il Tempo Della Cultura Moderna, ??(??):??, December 11, 1982. CODEN ???? ISSN ????

Wigner:1982:RDB

[Wig82k] Eugene P. Wigner. Remarks on Dr. Beilenson’s article. Journal of Civil Defense, 15(2):22–??, ???? 1982. CODEN ???? ISSN 0740-5537.

Wigner:1982:SCR

[Wig82l] Eugene P. Wigner. The story of the 1st chain reaction: Address at the American Nuclear Society Meeting, Nov. 16, 1982. In Anonymous, editor, American Nuclear Society Meeting, page ?? American Nuclear Society, ????, 1982. ISBN ???? LCCN ????

Wigner:1983:CCQ

[Wig83a] E. P. Wigner. Citation classic — on the quantum correction for thermodynamic equilibrium. Current Contents/Physical Chemical & Earth Sciences, ??(51):18, ???? 1983. ISSN 0163-2574.

Wigner:1983:DSP

[Wig83b] Eugene P. Wigner. Death or surrender? I prefer freedom. New York Tribune, ??(??):??, April 27, 1983. CODEN ???? ISSN ????

Wigner:1983:GDPa

[Wig83c] Eugene P. Wigner. The glorious days of physics. In Zichichi [Zic83b], pages 765–774. ISBN 0-306-41242-X. LCCN ????

Wigner:1983:GDPb

[Wig83d] Eugene P. Wigner. The glorious days of physics. In Meystre and Scully [MS83], pages 1–7. ISBN 0-306-41354-X. LCCN QC446.15 .N37 1981. Proceedings of the NATO Advanced Study Institute on Quantum Optics and Experimental General Relativity, held August 16–29, 1981, in Bad Windsheim, Federal Republic of Germany.

Wigner:1983:HCA

[Wig83e] Eugene P. Wigner. Honorary Chairman’s address. In Anonymous [Ano83a], pages 25–?? ISBN 0-89226-020-3 (set), 0-89226-021-1 (paperback : set). LCCN T14.5 .I58 1982.

Wigner:1983:ICL

[Wig83f] Eugene P. Wigner. Initial conditions, laws of nature, symmetries, and their history in physics. In García Doncel et al. [GHM^{+}87], pages xv + 678. ISBN 84-7488-148-8. LCCN QC174.17.S9 I57 1983.

Wigner:1983:IQM

[Wig83g] Eugene P. Wigner. The interpretation of quantum mechanics. In Wheeler and Zurek [WZ83], pages 260–314. ISBN 0-691-08315-0 , 0-691-08316-9 (paperback). LCCN QC174.125 .Q38 1983. US$60.00; US$19.50.

Wigner:1983:LD

[Wig83h] Eugene P. Wigner. The limitations of determinism. In Anonymous [Ano83a], pages 1365–1370. ISBN 0-89226-020-3 (set), 0-89226-021-1 (paperback : set). LCCN T14.5 .I58 1982.

Wigner:1983:PM

[Wig83i] Eugene P. Wigner. The problem of measurement. In Wheeler and Zurek [WZ83], pages 324–341. ISBN 0-691-08315-0 , 0-691-08316-9 (paperback). LCCN QC174.125 .Q38 1983. US$60.00; US$19.50.

Wigner:1983:QAI

[Wig83j] Eugene P. Wigner. Questions and answers: Interview with E. P. Wigner. Washington Post, 2C, ??(??):??, January 19, 1983. CODEN ???? ISSN ????

Wigner:1983:RQMa

[Wig83k] Eugene P. Wigner. Review of the quantum mechanical measurement problem. In Meystre and Scully [MS83], pages 43–63. ISBN 0-306-41354-X. LCCN QC446.15 .N37 1981. Proceedings of the NATO Advanced Study Institute on Quantum Optics and Experimental General Relativity, held August 16–29, 1981, in Bad Windsheim, Federal Republic of Germany.

Wigner:1983:SSP

[Wig83l] Eugene P. Wigner. Sarebbe stati piu difficile prevedere la crisi del chianti in America che l’accensione del fuoco nucleare. (Italian) [It would have been more difficult to predict the crisis in the chianti (??) America from the ignition of nuclear fire]. Il Tempo della Cultura Moderna, 17(??):8–??, January 1983. CODEN ???? ISSN ????

Wigner:1984:DFP

[Wig84a] E. P. Wigner. Distribution functions in physics: Fundamentals. Physics Reports, 106(3):121–167, April 1984. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0370157384901601.

Wigner:1984:CRS

[Wig84b] Eugene P. Wigner. Changes in the role of symmetry principles. In Zachary [Zac84], pages 591–597. ISBN 9971-966-87-5, 9971-966-88-3 (paperback). LCCN QC20.7.G76 I57 1984.

Wigner:1984:MS

[Wig84c] Eugene P. Wigner. The meaning of symmetry. In Zichichi [Zic84], pages 729–733. ISBN 0-306-41738-3. LCCN QC793.3.F5; 19845.

Wigner:1984:RRP

[Wig84d] Eugene P. Wigner. Reflection on the role and purpose of science. In Villee [Vil85], page 1984. ISBN 0-89226-040-8. LCCN AC5 .I696 1984.

Wigner:1984:RQM

[Wig84e] Eugene P. Wigner. Review of the quantum-mechanical measurement problem. In Kerr et al. [K^{+}84], pages 63–82. ISBN 0-12-404970-2. LCCN Q223 .S24 1984.

Wigner:1984:UUV

[Wig84f] Eugene P. Wigner. The use and ultimate validity of invariance principles. In Denardo et al. [DGW84], pages 1–12. ISBN 3-540-13335-6, 0-387-13335-6. LCCN ???? URL http://www.

springerlink.com/content/y24074280p7813l3.

Wigner:1985:ELNb

[Wig85a] Eugene P. Wigner. Events, laws of nature, and invariance principles. In Zichichi [Zic85], pages 699–711. ISBN 0-306-42203-4. LCCN QC793 .I555 1983. Proceedings of the Twenty-first Course of the International School of Subnuclear Physics, held August 3–14, 1983, in Erice, Trapani, Sicily, Italy.

Wigner:1985:MWN

[Wig85b] Eugene P. Wigner. My work for nuclear reactors. Princeton Packet, ??(??):??, July 22, 1985. CODEN ???? ISSN ????

Wigner:1985:SE

[Wig85c] Eugene P. Wigner. Symmetrie und Erhaltungssätze. (German) [Symmetry and conservation laws]. In Symmetry, Leitthemen, pages 46–60. Aulis/Deubner, Cologne, West Germany, 1985. Translation from the English and commentary by Lothar Jansen.

Wigner:1985:UEM

[Wig85d] Eugene P. Wigner. The unreasonable effectiveness of mathematics in the natural sciences [Comm. Pure Appl. Math. 13 (1960), 1–14; Zbl 102, 7]. In Mickens [Mic85], pages 1–14. ISBN 0-442-26077-6. LCCN QC20 .M36 1985.

Wightman:1986:PWI

[Wig86a] A. Wightman. E. P. Wigner: an introduction. In Greenberger [Gre86], pages xv–xvii. ISBN 0-89766-356-X. LCCN ????

Wigner:1986:NNP

[Wig86b] E. P. Wigner. The nonrelativistic nature of the present quantum mechanical measurement theory. Annals of the New York Academy of Sciences, 480(1):1–5, December 1986. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic).

Wigner:1986:R

[Wig86c] E. P. Wigner. Response. Annals of the New York Academy of Sciences, 480(1), December 1986. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic).

Wigner:1986:SPO

[Wig86d] E. P. Wigner. Some problems of our natural sciences. International Journal of Theoretical Physics, 25(5):467–476, ???? 1986. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575

(electronic). URL http://www.springerlink.com/content/
g711q41001620321.

Wigner:1986:MBE

[Wig86e] Eugene P. Wigner. Is mankind bright enough to survive? The World and I, 267(??):??, February 1986. CODEN ???? ISSN ????

Wigner:1986:NRN

[Wig86f] Eugene P. Wigner. The non-relativistic nature of the present quantum mechanical measurement theory. Annals of the New York Academy of Sciences, 480(??):1–5, ???? 1986. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic).

Wigner:1987:GPD

[Wig87a] E. P. Wigner. The general properties of the distribution function and remarks on its weakness. In Y. S. Kim and W. W. Zachary, editors, The Physics of Phase Space: Nonlinear Dynamics and Chaos, Geometric Quantization, and Wigner Function: Proceedings of the First International Conference on the Physics of Phase Space, Held at the University of Maryland, College Park, Maryland, May 20–23, 1986, volume 278 of Lecture Notes in Physics, pages 161–170. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 3-540-17894-5. ISSN 0075-8450 (print), 1616-6361 (electronic). URL http://

www.springerlink.com/content/y7qp023qh0064003.

Wigner:1987:ADM

[Wig87b] Eugene Wigner. Address delivered at Memorial Meeting in Tallahassee. In Taylor [Tay87], pages 40–42. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987. US$10.00. URL http://www.loc.

gov/catdir/enhancements/fy0745/87153334-d.html. Based

on the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Wigner:1987:RPD

[Wig87c] Eugene P. Wigner. Remembering Paul Dirac. In Kurşunoğlu and Wigner [KW87d], pages 57–65. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987. URL http://adsabs.

harvard.edu/abs/1987ragp.book.....D;   http://www.loc.
gov/catdir/description/cam023/86033409.html;   http://
www.loc.gov/catdir/toc/cam028/86033409.html.

Wigner:1988:RRP

[Wig88a] Eugene P. Wigner. Reflection on the role and purpose of science. International Journal on the Unity of the Sciences, 1(1): 5–11, ???? 1988. CODEN IJUSE5. ISSN 0896-2294.

Wigner:1985:RPA

[Wig88b] Eugene P. Wigner. Remembering Paul Adrien Maurice Dirac. In Zichichi [Zic88], pages 269–274. ISBN 0-306-43235-8. LCCN QC793.3.B5 I57 1985.

Wigner:1988:LER

[Wig88c] Wigner Jenő. … az ember érzi, hogy fontos dolgokat még nem tud: Beszélgetés Wigner Jenövel a József Attila Gimnáziumban. (Hungarian) [… people feel that things are not important to know: a discussion with Eugene Wigner at the József Attila High School]. Fizikai Szemle (Budapest), 38(5):180–183, ???? 1988. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1988:LNM

[Wig88d] Wigner Jenő. … nincs még átfogó elmélet [elöadás az ELTE ünnepi ülésén, 1987. november 17-én: Videofelvételröl átírta Marx György]. (Hungarian) [There is no comprehensive theory [Lecture at the celebration of Eötvös, 17 November 1987: Video recordings transcribed by George Marx]]. Fizikai Szemle (Budapest), 38(5): 174–175, ???? 1988. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1988:NMA

[Wig88e] Wigner Jenő. Nincs még átfogó elmélet. (Hungarian) [Still no comprehensive theory]. Fizikai Szemle (Budapest), 38(5):174– 175, May 1988. CODEN FISZA6. ISSN 0015-3257 (print),

1588-0540 (electronic). URL http://www.kfki.hu/fszemle/
archivum/fsz8805/tart8805.html.

Wigner:1989:URI

[Wig89a] E. Wigner. On unitary representations of the inhomogeneous Lorentz group. Nuclear Physics B, Proceedings Supplements, 6(1): 9–64, March 1989. CODEN NPBSE7. ISSN 0920-5632 (print),

1873-3832 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0920563289904027. Reprint of [Wig39b].

Wigner:1989:ALP

[Wig89b] Eugene P. Wigner. Accomplishments and limitations of physics. Nuclear Physics B, Proceedings Supplements, 6(1):7–8, March 1989. CODEN NPBSE7. ISSN 0920-5632 (print), 1873-3832

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0920563289904015. Spacetime symmetries (Col-

lege Park, MD, 1988).

Wigner:1989:SS

[Wig89c] Eugene P. Wigner. The scientist and society. In Bethe et al. [BDH^{+}89], pages 57–68. ISBN 9971-5-0937-7. LCCN QC71 .F74

1989. URL http://adsabs.harvard.edu/abs/1989liph.book.
....B.

Wigner:1990:ELT

[Wig90a] E. P. Wigner. Entropy and Lorentz transformations. Physics Letters A, 147(7):343–347, July 23, 1990. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0375960190905508.

Wigner:1990:UEM

[Wig90b] Eugene P. Wigner. The unreasonable effectiveness of mathematics in the natural sciences. In Mathematics and science, pages 291– 306. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1990.

Wigner:1991:UEM

[Wig91] Eugene P. Wigner. Unreasonable effectiveness of mathematics in the natural sciences. In Ferris and Fadiman [FF91], pages 526– 540. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Wigner:1992:MUP

[Wig92a] Eugene Wigner. Memoir of the uranium project. In The collected works of Eugene Paul Wigner: Part A, The scientific papers: Nuclear Energy [Wig92c], pages 41–?? ISBN 3-540-55343-6, 0-387-

55343-6. LCCN ???? URL http://www.gbv.de/dms/bowker/
toc/9780387553436.pdf.

Wigner:1992:UEM

[Wig92b] Eugene P. Wigner. The unreasonable effectiveness of mathematics in the natural sciences. In In the forest of symbols (Finnish), pages 264–283. Art House, Helsinki, Finland, 1992.

Wigner:1992:CWE

[Wig92c] Eugene Paul Wigner. The collected works of Eugene Paul Wigner: Part A, The scientific papers: Nuclear Energy, volume 5. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-55343-6, 0-387-55343-6. xiv + 808

pp. LCCN ???? URL http://www.gbv.de/dms/bowker/toc/
9780387553436.pdf.

Wigner:1992:FSC

[Wig92d] Wigner Jenő. A fizika szerepéről és céljáról. (Hungarian) [The role and purpose of physics]. Fizikai Szemle (Budapest), 42(11): 436–437, ???? 1992. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1992:SLA

[Wig92e] Wigner Jenő. Szilárd Leó [Angolból] ford. Tóth Eszter. (Hungarian) [Leo Szilard [English] rpm. Eszter Tóth]. Fizikai Szemle (Budapest), 42(11):406–407, ???? 1992. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1992:TFR

[Wig92f] Wigner Jenő. ‘Tudtam a fizikát.‘ [Riporter] John Walsh. [Angolból] ford. /Abonyi Iván/. (Hungarian) [‘I knew physics.‘ [Reporter] John Walsh. [English] rpm. Ivan Abonyi]. Természet Világa [Natural World], 123(10):464–467, ???? 1992. CODEN ???? ISSN ????

Wigner:1993:CWE

[Wig93] Eugene Paul Wigner. The collected works of Eugene Paul Wigner. Part A. The scientific papers. Vol. I. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-56560-4. xii + 717 pp. With a preface by Jagdish Mehra and Arthur S. Wightman. With a biographical sketch by Mehra, and annotation by Brian R. Judd and George W. Mackey. Edited by Wightman.

Wightman:1995:EPW

[Wig95a] A. S. Wightman. Eugene Paul Wigner (1902–1995). Notices of the American Mathematical Society, 42(7):769–771, July 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wigner:1995:CWE

[Wig95b] Eugene Paul Wigner. The collected works of Eugene Paul Wigner. Part B. Historical, philosophical, and socio-political papers. Vol. VI. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-56986-3. xii + 631 pp. Philosophical reflections and syntheses, Annotated by Gérard G. Emch, Edited and with a preface by Jagdish Mehra and Arthur S. Wightman.

Wigner:1995:KFS

[Wig95c] Wigner Jenő. A kezdõ feltételektõl a szimmetriákig. A természettörvényekrõl. lejegyezte fodor L. istván. (Hungarian) [The initial conditions of the ????. the laws of nature. transcribed by Stephen L. Fodor]. Élet és Tudomány [Life and Science], 50 (7):195–197, ???? 1995. CODEN ???? ISSN ????

Wigner:1995:ZFH

[Wig95d] Wigner Jenő. Üzenet a fiataloknak. (Hungarian) [Message to young people]. Fizikai Szemle (Budapest), 45(2):37–38, ???? 1995. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1996:ADM

[Wig96a] E. P. Wigner. Address delivered to the Memorial Meeting [for Paul Dirac]. In Wigner [Wig96m], pages 214–215. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:ABE

[Wig96b] E. P. Wigner. An appreciation on the 60th birthday of Edward Teller. In Wigner [Wig96m], pages 133–138. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:BJN

[Wig96c] E. P. Wigner. Biography of John von Neumann. In Wigner [Wig96m], pages 209–210. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:CRA

[Wig96d] E. P. Wigner. Concluding remarks (address at the Dirac Symposium). In Wigner [Wig96m], pages 195–196. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:EM

[Wig96e] E. P. Wigner. Einstein — A memoir. In Wigner [Wig96m], page 197. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:EF

[Wig96f] E. P. Wigner. Enrico Fermi (1901–1954). In Wigner [Wig96m], pages 115–119. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:EAE

[Wig96g] E. P. Wigner. Erinnerungen an Albert Einstein. (German) [Memories of Albert Einstein]. In Wigner [Wig96m], pages 198–200. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:JN

[Wig96h] E. P. Wigner. John von Neumann (1903–1957). In Wigner [Wig96m], pages 127–130. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:LS

[Wig96i] E. P. Wigner. Leo Szilard (1898–1964). In Wigner [Wig96m], pages 139–149. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:NLE

[Wig96j] E. P. Wigner. New light on Einstein letter — an interview with E. P. Wigner and A. M. Weinberg by D. Sundberg. In Wigner [Wig96m], pages 216–218. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:RPD

[Wig96k] E. P. Wigner. Remembering Paul Dirac. In Wigner [Wig96m], pages 219–230. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:TYK

[Wig96l] E. P. Wigner. Thirty years of knowing Albert Einstein. In Wigner [Wig96m], pages 201–208. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:CWE

[Wig96m] Eugene Paul Wigner, editor. The collected works of Eugene Paul Wigner. Part A. The scientific papers. Vol. II. Nuclear physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-56972-3. x + 574 pp. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1997:CWEc

[Wig97a] Eugene Paul Wigner. The collected works of Eugene Paul Wigner. Part A. The scientific papers. Vol. III. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-57293-7. xii + 576 pp. Part I. Particles and fields, Part II. Foundations of quantum mechanics, With a preface by Jagdish Mehra and Arthur S. Wightman, With annotations by Wightman and Abner Shimony, Edited by Wightman.

Wigner:1997:CWEd

[Wig97b] Eugene Paul Wigner. The collected works of Eugene Paul Wigner. Part A. The scientific papers. Vol. IV. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-56985-5. xii + 456 pp. Part I. Physical chemistry, Part II. Solid state physics, With a preface by Jagdish Mehra and Arthur S. Wightman, With annotations by Nandor Balazs and Walter Kohn, Edited by Wightman.

Wigner:1997:PRS

[Wig97c] Eugene Paul Wigner. Philosophical reflections and syntheses. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-63372-3. viii + 631 pp. Annotated and with an introduction by Gérard G. Emch. Edited by Jagdish Mehra and Arthur S. Wightman. Reprint of the 1995 original [ıt The collected works of Eugene Paul Wigner. Part B. Historical, philosophical and socio-political papers. Vol. VI, Springer, Berlin, 1995; MR1307827 (95m:01023)].

Wigner:1998:CWE

[Wig98] Eugene Paul Wigner. The collected works of Eugene Paul Wigner. Part B. Historical, philosophical and socio-political papers. Vol. VIII. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-57295-3. xii + 258 pp. Sociopolitical reflections and civil defense, With a preface by Jagdish Mehra and Arthur S. Wightman, With annotations by Conrad Chester, Edited by Mehra.

Wigner:1999:FP

[Wig99] Eugene P. Wigner. On the future of physics. Fizikai Szemle (Budapest), 49(5):174–175, ???? 1999. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:19xx:QMR

[Wigxx] Eugene P. Wigner. Are quantum mechanics and relativity theory consistent? pages 372–379. 19xx.

Wigner:2001:SDN

[Wig01a] Eugene P. Wigner. Safety device for a neutronic reactor. US Patent 6,226,341., May 1, 2001. URL http://www.google.com/ patents/US6226341. US Patent Application filed Mar 2, 1954.

Wigner:2001:CWE

[Wig01b] Eugene Paul Wigner. The collected works of Eugene Paul Wigner: Part B, Historical, philosophical, and socio-political papers: Historical and biographical reflections and syntheses, volume 7. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-57294-5. xii + 535 pp.

Wigner:2002:HLF

[Wig02a] Wigner Jenő. Hogyan lettem fizikus ? (1983. november 12. Magyar Club, Philadelphia). Melléklet: a visszaemlékezés hasonmáskiadása. (Hungarian) [How I became a physicist? (Hungarian Club, 12 November 1983, Philadelphia). Appendix: a facsimile edition of memoirs]. Magyar Tudomány [Hungarian Science], 47(11):1408–1412, ???? 2002. CODEN ???? ISSN ????

Wigner:2002:PML

[Wig02b] Wigner Jenő. Polányi Mihály élete, 1891. március 12.–1976. február 22., a Royal Society tagja (1944). (Hungarian) [Michael Polányi’s life, 12 March 1891–22 February 1976, a member of the Royal Society (1944)]. Polanyiana, 11(1–2):19–62, ???? 2002. CODEN ???? ISSN ????

Wigner:2002:WJE

[Wig02c] Wigner Jenő. Wigner Jenő emlékiratai Andrew Szanton lejegyzésében. (Hungarian) [Eugene Wigner, Memoirs, as told to Andrew Szanton]. Kairosz, Budapest, Hungary, 2002. ISBN ???? 336 pp. LCCN ????

Wigner:2007:RAA

[Wig07] Eugene P. Wigner. Roots of the atomic age. In Masters and Way [MW07], chapter 3, pages 11–15. ISBN 1-59558-227-4 (hardcover). LCCN UG1282.A8 O54 2007. URL http:/

/thenewpress.com/index.php?option=com_title&task=view_
title&metaproductid=1703;   http://www.loc.gov/catdir/
toc/ecip0718/2007020838.htm. Foreword by Niels Bohr. Intro-

duction by Arthur H. Compton. Reprint of [MW46].

Wigner:2009:LS

[Wig09] E. P. Wigner. The limits of science. Resonance, 14(10):1018– 1027, ???? 2009. CODEN RESOFE. ISSN 0971-8044 (print),

0973-712X (electronic). URL http://www.springerlink.com/
content/m881828214364335.

Winch:1976:IFW

[Win76] D. E. Winch. Integration formula for Wigner 3-j coefficients. Journal of Mathematical Physics, 17(7):1166–1170, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1166_s1.

Wigner:1976:OMP

[WJH76] Eugene P. Wigner, John Jewkes, and Rom Harré. Obituary: Michael Polányi (11 March 1891–22 February 1976). Nature, 261 (5555):83–84, May 6, 1976. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v261/n5555/pdf/261083a0.pdf.

Warren:1972:POI

[WK72] Russell E. Warren and William H. Klink. Position operators as “internal” symmetries. Journal of Mathematical Physics, 13 (3):312–316, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i3/p312_s1.

Wigner:1965:DQM

[WL65] E. P. Wigner and P. T. Landsberg. Does quantum mechanics exclude life? Nature, 205(4978):1307, March 27, 1965. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v205/
n4978/pdf/2051307a0.pdf.

Wigner:1972:CSI

[WL72] E. P. Wigner and David Lester. Cloud seeding. Science, 177 (4050):651, August 25, 1972. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1734734.

Wigner:1958:MSU

[WM58] Eugene P. Wigner and William A. McAdams. Method of separating uranium. US Patent 2,849,284., August 26, 1958. URL

http://www.google.com/patents/US2849284. US Patent Ap-

plication filed Mar 29, 1945.

Wigner:1995:FP

[WM95] Eugene P. Wigner and George Marx. On the future of physics. Acta Physica Hungarica A: Heavy Ion Physics, 1(2):87–90, ???? 1995. CODEN APHAC8. ISSN 1219-7580 (print), 1588-2675

(electronic). URL http://www.springerlink.com/content/

0576352584475568. English translation of [Wig88e].

Wigner:1995:PRS

[WMW95] Eugene Paul Wigner, Jagdish Mehra, and A. S. Wightman, editors. Philosophical Reflections and Syntheses. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-63372-3 (softcover), 3-540-56986-3 (hardcover). vii + 631 pp. LCCN QC21.2 .W49 1995. Annotated by Gérard G. Emch.

Wlodarz:1990:VCW

[WN90] Joachim J. Włodarz and Janusz Nowakowski. Variational calculations of the Wigner distribution function for selected anharmonic oscillators. International Journal of Quantum Chemistry, 38(2): 373–381, August 1990. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Witten:1984:HEP

[WNY^{+}84] E. Witten, Y. Nambu, C. N. Yang, G. Thooft, M. F. Atiyah, L. Susskind, M. K. Gaillard, Y. Neeman, S. Deser, C. Teitelboim, D. Gross, S. Adler, S. Weinberg, R. Brout, B. Dewitt, F. Englert, S. Fubini, T. Regge, E. P. Wigner, and J. A. Wheeler. Higher energy physics — what are the possibilities for extending our understanding of elementary-particles and their interactions to much greater energies? — Proceedings of the XVIIIth Solvay Conference on Physics at the University of Texas, Austin, Texas, USA, November 1982 — General Discussion. Physics Reports, 104(2–4): 201–236, ???? 1984. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270 (electronic).

Wigner:1958:NRS

[WO58] Eugene P. Wigner and Leo A. Ohlinger. Neutronic reactor shield and spacer construction. US Patent 2,861,034., November 18,

1958. URL http://www.google.com/patents/US2861034. US

Patent Application filed Sep 18, 1945.

Wolff:1967:TFY

[Wol67] Anthony Wolff. Twenty-five years with the bomb: A conversation with Nobel Prize-winner Eugene P. Wigner. Look Magazine, ?? (??):58, 60–61, December 26, 1967. URL http://library.ucsd.

edu/dc/object/bb5564714p.

Wolf:1971:RMC

[Wol71] Kurt Bernardo Wolf. Recursive method for the computation of the SO_{n}, SO_{n,1}, and ISO_{n}, representation matrix elements. Journal of Mathematical Physics, 12(2):197–206, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i2/
p197_s1.

Wolf:1972:RME

[Wol72] Kurt Bernardo Wolf. The U_{n,1} and IU_{n} representation matrix elements. Journal of Mathematical Physics, 13(10):1634–1638, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v13/i10/p1634_s1.

Wong:1971:UAT

[Won71] M. K. F. Wong. Unit adjoint tensor operators in SO(n). Journal of Mathematical Physics, 12(8):1530–1535, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i8/
p1530_s1.

Wong:1976:SMF

[Won76] M. K. F. Wong. On the structure of the multiplicity-free Wigner coefficients of U(n). Journal of Mathematical Physics, 17(8):1558– 1569, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i8/p1558_s1.

Wong:1979:MFW

[Won79] M. K. F. Wong. On the multiplicity-free Wigner and Racah coefficients of U(n). Journal of Mathematical Physics, 20(12): 2391–2397, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i12/p2391_s1.

Woolf:1980:SSP

[Woo80] Harry Woolf, editor. Some strangeness in the proportion: a centennial symposium to celebrate the achievements of Albert Einstein. Addison-Wesley, Reading, MA, USA, 1980. ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Wooten:1995:WER

[Woo95] J. W. Wooten. Eugene Wigner remembered. Physics Today, 48 (12):15, December 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Wigner:1957:ICD

[WOY57] Eugene P. Wigner, Leo A. Ohlinger, and Gale J. Young. Isotope conversion device. US Patent 2,815,321., December 3, 1957.

URL http://www.google.com/patents/US2815321. US Patent

Application filed Nov 13, 1945.

Wigner:1956:HWM

[WOYW56a] Eugene P. Wigner, Leo A. Ohlinger, Gale J. Young, and Alvin M. Weinberg. Heavy water moderated neutronic reactor. US Patent 2,770,591., November 13, 1956. URL http://www.google.com/ patents/US2770591. US Patent Application filed Jul 3, 1946.

Wigner:1956:R

[WOYW56b] Eugene P. Wigner, Leo A. Ohlinger, Gale J. Young, and Alvin M. Weinberg. Reactor. US Patent 2,736,696., February 28, 1956.

URL http://www.google.com/patents/US2736696. US Patent

Application filed Aug 29, 1945.

Wigner:1957:FMR

[WOYW57] Eugene P. Wigner, Leo A. Ohlinger, Gale J. Young, and Alvin M. Weinberg. Fluid moderated reactor. US Patent 2,810,689.,

October 22, 1957. URL http://www.google.com/patents/

US2810689. US Patent Application filed Nov 6, 1945.

Wigner:1959:MSC

[WOYW59] Eugene P. Wigner, Leo A. Ohlinger, Gale J. Young, and Alvin M. Weinberg. Means for shielding and cooling. US Patent 2,873,243.,

February 10, 1959. URL http://www.google.com/patents/

US2873243. US Patent Application filed Oct 8, 1946.

Wigner:1962:MAC

[WOYW62] Eugene P. Wigner, Leo A. Ohlinger, Gale J. Young, and Alvin M. Weinberg. Method and apparatus for conducting a nuclear chain reaction. US Patent 3,052,613., September 4, 1962. URL http:

//www.google.com/patents/US3052613. US Patent Application

filed Aug 29, 1945.

Wigner:1977:FVS

[WPG77] Eugene P. Wigner and E. Padányi-Gulyás. The future voction of science. Studies for a New Central Europe, ser. 4, ??(1–2):3–7, ???? 1977. CODEN ???? ISSN ????

Wigner:1933:CMS

[WS33] E. Wigner and F. Seitz. On the constitution of metallic sodium, I. Physical Review (2), 43(10):804–810, May 15, 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.43.804.

Wigner:1934:CMS

[WS34] E. Wigner and F. Seitz. On the constitution of metallic sodium. II. Physical Review (2), 46(6):509–524, September 15, 1934. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.46.509.

Wigner:1955:QTC

[WS55] Eugene P. Wigner and F. Seitz. Qualitative theory of cohesion in metals. In Frederick Seitz and David Turnbull, editors, Solid state physics: advances in research and applications, volume 51, pages 97–126. Academic Press, New York, USA, 1955. LCCN ????

Wigner:1992:REP

[WS92] Eugene Paul Wigner and Andrew Szanton. The recollections of Eugene P. Wigner as told to Andrew Szanton. Plenum Press, New York, NY, USA; London, UK, 1992. ISBN 0-306-44326-0. xxiv + 335 pp. LCCN QC16.W52 A3 1992.

Wigner:2003:REP

[WS03] Eugene Paul Wigner and Andrew Szanton. The recollections of Eugene P. Wigner as told to Andrew Szanton. Basic Books, New York, NY, USA, 2003. ISBN 0-7382-0886-8. xxiv + 335 pp. LCCN QC16.W52.

Wigner:1959:JFE

[WSC59] Eugene P. Wigner, Leo Szilard, and Edward C. Creutz. Jacketed fuel element. US Patent 2,872,401., February 3, 1959. URL http:

//www.google.com/patents/US2872401. US Patent Application

filed May 8, 1946, serial number 668,110.

Wigner:1961:MLI

[WSCF61] Eugene P. Wigner, Leo Szilard, Robert F. Christy, and Francis Lee Friedman. Massive leakage irradiator. US Patent 2,986,510., May

30, 1961. URL http://www.google.com/patents/US2986510.

US Patent Application filed May 14, 1946, serial number 669,524.

Wigner:1979:ANB

[WT79] Eugene P. Wigner and Edward Teller. America needs a better civil defense program: Congressional Record Senate, August 2, 1979. In USA Congressional Record. Proc. and Debates of the 96th Congress, First Session, page ?? US Government Printing Office, Washington, DC, USA, 1979.

Wu:1972:SWC

[Wu72] Alfred C. T. Wu. Structure of the Wigner 9-j coefficients in the Bargmann approach. Journal of Mathematical Physics, 13(1):84– 90, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i1/p84_s1.

Wu:1973:SAJ

[Wu73] Alfred C. T. Wu. Structure of Alisauskas–Jucys form of the 9-j coefficients. Journal of Mathematical Physics, 14(9):1222–1223, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i9/p1222_s1.

Wigner:1975:NRO

[WU75] Eugene P. Wigner and Harold C. Urey. Nuclear reactors offer the surest, safest way of meeting energy needs: Pro and con. U.S. Information Service, 9(3):??, ???? 1975. CODEN ???? ISSN ????

Wigner:1954:SLT

[WvN54] Eugene P. Wigner and John von Neumann. Significance of Loewner’s Theorem in the quantum theory of collisions. Annals of Mathematics (2), 59(3):418–433, May 1954. CODEN AN-MAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL

http://www.jstor.org/stable/1969710. Reprinted in [Tau62,

Paper 27].

Wigner:1928:SZM

[WW28] Eugene Wigner and Enos E. Witmer. Über die Struktur der zweiatomigen Molekelspektren nach der Quantenmechanik. (German) [On the structure of diatomic molecule spektra according to quantum mechanics]. Zeitschrift für Physik, 51(11–12):859–886, ???? 1928. CODEN ZEPYAA. ISSN ???? URL http://www.

springerlink.com/content/gw1k173441v71jw2.

Weisskopf:1930:NLS

[WW30a] V. Weisskopf and E. Wigner. Über die natürliche Linienbreite in der Strahlung des harmonischen Oszillators. (German) [On the natural linewidth of the radiation of the harmonic oscillator]. Zeitschrift für Physik, 65(1–2):18–29, ???? 1930. CODEN

ZEPYAA. ISSN ???? URL http://www.springerlink.com/
content/n235152287079172.

Weisskopf:1930:BNL

[WW30b] Victor Weisskopf and Eugene P. Wigner. Berechnung der natürlichen Linienbreite auf Grund der Diracschen Lichttheorie. (German) [Calculation of the natural linewidth on the basis of the Dirac theory of light]. Zeitschrift für Physik, 63(1–2):54–73, ???? 1930. CODEN ZEPYAA. ISSN ???? URL http://www.

springerlink.com/content/l33q5205r32537v7.

Way:1946:RFPa

[WW46a] K. Way and E. P. Wigner. Radiation from fission products. Report, Technical Information Division, Atomic Energy Commission, Oak Ridge, TN, USA, 1946. 4 pp. URL http://catalog.

hathitrust.org/Record/007841261.

Way:1946:RFPb

[WW46b] K. Way and E. P. Wigner. Radiation from fission products. Physical Review (2), 70(1–2):115, July 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v70/i1-2/p99_1.

Way:1948:RDF

[WW48] K. Way and E. P. Wigner. The rate of decay of fission products. Physical Review (2), 73(11):1318–1330, June 1, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.73.1318.

Way:1951:RDF

[WW51] K. Way and E. Wigner. Rate of decay of fission products. In Coryell and Sugarman [CS51], page ?? LCCN QD601 .C65. With special editorial assistance from R.A. Brightsen and others.

Weinberg:1958:PTN

[WW58] Alvin Martin Weinberg and Eugene Paul Wigner. The physical theory of neutron chain reactors. University of Chicago Press, Chicago, IL, USA and London, UK, 1958. xii + 801 pp. LCCN QC787.N8 W4.

Weinberg:1960:LRV

[WW60] Alvin M. Weinberg and Eugene P. Wigner. Longer range view of nuclear energy. Bulletin of the Atomic Scientists, 16(10):400– 403, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mei61].

Weinberg:1961:LEW

[WW61a] Alvin M. Weinberg and Eugene P. Wigner. Letter to the Editor: Why reactor development is economical. Bulletin of the Atomic Scientists, 17(4):120, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mei61].

Wigner:1961:MPP

[WW61b] Eugene P. Wigner and Alvin M. Weinberg. Means for producing plutonium chain reactions. US Patent 2,969,311., January 24,

1961. URL http://www.google.com/patents/US2969311. US

Patent Application filed Oct 8, 1946.

WheelerWigner:1962:SHIa

[WW62a] M. Wheeler Wigner and E. Wigner. The scientist: his increased responsibilities. Hungarian Quarterly, ??(??):39–??, April/July 1962. CODEN ???? ISSN ????

WheelerWigner:1962:SHIb

[WW62b] M. Wheeler Wigner and E. Wigner. The scientist: his increased responsibilities. Wellesley Alumnae Magazine, 46(??):80–??, January 1962. CODEN ???? ISSN ????

Wigner:1984:VBS

[WW84] Eugene P. Wigner and A. M. Weinberg. Values and benefits of science. Journal of the International Cultural Foundation, 2(4): ??, ???? 1984. CODEN ???? ISSN ????

Wigner:1986:NLE

[WW86] Eugene P. Wigner and A. M. Weinberg. New light on the Einstein letter. The Oak Ridger, 37(1):7–??, ???? 1986. CODEN ???? ISSN 0890-6009.

Weimar-Woods:1995:CLA

[WW95] Evelyn Weimar-Woods. Contractions of Lie algebras: generalized Inönü-Wigner contractions versus graded contractions. Journal of Mathematical Physics, 36(8):4519–4548, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Szilard:1942:ABC

[WWCS42] A. M. Weinberg, E. P. Wigner, R. F. Christy, and Leo Szilard. Approximate boundary conditions for diffusion equation at interface between two media. Report CP-189, US Atomic Energy Commission, Washington, DC, USA, July 10, 1942.

Wigner:1992:CWExx

[WWM98] Eugene Paul Wigner, Arthur S. Wightman, and Jagdish Mehra. The collected works of Eugene Paul Wigner. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992–1998. ISBN 3-540-55343-6 (Berlin: vol. 5), 0-387-55343-6 (New York: vol. 5), 3-540-57295-3 (Berlin: vol. 8). various pp.

LCCN QC19.3 .W54 1992. URL http://www.loc.gov/catdir/
enhancements/fy0818/92038376-d.html;   http://www.loc.
gov/catdir/enhancements/fy0818/92038376-t.html.

Wick:1952:IPE

[WWW52] G. C. Wick, Arthur S. Wightman, and E. P. Wigner. The intrinsic parity of elementary particles. Physical Review (2), 88 (1):101–105, October 1, 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.88.101.

Wick:1964:ERS

[WWW64] G. C. Wick, A. S. Wightman, and Wigner Jenő. Az elemi részek sajátparitása. (Hungarian) [The parity of elementary particles]. Magyar Fizikai Folyóirat, 12(3):327–335, ???? 1964. CODEN MGFFAC. ISSN 0025-0104.

Wick:1970:SRC

[WWW70] Gian-Carlo Wick, Arthur S. Wightman, and Eugene P. Wigner. Superselection rule for charge. Physical Review D (Particles and Fields), 1(12):3267–3269, June 15, 1970. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.1.
3267.

Wigner:1958:NRa

[WWY58a] Eugene P. Wigner, Alvin M. Weinberg, and Gale J. Young. Neutronic reactor. US Patent 2,830,944., April 15, 1958. URL http:

//www.google.com/patents/US2830944. US Patent Application

filed Aug 28, 1945.

Wigner:1958:SFM

[WWY58b] Eugene P. Wigner, Robert R. Williamson, and Gale J. Young. Shaped fissionable metal bodies. US Patent 2,856,340., October

14, 1958. URL http://www.google.com/patents/US2856340.

US Patent Application filed Jun 15, 1945.

Wigner:1958:RSD

[WY58] Eugene P. Wigner and Gale J. Young. Radiation shielding device. US Patent 2,853,624., September 23, 1958. URL http://www.

google.com/patents/US2853624. US Patent Application filed

May 22, 1945.

Wigner:1960:NRF

[WY60] Eugene P. Wigner and Gale J. Young. Nuclear reactor fuel charging/discharging system. Canadian Patent 604569., September

6, 1960. URL http://brevets-patents.ic.gc.ca/opic-cipo/
cpd/eng/patent/604569/summary.html.

Wigner:1963:ICD

[WY63] Eugene P. Wigner and Mutsuo M. Yanase. Information contents of distributions. Proceedings of the National Academy of Sciences of the United States of America, 49(6):910–918, June 15, 1963. CO-DEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

URL http://www.jstor.org/stable/71797.

Wigner:1964:PSN

[WY64] Eugene P. Wigner and Mutsuo M. Yanase. On the positive semidefinite nature of a certain matrix expression. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):397–406, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wigner:1973:AQM

[WY73] Eugene P. Wigner and M. M. Yanase. Analysis of the quantum mechanical measurement process. Annals of the Japan Association for the Philosophy of Science, 4(??):171–186, ???? 1973. CODEN ???? ISSN 0453-0691.

Yang:1974:RQS

[wYDK74] Janilla Jyy wen Yang, Russell Davidson, and John J. Kozak. On the relaxation to quantum-statistical equilibrium of the Wigner– Weisskopf atom in a one-dimensional radiation field. VI. Influence of the coupling function on the dynamics. Journal of Mathematical Physics, 15(4):491–501, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v15/i4/p491_s1.

Wigner:1961:MAC

[WYW61] Eugene P. Wigner, Gale J. Young, and Alvin M. Weinberg. Method and apparatus for controlling neutron density. US Patent 2,990,357., June 27, 1961. URL http://www.google.com/ patents/US2990357. US Patent Application filed May 14, 1946.

Wheeler:1983:QTM

[WZ83] John Archibald Wheeler and Wojciech Hubert Zurek, editors. Quantum theory and measurement. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1983. ISBN 0-691-08315-0 , 0-691-08316-9 (paperback). xxviii + 811 pp. LCCN QC174.125 .Q38 1983. US$60.00; US$19.50.

Yanagi:2010:URG

[Yan10] Kenjiro Yanagi. Uncertainty relation on generalized Wigner– Yanase–Dyson skew information. Linear Algebra and its Applications, 433(8–10):1524–1532, December 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zachary:1984:XIC

[Zac84] W. W. Zachary, editor. XIIIth International Colloquium on Group Theoretical Methods in Physics: College Park, Maryland, USA, 21–25 May 1984. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1984. ISBN 9971-966-87-5, 9971-966-88-3 (paperback). xvi + 620 pp. LCCN QC20.7.G76 I57 1984.

Zichichi:1984:SSS

[ZDF^{+}84] A. Zichichi, Paul A. M. Dirac, S. Ferrara, H. Kleinert, Andre Martin, Eugene P. Wigner, and Chen-Ning Yang. Special session on symmetries and gauge invariance. In Zichichi [Zic84], pages 725–745. ISBN 0-306-41738-3. LCCN QC793.3.F5; 19845. URL

http://usparc.ihep.su/spires/find/hep/www?irn=1397141.

Zhang:2003:WMS

[Zha03] Ping Zhang. Wigner measure and the semiclassical limit of Schrödinger–Poisson equations. SIAM Journal on Mathematical Analysis, 34(3):700–718, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://epubs.siam.

org/sam-bin/dbq/article/39340.

Zucker:1960:RBC

[ZHH60] Alexander Zucker, Frederick T. Howard, and Edith C. Halbert, editors. Reactions between complex nuclei: proceedings of the Second Conference on Reactions Between Complex Nuclei, May 2–4, 1960, Gatlinburg, Tennessee. Wiley, New York, NY, USA, 1960. LCCN QC793.9 .C66 1960.

Zichichi:1973:HPP

[Zic73] Antonino Zichichi, editor. Highlights in particle physics, volume 10 of Subnuclear series. Compositori, Bologna, Italy, 1973. xxiii + 879 pp. LCCN QC793 .I555 1972.

Zichichi:1978:UFC

[Zic78] Antonino Zichichi, editor. Understanding the fundamental constituents of matter: proceedings of the 1976 International School of Subnuclear Physics (NATO-MPI-MRST Advanced Study Institute) held in Erice, Trapani, Sicily, July 23–August 8, 1976 and sponsored by the Sicilian Regional Government and the Weizmann Institute of Science, volume 14 of The Subnuclear series. Plenum Press, New York, NY, USA; London, UK, 1978. ISBN 0-306-38183-4. LCCN QC793 .I555 1976.

Zichichi:1983:HEL

[Zic83a] Antonino Zichichi, editor. The high-energy limit: Proceedings of the Eighteenth Course of the International School of Subnuclear Physics, held July 31–August 11, 1980, in Erice, Trapani, Sicily, volume 18 of The Subnuclear series. Plenum Press, New York, NY, USA; London, UK, 1983. ISBN 0-306-41036-2. LCCN QC793 .I555 1980.

Zichichi:1983:UFI

[Zic83b] Antonino Zichichi, editor. The unity of the fundamental interactions: [proceedings of the nineteenth course of the International School of Subnuclear Physics, held July 31–August 11, 1981, in Erice, Trapani, Sicily], volume 19 of The subnuclear series.

Plenum Press, New York, NY, USA; London, UK, 1983. ISBN 0-306-41242-X. LCCN ????

Zichichi:1984:GIT

[Zic84] Antonino Zichichi, editor. Gauge interactions: theory and experiment, volume 20 of Subnuclear series. Plenum Press, New York, NY, USA; London, UK, 1984. ISBN 0-306-41738-3. LCCN QC793.3.F5; 19845.

Zichichi:1985:HFW

[Zic85] Antonino Zichichi, editor. How far are we from the gauge forces: Proceedings of the Twenty-first Course of the International School of Subnuclear Physics, held August 3–14, 1983, in Erice, Trapani, Sicily, Italy, volume 21 of The Subnuclear series. Plenum Press, New York, NY, USA; London, UK, 1985. ISBN 0-306-42203-4. LCCN QC793 .I555 1983. Proceedings of the Twenty-first Course of the International School of Subnuclear Physics, held August 3–14, 1983, in Erice, Trapani, Sicily, Italy.

Zichichi:1988:ONF

[Zic88] Antonino Zichichi, editor. Old and new forces of nature: Proceedings of the twenty-third course of the International School of Subnuclear Physics on Old and new forces of nature, held August 4–14, 1985, in Erice, Sicily, Italy, volume 23 of The Subnuclear series. Plenum Press, New York, NY, USA; London, UK, 1988. ISBN 0-306-43235-8. LCCN QC793.3.B5 I57 1985.

Zielinski:1988:SIW

[Zie88] Tomasz P. Zielinski. On a software implementation of the Wigner– Ville transform. Computer Physics Communications, 50(1–2): 269–272, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465588901361.

Zsadanyi:1976:PWJ

[Zsa76] Zsadanyi Oszkar. Párbeszéd Wigner Jenő Nobel-díjas professzorral. (Hungarian) [Dialogue with Nobel laureate Professor Eugene Wigner]. ????, ??(??):??, September 1, 1976. URL http:/

/library.ucsd.edu/dc/object/bb7851423b.

Zicovich-Wilson:2012:BWT

[ZWE12] Claudio M. Zicovich-Wilson and Alessandro Erba. Beyond Wigner’s theorems: the role of symmetry equivalences in quantum systems. International Journal of Quantum Chemistry, 112(21): 3543–3551, November 5, 2012. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zeng:1988:WOR

[ZY88] Gao Jian Zeng and Xiou Hua Yuan. Wigner operator and Racah operator of the Lie superalgebra OSp(1, 2). Journal of Mathematical Physics, 29(12):2553–2562, December 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i12/p2553_s1.